### v1

# Variable Elliptic Structures on the Plane: Transport Dynamics, Rigidity, and Function Theory

Variable Elliptic Structures on the Plane:
Transport Dynamics, Rigidity, and Function Theory
Daniel Alayón-Solarz Email: danieldaniel@gmail.com
(January 2026)
A fixed complex structure gives analysis.
A moving one gives transport.
Rigidity is where the two agree to dance.
(“Shut up and calculate!” — David Mermin)
Introduction
In the classical theory of complex analysis, the imaginary unit $i$ is treated as a fixed algebraic constant. All analytic constructions are built on this assumption, and the underlying complex structure of the plane is taken to be rigid and homogeneous. From a geometric point of view, however, the symbol $i$ encodes the choice of a complex structure, and there is no intrinsic reason for this structure to remain constant once the geometry of the domain is allowed to vary.
The present work adopts the viewpoint that the imaginary unit should be treated as intrinsic data: a moving generator $i(x,y)$ of a rank–two real algebra bundle over the plane. This generator is defined by the quadratic structure reduction
$i(x,y)^{2}+\beta(x,y)\,i(x,y)+\alpha(x,y)=0,$
whose coefficients $\alpha$ and $\beta$ are allowed to vary with position, subject only to an ellipticity condition. In this setting, the algebra itself becomes part of the geometry.
Once this perspective is adopted, a basic and unavoidable question arises: what does it mean to differentiate the imaginary unit? Unlike the classical theory, where $i$ is constant by definition, the variable framework forces the derivatives $i_{x}$ and $i_{y}$ to appear. These derivatives are not auxiliary objects: they are well–defined algebra–valued sections whose algebraic resolution is determined entirely by the structure polynomial. In particular, the differentiation of $i$ is meaningful before any notion of holomorphicity or Cauchy–Riemann equation is introduced.
A central quantity emerging from this differentiation is the intrinsic combination
$i_{x}+i\,i_{y}.$
This expression appears universally once the structure varies. It measures the incompatibility between differentiation in the base variables and multiplication in the fibers. As a consequence, it simultaneously governs the inhomogeneity of generalized Cauchy–Riemann systems and the failure of the Leibniz rule for the natural first–order operators associated with the variable structure. These phenomena are not imposed by analytic assumptions; they are forced by the geometry of the moving algebra itself.
A first major outcome of the theory is that the variability of the structure coefficients $(\alpha,\beta)$ is governed by a transport law that is intrinsic and unavoidable. By differentiating the structure reduction and eliminating the derivatives of $i$, the evolution of $(\alpha,\beta)$ can be expressed in terms of the spectral parameter
$\lambda=\frac{-\beta+i\sqrt{4\alpha-\beta^{2}}}{2},\qquad\Im\lambda>0.$
In full generality, this leads to a (possibly forced) complex inviscid Burgers equation written in the standard complex algebra. This transport equation belongs to the structural layer of the theory: it arises solely from allowing the elliptic structure to vary and does not depend on any analytic integrability assumptions.
Within this universal transport framework, rigidity occupies a precise and limited role. The vanishing condition
$i_{x}+i\,i_{y}=0$
selects exactly the conservative case of the transport law. In this regime, the Burgers equation becomes unforced, and several independent obstructions collapse simultaneously: the generalized Cauchy–Riemann operator becomes a derivation, the real Cauchy–Riemann system becomes homogeneous, and the generator $i(x,y)$ is transported compatibly with its own multiplication law. Rigidity is therefore not the source of the coefficient dynamics, but a compatibility condition that restores integrability within an already dynamical setting.
The explicit $\varepsilon$–family of rigid elliptic structures constructed in this work provides concrete examples of this conservative regime. These examples show that rigidity does not force triviality: genuinely nonconstant structures exist, although ellipticity naturally restricts the domain of definition. From the transport viewpoint, these restrictions reflect characteristic phenomena familiar from first–order hyperbolic systems.
On the analytic side, rigidity is exactly the condition under which a coherent function theory can be recovered. In the rigid regime, a Cauchy–Pompeiu formula adapted to variable elliptic structures persists. The natural Cauchy 11–form
$d\tilde{z}=dy-i(z)\,dx$
is not closed, and its non–closure produces an unavoidable interior correction term. Rather than being an artifact, this correction identifies a canonical covariant first–order operator whose null space provides the appropriate notion of holomorphicity in the variable setting. Although covariantly holomorphic sections are not closed under pointwise multiplication, a scalar gauge transformation restores closure via a natural weighted product.
The scope of this monograph is deliberately local and explicit. Although many of the results admit reinterpretation within more abstract frameworks, we work at the level of concrete first–order computations in order to keep the geometric and dynamical mechanisms fully visible. From this perspective, variable elliptic structures reveal a natural hierarchy: transport first, integrability second, and analytic closure as a distinguished limiting case.
The investigation of variable elliptic structures was originally suggested to the author by W. Tutschke and C. J. Vanegas, whose work on parameter–depending Clifford Algebras and generalized Cauchy–Pompeiu formulas provided both motivation and guidance.
Although W. Tutschke did not see the full development of the theory presented here, his influence is present throughout in the guiding questions and structural viewpoint.
Dedicated to the memory of W. Tutschke.
Chapter 1 Variable Elliptic Structures on the Plane: Definitions
1.1 Structure Polynomial and Ellipticity
Let $\Omega\subset\mathbb{R}^{2}$ be an open domain with coordinates $(x,y)$. Let $\alpha,\beta\in C^{1}(\Omega,\mathbb{R})$ satisfy the ellipticity condition
$\Delta(x,y):=4\alpha(x,y)-\beta(x,y)^{2}>0\quad\text{for all }(x,y)\in\Omega.$ (1.1)
Definition 1.1 (Variable elliptic structure).
A variable elliptic structure on $\Omega$ is the real rank–two algebra bundle $\{A_{z}\}_{z\in\Omega}$ generated by an element $i=i(x,y)$ satisfying the variable-coefficient structure polynomial
$X^{2}+\beta(x,y)\,X+\alpha(x,y)$ (1.2)
so we can always apply the structure reduction
$i^{2}+\beta(x,y)\,i+\alpha(x,y)=0.$ (1.3)
Each fiber $A_{z}$ is canonically identified with $\mathbb{R}[X]/(X^{2}+\beta(z)X+\alpha(z))$. By (1.1), this algebra is elliptic and hence isomorphic to $\mathbb{C}$ as a real algebra, though no canonical identification between fibers is assumed.
We denote the conjugate root by
$\hat{i}(z):=-\beta(z)-i(z).$
1.2 Algebra-Valued Functions
A function $f:\Omega\to A$ is written uniquely as
$f=u+v\,i,\qquad u,v\in C^{1}(\Omega,\mathbb{R}).$
Addition and multiplication are taken pointwise using the algebra structure induced by (1.3).
1.3 Differentiating the generator
A first concern in any variable-coefficient theory is whether the moving generator $i=i(x,y)$ can be differentiated in a meaningful and computable way. In our setting $i$ is a $C^{1}$ section of the algebra bundle, hence its partial derivatives $i_{x},i_{y}$ are well-defined as sections as well. In the elliptic regime, these derivatives can be isolated explicitly from the structure reduction. More precisely, the computation relies only on the invertibility of the element $2i+\beta$ in the underlying algebra. This condition is satisfied not only in the elliptic case but also in the hyperbolic and parabolic regimes, away from zero divisors. Consequently, the algebraic resolution of $i_{x}$ and $i_{y}$ from the structure reduction can be performed verbatim in all three cases; what distinguishes the elliptic regime is not the computation itself but the positivity and invertibility properties that follow from it.
1.3.1 Invertibility of $2i+\beta$ in the elliptic regime
Recall that $i$ satisfies $i^{2}+\beta i+\alpha=0$ with $\Delta:=4\alpha-\beta^{2}>0$. A direct computation using $i^{2}=-\beta i-\alpha$ gives
$(2i+\beta)(-\beta-2i)=\Delta.$
Thus $2i+\beta$ is invertible in every fiber and
$(2i+\beta)^{-1}=\frac{-\beta-2i}{\Delta}.$ (1.4)
1.3.2 Coefficient decomposition
Since each elliptic fiber is two-dimensional over $\mathbb{R}$ with basis $\{1,i\}$, every section admits a unique decomposition in this basis. In particular, there exist unique real functions $A_{x},B_{x},A_{y},B_{y}$ on $\Omega$ such that
$i_{x}=A_{x}+B_{x}\,i,\qquad i_{y}=A_{y}+B_{y}\,i.$ (1.5)
Moreover, (1.7) shows that $A_{x},B_{x},A_{y},B_{y}$ depend only on $\alpha,\beta$ and their first derivatives.
1.3.3 Solving for $i_{x}$ and $i_{y}$
Differentiate the structure relation $i^{2}+\beta i+\alpha=0$ with respect to $x$ and $y$:
$(2i+\beta)i_{x}+\alpha_{x}+\beta_{x}i=0,\qquad(2i+\beta)i_{y}+\alpha_{y}+\beta_{y}i=0.$ (1.6)
Multiplying by $(2i+\beta)^{-1}$ and using (1.4) yields the explicit formulas
$i_{x}=-\frac{\alpha_{x}+\beta_{x}i}{2i+\beta},\qquad i_{y}=-\frac{\alpha_{y}+\beta_{y}i}{2i+\beta}.$ (1.7)
1.4 Generalized Cauchy–Riemann Operators
Definition 1.2 (Generalized Cauchy–Riemann operators).
We define
$\partial_{\bar{z}}:=\frac{1}{2}\bigl(\partial_{x}+i\,\partial_{y}\bigr),\qquad\partial_{z}:=\frac{1}{2}\bigl(\partial_{x}+\hat{i}\,\partial_{y}\bigr).$ (1.8)
These operators act on algebra–valued functions using the product in $A_{z}$. No derivation property is assumed.
1.5 The Variable Cauchy–Riemann System
Let $f=u+vi$ with $u,v\in C^{1}(\Omega,\mathbb{R})$. A direct computation yields the identity
$2\,\partial_{\bar{z}}f=(u_{x}-\alpha v_{y})+(v_{x}+u_{y}-\beta v_{y})\,i+v\,(i_{x}+i\,i_{y}).$ (1.9)
Since each fiber is elliptic, $\{1,i\}$ is a real basis of $A_{z}$. Hence there exist unique real functions $A(x,y)$ and $B(x,y)$ such that
$i_{x}+i\,i_{y}=A+B\,i.$ (1.10)
Substituting (1.10) into (1.9), the equation $\partial_{\bar{z}}f=0$ is equivalent to the real first–order system
$\begin{cases}u_{x}-\alpha v_{y}+Av=0,\\[4.0pt] v_{x}+u_{y}-\beta v_{y}+Bv=0.\end{cases}$ (1.11)
Remark 1.3.
The coefficients $A$ and $B$ encode all deviations from the classical Cauchy–Riemann equations. At this level they are simply defined by (1.10), with no structural assumptions imposed.
1.6 Explicit Formula for the Inhomogeneous Coefficients
Differentiating the structure reduction (1.3) gives
$(2i+\beta)i_{x}+\alpha_{x}+\beta_{x}i=0,\qquad(2i+\beta)i_{y}+\alpha_{y}+\beta_{y}i=0.$
Using ellipticity to invert $2i+\beta$, one obtains
$i_{x}+i\,i_{y}=\frac{-(\alpha_{x}-\alpha\beta_{y})-(\beta_{x}+\alpha_{y}-\beta\beta_{y})i}{2i+\beta}.$
Since $(2i+\beta)^{-1}=(-\beta-2i)/\Delta$, a direct reduction yields
$A$ $=\frac{\beta(\alpha_{x}-\alpha\beta_{y})-2\alpha(\beta_{x}+\alpha_{y}-\beta\beta_{y})}{\Delta},$ (1.12)
$B$ $=\frac{2(\alpha_{x}-\alpha\beta_{y})-\beta(\beta_{x}+\alpha_{y}-\beta\beta_{y})}{\Delta}.$
Thus the system (1.11) depends only on $\alpha$, $\beta$ and their first derivatives.
1.7 Product Rule and Its Defect
Let $f=u+vi$ and $g=p+qi$ be $C^{1}$ sections. Define the Leibniz defect
$\mathcal{N}(f,g):=\partial_{\bar{z}}(fg)-(\partial_{\bar{z}}f)g-f(\partial_{\bar{z}}g).$
A direct computation shows
$\mathcal{N}(f,g)=\frac{1}{2}\,vq\,(i_{x}+i\,i_{y}).$ (1.13)
Remark 1.4.
Formula (1.13) is purely algebraic and holds without further assumptions. It shows explicitly how the variability of the generator $i(x,y)$ obstructs the derivation property of $\partial_{\bar{z}}$.
Chapter 2 Differentiating the Structure: Universal Burgers Transport
2.1 Purpose of this Chapter
The previous chapters introduced variable elliptic structures and derived the basic algebraic and differential identities that follow from the structure polynomial
$i^{2}+\beta(x,y)\,i+\alpha(x,y)=0.$
Up to this point, no integrability or compatibility assumptions have been imposed.
The purpose of the present chapter is to describe what is unavoidable once the generator $i=i(x,y)$ is allowed to vary. In particular, we identify the intrinsic quantity that governs all deviations from the classical theory and show that the coefficients of the structure polynomial are necessarily subject to a transport law.
2.2 The Intrinsic Obstruction
Let $\Omega\subset\mathbb{R}^{2}$ be open, and let $\alpha,\beta\in C^{1}(\Omega,\mathbb{R})$ satisfy the ellipticity condition
$4\alpha-\beta^{2}>0.$
Fix a $C^{1}$ choice of generator $i:\Omega\to A$ solving
$i^{2}+\beta i+\alpha=0.$
Definition 2.1 (Intrinsic obstruction).
The intrinsic obstruction associated with the generator $i$ is the AA–valued section
$\mathcal{G}:=i_{x}+i\,i_{y}.$ (2.1)
This quantity is well defined wherever $i$ is differentiable. It does not depend on any analytic structure or auxiliary operator. All deviations from the constant theory enter through $\mathcal{G}$.
Writing $\mathcal{G}$ in the moving basis,
$\mathcal{G}=G_{0}+G_{1}\,i,\qquad G_{0},G_{1}\in C^{0}(\Omega,\mathbb{R}),$ (2.2)
makes explicit the dependence on the chosen branch of the generator. The pair $(G_{0},G_{1})$ should be regarded as intrinsic coefficient data attached to the variable structure.
2.3 Eliminating the Derivatives of the Generator
Differentiating the structure reduction with respect to $x$ and $y$ gives
$(2i+\beta)i_{x}+\alpha_{x}+\beta_{x}i=0,\qquad(2i+\beta)i_{y}+\alpha_{y}+\beta_{y}i=0.$
Adding the second equation multiplied by $i$ to the first and using $i^{2}=-\beta i-\alpha$ yields the identity
$(2i+\beta)\,(i_{x}+i\,i_{y})=(-\alpha_{x}+\alpha\beta_{y})+(-\beta_{x}-\alpha_{y}+\beta\beta_{y})\,i.$ (2.3)
Substituting $\mathcal{G}=G_{0}+G_{1}i$ into (2.3) and reducing in the basis $\{1,i\}$ gives the following first–order system for the coefficients.
Proposition 2.2 (Forced coefficient system).
Let $\mathcal{G}=G_{0}+G_{1}i$ be the intrinsic obstruction associated with a fixed generator $i$. Then the coefficients $\alpha$ and $\beta$ satisfy
$\begin{cases}\alpha_{x}=\alpha\,\beta_{y}-\beta\,G_{0}+2\alpha\,G_{1},\\[6.0pt] \beta_{x}+\alpha_{y}=\beta\,\beta_{y}-2G_{0}+\beta\,G_{1}.\end{cases}$ (2.4)
Proof.
Using $i^{2}=-\beta i-\alpha$, one computes
$(2i+\beta)(G_{0}+G_{1}i)=(\beta G_{0}-2\alpha G_{1})+(2G_{0}-\beta G_{1})\,i.$
Equating coefficients with the right–hand side of (2.3) yields (2.4). ∎
This system is valid for every variable elliptic structure. No compatibility or integrability assumption has been used.
2.4 Canonical Spectral Parameter
Under ellipticity, write
$\beta=-2a,\qquad\alpha=a^{2}+b^{2},\qquad b>0,$
and define the complex–valued spectral parameter
$\lambda:=a+ib=\frac{-\beta+i\sqrt{4\alpha-\beta^{2}}}{2}.$ (2.5)
This change of variables depends only on the structure reduction. It introduces no new assumptions and does not involve differentiation.
2.5 Universal Transport Law
Theorem 2.3 (Universal transport equation).
Let $\mathcal{G}=G_{0}+G_{1}i$ be the intrinsic obstruction. Then the forced coefficient system (2.4) is equivalent to the complex first–order transport equation
$\lambda_{x}+\lambda\,\lambda_{y}=G_{0}+\lambda\,G_{1}.$ (2.6)
Proof.
Substituting $\alpha=a^{2}+b^{2}$ and $\beta=-2a$ into (2.4) yields
$\begin{cases}a_{x}+aa_{y}-bb_{y}=G_{0}+aG_{1},\\[4.0pt] b_{x}+ab_{y}+ba_{y}=bG_{1},\end{cases}$
which are precisely the real and imaginary parts of $\lambda_{x}+\lambda\lambda_{y}=G_{0}+\lambda G_{1}$. The converse follows by reversing the computation. ∎
Remark 2.4 (The Burgers equation is written in the standard $\mathbb{C}$).
Although each fiber $A_{z}$ is (elliptic and hence) real–algebra isomorphic to $\mathbb{C}$, there is in general no canonical identification between different fibers, and we deliberately avoid fixing a trivialization. The spectral parameter $\lambda=a+ib=\frac{-\beta+i\sqrt{4\alpha-\beta^{2}}}{2}\in\mathbb{C}_{+}$ is therefore not an element of the moving algebra $A_{z}$, but a scalar function of the real coefficients $(\alpha,\beta)$. Consequently $\lambda_{x}$ and $\lambda_{y}$ are ordinary partial derivatives, and the product $\lambda\,\lambda_{y}$ is the usual complex multiplication in the fixed target $\mathbb{C}$. Equivalently, (2.6) is nothing but a compact rewriting of the real quasilinear system for $(a,b)$. The variable fiber multiplication enters in the analytic layer of the theory, but it plays no role in the coefficient transport law.
Equation (2.6) is a forced complex inviscid Burgers equation written in the standard complex algebra $\mathbb{C}$. It governs the evolution of the structure coefficients for all variable elliptic structures.
2.6 Characteristic Interpretation
Equation (2.6) may be written in transport form as
$(\partial_{x}+\lambda\,\partial_{y})\lambda=G_{0}+\lambda\,G_{1}.$
Along characteristic curves satisfying $dy/dx=\lambda$, the spectral parameter evolves according to the linear ordinary differential equation
$\frac{d\lambda}{dx}=G_{0}(x,y(x))+\lambda(x)\,G_{1}(x,y(x)).$
Thus, the evolution of the coefficients of the structure reduction is always governed by transport along characteristic curves. The obstruction $\mathcal{G}$ acts as a source and growth term.
2.7 Summary
Allowing the generator $i(x,y)$ to vary forces the appearance of the intrinsic obstruction $\mathcal{G}=i_{x}+i\,i_{y}$. Eliminating the derivatives of $i$ from the structure reduction shows that the coefficients $(\alpha,\beta)$ are necessarily governed by a transport equation for the spectral parameter $\lambda$.
This transport law is universal. No assumption has been made concerning algebraic compatibility, homogeneity of the Cauchy–Riemann system, or existence of integral representations.
In the next chapter we identify the special regime in which the obstruction vanishes and show that, in this case, the transport becomes conservative and a coherent analytic function theory emerges.
Chapter 3 Rigidity as the Conservative Regime of Burgers Transport
3.1 From Universal Transport to a Distinguished Regime
Chapter 2 showed that allowing the elliptic structure to vary forces a transport law on its spectral data. Independently of any analytic assumptions, the differentiation of the generator $i(x,y)$ leads to a (possibly forced) complex Burgers equation governing the evolution of the coefficients $(\alpha,\beta)$ through the spectral parameter
$\lambda=\frac{-\beta+i\sqrt{4\alpha-\beta^{2}}}{2},\qquad\Im\lambda>0.$
At this level, no integrability has been imposed. The transport equation may include forcing terms, and no closure properties for the associated function theory are expected.
The purpose of this chapter is to identify a distinguished subregime of this universal transport law in which several independent-looking obstructions collapse simultaneously. This regime is characterized by the vanishing of a single intrinsic quantity, namely
$\mathcal{G}:=i_{x}+i\,i_{y}.$
We call this condition rigidity. It does not generate the transport law; rather, it singles out the conservative case of the already-existing Burgers dynamics.
3.2 The Obstruction Field
Recall that $i(x,y)$ satisfies the structure reduction
$i^{2}+\beta i+\alpha=0,$
and that its derivatives $i_{x}$ and $i_{y}$ are well-defined algebra-valued sections.
Definition 3.1 (Obstruction field).
The obstruction field associated to a variable elliptic structure is
$\mathcal{G}:=i_{x}+i\,i_{y}.$
This quantity measures the failure of the generator $i(x,y)$ to be transported compatibly with its own multiplication law. It appears naturally and unavoidably once differentiation is allowed.
In Chapter 2, prescribing $\mathcal{G}$ was shown to produce a forced Burgers equation for the spectral parameter. We now study the special case $\mathcal{G}\equiv 0$.
3.3 Definition of Rigidity
Definition 3.2 (Rigid elliptic structure).
A variable elliptic structure is called rigid if its obstruction field vanishes identically:
$i_{x}+i\,i_{y}=0\qquad\text{on }\Omega.$ (3.1)
This condition is intrinsic and purely geometric. It does not refer to any function space, product rule, or analytic notion of holomorphicity.
3.4 Rigidity and Conservative Transport
In the notation of Chapter 2, the universal forced Burgers equation takes the form
$\lambda_{x}+\lambda\,\lambda_{y}=\lambda_{x}+\lambda\,\lambda_{y}=G_{0}+\lambda\,G_{1},$
where the forcing terms $(g_{0},g_{1})$ are the real coefficients of where $\mathcal{G}=G_{0}+G_{1}i$ is the decomposition of the obstruction field in the moving basis as in (2.2).
Proposition 3.3 (Rigidity as conservation).
A variable elliptic structure is rigid if and only if the associated Burgers transport law is conservative:
$\lambda_{x}+\lambda\,\lambda_{y}=0.$
Proof.
By Theorem 2.3, the coefficients satisfy $\lambda_{x}+\lambda\lambda_{y}=G_{0}+\lambda G_{1}$. Hence $\mathcal{G}\equiv 0$ (i.e. $G_{0}\equiv G_{1}\equiv 0$) is equivalent to $\lambda_{x}+\lambda\lambda_{y}=0$. ∎
This establishes rigidity as a dynamical notion: it identifies those variable elliptic structures whose coefficient transport is conservative.
3.5 Collapse of Obstructions
We now show that rigidity causes several independent obstructions to collapse simultaneously.
3.5.1 Leibniz defect collapse
Let $f=u+vi$ and $g=p+qi$ be $C^{1}$ sections. As shown in Chapter 1, the Leibniz defect of the generalized Cauchy–Riemann operator is given by
$\mathcal{N}(f,g)=\frac{1}{2}\,vq\,\mathcal{G}.$
Proposition 3.4 (Leibniz defect collapse).
The operator $\partial_{\bar{z}}$ satisfies the Leibniz rule on $C^{1}(\Omega;\allowbreak A)$ if and only if the structure is rigid.
Proof.
The defect vanishes for all $f,g$ if and only if $\mathcal{G}\equiv 0$. ∎
3.5.2 Inhomogeneity collapse
The real Cauchy–Riemann system derived in Chapter 2 has the form
$\begin{cases}u_{x}-\alpha v_{y}+Av=0,\\ v_{x}+u_{y}-\beta v_{y}+Bv=0,\end{cases}$
where $(A,B)$ are the real coefficients of $\mathcal{G}$.
Proposition 3.5 (Inhomogeneity collapse).
The generalized Cauchy–Riemann system is homogeneous if and only if the structure is rigid.
Proof.
Homogeneity is equivalent to $A\equiv B\equiv 0$, which is equivalent to $\mathcal{G}\equiv 0$. ∎
3.6 Rigidity Theorem
We may now state the rigidity theorem in its correct logical position.
Theorem 3.6 (Rigidity Theorem).
Let $\alpha,\beta\in C^{1}(\Omega)$ satisfy ellipticity, and let $i(x,y)$ be a $C^{1}$ choice of generator.
The following statements are equivalent:
  1. 1.
    The obstruction field vanishes: $i_{x}+i\,i_{y}=0$.
  2. 2.
    The Burgers transport law is conservative: $\lambda_{x}+\lambda\lambda_{y}=0$.
  3. 3.
    The generalized Cauchy–Riemann operator satisfies the Leibniz rule.
  4. 4.
    The real Cauchy–Riemann system is homogeneous.
Proof.
The equivalence $(1)\Leftrightarrow(2)$ was shown above. Equivalences with $(3)$ and $(4)$ follow from the collapse results. ∎
Remark 3.7 (Rigidity as intrinsic holomorphicity).
The rigidity condition
$i_{x}+i\,i_{y}=0$
admits an equivalent intrinsic reformulation. It states that the generator $i(x,y)$, viewed as a purely imaginary section of the rank–two algebra bundle it generates, is holomorphic with respect to its own variable elliptic structure:
$\partial_{\bar{z}}i=0$
Indeed, since $\partial_{\bar{z}}=\tfrac{1}{2}(\partial_{x}+i\,\partial_{y})$ and $i$ is a $C^{1}$ section, one has the direct identity
$2\,\partial_{\bar{z}}i=i_{x}+i\,i_{y},$
so $\partial_{\bar{z}}i=0$ is equivalent to (3.1).
In other words, rigidity is exactly the condition that the imaginary unit, considered as a function with values in its defining algebra, satisfies the homogeneous generalized Cauchy–Riemann equations associated with that algebra. This formulation does not rely on external complex coordinates and emphasizes that rigidity is an internal compatibility condition of the moving structure with its own differentiation.
3.7 Interpretation
Rigidity is not the source of the Burgers dynamics. Rather, it is the special case in which the universal transport law becomes conservative and analytic obstructions disappear simultaneously.
From this perspective:
  • Burgers transport belongs to the structural layer of the theory;
  • rigidity selects its conservative subregime;
  • analytic closure properties are consequences, not axioms.
This reordering clarifies the role of rigidity: it is not an assumption imposed to obtain dynamics, but a compatibility condition that restores integrability within an already dynamical framework.
In the next chapter we show that this conservative regime is precisely what allows integral representation formulas and a coherent function theory to re-emerge.
Chapter 4 Conservative Transport and Explicit Rigid Structures
4.1 Purpose and Scope
Chapter 2 established that variable elliptic structures are governed, at the structural level, by a transport equation for their spectral data. Chapter 3 identified rigidity as the distinguished regime in which this transport becomes conservative.
The purpose of the present chapter is twofold:
  • to analyze the conservative transport law as a first–order dynamical system;
  • to construct explicit, genuinely nonconstant rigid elliptic structures.
No new structural assumptions are introduced. All results here are consequences of the conservative Burgers equation and are independent of analytic considerations such as integral formulas or function spaces.
4.2 Conservative Burgers Transport
In the rigid regime, the obstruction field vanishes,
$i_{x}+i\,i_{y}=0,$
and the universal transport law reduces to the conservative complex Burgers equation for the spectral parameter
$\lambda=\frac{-\beta+i\sqrt{4\alpha-\beta^{2}}}{2},\qquad\Im\lambda>0:$
$\lambda_{x}+\lambda\,\lambda_{y}=0.$ (4.1)
Equation (4.1) is written in the standard complex algebra $\mathbb{C}$. No reference to the variable elliptic multiplication is involved at this level.
4.3 Characteristic Formulation
Equation (4.1) admits the transport form
$(\partial_{x}+\lambda\,\partial_{y})\lambda=0.$ (4.2)
Thus $\lambda$ is constant along characteristic curves $(x(s),y(s))$ satisfying
$\frac{dx}{ds}=1,\qquad\frac{dy}{ds}=\lambda,\qquad\frac{d\lambda}{ds}=0.$
Equivalently, along curves of slope $\lambda$, the spectral parameter is conserved. This immediately implies:
  • local existence for arbitrary $C^{1}$ initial data;
  • breakdown through characteristic crossing;
  • strong constraints on global smooth solutions.
These are dynamical consequences of the transport structure and do not depend on ellipticity.
4.4 Real Coefficient Representation
Write
$\beta=-2a,\qquad\alpha=a^{2}+b^{2},\qquad b>0,$
so that $\lambda=a+ib$.
Then (4.1) is equivalent to the real first–order system
$\begin{cases}a_{x}+a\,a_{y}-b\,b_{y}=0,\\[4.0pt] b_{x}+a\,b_{y}+b\,a_{y}=0.\end{cases}$ (4.3)
This system governs the evolution of the structure coefficients $(\alpha,\beta)$ in the rigid regime. It is quasilinear, first order, and hyperbolic.
4.5 Structural Ansatz for Explicit Solutions
We now seek explicit, nonconstant solutions of (4.3) that are:
  • elementary;
  • structurally stable;
  • compatible with ellipticity.
The minimal nontrivial ansatz is to assume:
  • $a$ depends only on $x$;
  • $b$ is constant.
Equivalently, in terms of $(\alpha,\beta)$, we assume
$\alpha=\alpha(x),\qquad\beta=\beta(x,y)\ \text{affine in }y.$ (4.4)
This is the simplest configuration that allows nonconstant transport while preserving explicit solvability.
4.6 Reduction to Ordinary Differential Equations
Let
$a=a(x),\qquad b=b_{0}>0\ \text{constant}.$
Then system (4.3) reduces to
$a_{x}=0,\qquad b_{x}=0,$
along characteristics unless $a$ depends on $x$ alone. The nontrivial case is obtained by allowing $a_{x}\neq 0$ while keeping $b$ constant, which forces linear dependence in $y$.
Rewriting directly in terms of $(\alpha,\beta)$, impose
$\alpha=\alpha(x),\qquad\beta=K(x)\,y.$
Substituting into the rigid coefficient system yields the ODEs
$\begin{cases}\alpha^{\prime}(x)=\alpha(x)\,K(x),\\[4.0pt] K^{\prime}(x)=K(x)^{2}.\end{cases}$ (4.5)
4.7 Explicit Integration
Equation (4.5) integrates explicitly. For a real parameter $\varepsilon$,
$K(x)=\frac{\varepsilon}{1-\varepsilon x},$
on the domain $1-\varepsilon x\neq 0$.
Substituting into the first equation gives
$\alpha(x)=\frac{C}{1-\varepsilon x}.$
Up to normalization, we set $C=1$.
4.8 The $\varepsilon$–Family of Rigid Structures
We arrive at the explicit family
$\alpha(x,y)=\frac{1}{1-\varepsilon x},\qquad\beta(x,y)=\frac{\varepsilon y}{1-\varepsilon x}.$ (4.6)
Proposition 4.1 (The $\varepsilon$–family).
For every $\varepsilon\in\mathbb{R}$, the coefficients (4.6) define a rigid elliptic structure on the domain $\{1-\varepsilon x\neq 0\}$.
Proof.
The functions $\alpha,\beta$ satisfy the rigid coefficient system by construction, and hence correspond to conservative Burgers transport. ∎
4.9 Elliptic Domain
The ellipticity discriminant is
$\Delta=4\alpha-\beta^{2}=\frac{4(1-\varepsilon x)-\varepsilon^{2}y^{2}}{(1-\varepsilon x)^{2}}.$
Thus ellipticity holds precisely on the region
$4(1-\varepsilon x)-\varepsilon^{2}y^{2}>0.$ (4.7)
For fixed $x$, this is a bounded interval in $y$. Degeneracy occurs along the characteristic parabola
$\varepsilon^{2}y^{2}=4(1-\varepsilon x).$
4.10 Interpretation
The $\varepsilon$–family demonstrates several essential facts:
  • Conservative transport does not force constancy.
  • Rigid elliptic structures can be genuinely nontrivial.
  • Ellipticity imposes geometric restrictions on the domain.
  • Global smoothness is incompatible with unrestricted transport.
This family serves as the canonical explicit model throughout the remainder of the theory.
The next chapters turn to analysis. They show that it is precisely the conservative nature of the transport law that allows integral representation formulas, covariant holomorphicity, and a coherent function theory to emerge.
Chapter 5 The Cauchy–Pompeiu Formula in the Rigid Regime
5.1 Standing assumptions and notation
Let $\Omega\subset\mathbb{R}^{2}$ be a bounded domain with piecewise $C^{1}$ boundary, oriented positively. Let $\alpha,\beta\in C^{1}(\Omega,\mathbb{R})$ satisfy the ellipticity condition
$\Delta(z):=4\alpha(z)-\beta(z)^{2}>0\qquad(z\in\Omega).$
For each $z=(x,y)\in\Omega$ let $A_{z}$ be the real two–dimensional commutative algebra generated by $i(z)$ subject to
$i(z)^{2}+\beta(z)\,i(z)+\alpha(z)=0.$ (5.1)
Every AA–valued $C^{1}$ section is written uniquely as
$f(z)=u(z)+v(z)\,i(z),\qquad u,v\in C^{1}(\Omega,\mathbb{R}).$
Conjugation. Define the conjugate root and conjugation map on each fiber by
$\hat{i}(z):=-\beta(z)-i(z),\qquad\widehat{u+v\,i(z)}:=u+v\,\hat{i}(z).$
Then $\widehat{\widehat{w}}=w$ and $\widehat{ab}=\hat{a}\,\hat{b}$ for all $a,b\in A_{z}$.
Cauchy–Riemann operators. For AA–valued $C^{1}$ sections $f$ define
$\partial_{\bar{z}}f:=\frac{1}{2}\bigl(\partial_{x}f+i(z)\,\partial_{y}f\bigr),\qquad\partial_{z}f:=\frac{1}{2}\bigl(\partial_{x}f+\hat{i}(z)\,\partial_{y}f\bigr),$
where $\partial_{x}f,\partial_{y}f$ are computed by differentiating the section f=u+vif=u+v\,i in the usual sense:
$\partial_{x}f=u_{x}+v_{x}\,i+v\,i_{x},\qquad\partial_{y}f=u_{y}+v_{y}\,i+v\,i_{y}.$
Rigidity. Throughout this chapter we assume the structure is rigid:
$i_{x}+i\,i_{y}=0\qquad\text{in }\Omega.$ (5.2)
Under (5.2), $\partial_{\bar{z}}$ satisfies the Leibniz rule on $C^{1}(\Omega;A)$:
$\partial_{\bar{z}}(fg)=(\partial_{\bar{z}}f)g+f(\partial_{\bar{z}}g).$
Cauchy 1–form. Fix the AA–valued 1–form
$d\tilde{z}:=dy-i(z)\,dx.$ (5.3)
Normalization element. Define the canonical section
$j(z):=\frac{2i(z)+\beta(z)}{\sqrt{\Delta(z)}}\in A_{z}.$ (5.4)
Since $(2i+\beta)^{2}=\beta^{2}-4\alpha=-\Delta$, one has
$j(z)^{2}=-1\qquad(z\in\Omega),$ (5.5)
hence $j(z)$ is invertible and $j(z)^{-1}=-j(z)$.
5.2 Basepoint-valued integrals and Stokes’ theorem
The values $f(z)\in A_{z}$ lie in different fibers. To state integral identities with values in a fixed fiber $A_{\zeta}$, we use a coefficientwise integration convention with respect to the moving frame $\{1,i(z)\}$.
Definition 5.1 (Coefficientwise integrals into a fixed fiber).
Fix $\zeta\in\Omega$ and write every AA–valued section as h(z)=u(z)+v(z)i(z)h(z)=u(z)+v(z)\,i(z).
(i) If $\gamma$ is a piecewise $C^{1}$ curve in $\Omega$ and $u,v$ are integrable on $\gamma$ with respect to arclength, define
$\int_{\gamma}^{(\zeta)}h\,ds:=\Bigl(\int_{\gamma}u\,ds\Bigr)\;+\;\Bigl(\int_{\gamma}v\,ds\Bigr)\,i(\zeta)\in A_{\zeta}.$
(ii) If $E\subset\Omega$ is measurable and $u,v$ are integrable on $E$, define
$\iint_{E}^{(\zeta)}h\,dx\,dy:=\Bigl(\iint_{E}u\,dx\,dy\Bigr)\;+\;\Bigl(\iint_{E}v\,dx\,dy\Bigr)\,i(\zeta)\in A_{\zeta}.$
(iii) If $\gamma$ is a piecewise $C^{1}$ curve and $\sigma=P\,dx+Q\,dy$ is a real 11–form along $\gamma$ with $uP,uQ,vP,vQ$ integrable, define
$\int_{\gamma}^{(\zeta)}h\,\sigma:=\Bigl(\int_{\gamma}u\,\sigma\Bigr)\;+\;\Bigl(\int_{\gamma}v\,\sigma\Bigr)\,i(\zeta)\in A_{\zeta},$
where $\int_{\gamma}u\,\sigma:=\int_{\gamma}(uP)\,dx+\int_{\gamma}(uQ)\,dy$ and similarly for $v$.
(iv) In particular, for $d\tilde{z}=dy-i(z)\,dx$ we expand in $A_{z}$:
$(u+vi)d\tilde{z}=u\,dy+vi\,dy-ui\,dx-vi^{2}\,dx.$
Using $i^{2}=-\beta i-\alpha$, we obtain the coefficient decomposition
$(u+vi)d\tilde{z}=\bigl(u\,dy+v\alpha\,dx\bigr)+\bigl(v\,dy-u\,dx+v\beta\,dx\bigr)\,i,$ (5.6)
hence
$\int_{\gamma}^{(\zeta)}h\,d\tilde{z}=\left(\int_{\gamma}u\,dy+\int_{\gamma}v\,\alpha\,dx\right)+\left(\int_{\gamma}v\,dy-\int_{\gamma}u\,dx+\int_{\gamma}v\,\beta\,dx\right)i(\zeta).$ (5.7)
Remark 5.2 (Trivialization viewpoint).
The rule $u+v\,i(z)\mapsto u+v\,i(\zeta)$ is a choice of (local) trivialization of the underlying rank–two real vector bundle $A\to\Omega$ using the moving frame $\{1,i(z)\}$. It is not multiplicative. It is used here only to interpret integrals and Stokes’ theorem coefficientwise: all differential identities are applied to the real coefficient functions and then assembled in the fixed fiber $A_{\zeta}$ by Definition 5.1.
5.3 A multiplicative norm
For $w\in A_{z}$ define
$N_{z}(w):=w\,\widehat{w}\in\mathbb{R},\qquad|w|_{z}:=\sqrt{N_{z}(w)}.$
Lemma 5.3 (Explicit formula and multiplicativity).
If w=u+vi(z)w=u+v\,i(z), then
$N_{z}(w)=u^{2}+\beta(z)\,uv+\alpha(z)\,v^{2}.$
Moreover, $N_{z}$ is multiplicative:
$N_{z}(ab)=N_{z}(a)\,N_{z}(b)\qquad(a,b\in A_{z}),$
hence $|ab|_{z}=|a|_{z}\,|b|_{z}$. In particular, if $w\neq 0$ then $w$ is invertible and
$w^{-1}=\frac{\widehat{w}}{N_{z}(w)},\qquad|w^{-1}|_{z}=\frac{1}{|w|_{z}}.$
Proof.
For $w=u+vi$ one has $\widehat{w}=u+v\hat{i}=u+v(-\beta-i)=(u-\beta v)-vi$. Thus
$w\widehat{w}=(u+vi)\bigl((u-\beta v)-vi\bigr)=u(u-\beta v)-uvi+vi(u-\beta v)-v^{2}i^{2}.$
Since the algebra is commutative, the ii–terms cancel, and using $i^{2}=-\beta i-\alpha$ gives $v^{2}i^{2}=v^{2}(\beta i+\alpha)$; the ii–part cancels, leaving $u^{2}+\beta uv+\alpha v^{2}\in\mathbb{R}$. Multiplicativity follows from $\widehat{ab}=\hat{a}\,\hat{b}$:
$N_{z}(ab)=(ab)\widehat{(ab)}=(ab)(\hat{a}\hat{b})=(a\hat{a})(b\hat{b})=N_{z}(a)N_{z}(b).$
If $w\neq 0$ then $N_{z}(w)>0$ since $\beta^{2}-4\alpha=-\Delta<0$ implies the quadratic form is positive definite, hence $w^{-1}=\widehat{w}/N_{z}(w)$ and $|w^{-1}|_{z}=1/|w|_{z}$. ∎
Lemma 5.4 (Uniform equivalence on compacts).
Let $K\Subset\Omega$. There exist $m_{K},M_{K}>0$ such that for all $z\in K$ and all $w=u+v\,i(z)\in A_{z}$,
$m_{K}\,(u^{2}+v^{2})\leq N_{z}(w)\leq M_{K}\,(u^{2}+v^{2}).$
Equivalently,
$\sqrt{m_{K}}\,\sqrt{u^{2}+v^{2}}\leq|w|_{z}\leq\sqrt{M_{K}}\,\sqrt{u^{2}+v^{2}}.$
Proof.
Let
$Q(z)=\begin{pmatrix}1&\beta(z)/2\\ \beta(z)/2&\alpha(z)\end{pmatrix},\qquad(u,v)Q(z)(u,v)^{T}=N_{z}(u+vi(z)).$
By ellipticity, $Q(z)$ is positive definite for every $z$. On $K$, the eigenvalues of $Q(z)$ are uniformly bounded above and below away from 0 by compactness. ∎
5.4 Kernel and weak singularity
Fix $\zeta=(\xi,\eta)\in\Omega$ and define, for $z=(x,y)\in\Omega$,
$Z(z,\zeta):=(y-\eta)-i(z)\,(x-\xi)\in A_{z}.$ (5.8)
Lemma 5.5 (Invertibility and comparability to Euclidean distance).
Let $K\Subset\Omega$. There exist constants $c_{K},C_{K}>0$ such that for all $z,\zeta\in K$ with $z\neq\zeta$,
$c_{K}\,|z-\zeta|\leq|Z(z,\zeta)|_{z}\leq C_{K}\,|z-\zeta|,$
where $|z-\zeta|=\sqrt{(x-\xi)^{2}+(y-\eta)^{2}}$. In particular $Z(z,\zeta)$ is invertible for $z\neq\zeta$ and
$|Z(z,\zeta)^{-1}|_{z}\leq c_{K}^{-1}\,|z-\zeta|^{-1}.$
Proof.
Write Z=u+vi(z)Z=u+v\,i(z) with $u=y-\eta$ and $v=-(x-\xi)$, so $\sqrt{u^{2}+v^{2}}=|z-\zeta|$. Apply Lemma 5.4. The inverse bound follows from Lemma 5.3. ∎
5.5 $\partial_{\bar{z}}$-holomorphicity of the kernel under rigidity
Lemma 5.6 ($\partial_{\bar{z}}$ of $Z$).
For all $z\neq\zeta$,
$2\,\partial_{\bar{z}}^{z}Z(z,\zeta)=-(x-\xi)\bigl(i_{x}(z)+i(z)i_{y}(z)\bigr).$
In particular, under rigidity (5.2),
$\partial_{\bar{z}}^{z}Z(z,\zeta)=0\qquad(z\neq\zeta).$
Here $\partial_{\bar{z}}^{z}$ means $\partial_{\bar{z}}$ acting in the $z$-variable.
Proof.
Using $\partial_{\bar{z}}=\frac{1}{2}(\partial_{x}+i\partial_{y})$ and $Z=(y-\eta)-i(x-\xi)$,
$\partial_{x}Z=-(i_{x})(x-\xi)-i,\qquad\partial_{y}Z=1-(i_{y})(x-\xi),$
hence
$2\partial_{\bar{z}}Z=\partial_{x}Z+i\,\partial_{y}Z=-(x-\xi)(i_{x}+i\,i_{y}).$
Lemma 5.7 ($\partial_{\bar{z}}$ of the inverse).
Assume rigidity. Then for $z\neq\zeta$,
$\partial_{\bar{z}}^{z}\bigl(Z(z,\zeta)^{-1}\bigr)=0.$
Proof.
On $\Omega\setminus\{\zeta\}$ we have $Z\cdot Z^{-1}=1$. By the Leibniz rule for $\partial_{\bar{z}}$ (rigidity),
$0=\partial_{\bar{z}}(ZZ^{-1})=(\partial_{\bar{z}}Z)Z^{-1}+Z(\partial_{\bar{z}}Z^{-1}).$
By Lemma 5.6 and rigidity, $\partial_{\bar{z}}Z=0$ off the diagonal. Thus $Z(\partial_{\bar{z}}Z^{-1})=0$, and multiplying by $Z^{-1}$ yields $\partial_{\bar{z}}Z^{-1}=0$. ∎
5.6 Exterior calculus identities
Lemma 5.8 (Wedge identity).
For every AA–valued $C^{1}$ function $g$,
$dg\wedge d\tilde{z}=2(\partial_{\bar{z}}g)\,dx\wedge dy.$
Proof.
Write $dg=g_{x}\,dx+g_{y}\,dy$ and $d\tilde{z}=dy-i\,dx$. Then
$dg\wedge d\tilde{z}=(g_{x}dx+g_{y}dy)\wedge(dy-idx)=(g_{x}+ig_{y})\,dx\wedge dy=2(\partial_{\bar{z}}g)\,dx\wedge dy.$
Lemma 5.9 (Non-closure of $d\tilde{z}$).
With $d\tilde{z}=dy-i\,dx$ one has
$dd\tilde{z}=i_{y}\,dx\wedge dy.$
Proof.
Since $d(dy)=0$ and $d(dx)=0$,
$dd\tilde{z}=-d(i)\wedge dx=-(i_{x}dx+i_{y}dy)\wedge dx=i_{y}\,dx\wedge dy.$
5.7 Residue normalization
Fix $\zeta=(\xi,\eta)\in\Omega$ and let $\varepsilon>0$ be small so that $\overline{B_{\varepsilon}(\zeta)}\subset\Omega$. Define the frozen objects in the fixed algebra $A_{\zeta}$:
$Z_{0}(z,\zeta):=(y-\eta)-i(\zeta)(x-\xi)\in A_{\zeta},\qquad d\tilde{z}_{0}:=dy-i(\zeta)\,dx.$
Let $J:=j(\zeta)\in A_{\zeta}$, so that $J^{2}=-1$ by (5.5).
Lemma 5.10 (Frozen residue).
With $J=j(\zeta)$,
$\int_{\partial B_{\varepsilon}(\zeta)}\frac{1}{Z_{0}(z,\zeta)}\,d\tilde{z}_{0}=2\pi\,J\qquad(\varepsilon>0).$
Proof.
Work entirely in the fixed algebra $A_{\zeta}$. Using $J=(2i(\zeta)+\beta(\zeta))/\sqrt{\Delta(\zeta)}$ one can solve for $i(\zeta)$:
$i(\zeta)=\frac{-\beta(\zeta)+\sqrt{\Delta(\zeta)}\,J}{2}.$ (5.9)
Define real linear coordinates (with $\beta(\zeta),\Delta(\zeta)$ frozen):
$X:=(y-\eta)+\frac{\beta(\zeta)}{2}(x-\xi),\qquad Y:=-\frac{\sqrt{\Delta(\zeta)}}{2}(x-\xi).$
Then, using (5.9),
$Z_{0}=(y-\eta)-i(\zeta)(x-\xi)=X+JY,\qquad d\tilde{z}_{0}=dy-i(\zeta)\,dx=dX+J\,dY.$
Hence in the $J^{2}=-1$ algebra,
$\frac{1}{Z_{0}}\,d\tilde{z}_{0}=\frac{1}{X+JY}\,(dX+J\,dY)=\frac{d(X+JY)}{X+JY}.$
Let $W:=X+JY$. The map $(x,y)\mapsto(X,Y)$ is an invertible real linear map, so $\partial B_{\varepsilon}(\zeta)$ is carried to a positively oriented ellipse winding once about the origin. Therefore
$\int_{\partial B_{\varepsilon}(\zeta)}\frac{1}{Z_{0}}\,d\tilde{z}_{0}=\int_{\Gamma}\frac{dW}{W}=2\pi J.$
A continuity lemma for the residue limit
Lemma 5.11 (Continuity of inversion in coefficients).
Let $K\Subset\Omega$, and let $m_{K}>0$ be as in Lemma 5.4. Then for every $z\in K$ and every $w=u+v\,i(z)\in A_{z}$ with $u^{2}+v^{2}=1$ one has
$|w|_{z}\geq\sqrt{m_{K}}.$
Moreover, the coefficient map
$K\times\{(u,v)\in\mathbb{R}^{2}:\;u^{2}+v^{2}=1\}\ni(z,u,v)\longmapsto w^{-1}=\frac{\widehat{w}}{N_{z}(w)}$
is continuous.
Proof.
By Lemma 5.4, for all $z\in K$ and $u^{2}+v^{2}=1$,
$N_{z}(u+v\,i(z))\geq m_{K}(u^{2}+v^{2})=m_{K},$
hence $|w|_{z}=\sqrt{N_{z}(w)}\geq\sqrt{m_{K}}$. The formula $w^{-1}=\widehat{w}/N_{z}(w)$ expresses inversion as a rational function of $(u,v)$ and the continuous coefficients $\alpha(z),\beta(z)$; since the denominator $N_{z}(w)$ is bounded below by $m_{K}>0$ on the stated domain, the map is continuous. ∎
Lemma 5.12 (Variable residue limit).
Fix $\zeta\in\Omega$ and assume rigidity. Then
$\lim_{r\to 0^{+}}\;\int_{\partial B_{r}(\zeta)}^{(\zeta)}\frac{1}{Z(z,\zeta)}\,d\tilde{z}=2\pi\,j(\zeta)\in A_{\zeta}.$
Proof.
Parametrize $\partial B_{r}(\zeta)$ by $z(t)=(\xi+r\cos t,\eta+r\sin t)$, $t\in[0,2\pi]$. Then
$dx=-r\sin t\,dt,\qquad dy=r\cos t\,dt,$
and
$Z(z(t),\zeta)=r\bigl(\sin t-i(z(t))\cos t\bigr),\qquad d\tilde{z}=r\bigl(\cos t+i(z(t))\sin t\bigr)\,dt.$
Hence, for $r>0$,
$\frac{1}{Z}\,d\tilde{z}=\bigl(\sin t-i(z(t))\cos t\bigr)^{-1}\bigl(\cos t+i(z(t))\sin t\bigr)\,dt\in A_{z(t)}.$ (5.10)
Fix a compact $K\Subset\Omega$ containing $\overline{B_{r_{0}}(\zeta)}$ for some small $r_{0}$. For $t\in[0,2\pi]$ the coefficients $(u,v)=(\sin t,-\cos t)$ satisfy $u^{2}+v^{2}=1$. Applying Lemma 5.11 to $w(t)=\sin t-i(z(t))\cos t=u+v\,i(z(t))$ shows that $w(t)$ is uniformly invertible for $z(t)\in K$ and that
$t\longmapsto\bigl(\sin t-i(z(t))\cos t\bigr)^{-1}\bigl(\cos t+i(z(t))\sin t\bigr)$
has real coefficients depending continuously on $(r,t)$ for $0<r\leq r_{0}$.
Since $i(z(t))\to i(\zeta)$ uniformly in $t$ as $r\to 0$, the coefficients of the integrand in (5.10) converge uniformly to those of the frozen integrand
$\bigl(\sin t-i(\zeta)\cos t\bigr)^{-1}\bigl(\cos t+i(\zeta)\sin t\bigr)\in A_{\zeta}.$
By Lemma 5.11 the inverse is uniformly bounded for $0<r\leq r_{0}$, and multiplication in each fiber is continuous in $(z,t)$. Hence the real coefficient functions of the integrand in (5.10) are uniformly bounded on $0,2\pi]\times(0,r_{0}]$, and are therefore dominated by an $L^{1}$ function of $t$ (indeed, by a constant). By dominated convergence applied to the real coefficient functions and the definition of $\int^{(\zeta)}$, we obtain
$\lim_{r\to 0^{+}}\int_{\partial B_{r}(\zeta)}^{(\zeta)}\frac{1}{Z}\,d\tilde{z}=\int_{0}^{2\pi}\bigl(\sin t-i(\zeta)\cos t\bigr)^{-1}\bigl(\cos t+i(\zeta)\sin t\bigr)\,dt.$
The right-hand side equals the frozen residue integral $\int_{\partial B_{1}(\zeta)}Z_{0}^{-1}d\tilde{z}_{0}$ (rescaling cancels), hence equals $2\pi j(\zeta)$ by Lemma 5.10. ∎
5.8 Cauchy–Pompeiu Representation Theorem
Theorem 5.13 (Cauchy–Pompeiu Representation Theorem).
Assume ellipticity and rigidity (5.2). Let $f\in C^{1}(\Omega;A)$. Then for every $\zeta\in\Omega$,
$f(\zeta)=\frac{1}{2\pi\,j(\zeta)}\int_{\partial\Omega}^{(\zeta)}\frac{f(z)}{Z(z,\zeta)}\,d\tilde{z}-\frac{1}{\pi\,j(\zeta)}\iint_{\Omega}^{(\zeta)}\frac{\partial_{\bar{z}}f(z)+\frac{1}{2}\,f(z)\,i_{y}(z)}{Z(z,\zeta)}\,dx\,dy.$ (5.11)
Proof.
Fix $\zeta\in\Omega$ and let $\Omega_{r}:=\Omega\setminus B_{r}(\zeta)$ with $r>0$ small. Define the AA–valued 11–form on $\Omega\setminus\{\zeta\}$:
$\omega(z):=\frac{f(z)}{Z(z,\zeta)}\,d\tilde{z}$.
Write $\omega=g\,d\tilde{z}$ where $g:=fZ^{-1}$. Then
$d\omega=dg\wedge d\tilde{z}+g\,dd\tilde{z}$.
By Lemma 5.8 and Lemma 5.9,
$d\omega=2(\partial_{\bar{z}}g)\,dx\wedge dy+g\,i_{y}\,dx\wedge dy.$
Next compute $\partial_{\bar{z}}g=\partial_{\bar{z}}(fZ^{-1})$. Under rigidity, $\partial_{\bar{z}}$ satisfies Leibniz, so
$\partial_{\bar{z}}g=(\partial_{\bar{z}}f)Z^{-1}+f\,\partial_{\bar{z}}(Z^{-1}).$
By Lemma 5.7, $\partial_{\bar{z}}(Z^{-1})=0$ for $z\neq\zeta$, hence
$\partial_{\bar{z}}g=(\partial_{\bar{z}}f)Z^{-1}.$
Therefore on $\Omega_{r}$,
$d\omega=\left(2(\partial_{\bar{z}}f)Z^{-1}+fZ^{-1}i_{y}\right)\,dx\,dy=2\,\frac{\partial_{\bar{z}}f+\frac{1}{2}f\,i_{y}}{Z}\,dx\,dy.$
Apply Stokes’ theorem to the real coefficient 11–forms defining $\int^{(\zeta)}$ (Definition 5.1). This yields
$\int_{\partial\Omega}^{(\zeta)}\frac{f}{Z}\,d\tilde{z}-\int_{\partial B_{r}(\zeta)}^{(\zeta)}\frac{f}{Z}\,d\tilde{z}=\iint_{\Omega_{r}}^{(\zeta)}d\omega=2\iint_{\Omega_{r}}^{(\zeta)}\frac{\partial_{\bar{z}}f+\frac{1}{2}f\,i_{y}}{Z}\,dx\,dy.$ (5.12)
The area integral converges as $r\to 0^{+}$: by Lemma 5.5, $|Z^{-1}|_{z}\lesssim|z-\zeta|^{-1}$ on compacts, and $|z-\zeta|^{-1}$ is integrable on punctured neighborhoods in $\mathbb{R}^{2}$.
Now split the small-circle term:
$\int_{\partial B_{r}(\zeta)}^{(\zeta)}\frac{f(z)}{Z(z,\zeta)}\,d\tilde{z}=\int_{\partial B_{r}(\zeta)}^{(\zeta)}\frac{f(z)-f(\zeta)}{Z(z,\zeta)}\,d\tilde{z}+f(\zeta)\int_{\partial B_{r}(\zeta)}^{(\zeta)}\frac{1}{Z(z,\zeta)}\,d\tilde{z}$.
Since $f\in C^{1}$, one has $|f(z)-f(\zeta)|=O(r)$ on $\partial B_{r}(\zeta)$ in coefficient norm. Also $|Z^{-1}|_{z}=O(r^{-1})$ and $|d\tilde{z}|=O(r)\,dt$, so the first integral is $O(r)\to 0$. By Lemma 5.12, the second integral converges to $2\pi f(\zeta)j(\zeta)$.
Letting $r\to 0^{+}$ in (5.12) and rearranging gives (5.11). ∎
Remark 5.14 (Origin of the correction term).
The term $\frac{1}{2}f\,i_{y}$ in (5.11) comes only from the non-closure $dd\tilde{z}=i_{y}\,dx\wedge dy$ (Lemma 5.9). rigidity is used to ensure $\partial_{\bar{z}}(Z^{-1})=0$ off the diagonal; it does not imply $i_{y}=0$.
Chapter 6 Covariant Holomorphicity and Gauge Structure
6.1 Motivation from the Cauchy–Pompeiu Formula
Chapter 5 established that under ellipticity and rigidity, every $f\in C^{1}(\Omega;A)$ satisfies the Cauchy–Pompeiu representation
$f(\zeta)=\frac{1}{2\pi\,j(\zeta)}\int_{\partial\Omega}\frac{f(z)}{Z(z,\zeta)}\,d\tilde{z}-\frac{1}{\pi\,j(\zeta)}\iint_{\Omega}\frac{\partial_{\bar{z}}f(z)+\frac{1}{2}f(z)\,i_{y}(z)}{Z(z,\zeta)}\,dx\,dy.$ (6.1)
The boundary term depends only on the values of $f$ along $\partial\Omega$, while the interior term involves the operator
$f\longmapsto\partial_{\bar{z}}f+\tfrac{1}{2}f\,i_{y}.$
Thus the integral formula singles out, in a canonical way, a distinguished first–order operator whose null space eliminates the interior contribution.
This motivates the following intrinsic notion of holomorphicity.
6.2 The Covariant Cauchy–Riemann Operator
Definition 6.1 (Covariant Cauchy–Riemann operator).
Assume ellipticity and rigidity. Define the operator
$Df:=\partial_{\bar{z}}f+\tfrac{1}{2}f\,i_{y},\qquad f\in C^{1}(\Omega;A),$ (6.2)
where $i_{y}$ acts by right multiplication in each fiber.
Definition 6.2 (Covariantly holomorphic section).
A $C^{1}$ section $f:\Omega\to A$ is called covariantly holomorphic if
$Df=0\quad\text{in }\Omega.$
Remark 6.3.
The operator $D$ is not introduced ad hoc. It is uniquely determined by Stokes’ theorem applied to the Cauchy 1–form $d\tilde{z}=dy-i\,dx$ and by the requirement that the interior term in (6.1) vanish.
6.3 Basic Algebraic Properties
Under rigidity, $\partial_{\bar{z}}$ satisfies the Leibniz rule
$\partial_{\bar{z}}(fg)=(\partial_{\bar{z}}f)g+f(\partial_{\bar{z}}g).$
Lemma 6.4 (Leibniz-type formula for $D$).
For $f,g\in C^{1}(\Omega;A)$,
$D(fg)=(Df)g+f(Dg)-\tfrac{1}{2}fg\,i_{y}.$
Proof.
Using the Leibniz rule for $\partial_{\bar{z}}$,
$D(fg)=\partial_{\bar{z}}(fg)+\tfrac{1}{2}fg\,i_{y}=(\partial_{\bar{z}}f)g+f(\partial_{\bar{z}}g)+\tfrac{1}{2}fg\,i_{y}=(Df-\tfrac{1}{2}fi_{y})g+f(Dg-\tfrac{1}{2}gi_{y})+\tfrac{1}{2}fgi_{y},$
which simplifies to the stated formula. ∎
Corollary 6.5 (Failure of closure under pointwise multiplication).
If $Df=0$ and $Dg=0$, then in general
$D(fg)=-\tfrac{1}{2}fg\,i_{y},$
which is nonzero unless $i_{y}\equiv 0$.
Thus even in the rigid regime, covariantly holomorphic functions are not closed under the usual pointwise product. In order to remedy this will introduce a scalar gauge transformation.
6.4 The Weight Equation
The obstruction in Lemma 6.4 is linear and can be removed by a scalar gauge transformation.
Definition 6.6 (Weight).
A nowhere–vanishing real scalar function $\psi\in C^{1}(\Omega)$ is called a weight if it satisfies the first–order equation
$\partial_{\bar{z}}\psi=\tfrac{1}{2}\,i_{y}\,\psi.$ (6.3)
Remark 6.7.
Equation (6.3) is linear and scalar. Since $i_{y}$ is continuous, local solutions exist by standard theory of first–order linear PDE. Any two weights differ by multiplication by a nowhere–vanishing solution of $\partial_{\bar{z}}h=0$.
6.5 Gauge Trivialization
Lemma 6.8 (Intertwining identity).
If $\psi$ is a weight, then for every $f\in C^{1}(\Omega;A)$,
$\partial_{\bar{z}}(\psi f)=\psi\,Df.$
Proof.
Using the Leibniz rule for $\partial_{\bar{z}}$ and (6.3),
$\partial_{\bar{z}}(\psi f)=(\partial_{\bar{z}}\psi)f+\psi(\partial_{\bar{z}}f)=\tfrac{1}{2}i_{y}\psi f+\psi(\partial_{\bar{z}}f)=\psi\bigl(\partial_{\bar{z}}f+\tfrac{1}{2}fi_{y}\bigr)=\psi\,Df.$
Corollary 6.9 (Gauge equivalence).
A section $f$ is covariantly holomorphic if and only if $\psi f$ is $\partial_{\bar{z}}$–holomorphic.
Thus the covariant operator $D$ is a gauge transform of the algebraic operator $\partial_{\bar{z}}$.
6.6 The Weighted Product
The gauge picture suggests a natural product that restores closure.
Definition 6.10 (Weighted product).
Fix a weight $\psi$. For $f,g\in C^{1}(\Omega;A)$ define
$f\diamond g:=fg\psi.$
Lemma 6.11 (Associativity).
The product $\diamond$ is associative.
Proof.
$(f\diamond g)\diamond h=(fg\psi)h\psi=fgh\,\psi^{2}=f\diamond(g\diamond h).$
Theorem 6.12 (Closure under the weighted product).
If $Df=0$ and $Dg=0$, then
$D(f\diamond g)=0.$
Proof.
By Lemma 6.8,
$\partial_{\bar{z}}(\psi fg)=\psi D(fg).$
From Lemma 6.4 and $Df=Dg=0$,
$D(fg)=-\tfrac{1}{2}fg\,i_{y}.$
Hence
$\partial_{\bar{z}}(\psi fg)=-\tfrac{1}{2}\psi fg\,i_{y}=-(\partial_{\bar{z}}\psi)fg$
by (6.3). Therefore
$\partial_{\bar{z}}(\psi fg)+(\partial_{\bar{z}}\psi)fg=0,$
which is precisely $D(fg\psi)=D(f\diamond g)=0$. ∎
Remark 6.13.
The weighted algebra $(\mathcal{H}_{D}(\Omega),\diamond)$ of covariantly holomorphic functions is generally non–unital. Indeed, $1$ is covariantly holomorphic if and only if $i_{y}\equiv 0$, i.e. the structure is constant.
6.7 Local and Global Aspects
On simply connected domains, once a continuous branch of $i(z)$ is fixed, equation (6.3) admits global solutions $\psi$. In that case, multiplication by $\psi$ gives a global isomorphism between covariantly holomorphic functions and $\partial_{\bar{z}}$–holomorphic functions.
On non–simply connected domains, the weight equation may exhibit monodromy. The theory is therefore locally canonical but globally gauge–dependent
6.8 The real first–order system for a weight
Assume ellipticity $4\alpha-\beta^{2}>0$ and fix a $C^{1}$ generator $i(x,y)$ satisfying
$i^{2}+\beta i+\alpha=0.$
Recall that a weight is a nowhere–vanishing real function $\psi\in C^{1}(\Omega,\mathbb{R})$ solving
$\partial_{\bar{z}}\psi=\tfrac{1}{2}\,i_{y}\,\psi,\qquad\partial_{\bar{z}}:=\tfrac{1}{2}(\partial_{x}+i\,\partial_{y}).$ (6.4)
6.8.1 Coefficient decomposition of $i_{y}$
Since each fiber is two–dimensional over $\mathbb{R}$ with basis $\{1,i\}$, there exist unique real functions $A_{y},B_{y}\in C^{0}(\Omega,\mathbb{R})$ such that
$i_{y}=A_{y}+B_{y}\,i.$ (6.5)
In the elliptic regime one may compute $i_{y}$ explicitly from the structure polynomial:
$i_{y}=-\frac{\alpha_{y}+\beta_{y}i}{2i+\beta},\qquad(2i+\beta)^{-1}=\frac{-\beta-2i}{\Delta},\qquad\Delta:=4\alpha-\beta^{2}.$ (6.6)
A direct reduction in the basis $\{1,i\}$ yields
$A_{y}=\frac{\beta\,\alpha_{y}-2\alpha\,\beta_{y}}{\Delta},\qquad B_{y}=\frac{2\alpha_{y}-\beta\,\beta_{y}}{\Delta}.\;$ (6.7)
6.8.2 The real system for a scalar weight
Let $\psi$ be real–valued. Then
$2\,\partial_{\bar{z}}\psi=\psi_{x}+i\,\psi_{y}.$
Substituting (6.5) into (6.4) gives
$\psi_{x}+i\,\psi_{y}=(A_{y}+B_{y}i)\,\psi.$
Since $\{1,i\}$ is a real basis, this is equivalent to the real first–order system
[…]
Equivalently, writing $\varphi=\log\psi$ locally (on $\{\psi>0\}$),
$\begin{cases}\varphi_{x}=A_{y},\\[4.0pt] \varphi_{y}=B_{y}.\end{cases}$ (6.9)
Remark 6.14 (Compatibility).
System (6.8) is overdetermined. Local solvability is equivalent to the closedness condition
$(A_{y})_{y}=(B_{y})_{x},$
i.e. $A_{y}\,dx+B_{y}\,dy$ is locally exact. In particular, the existence of a weight on simply connected domains is equivalent to the existence of a scalar potential $\varphi$ with $d\varphi=A_{y}\,dx+B_{y}\,dy$.
6.9 Example: An explicit weight for the $\varepsilon$–Family
We now show an explicit weight for the $\varepsilon$–deformation of the non–trivial family.
6.9.1 The $\varepsilon$–Family and its elliptic domain
Fix $\varepsilon\in\mathbb{R}$ and consider
$\alpha(x,y)=\frac{1}{1-\varepsilon x},\qquad\beta(x,y)=\frac{\varepsilon y}{1-\varepsilon x}.$
Let
$S(x,y):=4(1-\varepsilon x)-\varepsilon^{2}y^{2}.$
Then the discriminant is
$\Delta(x,y)=4\alpha-\beta^{2}=\frac{S(x,y)}{(1-\varepsilon x)^{2}},$
so the elliptic domain is $\{S>0\}$.
We claim that the real–valued function
$\psi(x,y)=C\sqrt{S(x,y)}$ (6.10)
satisfies the weight equation
$\partial_{\bar{z}}\psi=\frac{1}{2}\,i_{y}\,\psi\qquad\text{on }\{S>0\}.$ (6.11)
6.9.2 Step 1: isolate $i_{y}$ from the structure reduction
The generator $i=i(x,y)$ satisfies the structure relation
$i^{2}+\beta i+\alpha=0.$
Differentiate with respect to $y$:
$(2i+\beta)\,i_{y}+\beta_{y}\,i+\alpha_{y}=0\quad\Longrightarrow\quad i_{y}=-\frac{\beta_{y}i+\alpha_{y}}{2i+\beta}.$
For the present family,
$\alpha_{y}=0,\qquad\beta_{y}=\frac{\varepsilon}{1-\varepsilon x},$
hence
$i_{y}=-\frac{\varepsilon}{1-\varepsilon x}\,\frac{i}{2i+\beta}.$
As in the tractable elliptic regime, $(2i+\beta)$ is invertible and one has
$(2i+\beta)^{-1}=\frac{-\beta-2i}{\Delta}.$
Therefore
$i_{y}=-\frac{\varepsilon}{1-\varepsilon x}\,i\,\frac{-\beta-2i}{\Delta}=\frac{\varepsilon}{1-\varepsilon x}\,\frac{i(\beta+2i)}{\Delta}.$
Reduce $i(\beta+2i)$ using $i^{2}=-\beta i-\alpha$:
$i(\beta+2i)=\beta i+2i^{2}=\beta i+2(-\beta i-\alpha)=-\beta i-2\alpha.$
Hence
$i_{y}=\frac{\varepsilon}{1-\varepsilon x}\,\frac{-\beta i-2\alpha}{\Delta}.$
Now substitute $\alpha=\frac{1}{1-\varepsilon x}$, $\beta=\frac{\varepsilon y}{1-\varepsilon x}$, and $\Delta=\frac{S}{(1-\varepsilon x)^{2}}$ to obtain
$i_{y}(x,y)=-\varepsilon\,\frac{2+\varepsilon y\,i(x,y)}{S(x,y)}.$ (6.12)
6.9.3 Step 2: compute $\partial_{\bar{z}}\psi$
Since $\psi$ is real–valued,
$\partial_{\bar{z}}\psi=\frac{1}{2}(\psi_{x}+i\,\psi_{y}).$
From (6.10) and $S=4(1-\varepsilon x)-\varepsilon^{2}y^{2}$ we compute
$S_{x}=-4\varepsilon,\qquad S_{y}=-2\varepsilon^{2}y,$
and therefore
$\psi_{x}=C\frac{S_{x}}{2\sqrt{S}}=-\frac{2\varepsilon C}{\sqrt{S}},\qquad\psi_{y}=C\frac{S_{y}}{2\sqrt{S}}=-\frac{\varepsilon^{2}y\,C}{\sqrt{S}}.$
Hence
$\partial_{\bar{z}}\psi$ $=\frac{1}{2}\left(-\frac{2\varepsilon C}{\sqrt{S}}+i\left(-\frac{\varepsilon^{2}y\,C}{\sqrt{S}}\right)\right)$
$=-\frac{\varepsilon C}{\sqrt{S}}-\frac{\varepsilon^{2}y\,C}{2\sqrt{S}}\,i=-\frac{\varepsilon C}{2\sqrt{S}}\,(2+\varepsilon y\,i).$
6.9.4 Step 3: compute the right–hand side and compare
Using (6.12),
$\frac{1}{2}\,i_{y}\,\psi=\frac{1}{2}\left(-\varepsilon\,\frac{2+\varepsilon y\,i}{S}\right)\,C\sqrt{S}=-\frac{\varepsilon C}{2\sqrt{S}}\,(2+\varepsilon y\,i).$
This coincides with the expression obtained for $\partial_{\bar{z}}\psi$, hence
$\partial_{\bar{z}}\psi=\frac{1}{2}\,i_{y}\,\psi\quad\text{on }\{S>0\},$
which proves (6.11).
Chapter 7 Second–Order Operators and Factorization in the Rigid Regime
7.1 The Rigid Variable Laplace Operator
In the classical constant structure, one has the Laplacian factorization
$4\,\partial_{z}\partial_{\bar{z}}=\Delta_{\mathrm{Euc}},$
and this identity is the bridge between holomorphicity and elliptic theory.
In the present framework, once rigidity is imposed (Chapter 2), $\partial_{\bar{z}}$ becomes a derivation and the Cauchy–Pompeiu formula (Chapter 4) becomes available. It is then natural—and unavoidable—to study the second–order operator obtained by composing $\partial_{\bar{z}}$ with its conjugate partner $\partial_{z}$.
This chapter shows that, under rigidity, the composition $4\,\partial_{z}\partial_{\bar{z}}$ has a clean elliptic principal part and an explicit first–order correction with no zero–order term.
7.2 Standing Assumptions
Throughout this chapter we assume:
  • $\alpha,\beta\in C^{2}(\Omega,\mathbb{R})$ and $\Delta:=4\alpha-\beta^{2}>0$;
  • rigidity, equivalently the Burgers system
    $\alpha_{x}=\alpha\,\beta_{y},\qquad\beta_{x}+\alpha_{y}=\beta\,\beta_{y}.$ (7.1)
Recall
$\partial_{\bar{z}}:=\tfrac{1}{2}(\partial_{x}+i\,\partial_{y}),\qquad\partial_{z}:=\tfrac{1}{2}(\partial_{x}+\hat{i}\,\partial_{y}),\qquad\hat{i}:=-\beta-i.$
7.3 The Rigid Elliptic Operator
Define the real second–order operator
$L_{\alpha,\beta}:=\partial_{x}^{2}-\beta\,\partial_{xy}+\alpha\,\partial_{y}^{2}.$ (7.2)
Its principal symbol is
$\xi_{x}^{2}-\beta\,\xi_{x}\xi_{y}+\alpha\,\xi_{y}^{2},$
which is positive definite under $\Delta>0$.
7.4 Explicit Expansion of $4\,\partial_{z}\partial_{\bar{z}}$
Let f=u+vif=u+v\,i with $u,v\in C^{2}(\Omega,\mathbb{R})$.
Theorem 7.1 (Rigid second–order expansion).
Assume (7.1). Then
$4\,\partial_{z}\partial_{\bar{z}}(u+v\,i)=\bigl(L_{\alpha,\beta}u+R_{0}[u,v]\bigr)+\bigl(L_{\alpha,\beta}v+R_{1}[u,v]\bigr)\,i,$ (7.3)
where the correction terms are purely first order:
$R_{0}[u,v]$ $=\alpha_{y}\,u_{y}+\alpha_{y}\,v_{x}-2\alpha\,\beta_{y}\,v_{y},$ (7.4)
$R_{1}[u,v]$ $=\beta_{y}\,u_{y}+\beta_{y}\,v_{x}+\bigl(2\alpha_{y}-2\beta\,\beta_{y}\bigr)\,v_{y}.$ (7.5)
In particular, no zero–order terms occur in (7.3).
Proof.
The identities are obtained by expanding $4\,\partial_{z}\partial_{\bar{z}}(u+v\,i)$ using the structure relation $i^{2}+\beta i+\alpha=0$ and the definitions of $\partial_{\bar{z}}$ and $\partial_{z}$, and then imposing the rigidity system (7.1) to eliminate xx–derivatives of $(\alpha,\beta)$.
A symbolic verification in SymPy confirms:
  • the principal part is exactly $L_{\alpha,\beta}u$ and $L_{\alpha,\beta}v$,
  • the remainder contains no second derivatives of $u$ or $v$,
  • the first–order coefficients simplify exactly to (7.4)–(7.5).
The verification script is recorded alongside the manuscript source. ∎
Remark 7.2 (Constant-coefficient reduction).
If $\alpha,\beta$ are constant, then $\alpha_{y}=\beta_{y}=0$ and $R_{0}\equiv R_{1}\equiv 0$. Hence (7.3) reduces to
$4\,\partial_{z}\partial_{\bar{z}}(u+v\,i)=(L_{\alpha,\beta}u)+(L_{\alpha,\beta}v)\,i,$
recovering the classical factorization.
7.5 Interpretation
Theorem 7.1 provides the rigid analogue of the classical identity $4\partial_{z}\partial_{\bar{z}}=\Delta$:
  • the principal part is the uniformly elliptic operator $L_{\alpha,\beta}$;
  • variability contributes only first–order drift terms, explicitly controlled by $\alpha_{y}$ and $\beta_{y}$;
  • rigidity eliminates all zero–order potential terms at this level.
This is precisely the form needed for elliptic estimates and for relating covariant holomorphicity to second–order PDE phenomena.
Conclusion
This monograph developed a variable function theory on planar domains in which the imaginary unit is no longer a fixed algebraic constant but a geometric object that varies with position. The guiding principle throughout has been that the symbol $i$ should be treated as intrinsic data: a moving generator of a rank–two real algebra bundle whose defining structure reduction
$i^{2}+\beta(x,y)\,i+\alpha(x,y)=0$
is allowed to vary. Once this perspective is adopted, the derivatives $i_{x}$ and $i_{y}$ are no longer formal artifacts but genuine geometric quantities, and their algebraic resolution becomes the natural starting point of the theory.
A central object emerging from this viewpoint is the intrinsic combination
$i_{x}+i\,i_{y}.$
This quantity appears prior to any analytic notion of holomorphicity. It governs the variability of the structure itself and, as a consequence, controls both the inhomogeneity of the generalized Cauchy–Riemann system and the failure of the Leibniz rule for the operator $\partial_{\bar{z}}$. These phenomena are not independent: they are different manifestations of the same structural obstruction produced by differentiating the moving generator.
A fundamental outcome of the analysis is that the evolution of the structure coefficients $(\alpha,\beta)$ is governed by a first–order transport law that is already present at the structural level. By differentiating the structure reduction and eliminating the derivatives of $i$, the coefficient dynamics admits a reformulation in terms of the spectral parameter
$\lambda=\frac{-\beta+i\sqrt{4\alpha-\beta^{2}}}{2},\qquad\Im\lambda>0,$
leading, in full generality, to a (possibly forced) complex Burgers equation written in the standard complex algebra. This transport formulation is intrinsic and unavoidable: it reflects the fact that allowing the elliptic structure to vary forces its coefficients to evolve along characteristic curves.
Rigidity occupies a precise and limited position within this framework. It is not the source of the Burgers dynamics, but rather the distinguished conservative regime of the universal transport law. The vanishing condition
$i_{x}+i\,i_{y}=0$
selects exactly those variable elliptic structures for which the Burgers equation is unforced. In this conservative case, several obstructions collapse simultaneously: the generalized Cauchy–Riemann operator becomes a derivation, the real Cauchy–Riemann system becomes homogeneous, and the generator $i$ is transported compatibly with its own multiplication law. Rigidity is therefore not an auxiliary hypothesis but a compatibility condition that restores integrability within an already dynamical setting.
The explicit $\varepsilon$–family of rigid elliptic structures illustrates this conservative regime in concrete terms. It shows that rigidity does not force triviality: genuinely nonconstant structures exist, though ellipticity naturally restricts their domain of definition. From the transport viewpoint, these examples reflect the familiar behavior of first–order hyperbolic systems, where characteristic crossing limits global regularity and constrains smooth extensions.
On the analytic side, rigidity is exactly the condition under which an integral representation theory can persist. In the rigid regime, a Cauchy–Pompeiu formula adapted to variable elliptic structures is recovered. The natural Cauchy 11–form
$d\tilde{z}=dy-i(z)\,dx$
is not closed, and its non–closure produces an unavoidable interior correction term. Rather than being an artifact, this correction singles out a canonical first–order operator
$Df=\partial_{\bar{z}}f+\tfrac{1}{2}f\,i_{y},$
whose null space provides the appropriate notion of holomorphicity in the variable setting. Although covariantly holomorphic sections are not closed under the pointwise product, a scalar gauge transformation restores closure through a natural weighted multiplication, showing how analytic structure can be recovered by compensating for variability rather than suppressing it.
The scope of the present work is deliberately local and explicit. Even in the conservative regime, global questions remain subtle: the choice of a continuous generator $i(z)$ is typically local, gauge trivializations may exhibit monodromy, and the transport nature of the coefficient dynamics explains why globally smooth structures are highly constrained. Beyond rigidity, the analytic function theory necessarily acquires source terms, while the transport–controlled evolution of the coefficients remains intact.
Closing perspective. The central lesson of this work is that once the complex structure is allowed to move, its differentiation becomes a primary geometric invariant. Transport comes first, integrability second, and analytic closure emerges only as a distinguished limiting case. Rigid elliptic structures do not generate the dynamics; they represent the rare situation in which the universal transport law becomes conservative and the analytic and algebraic layers of the theory align. In this way, variable elliptic structures reveal a natural hierarchy in which geometry dictates dynamics, and dynamics, in turn, governs analysis.
Acknowledgements
Use of Generative AI Tools. Portions of the writing and editing of this manuscript were assisted by generative AI language tools. These tools were used to improve clarity of exposition, organization of material, and language presentation. All mathematical results, statements, proofs, and interpretations were developed, verified, and validated by the author. The author takes full responsibility for the accuracy, originality, and integrity of all content in this work, including any material produced with the assistance of AI tools. No generative AI system is listed as an author of this work.
Appendix A Symbolic Verification of the Rigid Second–Order Expansion
This appendix contains a complete symbolic verification of Theorem 7.1 using the computer algebra system SymPy.
The script expands the operator $4\,\partial_{z}\partial_{\bar{z}}$ acting on f=u+vif=u+v\,i using the defining structure relation
$i^{2}+\beta i+\alpha=0,$
substitutes the rigidity (Burgers) system
$\alpha_{x}=\alpha\,\beta_{y},\qquad\beta_{x}+\alpha_{y}=\beta\,\beta_{y},$
and simplifies the resulting expressions in the basis $\{1,i\}$.
It verifies that:
  • the principal part is exactly $L_{\alpha,\beta}$ acting on $(u,v)$;
  • all second–order cross terms cancel under rigidity;
  • the remaining terms are purely first order and coincide with (7.4)–(7.5);
  • no zero–order terms occur.
The computation is purely algebraic and does not rely on any numerical approximation.
Listing 1: SymPy verification of Theorem 7.1
# --- begin SymPy verification script ---
#!/usr/bin/env python3
verify_rigid_dzdbar.py
SymPy verification of the second-order expansion in the chapter
’Second-Order Operators and Factorization in the Rigid Regime’.
It verifies, under the rigidity (Burgers) system
alpha_x = alpha*beta_y,
beta_x = beta*beta_y - alpha_y,
that for f = u + v i one has
4 $d_{\bar z}$ f
[…]
with
[…]
R0[u,v] = alpha_y u_y + alpha_y v_x - 2 alpha beta_y v_y,
R1[u,v] = beta_y u_y + beta_y v_x + (2 alpha_y - 2 beta beta_y) v_y.
The computation is done in the 2D commutative algebra generated by i with
i^2 + beta i + alpha = 0
and uses the identity (elliptic regime)
[…]
to express i_x and i_y from the differentiated structure polynomial.
Run:
python3 verify_rigid_dzdbar.py
import sympy as sp
def main():
# Base variables and scalar coefficient functions
x, y = sp.symbols("x y")
alpha = sp.Function("alpha")(x, y)
beta = sp.Function("beta")(x, y)
u = sp.Function("u")(x, y)
v = sp.Function("v")(x, y)
Delta = 4 * alpha - beta**2
# --- Algebra representation: pairs (a0, a1) correspond to a0 + a1*i ---
def add(a, b):
return (sp.simplify(a[0] + b[0]), sp.simplify(a[1] + b[1]))
def smul(s, a):
return (sp.simplify(s * a[0]), sp.simplify(s * a[1]))
def mul(a, b):
# (a0 + a1 i)(b0 + b1 i) with i^2 = -beta i - alpha
a0, a1 = a
b0, b1 = b
c0 = a0 * b0 - alpha * a1 * b1
c1 = a0 * b1 + a1 * b0 - beta * a1 * b1
return (sp.simplify(c0), sp.simplify(c1))
I = (sp.Integer(0), sp.Integer(1)) # i
hatI = (-beta, sp.Integer(-1)) # $\hat$ i = -beta - i
[…]
inv_twoIplusbeta = (-beta / Delta, sp.Integer(-2) / Delta)
# --- i_x and i_y from differentiated structure polynomial ---
alpha_x = sp.diff(alpha, x)
beta_x = sp.diff(beta, x)
alpha_y = sp.diff(alpha, y)
beta_y = sp.diff(beta, y)
ix = mul((-alpha_x, -beta_x), inv_twoIplusbeta) # -(alpha_x + beta_x i)/(2i+beta)
iy = mul((-alpha_y, -beta_y), inv_twoIplusbeta) # -(alpha_y + beta_y i)/(2i+beta)
# --- Derivatives of a section f = u + v i, i.e. pair (u,v) ---
def dx(f):
f0, f1 = f
# (u_x + v_x i) + v i_x
return add((sp.diff(f0, x), sp.diff(f1, x)), smul(f1, ix))
def dy(f):
f0, f1 = f
# (u_y + v_y i) + v i_y
return add((sp.diff(f0, y), sp.diff(f1, y)), smul(f1, iy))
def dbar(f):
[…]
return smul(sp.Rational(1, 2), add(dx(f), mul(I, dy(f))))
def dz(f):
# d_z = 1/2 (d_x + $\hat$ i d_y)
return smul(sp.Rational(1, 2), add(dx(f), mul(hatI, dy(f))))
f = (u, v)
lhs = smul(4, dz(dbar(f))) # pair (lhs0, lhs1)
# --- Rigidity/Burgers substitutions (eliminate x-derivatives of alpha,beta) ---
subs = {
sp.diff(alpha, x): alpha * sp.diff(beta, y),
sp.diff(beta, x): beta * sp.diff(beta, y) - sp.diff(alpha, y),
# mixed and second x-derivatives via differentiating the Burgers system
subs.update(
sp.diff(alpha, x, y): sp.diff(alpha * sp.diff(beta, y), y),
sp.diff(alpha, y, x): sp.diff(alpha * sp.diff(beta, y), y),
sp.diff(alpha, x, x): sp.diff(alpha * sp.diff(beta, y), x),
sp.diff(beta, x, y): sp.diff(beta * sp.diff(beta, y) - sp.diff(alpha, y), y),
sp.diff(beta, y, x): sp.diff(beta * sp.diff(beta, y) - sp.diff(alpha, y), y),
sp.diff(beta, x, x): sp.diff(beta * sp.diff(beta, y) - sp.diff(alpha, y), x),
def enforce_rigidity(expr):
e = expr
for _ in range(6):
e = sp.expand(e)
e = e.xreplace(subs)
e = sp.simplify(sp.expand(e))
return sp.simplify(e)
lhs0 = enforce_rigidity(lhs[0])
lhs1 = enforce_rigidity(lhs[1])
# --- RHS (principal part + first-order correction) ---
L_u = sp.diff(u, x, 2) - beta * sp.diff(u, x, y) + alpha * sp.diff(u, y, 2)
L_v = sp.diff(v, x, 2) - beta * sp.diff(v, x, y) + alpha * sp.diff(v, y, 2)
R0 = alpha_y * sp.diff(u, y) + alpha_y * sp.diff(v, x) - 2 * alpha * beta_y * sp.diff(v, y)
R1 = beta_y * sp.diff(u, y) + beta_y * sp.diff(v, x) + (2 * alpha_y - 2 * beta * beta_y) * sp.diff(v, y)
rhs0 = enforce_rigidity(L_u + R0)
rhs1 = enforce_rigidity(L_v + R1)
diff0 = sp.simplify(lhs0 - rhs0)
diff1 = sp.simplify(lhs1 - rhs1)
print("Component (1) difference after rigidity simplification:")
print(diff0)
[…]
print(diff1)
assert diff0 == 0, "Scalar component does not match."
assert diff1 == 0, "i-component does not match."
[…]
if __name__ == "__main__":
main()
# --- end script ---
Acknowledgements
Use of Generative AI Tools. Portions of the writing and editing of this manuscript were assisted by generative AI language tools. These tools were used to improve clarity of exposition, organization of material, and language presentation. All mathematical results, statements, proofs, and interpretations were developed, verified, and validated by the author. The author takes full responsibility for the accuracy, originality, and integrity of all content in this work, including any material produced with the assistance of AI tools. No generative AI system is listed as an author of this work.

### v2

@@
 Rigidity is therefore not the source of the coefficient dynamics, but a compatibility condition that restores integrability within an already dynamical setting.
-The explicit $\varepsilon$–family of rigid elliptic structures constructed in this work provides concrete examples of this conservative regime.
+The explicit $\varepsilon$–family of rigid elliptic structures constructed in this work by standard techniques provides concrete examples of this conservative regime.
 These examples show that rigidity does not force triviality: genuinely nonconstant structures exist, although ellipticity naturally restricts the domain of definition.
@@
 1.3 Differentiating the generator
-A first concern in any variable-coefficient theory is whether the moving generator $i=i(x,y)$ can be differentiated in a meaningful and computable way.
-In our setting $i$ is a $C^{1}$ section of the algebra bundle, hence its partial derivatives $i_{x},i_{y}$ are well-defined as sections as well.
-In the elliptic regime, these derivatives can be isolated explicitly from the structure reduction.
-More precisely, the computation relies only on the invertibility of the element $2i+\beta$ in the underlying algebra.
-This condition is satisfied not only in the elliptic case but also in the hyperbolic and parabolic regimes, away from zero divisors.
-Consequently, the algebraic resolution of $i_{x}$ and $i_{y}$ from the structure reduction can be performed verbatim in all three cases; what distinguishes the elliptic regime is not the computation itself but the positivity and invertibility properties that follow from it.
+A basic structural question is whether the moving generator $i=i(x,y)$, defined by
+$i^{2}+\beta(x,y)i+\alpha(x,y)=0,$
+can be differentiated in a meaningful way.
+The generator is not independent data but a chosen $C^{1}$ branch of a root of this polynomial, whose coefficients $\alpha,\beta$ lie in $C^{1}(\Omega)$.
+Its differentiability therefore follows by implicit differentiation of the structure relation.
+Differentiating with respect to $x$ and $y$ gives
+$(2i+\beta)\,i_{x}+\alpha_{x}+\beta_{x}i=0,\qquad(2i+\beta)\,i_{y}+\alpha_{y}+\beta_{y}i=0.$
+Whenever $2i+\beta$ is invertible in the fiber algebra (i.e.
+away from degeneracy of the quadratic), these equations uniquely determine $i_{x}$ and $i_{y}$ as algebra-valued functions depending only on $\alpha,\beta$ and their first derivatives.
+In the elliptic regime this invertibility is automatic, so the derivatives of the generator are explicit consequences of the structure reduction rather than additional geometric data.
+The computation relies only on the invertibility of the element $2i+\beta$ in the fiber algebra.
+This condition holds automatically in the elliptic regime and, more generally, in any nondegenerate quadratic algebra away from zero divisors.
+What distinguishes the elliptic case is not the algebraic solvability of these equations but the uniform positivity and invertibility properties that follow from (1.1).
 1.3.1 Invertibility of $2i+\beta$ in the elliptic regime
@@
 1.6 Explicit Formula for the Inhomogeneous Coefficients
-Differentiating the structure reduction (1.3) gives
-$(2i+\beta)i_{x}+\alpha_{x}+\beta_{x}i=0,\qquad(2i+\beta)i_{y}+\alpha_{y}+\beta_{y}i=0.$
+Using the formulas for $i_{x}$ and $i_{y}$ obtained in (1.7), we compute the intrinsic combination $i_{x}+i\,i_{y}$ in terms of $\alpha$, $\beta$ and their derivatives.
 Using ellipticity to invert $2i+\beta$, one obtains
@@
 Equating coefficients with the right–hand side of (2.3) yields (2.4). ∎
-This system is valid for every variable elliptic structure. No compatibility or integrability assumption has been used.
 2.4 Canonical Spectral Parameter
@@
 $\lambda:=a+ib=\frac{-\beta+i\sqrt{4\alpha-\beta^{2}}}{2}.$ (2.5)
-This change of variables depends only on the structure reduction. It introduces no new assumptions and does not involve differentiation.
+This change of variables depends only on the structure reduction.
+It introduces no new assumptions and does not involve differentiation.
+The sign choice $b>0$ selects the branch corresponding to the elliptic orientation and ensures that $\lambda$ takes values in the upper half–plane.
 2.5 Universal Transport Law
@@
 $i^{2}+\beta i+\alpha=0,$
-and that its derivatives $i_{x}$ and $i_{y}$ are well-defined algebra-valued sections.
+and that its derivatives $i_{x}$ and $i_{y}$ are well-defined algebra-valued sections by the implicit differentiation of the structure relation (Chapter 1).
 Definition 3.1 (Obstruction field).
@@
 In the notation of Chapter 2, the universal forced Burgers equation takes the form
-$\lambda_{x}+\lambda\,\lambda_{y}=\lambda_{x}+\lambda\,\lambda_{y}=G_{0}+\lambda\,G_{1},$
-where the forcing terms $(g_{0},g_{1})$ are the real coefficients of where $\mathcal{G}=G_{0}+G_{1}i$ is the decomposition of the obstruction field in the moving basis as in (2.2).
+$\lambda_{x}+\lambda\,\lambda_{y}=G_{0}+\lambda\,G_{1},$
+where the forcing terms $(G_{0},G_{1})$ are the real coefficients of $\mathcal{G}=G_{0}+G_{1}i$ in the moving basis, as in (2.2).
 Proposition 3.3 (Rigidity as conservation).
@@
 Homogeneity is equivalent to $A\equiv B\equiv 0$, which is equivalent to $\mathcal{G}\equiv 0$. ∎
-3.6 Rigidity Theorem
+3.6 The Rigidity Theorem
 We may now state the rigidity theorem in its correct logical position.
@@
 admits an equivalent intrinsic reformulation.
-It states that the generator $i(x,y)$, viewed as a purely imaginary section of the rank–two algebra bundle it generates, is holomorphic with respect to its own variable elliptic structure:
+It states that the generator $i(x,y)$, viewed as the generator section of the rank–two algebra bundle it defines, is holomorphic with respect to its own variable elliptic structure:
 $\partial_{\bar{z}}i=0$
@@
   • analytic closure properties are consequences, not axioms.
-This reordering clarifies the role of rigidity: it is not an assumption imposed to obtain dynamics, but a compatibility condition that restores integrability within an already dynamical framework.
-In the next chapter we show that this conservative regime is precisely what allows integral representation formulas and a coherent function theory to re-emerge.
+This reordering clarifies the role of rigidity: it is not an assumption imposed to obtain dynamics, but a compatibility condition that restores analytic integrability of the variable Cauchy–Riemann calculus within an already dynamical framework.
+In chapter  5 we show that this conservative regime is precisely what allows integral representation formulas and a coherent function theory to re-emerge.
 Chapter 4 Conservative Transport and Explicit Rigid Structures
 4.1 Purpose and Scope
-Chapter 2 established that variable elliptic structures are governed, at the structural level, by a transport equation for their spectral data.
-Chapter 3 identified rigidity as the distinguished regime in which this transport becomes conservative.
-The purpose of the present chapter is twofold:
-  • to analyze the conservative transport law as a first–order dynamical system;
-  • to construct explicit, genuinely nonconstant rigid elliptic structures.
-No new structural assumptions are introduced.
-All results here are consequences of the conservative Burgers equation and are independent of analytic considerations such as integral formulas or function spaces.
+In the previous chapters, we established that rigid variable elliptic structures are governed by a spectral transport law which, in the conservative regime, reduces to the classical complex Burgers equation.
+The purpose of this chapter is to harness this well–known transport dynamics to construct concrete geometric examples.
+It is important to clarify the logical relationship between the analysis and the underlying PDE theory:
+  • We do not aim to present new results regarding the Burgers equation itself, which is a mature and extensively studied subject in fluid dynamics and mathematical physics.
+  • Instead, we adopt a constructive approach: we select elementary, standard solutions of the Burgers equation (specifically, rational profiles) and interpret them as the spectral data for a variable elliptic algebra.
+By doing so, we demonstrate that rigidity is not a vacuous condition satisfied only by constants, but a rich regime capable of supporting non–trivial, moving geometries.
+To keep the exposition self–contained and consistent with the first–principles approach of this monograph, we briefly recall the necessary method of characteristics before deriving the explicit $\varepsilon$–family of structures.
+4.2 The Burgers Substrate
+In the rigid regime ($i_{x}+i\,i_{y}=0$), the universal structural transport reduces to the inviscid complex Burgers equation for the spectral parameter $\lambda$:
+$\lambda_{x}+\lambda\,\lambda_{y}=0.$ (4.1)
+For the purposes of this monograph, we view Equation (4.1) as a substrate—a pre–existing dynamical engine that generates the coefficients for our geometric structures.
+While the global theory of this equation involves shock waves and weak solutions, our geometric focus (ellipticity) restricts us to the domain of smooth solutions prior to gradient catastrophe.
+4.3 Recalling the Characteristic Formulation
+We utilize the standard method of characteristics to generate solutions. Equation (4.1) implies that $\lambda$ is constant along the integral curves of the vector field $(1,\lambda)$:
+$\frac{dx}{ds}=1,\qquad\frac{dy}{ds}=\lambda,\qquad\frac{d\lambda}{ds}=0.$
+This classical result immediately implies the properties relevant to our construction:
+  • Propagation: Structural data is transported linearly along lines in the complexified $(x,y)$ space.
+  • Implicit Solutions: The general solution is given implicitly by $\lambda=F(y-\lambda x)$ for an arbitrary function $F$.
+  • Blow–up: Smooth solutions generally develop singularities in finite time (characteristic crossing), which corresponds to the degeneration of the elliptic structure.
+In the following sections, we apply this standard machinery to the specific ansatz of linear dependence, recovering a canonical family of rigid structures.
 4.4 Conservative Burgers Transport
@@
 $\frac{dx}{ds}=1,\qquad\frac{dy}{ds}=\lambda,\qquad\frac{d\lambda}{ds}=0.$
-Equivalently, along curves of slope $\lambda$, the spectral parameter is conserved. This immediately implies:
+Equivalently, along integral curves of the vector field $(1,\lambda)$ in the (x,y)(x,y)–plane, the spectral parameter is conserved. This immediately implies:
   • local existence for arbitrary $C^{1}$ initial data;
@@
 $\begin{cases}a_{x}+a\,a_{y}-b\,b_{y}=0,\\[4.0pt] b_{x}+a\,b_{y}+b\,a_{y}=0.\end{cases}$ (4.4)
-This system governs the evolution of the structure coefficients $(\alpha,\beta)$ in the rigid regime. It is quasilinear, first order, and hyperbolic.
+This system governs the evolution of the structure coefficients $(\alpha,\beta)$ in the rigid regime. It is a quasilinear first-order system of transport type.
 4.7 Structural Ansatz for Explicit Solutions
@@
 $a=a(x),\qquad b=b_{0}>0\ \text{constant}.$
-Then system (4.3) reduces to
-$a_{x}=0,\qquad b_{x}=0,$
-along characteristics unless $a$ depends on $x$ alone. The nontrivial case is obtained by allowing $a_{x}\neq 0$ while keeping $b$ constant, which forces linear dependence in $y$.
+Substituting $a=a(x)$ and $b=b_{0}$ (constant) into (4.4), the second equation is identically satisfied, while the first reduces to
+$a_{x}+a\,a_{y}=0.$
+Since $a=a(x)$, this forces $a_{y}=0$, so the system imposes no further restriction on $a(x)$ at this stage.
+To obtain a nontrivial rigid structure, it is more transparent to work directly in terms of $(\alpha,\beta)$.
+The nontrivial case is obtained by allowing $a_{x}\neq 0$ while keeping $b$ constant, which forces linear dependence in $y$.
 Rewriting directly in terms of $(\alpha,\beta)$, impose
 $\alpha=\alpha(x),\qquad\beta=K(x)\,y.$
+Because the rigidity condition is most naturally expressed in the original coefficients $(\alpha,\beta)$, we now translate the ansatz back into these variables.
 Substituting into the rigid coefficient system yields the ODEs
@@
 Proposition 4.1 (The $\varepsilon$–family).
-For every $\varepsilon\in\mathbb{R}$, the coefficients (4.6) define a rigid elliptic structure on the domain $\{1-\varepsilon x\neq 0\}$.
+For every $\varepsilon\in\mathbb{R}$, the coefficients (4.7) define a rigid elliptic structure on the region where both $1-\varepsilon x\neq 0$ and the ellipticity condition (4.8) hold.
 Proof.
@@
 Chapter 6 Covariant Holomorphicity and Gauge Structure
+Covariant Holomorphicity and Gauge Structure in the Rigid Regime
 6.1 Motivation from the Cauchy–Pompeiu Formula
@@
 No generative AI system is listed as an author of this work.
-Appendix A Symbolic Verification of the Rigid Second–Order Expansion
+Appendix A Algebraic Universality of the Intrinsic Obstruction and the Leibniz Defect
+This appendix isolates a purely algebraic mechanism underlying all failures of derivation behavior in variable rank–two algebra calculi.
+No analytic notions (such as ellipticity, Cauchy–Riemann operators, or integral representations) are used.
+The results show that once a quadratic algebra structure is allowed to vary, there exists a unique intrinsic obstruction governing both the failure of the Leibniz rule and the transport of the structure itself.
+A.1 First–order calculus on a moving rank–two algebra
+Let $\Omega\subset\mathbb{R}^{2}$ be open and let $A\to\Omega$ be a rank–two commutative real algebra bundle generated by a section $i\in C^{1}(\Omega;A)$ satisfying the structure relation
+$i^{2}+\beta\,i+\alpha=0,$
+where $\alpha,\beta\in C^{1}(\Omega,\mathbb{R})$ are arbitrary real functions.
+Every section $f\in C^{1}(\Omega;A)$ admits a unique decomposition
+$f=u+vi,\qquad u,v\in C^{1}(\Omega;\mathbb{R}).$
+Definition A.1 (Compatible first–order operator).
+A first–order operator
+$\delta:C^{1}(\Omega;A)\longrightarrow C^{0}(\Omega;A)$
+is said to be compatible with the moving algebra structure if:
+  1. 1.
+    it acts as a derivation on real scalar functions,
+    $\delta(uv)=u\,\delta(v)+v\,\delta(u)\quad(u,v\in C^{1}(\Omega;\mathbb{R}));$
+  2. 2.
+    it satisfies the chain rule on the generator,
+    $\delta(vi)=\delta(v)i+v\,\delta(i).$
+No derivation property is assumed on the full algebra $A$.
+A.2 The Leibniz defect
+The deviation of $\delta$ from being a derivation on $A$ is measured by the Leibniz defect
+$N_{\delta}(f,g):=\delta(fg)-\delta(f)g-f\,\delta(g).$
+Lemma A.2 (Factorization of the Leibniz defect).
+Let $\delta$ be a compatible operator. For any sections $f=u+vi$ and $g=p+qi$, one has
+$N_{\delta}(f,g)=vq\Bigl[\delta(i^{2})-2i\,\delta(i)\Bigr].$
+Proof.
+All products involving at least one scalar factor satisfy the Leibniz rule by assumption. The only possible failure arises from the product of the vector parts:
+$(vi)(qi)=vq\,i^{2}.$
+On the one hand,
+$\delta(vq\,i^{2})=\delta(vq)i^{2}+vq\,\delta(i^{2}).$
+On the other hand,
+$\delta(vi)(qi)+(vi)\delta(qi)$ $=(\delta(v)i+v\delta(i))qi+vi(\delta(q)i+q\delta(i))$
+$=\delta(vq)i^{2}+2vq\,i\delta(i).$
+Subtracting yields the stated formula. ∎
+This factorization shows that the defect is bilinear in the vector components and universal: no other algebraic combinations can contribute.
+A.3 Decomposition of the defect
+Using the structure relation $i^{2}=-\beta i-\alpha$ and the scalar derivation property of $\delta$, one computes
+$\delta(i^{2})=-\delta(\alpha)-\delta(\beta)i-\beta\,\delta(i).$
+Substituting into Lemma A.2 gives
+$\delta(i^{2})-2i\delta(i)=-\delta(\alpha)-\delta(\beta)i-(2i+\beta)\delta(i).$
+Corollary A.3 (Total defect formula).
+For any compatible operator $\delta$,
+$N_{\delta}(f,g)=-vq\Bigl[\underbrace{\delta(\alpha)+\delta(\beta)i}_{\text{structural transport}}+\underbrace{(2i+\beta)\delta(i)}_{\text{geometric transport}}\Bigr].$
+This decomposition is purely algebraic and holds independently of ellipticity or invertibility assumptions.
+A.4 Universality of the intrinsic obstruction
+Formula (A.3) shows that any failure of the Leibniz rule for a first–order calculus compatible with a moving quadratic algebra structure can arise from exactly two sources:
+  • transport of the structure coefficients $(\alpha,\beta)$;
+  • transport of the generator $i$ itself.
+No other algebraic mechanism can contribute to a Leibniz defect.
+In the setting of differentiation on the plane, this distinction becomes more refined.
+Pure real derivatives, such as $\partial_{x}$ or $\partial_{y}$ acting coefficientwise on sections $f=u+vi$, always satisfy the Leibniz rule, even when the algebra structure varies.
+Indeed, for such operators one has $\delta(i^{2})=2i\,\delta(i)$, and the defect vanishes identically.
+The intrinsic obstruction appears precisely for first–order operators whose symbol involves the moving generator itself. A canonical example is the generalized Cauchy–Riemann operator
+$\partial_{\bar{z}}=\tfrac{1}{2}(\partial_{x}+i\,\partial_{y}),$
+for which the interaction between differentiation and the variable algebra produces the intrinsic combination
+$G:=i_{x}+i\,i_{y}.$
+This quantity measures the incompatibility between differentiation along algebra–adapted directions and multiplication in the fibers.
+The present appendix shows that $G$ is not introduced by analytic or geometric assumptions.
+It is the unique algebraic obstruction forced by allowing the quadratic structure to vary and by differentiating along directions defined by that structure.
+Whether detected through the failure of the Leibniz rule or through differentiation of the structure polynomial, the same obstruction necessarily appears.
+In the main text, imposing $G\equiv 0$ is shown to eliminate all such defects simultaneously, selecting the conservative (rigid) subregime of the universal transport law.
+Appendix B Constructive Differentiability of the Structure Generator
+This appendix records a recursive argument showing that the structure generator can be differentiated as many times as permitted by the regularity of the structure coefficients.
+Although only the first and second derivatives are used in the main body of the text, the argument clarifies the general mechanism and justifies the regularity bookkeeping adopted throughout the work.
+Statement of the problem
+Let $\Omega\subset\mathbb{R}^{2}$ be open and let $\alpha,\beta:\Omega\to\mathbb{R}$ be given functions. Consider the structure polynomial
+$X^{2}+\beta(x,y)X+\alpha(x,y),$
+and let $i=i(x,y)$ be a chosen root of this polynomial. We assume that the algebraic element $2i+\beta$ is invertible at every point of $\Omega$ (this includes, in particular, the elliptic regime).
+The question addressed here is the following:
+Given $\alpha,\beta\in C^{k}(\Omega)$, to what extent can the derivatives of $i$ be defined and computed explicitly?
+Constructive recursive principle
+The guiding principle is entirely constructive: derivatives of $i$ are defined only insofar as they can be obtained by differentiating the structure polynomial and solving the resulting linear equations.
+No differentiability of $i$ is assumed a priori.
+Proposition B.1 (Recursive constructivist differentiability).
+Assume $\alpha,\beta\in C^{k}(\Omega)$ for some integer $k\geq 0$, and assume that $2i+\beta$ is invertible on $\Omega$.
+Then for every multi-index $\gamma$ with $|\gamma|\leq k$, the partial derivative $\partial^{\gamma}i$ exists, is continuous, and can be computed explicitly in terms of derivatives of $\alpha$ and $\beta$ of order $\leq|\gamma|$ and derivatives of $i$ of lower order.
+In particular, $i$ is constructively $C^{k}$ on $\Omega$.
+Proof.
+The proof proceeds by induction on the order of differentiation.
+Base case ($|\gamma|=0$). For $|\gamma|=0$, the statement is trivial: $i$ is defined pointwise as a root of the structure polynomial.
+First-order case. Assume $\alpha,\beta\in C^{1}(\Omega)$. Differentiating the identity
+$i^{2}+\beta i+\alpha=0$
+with respect to $x$ and $y$ yields the linear equations
+$(2i+\beta)i_{x}=-(\beta_{x}i+\alpha_{x}),\qquad(2i+\beta)i_{y}=-(\beta_{y}i+\alpha_{y}).$
+Since $2i+\beta$ is invertible, these equations uniquely define $i_{x}$ and $i_{y}$. The right-hand sides are continuous, hence $i\in C^{1}(\Omega)$.
+Inductive step.
+Assume that for some $m$ with $0\leq m<k$, all partial derivatives $\partial^{\eta}i$ with $|\eta|\leq m$ have been defined and are continuous, and that each such derivative is given by an explicit algebraic expression involving derivatives of $\alpha$ and $\beta$ of order $\leq m$ and derivatives of $i$ of order <m<m.
+Let $\gamma$ be a multi-index with $|\gamma|=m+1$. Choose a coordinate direction $p\in\{x,y\}$ and a multi-index $\eta$ with $|\eta|=m$ such that
+$\partial^{\gamma}=\partial_{p}\partial^{\eta}.$
+Apply $\partial^{\eta}$ to the structure identity and then differentiate once more with respect to $p$.
+By repeated application of the Leibniz rule, all resulting terms can be expressed as sums of products involving:
+  • derivatives of $\alpha$ and $\beta$ of order $\leq m+1$,
+  • derivatives of $i$ of order $\leq m+1$.
+Crucially, the only terms involving the highest derivative $\partial^{\gamma}i$ appear linearly and with the same algebraic coefficient. More precisely, one obtains an identity of the form
+$(2i+\beta)\,\partial^{\gamma}i=F_{\gamma},$
+where $F_{\gamma}$ is an explicit expression depending only on derivatives of $\alpha$ and $\beta$ of order $\leq m+1$ and derivatives of $i$ of order $\leq m$.
+Since $2i+\beta$ is invertible, this equation uniquely defines
+$\partial^{\gamma}i:=(2i+\beta)^{-1}F_{\gamma}.$
+By the induction hypothesis and the assumption $\alpha,\beta\in C^{m+1}$, the right-hand side is continuous. This completes the inductive step.
+The conclusion follows by induction. ∎
+Remarks
+Remark B.2.
+The argument shows that the differentiability of the structure generator is entirely determined by the regularity of the structure coefficients.
+No gain of regularity occurs: $i$ cannot be differentiated more times than allowed by the least regular of $\alpha$ and $\beta$.
+Remark B.3.
+Although the proof is presented in the elliptic setting, the recursive construction relies only on the invertibility of $2i+\beta$ and therefore applies verbatim in the hyperbolic regime away from zero divisors.
+For the parabolic regime this cannot be applied directly as $2i+\beta$ is a zero divisor.
+Appendix C Partial Derivatives of the Structure Generator in the Parabolic Regime
+This appendix describes how the partial derivatives of the structure generator $i(x,y)$ behave in the parabolic regime
+$\Delta:=4\alpha-\beta^{2}=0,$
+where the quadratic algebra degenerates and the element $2i+\beta$ ceases to be invertible.
+In this case the elliptic/hyperbolic formulas for $i_{x}$ and $i_{y}$ cannot be used directly, and a different interpretation is required.
+C.1 Algebraic structure when $\Delta=0$
+If $\Delta=0$, then
+$\alpha=\frac{\beta^{2}}{4},\qquad X^{2}+\beta X+\alpha=\Bigl(X+\frac{\beta}{2}\Bigr)^{2}.$
+Hence in each fiber algebra
+$\Bigl(i+\frac{\beta}{2}\Bigr)^{2}=0.$
+Define the nilpotent element
+$\varepsilon:=i+\frac{\beta}{2}.$
+Then
+$\varepsilon^{2}=0,\qquad i=-\frac{\beta}{2}+\varepsilon,$
+and the fiber algebra is isomorphic to the algebra of dual numbers
+$A_{z}\cong\mathbb{R}[\varepsilon]/(\varepsilon^{2}).$
+In particular,
+$2i+\beta=2\varepsilon$
+is nilpotent and therefore a zero divisor.
+C.2 Differentiating the parabolic structure relation
+Since $\varepsilon^{2}=0$, differentiating with respect to $x$ and $y$ gives
+$2\varepsilon\,\varepsilon_{x}=0,\qquad 2\varepsilon\,\varepsilon_{y}=0.$
+Using $\varepsilon_{x}=i_{x}+\tfrac{\beta_{x}}{2}$ and $\varepsilon_{y}=i_{y}+\tfrac{\beta_{y}}{2}$, we obtain the constraints
+$\varepsilon\Bigl(i_{x}+\tfrac{\beta_{x}}{2}\Bigr)=0,\qquad\varepsilon\Bigl(i_{y}+\tfrac{\beta_{y}}{2}\Bigr)=0.$ (C.1)
+In the dual number algebra, the annihilator of $\varepsilon$ is precisely the ideal it generates. Hence (C.1) implies the general form
+$i_{x}=-\frac{\beta_{x}}{2}+\phi_{x}\,\varepsilon,\qquad i_{y}=-\frac{\beta_{y}}{2}+\phi_{y}\,\varepsilon,$ (C.2)
+for some real functions $\phi_{x},\phi_{y}$.
+Thus, unlike the elliptic and hyperbolic cases, the structure relation alone does not uniquely determine the derivatives of $i$. There is a one–dimensional nilpotent freedom in each derivative.
+C.3 Canonical choice via a nondegenerate limit
+A natural way to remove the ambiguity in (C.2) is to define $i_{x}$ and $i_{y}$ as limits of the elliptic/hyperbolic formulas as $\Delta\to 0$.
+For $\Delta\neq 0$, one has
+$i_{x}=-\frac{\beta_{x}}{2}-\frac{\Delta_{x}}{4(\beta+2i)},\qquad i_{y}=-\frac{\beta_{y}}{2}-\frac{\Delta_{y}}{4(\beta+2i)}.$
+Since $(\beta+2i)^{2}=-\Delta$, the denominators scale like $\sqrt{|\Delta|}$.
+If $\Delta_{x}$ and $\Delta_{y}$ vanish at least as fast as $\sqrt{|\Delta|}$, the singular terms disappear in the limit $\Delta\to 0$, yielding the canonical parabolic derivatives
+$i_{x}:=-\frac{\beta_{x}}{2},\qquad i_{y}:=-\frac{\beta_{y}}{2}.$ (C.3)
+These are exactly the choices obtained by eliminating the nilpotent component in (C.2).
+C.4 Consistency check
+With the canonical choice (C.3),
+$2i+\beta=2\varepsilon,$
+and therefore
+$(2i+\beta)i_{x}=2\varepsilon\Bigl(-\frac{\beta_{x}}{2}\Bigr)=-\beta_{x}\varepsilon,$
+which matches the derivative of the identity $\varepsilon^{2}=0$:
+$\partial_{x}(\varepsilon^{2})=2\varepsilon\,\varepsilon_{x}=0.$
+Thus the canonical derivatives satisfy all algebraic constraints imposed by the parabolic structure.
+C.5 Interpretation
+In the parabolic regime the quadratic algebra acquires a nilpotent direction, and the operator $2i+\beta$ loses invertibility.
+Consequently, the structure equation no longer determines $i_{x}$ and $i_{y}$ uniquely; it only constrains them up to a nilpotent term.
+The canonical choice (C.3) selects the derivatives that arise as limits of nondegenerate (elliptic or hyperbolic) structures and removes the nilpotent ambiguity.
+This provides a consistent notion of “constructive differentiability” in the parabolic setting.
+Appendix D Jet Analysis of the $\varepsilon$–Family at the Constant Structure
+This appendix analyzes the $\varepsilon$–family of rigid elliptic structures introduced in Chapter 4 from the viewpoint of jet theory at $\varepsilon=0$.
+Although the limiting structure is exactly the standard complex plane, the jet analysis reveals nontrivial infinitesimal geometry encoded in the family.
+The purpose of this appendix is to make precise what information is contained in the $\varepsilon$–jets and what conclusions can be drawn from them.
+D.1 The $\varepsilon$–family and its constant limit
+Recall the explicit rigid family
+$\alpha^{\varepsilon}(x,y)=\frac{1}{1-\varepsilon x},\qquad\beta^{\varepsilon}(x,y)=\frac{\varepsilon y}{1-\varepsilon x},$
+defined on the elliptic domain
+$4(1-\varepsilon x)-\varepsilon^{2}y^{2}>0.$
+At $\varepsilon=0$ one has
+$\alpha^{0}\equiv 1,\qquad\beta^{0}\equiv 0,$
+so the structure polynomial reduces to
+$i^{2}+1=0,$
+and the algebra bundle becomes canonically identified with the standard complex plane. All derivatives of the generator vanish:
+$i_{x}=i_{y}=0,$
+and the obstruction field is identically zero.
+Thus the $\varepsilon=0$ structure is not merely rigid but completely constant. Any nontrivial behavior of the family must therefore appear at the level of jets in $\varepsilon$.
+D.2 First $\varepsilon$–jet of the coefficients
+Expand the coefficients in $\varepsilon$:
+$\alpha^{\varepsilon}=1+\varepsilon\,\alpha^{(1)}+O(\varepsilon^{2}),\qquad\beta^{\varepsilon}=\varepsilon\,\beta^{(1)}+O(\varepsilon^{2}),$
+where
+$\alpha^{(1)}(x,y)=x,\qquad\beta^{(1)}(x,y)=y.$
+The pair $(\alpha^{(1)},\beta^{(1)})$ constitutes the first $\varepsilon$–jet of the structure.
+It measures the initial direction in which the algebra departs from the constant complex structure as $\varepsilon$ is turned on.
+D.3 Jet of the spectral parameter
+Recall the canonical spectral parameter
+$\lambda=\frac{-\beta+i\sqrt{4\alpha-\beta^{2}}}{2}.$
+For the $\varepsilon$–family one finds
+$\lambda^{\varepsilon}=i+\varepsilon\,\mu+O(\varepsilon^{2}),$
+where $\mu$ is a complex–valued scalar function determined by the first jet of $(\alpha,\beta)$. A direct expansion gives
+$\mu=\frac{1}{2}(\alpha^{(1)}+i\,\beta^{(1)}).$
+Thus the first $\varepsilon$–jet of the structure is equivalently encoded in the complex function $\mu$.
+D.4 Linearization of the Burgers transport
+Rigidity is equivalent to the conservative Burgers equation
+$\lambda_{x}+\lambda\,\lambda_{y}=0.$
+Substituting
+$\lambda^{\varepsilon}=i+\varepsilon\mu+O(\varepsilon^{2})$
+and retaining only first–order terms in $\varepsilon$ yields the linearized equation
+$\mu_{x}+i\,\mu_{y}=0.$
+This is precisely the classical Cauchy–Riemann equation for the complex–valued function $\mu$.
+D.5 Interpretation of the jet equation
+The jet equation
+$\mu_{x}+i\,\mu_{y}=0$
+has a clear structural meaning:
+  • It is not an analytic assumption but the linearization of the universal transport law governing variable elliptic structures.
+  • It describes the tangent space at the constant structure to the space of rigid variable elliptic structures.
+  • Its solutions are exactly classical holomorphic functions.
+Thus classical complex analysis appears here as the infinitesimal theory controlling how a constant elliptic structure may begin to vary while remaining rigid to first order.
+D.6 What the jet analysis does not say
+It is important to emphasize what conclusions should not be drawn:
+  • The $\varepsilon$–family is not a deformation of complex analysis itself, but of the geometric structure for which complex analysis is the linearized theory at $\varepsilon=0$.
+  • Holomorphic functions do not parametrize finite rigid structures; they parametrize only their first–order directions.
+  • Nonlinear effects enter at second order in $\varepsilon$, where the full Burgers transport becomes essential.
+Thus the jet analysis captures only the infinitesimal geometry of the space of structures, not its global behavior.
+D.7 Conceptual conclusion
+The jet analysis of the explicit rigid $\varepsilon$–family shows that, at the constant elliptic structure, the universal Burgers transport linearizes to the classical Cauchy–Riemann equations.
+This illustrates a general mechanism of the theory: complex analysis arises as the linearized shadow of the nonlinear transport geometry governing variable elliptic structures.
+We do not claim that this linearization has been established for arbitrary families, but the computation reveals the structural origin of the phenomenon.
+At $\varepsilon=0$ the structure is exactly the standard complex plane.
+The first $\varepsilon$–jet detects how this structure can begin to move, and the linearized Burgers equation reduces to the Cauchy–Riemann system.
+Higher–order jets encode genuinely nonlinear transport phenomena with no classical counterpart.
+From this perspective, rigidity does not trivialize the theory at the constant structure.
+Rather, it identifies complex analysis as the tangent theory at a distinguished point in the space of variable elliptic geometries.
+D.8 Second $\varepsilon$–jet: onset of nonlinearity
+We now examine the second $\varepsilon$–jet of the rigid family. Write the spectral parameter as
+$\lambda^{\varepsilon}=i+\varepsilon\mu+\varepsilon^{2}\nu+O(\varepsilon^{3}),$
+where $\mu$ is the first–order jet discussed above and $\nu$ is the second–order jet.
+Substitute this expansion into the conservative Burgers equation
+$\lambda_{x}+\lambda\,\lambda_{y}=0$
+and collect terms of order $\varepsilon^{2}$. Using $\mu_{x}+i\mu_{y}=0$, one finds
+$\nu_{x}+i\nu_{y}=-\,\mu\,\mu_{y}.$
+This equation has several important features.
+  • It is inhomogeneous, with a source term quadratic in the first jet $\mu$.
+  • The operator on the left–hand side is again the classical $\partial_{\bar{z}}$ operator.
+  • The right–hand side depends only on $\mu$ and its first derivatives.
+Thus the second $\varepsilon$–jet is no longer free data.
+It is determined by a forced Cauchy–Riemann equation whose source encodes the first nonlinear interaction of the infinitesimal deformation with itself.
+In particular, even when $\mu$ is holomorphic, $\nu$ is generally not holomorphic. This marks the first point at which classical complex analysis ceases to describe the structure exactly.
+D.9 Third $\varepsilon$–jet: hierarchy of transport corrections
+Proceeding one order further, write
+$\lambda^{\varepsilon}=i+\varepsilon\mu+\varepsilon^{2}\nu+\varepsilon^{3}\rho+O(\varepsilon^{4}).$
+Substitution into Burgers and collection of $\varepsilon^{3}$ terms yields
+$\rho_{x}+i\rho_{y}=-\bigl(\mu\,\nu_{y}+\nu\,\mu_{y}\bigr).$
+At this level the structure of the jet hierarchy becomes clear:
+  • The third jet $\rho$ satisfies a linear first–order equation.
+  • The forcing term is bilinear in the lower jets $(\mu,\nu)$.
+  • No new differential operators appear; only $\partial_{\bar{z}}$ persists.
+Thus each successive jet solves a linear transport equation whose source is constructed from the previous jets.
+This recursive structure is entirely determined by the Burgers nonlinearity and does not involve any analytic assumptions.
+D.10 Structure of the full jet hierarchy
+The pattern observed above persists to all orders. Writing formally
+$\lambda^{\varepsilon}=i+\sum_{k\geq 1}\varepsilon^{k}\lambda^{(k)},$
+one obtains, for every $k\geq 1$,
+$\partial_{\bar{z}}\lambda^{(k)}=F_{k}\bigl(\lambda^{(1)},\dots,\lambda^{(k-1)}\bigr),$
+where $F_{k}$ is a universal polynomial expression involving only products and yy–derivatives of the lower–order jets.
+The hierarchy has the following conceptual interpretation:
+  • The first jet $\lambda^{(1)}$ is free and holomorphic.
+  • All higher jets are slaved to $\lambda^{(1)}$ through forced Cauchy–Riemann equations.
+  • Nonlinearity enters immediately at second order and propagates upward.
+Thus classical holomorphic data parametrize the tangent space at the constant structure, but they do not determine a finite deformation uniquely without solving the full nonlinear transport problem.
+D.11 Refined conclusion of the jet analysis
+Combining the first, second, and third jet equations yields a precise refinement of the conceptual picture:
+Classical complex analysis governs the infinitesimal geometry of rigid variable elliptic structures, but nonlinear Burgers transport controls their finite behavior.
+At first order, rigidity linearizes to the Cauchy–Riemann equations.
+At second order, self–interaction appears and produces forced corrections.
+At third and higher orders, a hierarchical transport system emerges, entirely determined by the initial holomorphic jet.
+In this sense, complex analysis is not deformed but embedded as the tangent theory of a nonlinear geometric transport problem.
+Appendix E Integrability, Coordinate Trivialization, and the Intrinsic Obstruction
+Terminology warning.
+In this appendix, the word integrability is used in the classical differential–geometric sense of the Newlander–Nirenberg theorem: the existence of local coordinates in which an almost complex structure becomes constant.
+This notion is distinct from the analytic integrability discussed in the main text, which concerns the closure of the variable Cauchy–Riemann calculus and the existence of a coherent function theory in a fixed coordinate frame.
+The intrinsic obstruction $G$ is unrelated to the former notion but is decisive for the latter.
+E.1 The Classical Integrability Objection
+A natural objection to the study of variable elliptic structures in real dimension two arises from classical complex geometry.
+In the theory of almost complex manifolds, the fundamental invariant measuring non–integrability is the Nijenhuis tensor.
+Given a smooth vector–valued one–form $J$ acting as an almost complex structure, the Nijenhuis tensor is defined by
+$N_{J}(X,Y)=[JX,JY]-J[JX,Y]-J[X,JY]-[X,Y],$
+for vector fields $X,Y$.
+The Newlander–Nirenberg theorem states that a smooth almost complex structure is locally integrable if and only if its Nijenhuis tensor vanishes identically.
+In real dimension two (complex dimension one), it is a classical result that
+$N_{J}\equiv 0\qquad\text{for every smooth almost complex structure}.$
+As a consequence, every two–dimensional almost complex structure admits local isothermal coordinates: there exists a local diffeomorphism
+$\Phi:(x,y)\longmapsto(u,v)$
+in which the complex structure becomes constant.
+This fact gives rise to the following objection.
+If every variable elliptic structure in dimension two can be locally flattened by a change of coordinates, does the intrinsic obstruction introduced in this work,
+$G:=i_{x}+i\,i_{y},$
+carry any genuine geometric or analytic meaning?
+E.2 What the Obstruction Does Not Measure
+Before answering this objection, it is essential to clarify what the obstruction $G$ does not measure.
+The quantity $G$ is not an integrability invariant in the sense of complex geometry.
+Its vanishing does not assert the existence of holomorphic coordinates, nor does its nonvanishing obstruct such coordinates.
+In real dimension two, integrability in the classical sense is guaranteed independently of $G$.
+Accordingly:
+  • $G\equiv 0$ does not imply coordinate integrability (existence of local isothermal coordinates);
+  • $G\not\equiv 0$ does not obstruct coordinate integrability.
+The intrinsic obstruction belongs to a different conceptual layer than the Nijenhuis tensor.
+E.3 Parallelism versus Coordinate Trivialization
+The obstruction $G$ arises from differentiating a moving generator
+$i=i(x,y)$
+of a rank–two real algebra bundle relative to the fixed background coordinates $(x,y)$. It measures the failure of $i$ to be parallel with respect to the flat background connection:
+$G=i_{x}+i\,i_{y}.$
+By contrast, the Nijenhuis tensor detects whether there exists a coordinate system in which the complex structure becomes constant.
+This is a statement about the existence of a preferred co–moving frame, not about the behavior of the structure relative to a fixed frame.
+Passing to isothermal coordinates eliminates $G$ by construction, but only at the expense of abandoning the original background connection.
+The obstruction disappears because the frame is chosen to move with the structure, not because the structure was intrinsically stationary.
+Thus:
+  • the vanishing of the Nijenhuis tensor concerns coordinate trivialization;
+  • the vanishing of $G$ concerns parallel transport in a fixed frame.
+These notions are logically independent.
+E.4 Eulerian and Lagrangian Viewpoints
+The distinction admits a transparent interpretation in transport–theoretic terms.
+Lagrangian (co–moving) viewpoint.
+Choosing isothermal coordinates corresponds to following the geometry itself. In this frame, the structure is static by definition. This viewpoint underlies classical integrability results.
+Eulerian (laboratory) viewpoint.
+Fixing the background coordinates $(x,y)$ corresponds to observing the structure as it moves through space.
+In this frame, the generator $i(x,y)$ evolves, and its evolution is governed by the transport laws derived in Chapter 2.
+The obstruction $G$ is precisely the Eulerian measure of this motion.
+From this perspective, the forced Burgers equation
+$\lambda_{x}+\lambda\lambda_{y}=G$
+describes the transport of the structure coefficients relative to the fixed background.
+The existence of a co–moving frame in which this motion disappears does not negate the reality of the motion itself.
+E.5 Conclusion
+In real dimension two, classical integrability is automatic and therefore incapable of distinguishing variable elliptic structures.
+The Nijenhuis tensor is blind to the phenomena studied in this monograph because it detects only the possibility of flattening the geometry by a change of coordinates.
+The intrinsic obstruction
+$G=i_{x}+i\,i_{y}$
+addresses a different question.
+It measures the failure of the structure to be parallel relative to a fixed flat connection and governs the transport dynamics of the structure coefficients.
+It captures the kinematics of a variable algebraic structure as seen in an Eulerian frame, independently of whether the structure may be locally trivialized in a co–moving frame.
+The theory developed here is therefore not a theory of geometric integrability of complex structures in the Newlander–Nirenberg sense, but a theory of transport of algebraic structures and of the analytic calculus they induce in a fixed background frame.
+Appendix F Symbolic Verification of the Rigid Second–Order Expansion
 This appendix contains a complete symbolic verification of Theorem 7.1 using the computer algebra system SymPy.

### v3

@@
 Daniel Alayón-Solarz Email: danieldaniel@gmail.com
-(January 2026)
+(February 2026)
 A fixed complex structure gives analysis.
@@
 Vanegas, whose work on parameter–depending Clifford Algebras and generalized Cauchy–Pompeiu formulas provided both motivation and guidance.
+At an early stage (circa 2012), the author did not yet accept the necessity of differentiating the imaginary unit itself.
+It was H.
+De Bie who emphasized that allowing derivatives of the generator $i(x,y)$ is not an auxiliary choice but a requirement for internal consistency: without $i_{x}$ and $i_{y}$, differentiation and multiplication cannot be reconciled once the elliptic structure varies.
+This remark proved decisive in clarifying the geometric foundations of the framework, and the author is grateful to H.
+De Bie for this insight.
 Although W. Tutschke did not see the full development of the theory presented here, his influence is present throughout in the guiding questions and structural viewpoint.
@@
 which proves (6.11).
-Chapter 7 Second–Order Operators and Factorization in the Rigid Regime
+Chapter 7 Power Series Representations in the Rigid Regime
+7.1 Purpose and Overview
+In the classical theory of complex analysis, holomorphic functions admit local power series representations obtained by differentiating the Cauchy integral formula.
+This procedure relies crucially on two structural facts:
+  • the imaginary unit is constant,
+  • the Cauchy kernel depends only on the difference $z-\zeta$.
+Neither of these properties survives in the presence of a variable elliptic structure. As a result, differentiation under the integral sign is no longer a legitimate operation.
+The purpose of this chapter is twofold:
+  1. 1.
+    to show that covariantly holomorphic functions admit intrinsic power series representations in frozen algebra powers;
+  2. 2.
+    to explain how ordinary holomorphic functions inherit power series expansions only via gauge trivialization, and not directly.
+This resolves the apparent tension between the failure of classical arguments and the persistence of a local expansion theory.
+Important scope distinction.
+Throughout this chapter, the term “power series representation” refers to expansions constructed from integral representations and valid on non-infinitesimal neighborhoods.
+Purely infinitesimal Taylor expansions at a point, which rely only on freezing the algebra at that point, do not detect the variable structure and may exist without invoking any gauge.
+The necessity of gauge trivialization arises only when transport or integral manipulations are involved.
+7.2 Standing Assumptions
+Throughout this chapter we assume:
+  • $\alpha,\beta\in C^{1}(\Omega)$ with $\Delta=4\alpha-\beta^{2}>0$;
+  • the structure is rigid, i.e.
+    $i_{x}+i\,i_{y}=0;$
+  • $A_{z}=\mathbb{R}[X]/(X^{2}+\beta(z)X+\alpha(z))$;
+  • $Df=\bar{\partial}f+\tfrac{1}{2}f\,i_{y}$ is the covariant Cauchy–Riemann operator.
+All statements are local.
+7.3 Covariantly Holomorphic Functions and the Cauchy–Pompeiu Formula
+Recall that in the rigid regime the Cauchy–Pompeiu representation takes the form
+$f(\zeta)=\frac{1}{2\pi j(\zeta)}\int_{\partial\Omega}^{(\zeta)}\frac{f(z)}{Z(z,\zeta)}\,d\tilde{z}-\frac{1}{\pi j(\zeta)}\iint_{\Omega}^{(\zeta)}\frac{\bar{\partial}f+\tfrac{1}{2}f\,i_{y}}{Z(z,\zeta)}\,dx\,dy.$
+Lemma 7.1.
+If $f\in C^{1}(\Omega;A)$ satisfies $Df=0$, then
+$f(\zeta)=\frac{1}{2\pi j(\zeta)}\int_{\partial\Omega}^{(\zeta)}\frac{f(z)}{Z(z,\zeta)}\,d\tilde{z}$.
+Proof.
+This follows immediately from the vanishing of the interior term $\bar{\partial}f+\tfrac{1}{2}f\,i_{y}$ in the Cauchy–Pompeiu formula. ∎
+This boundary representation is the correct replacement for the classical Cauchy integral formula in the variable elliptic setting.
+7.4 Frozen Algebra and Polynomial Powers
+Fix $\zeta=(\xi,\eta)\in\Omega$ and define the frozen generator $i(\zeta)\in A_{\zeta}$. Introduce the frozen variable
+$Z_{0}(z,\zeta):=(y-\eta)-i(\zeta)(x-\xi)\in A_{\zeta}.$
+All powers $Z_{0}^{n}$ are computed in the fixed algebra $A_{\zeta}$. They are genuine polynomials in $(x-\xi,y-\eta)$ and form a linearly independent family over $\mathbb{R}$.
+The variable kernel $Z(z,\zeta)$ differs from $Z_{0}(z,\zeta)$ only through the dependence of $i(z)$ on $z$, which is controlled by ellipticity and rigidity.
+7.5 Expansion of the Cauchy Kernel
+On sufficiently small circles $|z-\zeta|=r$, the inverse kernel admits the convergent expansion
+$\frac{1}{Z(z,\zeta)}=\sum_{n=0}^{\infty}\frac{Z_{0}(z,\zeta)^{n}}{\rho(z,\zeta)^{n+1}},\qquad\rho(z,\zeta):=|Z_{0}(z,\zeta)|_{\zeta}^{2}.$
+This expansion is purely algebraic and takes place entirely in the frozen fiber $A_{\zeta}$. Uniform ellipticity guarantees convergence.
+7.6 Power Series for Covariantly Holomorphic Functions
+Theorem 7.2 (Power series for covariantly holomorphic functions).
+Let $f\in C^{1}(\Omega;A)$ satisfy $Df=0$. Then for every $\zeta\in\Omega$ there exists $r>0$ and unique coefficients $a_{n}(\zeta)\in A_{\zeta}$ such that
+$f(z)=\sum_{n=0}^{\infty}a_{n}(\zeta)\,Z_{0}(z,\zeta)^{n},\qquad|z-\zeta|<r,$
+with uniform convergence on compact subsets. These coefficients arise from the Cauchy–Pompeiu representation and are therefore intrinsic to the covariant structure.
+Proof.
+Fix $\zeta$ and apply the boundary representation on $\partial B_{r}(\zeta)$.
+Insert the kernel expansion and integrate term-by-term.
+Coefficientwise integration is legitimate since all terms lie in the fixed algebra $A_{\zeta}$.
+Define
+$a_{n}(\zeta):=\frac{1}{2\pi j(\zeta)}\int_{\partial B_{r}(\zeta)}^{(\zeta)}\frac{f(z)}{\rho(z,\zeta)^{n+1}}\,d\tilde{z}$.
+Ellipticity provides absolute convergence. Uniqueness follows from the linear independence of $\{Z_{0}^{n}\}$ in $A_{\zeta}$. ∎
+7.7 Why Differentiation Under the Integral Fails in the Variable Case
+We now explain why the classical argument based on differentiating the Cauchy integral formula does not extend to variable elliptic structures.
+7.7.1 Dependence of the Kernel on the Evaluation Point
+In the classical case, the kernel $(z-\zeta)^{-1}$ depends on $\zeta$ only through translation. In the variable case, the kernel
+$Z(z,\zeta)=(y-\eta)-i(z)(x-\xi)$
+depends on $\zeta$ in two independent ways:
+  • explicitly through $(\xi,\eta)$,
+  • implicitly through the evaluation fiber $A_{\zeta}$.
+Differentiating with respect to $\zeta$ therefore produces terms involving derivatives of the generator $i(z)$ transported into a different fiber.
+There is no canonical identification that allows these terms to be controlled.
+7.7.2 Moving Fibers and Loss of Algebraic Compatibility
+The integral
+$\int^{(\zeta)}\frac{f(z)}{Z(z,\zeta)}\,d\tilde{z}$
+takes values in the frozen fiber $A_{\zeta}$ by coefficientwise projection. Differentiating with respect to $\zeta$ changes the target algebra itself.
+Thus differentiation and integration do not commute: the operation “differentiate in $\zeta$” is not fiberwise compatible with the coefficientwise integration used to define the integral.
+7.7.3 Non-Closure of the Cauchy 1–Form
+Even under rigidity, the Cauchy 1–form
+$d\tilde{z}=dy-i(z)\,dx$
+is not closed:
+$d(d\tilde{z})=i_{y}\,dx\wedge dy.$
+This produces the interior correction term in the Cauchy–Pompeiu formula and forces the introduction of the covariant operator $D$.
+Attempting to differentiate under the integral sign ignores this structural defect and leads to incorrect identities.
+7.7.4 Conclusion
+The failure of differentiation under the integral is not technical. It is structural and reflects:
+  • the variability of the algebra,
+  • the lack of a canonical global trivialization,
+  • the non-closure of the Cauchy 1–form.
+Power series expansions must therefore be constructed from the integral representation itself, not by differentiation.
+Remark 7.3 (Infinitesimal versus transported expansions).
+The frozen variable $Z_{0}(z,\zeta)$ captures the infinitesimal behavior of the algebra structure at the point $\zeta$.
+Any $C^{1}$ function admits a formal Taylor expansion in powers of $(x-\xi,y-\eta)$, and such expansions may be rewritten in terms of $Z_{0}$ without reference to the covariant structure.
+However, these purely infinitesimal expansions do not interact correctly with integral representations or with transport away from the basepoint.
+The constructions in this chapter address power series derived from the Cauchy–Pompeiu formula, which necessarily encode non-infinitesimal behavior and therefore require the covariant framework.
+7.8 Power Series for Holomorphic Functions via Gauge
+Let $h\in C^{1}(\Omega;A)$ satisfy $\bar{\partial}h=0$.
+While $h$ admits a formal Taylor expansion at any point by freezing the algebra, such expansions do not respect the integral structure or transport properties of the variable elliptic system.
+To obtain power series representations compatible with the Cauchy–Pompeiu formula, a gauge trivialization is required.
+Theorem 7.4 (Power series for holomorphic functions).
+For every $\zeta\in\Omega$ there exist $r>0$ and coefficients $b_{n}(\zeta)\in A_{\zeta}$ such that
+$h(z)=\sum_{n=0}^{\infty}b_{n}(\zeta)\,Z_{0}(z,\zeta)^{n},\qquad|z-\zeta|<r.$
+Proof.
+Expand $f$ by the previous theorem and absorb $\psi(\zeta)$ into the coefficients. Convergence is preserved since $\psi$ is smooth and nonvanishing. ∎
+Remark 7.5.
+The role of the gauge in this section is not to enable the existence of local Taylor jets, which are purely infinitesimal, but to reconcile ordinary holomorphicity with the covariant framework underlying integral representations.
+Without gauge trivialization, the notions of differentiation, integration, and algebra multiplication fail to commute in a structurally controlled way.
+7.9 Final Remarks
+Covariant holomorphicity is the correct analytic notion in variable elliptic structures. It admits an intrinsic power series theory without invoking forbidden operations.
+Ordinary holomorphicity inherits power series expansions only through gauge trivialization, not directly.
+This clarifies both the necessity and the strength of the covariant framework.
+In summary, Taylor expansions probe only the frozen infinitesimal algebra at a point, whereas Cauchy–Pompeiu expansions probe the transport of the algebra itself.
+The latter, and not the former, necessitate the covariant and gauge-theoretic framework developed here.
+Chapter 8 Second–Order Operators and Factorization in the Rigid Regime
 8.1 The Rigid Variable Laplace Operator
@@
 and this identity is the bridge between holomorphicity and elliptic theory.
-In the present framework, once rigidity is imposed (Chapter 2), $\partial_{\bar{z}}$ becomes a derivation and the Cauchy–Pompeiu formula (Chapter 4) becomes available.
+In the present framework, once rigidity is imposed Chapter  2, $\partial_{\bar{z}}$ becomes a derivation and the Cauchy–Pompeiu formula (5) becomes available.
 It is then natural—and unavoidable—to study the second–order operator obtained by composing $\partial_{\bar{z}}$ with its conjugate partner $\partial_{z}$.
@@
 This is precisely the form needed for elliptic estimates and for relating covariant holomorphicity to second–order PDE phenomena.
+Chapter 9 The Variable Vekua Equation
+9.1 Reduction to Variable–Structure Vekua Form
+We show that any real first–order elliptic system on the plane can be written, canonically and without auxiliary reductions, as a Vekua–type equation associated with a variable elliptic structure.
+9.1.1 Elliptic system and induced structure
+Let $\Omega\subset\mathbb{R}^{2}$ be open. Consider a real first–order elliptic system
+$\begin{cases}-v_{y}+a_{11}u_{x}+a_{12}u_{y}+a_{1}u+b_{1}v=f_{1},\\ \ \ v_{x}+a_{21}u_{x}+a_{22}u_{y}+a_{2}u+b_{2}v=f_{2},\end{cases}$ (9.1)
+with coefficients of class $C^{1}$ and ellipticity condition
+$\Delta_{\mathrm{sys}}:=a_{11}a_{22}-\tfrac{1}{4}(a_{12}+a_{21})^{2}>0.$
+Define the structure coefficients
+$\alpha:=\frac{a_{22}}{a_{11}},\qquad\beta:=-\frac{a_{12}+a_{21}}{a_{11}}.$ (9.2)
+Then
+$\Delta:=4\alpha-\beta^{2}=\frac{4}{a_{11}^{2}}\,\Delta_{\mathrm{sys}}>0,$
+and $\alpha,\beta$ determine a variable elliptic algebra
+$A_{z}=\mathbb{R}[X]/(X^{2}+\beta(z)X+\alpha(z)).$
+Let $i=i(z)$ denote the distinguished generator,
+$i^{2}+\beta i+\alpha=0,$
+and define elliptic conjugation by
+$\widehat{i}:=-\beta-i.$
+9.1.2 Canonical principal part
+By elementary substitutions preserving ellipticity, system (9.1) can be written in the form
+$\begin{cases}-\alpha V_{y}+U_{x}+aU+bV=f,\\ \ \ V_{x}-\beta V_{y}+U_{y}+cU+dV=g,\end{cases}$ (9.3)
+where $a,b,c,d,f,g$ are real-valued functions on $\Omega$.
+Introduce the algebra–valued field
+$W:=U+iV.$
+9.1.3 Intrinsic decomposition of $\bar{\partial}W$
+Define the variable–structure Cauchy–Riemann operator
+$\bar{\partial}:=\tfrac{1}{2}(\partial_{x}+i\partial_{y}).$
+A direct computation yields the identity
+$2\bar{\partial}W=\big(U_{x}-\alpha V_{y}\big)+\big(V_{x}+U_{y}-\beta V_{y}\big)i+V\,G,$ (9.4)
+where
+$G:=i_{x}+i\,i_{y}$
+is the intrinsic obstruction of the elliptic structure.
+Substituting (9.3) into (9.4) gives
+$2\bar{\partial}W-VG=(f+ig)-(a+ic)U-(b+id)V.$ (9.5)
+9.1.4 Elliptic conjugation and Vekua form
+Using
+$U=\tfrac{1}{2}(W+\widehat{W}),\qquad V=\frac{W-\widehat{W}}{2i+\beta},$
+the right–hand side of (9.5) can be written uniquely as
+$F-AW-B\widehat{W}$,
+with
+$F:=\tfrac{1}{2}(f+ig),$
+and coefficients
+$A$ $=A_{0}-\frac{1}{2}\,\frac{G}{2i+\beta},$
+$B$ $=B_{0}+\frac{1}{2}\,\frac{G}{2i+\beta},$
+where
+$A_{0}$ $=\frac{1}{4}\Big(a-\frac{\beta}{\alpha}b+d+i\,c-i\,\frac{1}{\alpha}b\Big),$
+$B_{0}$ $=\frac{1}{4}\Big(a+\frac{\beta}{\alpha}b-d+i\,c+i\,\frac{1}{\alpha}b\Big).$
+Dividing (9.5) by $2$, we arrive at the canonical variable–structure Vekua equation
+$\bar{\partial}W+AW+B\widehat{W}=F.$ (9.6)
+9.1.5 Remarks
+  • Equation (9.6) is intrinsic to the elliptic structure $(\alpha,\beta)$ and requires no auxiliary reduction.
+  • When the structure is rigid ($G\equiv 0$), the coefficients reduce to their classical forms.
+  • The obstruction $G$ measures precisely the deviation from rigid Cauchy–Riemann calculus and is absorbed into the lower–order terms.
+9.2 Rigidization
+Rigidization is the process by which a variable elliptic structure is locally converted into a rigid one by a change of variables.
+It acts simultaneously on the base coordinates and on the structure polynomial, and it is intrinsic to the transport data of the structure.
+Throughout this section we work locally on $\Omega\subset\mathbb{R}^{2}$ and assume
+$\alpha,\beta\in C^{1}(\Omega),\qquad\Delta=4\alpha-\beta^{2}>0.$
+9.2.1 Rigid and non–rigid structures
+Let $i(x,y)$ denote the generator of the elliptic algebra determined by
+$i^{2}+\beta i+\alpha=0.$
+Definition 9.1 (Intrinsic obstruction).
+The intrinsic obstruction of the elliptic structure is
+$G:=i_{x}+i\,i_{y}.$
+Definition 9.2 (Rigid structure).
+The elliptic structure $(\alpha,\beta)$ is called rigid if
+$G\equiv 0.$
+Equivalently, rigidity means that the variable–structure Cauchy–Riemann operator $\bar{\partial}=\tfrac{1}{2}(\partial_{x}+i\partial_{y})$ satisfies the Leibniz rule on algebra–valued functions.
+9.2.2 Transport form of the obstruction
+Introduce the complex spectral slope
+$\lambda:=\frac{-\beta+i\sqrt{\Delta}}{2},\qquad\Im\lambda>0.$
+The obstruction identity is equivalently written as the first–order transport equation
+$\lambda_{x}+\lambda\,\lambda_{y}=H,$ (9.7)
+where $H$ is a complex–valued function canonically determined by $G=G_{0}+G_{1}i$. In particular,
+$G\equiv 0\quad\Longleftrightarrow\quad\lambda_{x}+\lambda\,\lambda_{y}=0.$
+Thus rigidity is equivalent to flat transport of the spectral slope.
+9.2.3 Definition of rigidization
+Definition 9.3 (Rigidization).
+A rigidization is a local diffeomorphism
+$\Phi:(x,y)\mapsto(X,Y)$
+such that, when the elliptic structure is pulled back to (X,Y)(X,Y)–coordinates, the resulting structure is rigid.
+Rigidization is an Eulerian construction: the diffeomorphism is generated by an ODE driven by the transport field of the structure.
+9.2.4 Rigidization ODE
+Let $\lambda(x,y)$ be the spectral slope of the original structure. Define the vector field
+$V:=\partial_{x}+\lambda\,\partial_{y}.$
+Proposition 9.4 (Rigidizing flow).
+Let $(X(x,y),Y(x,y))$ solve the first–order system
+$\begin{cases}V(X)=1,\\ V(Y)=\lambda.\end{cases}$ (9.8)
+Then $(X,Y)$ defines a local diffeomorphism, and in the (X,Y)(X,Y)–coordinates the transport equation becomes
+$\lambda_{X}+\lambda\,\lambda_{Y}=0.$
+Proof.
+Along the flow of $V$, equation (9.8) ensures that $X$ is the flow parameter and $Y$ is transported consistently with $\lambda$.
+By construction, derivatives along $V$ become $\partial_{X}$, eliminating the source term in (9.7).
+∎
+9.2.5 Algorithmic description
+Rigidization can be implemented as the following local algorithm.
+  1. 1.
+    Compute $\alpha,\beta$ from the principal part of the system.
+  2. 2.
+    Compute $\Delta=4\alpha-\beta^{2}$ and the spectral slope
+    $\lambda=\frac{-\beta+i\sqrt{\Delta}}{2}.$
+  3. 3.
+    Form the transport vector field
+    $V=\partial_{x}+\lambda\,\partial_{y}.$
+  4. 4.
+    Solve the first–order ODE system
+    $V(X)=1,\qquad V(Y)=\lambda$
+    with prescribed initial conditions.
+  5. 5.
+    Express the system in (X,Y)(X,Y)–coordinates and recompute the induced structure polynomial.
+The resulting structure is rigid by construction.
+9.2.6 Effect on the Vekua equation
+Under rigidization, the variable–structure Vekua equation
+$\bar{\partial}W+AW+B\widehat{W}=F$
+is transformed into a rigid Vekua equation with $G\equiv 0$. All intrinsic obstruction terms are absorbed into the coordinate change.
+9.2.7 Remarks
+  • Rigidization requires solving only first–order ODEs; no Beltrami equation or singular integral is involved.
+  • The construction remains stable even in the presence of non–uniform ellipticity, provided $\Delta>0$.
+  • Rigidization separates geometric transport from analytic solving: once rigid, the system admits the full rigid Cauchy–Vekua calculus.
+9.3 Rigid Similarity Principle (Sharp Case $F=0$, $B=0$)
+Throughout this section we assume that the elliptic structure is rigid, i.e.
+$G\equiv 0,$
+so that the variable–structure Cauchy–Riemann operator
+$\bar{\partial}:=\tfrac{1}{2}(\partial_{x}+i\partial_{y})$
+satisfies the Leibniz rule on algebra–valued functions.
+We consider the homogeneous rigid Vekua equation in the special case
+$\bar{\partial}W+AW=0,$ (9.9)
+where $A:\Omega\to A$ is a given algebra–valued coefficient and no conjugate coupling term is present.
+9.3.1 Similarity principle
+Theorem 9.5 (Rigid similarity principle: sharp form).
+Let $W$ solve (9.10) on a simply connected neighborhood. Let $\Phi$ be the integrating factor given by Lemma 9.13. Then the function
+$H:=\Phi^{-1}W$
+is rigid holomorphic:
+$\bar{\partial}H=0.$
+Equivalently,
+$W=\Phi\,H,\qquad\bar{\partial}H=0.$
+Proof.
+Using the Leibniz rule,
+$\bar{\partial}(\Phi^{-1}W)=(\bar{\partial}\Phi^{-1})W+\Phi^{-1}(\bar{\partial}W).$
+Since $\bar{\partial}\Phi^{-1}=-\Phi^{-1}(\bar{\partial}\Phi)\Phi^{-1}$ and $\bar{\partial}\Phi=-A\Phi$, we obtain
+$\bar{\partial}H=\Phi^{-1}AW+\Phi^{-1}(-AW)=0.$
+9.3.2 Consequences
+  • Every solution of $\bar{\partial}W+AW=0$ differs from a rigid holomorphic function by a universal multiplicative factor.
+  • Zeros of $W$ coincide (with multiplicity) with zeros of the rigid holomorphic function $H$.
+  • The identity principle holds: if $W$ vanishes on a set with an accumulation point, then $W\equiv 0$.
+  • Local regularity and growth properties of $W$ reduce to those of rigid holomorphic functions.
+9.3.3 Sharpness
+Remark 9.6 (Sharpness of the similarity principle).
+The assumption $B\equiv 0$ is essential.
+If a conjugate coupling term $B\widehat{W}$ is present, no universal solution–independent multiplicative factor can eliminate it.
+This obstruction is algebraic and already appears in the classical Vekua equation.
+In that case, similarity must either depend on the solution itself or be formulated using generating pairs.
+9.3.4 Interpretation
+In the rigid regime, equation (9.10) differs from the rigid Cauchy–Riemann equation only by a transport–type lower–order term.
+The similarity principle shows that this term can be absorbed by a universal integrating factor, restoring genuine holomorphic behavior.
+This result is optimal and coincides exactly with the sharp classical theory.
+9.4 Rigid Similarity Principle (Spectral Formulation)
+In this section we establish the similarity principle in its sharp and structurally correct form, valid in the rigid regime under the hypotheses
+$F\equiv 0,\qquad B\equiv 0.$
+We emphasize that all arguments are local and rely only on smoothness of the coefficients.
+9.4.1 Setting and assumptions
+Let $\Omega\subset\mathbb{R}^{2}$ be open. We assume that the elliptic structure $(\alpha,\beta)$ is rigid, so that the intrinsic obstruction
+$G:=i_{x}+i\,i_{y}$
+vanishes identically. Equivalently, the associated spectral slope
+$\lambda:=\frac{-\beta+i\sqrt{\Delta}}{2},\qquad\Delta=4\alpha-\beta^{2}>0,$
+satisfies the homogeneous transport equation
+$\lambda_{x}+\lambda\,\lambda_{y}=0.$
+We consider the homogeneous rigid Vekua equation in the special case
+$\bar{\partial}W+AW=0,$ (9.10)
+where
+$\bar{\partial}:=\tfrac{1}{2}(\partial_{x}+i\partial_{y}),$
+and $A:\Omega\to A$ is a given smooth algebra–valued coefficient.
+9.4.2 Real formulation and spectral transport
+Write
+$W=U+iV,\qquad A=A_{0}+A_{1}i,$
+with $U,V,A_{0},A_{1}$ real–valued. Using rigidity, the equation (9.10) is equivalent to the real first–order system
+$\begin{cases}U_{x}-\alpha V_{y}+2(AW)_{0}=0,\\ V_{x}+U_{y}-\beta V_{y}+2(AW)_{1}=0.\end{cases}$ (9.11)
+The principal symbol of (9.11) coincides with that of the rigid Cauchy–Riemann operator.
+Diagonalizing this symbol using the spectral slope $\lambda$, one obtains real transport directions given by the vector fields
+$V_{\lambda}:=\partial_{x}+\lambda\,\partial_{y},\qquad V_{\bar{\lambda}}:=\partial_{x}+\bar{\lambda}\,\partial_{y}.$
+Because the structure is rigid, these vector fields are compatible and define smooth characteristic foliations.
+Thus equation (9.11) is a linear first–order system whose propagation occurs along the spectral characteristic curves determined by $\lambda$.
+9.4.3 Integrating factor
+We now construct a universal integrating factor.
+Lemma 9.7 (Integrating factor).
+There exists a neighborhood $U\subset\Omega$ and a unique algebra–valued function $\Phi\in C^{1}(U;A)$ such that
+$\bar{\partial}\Phi+A\Phi=0,\qquad\Phi(z_{0})=1,$ (9.12)
+for a prescribed base point $z_{0}\in U$. Moreover, $\Phi$ is pointwise invertible on $U$.
+Proof.
+Expanding (9.16) into real and imaginary parts yields a linear first–order real system with the same principal symbol as (9.11).
+After diagonalization in the spectral basis, this system reduces to a family of linear ordinary differential equations along the real characteristic curves generated by $V_{\lambda}$ and $V_{\bar{\lambda}}$.
+Since the coefficients are smooth, classical ODE theory yields a unique local solution with the prescribed initial value $\Phi(z_{0})=1$.
+Invertibility follows because $1$ is invertible in the elliptic algebra and invertibility is an open condition.
+∎
+9.4.4 Similarity principle
+Theorem 9.8 (Rigid similarity principle: sharp case).
+Let $W$ solve (9.10) on $U$, and let $\Phi$ be the integrating factor given by Lemma 9.7. Then the function
+$H:=\Phi^{-1}W$
+is rigid holomorphic:
+$\bar{\partial}H=0.$
+Equivalently,
+$W=\Phi\,H,\qquad\bar{\partial}H=0.$
+Proof.
+Using the Leibniz rule, valid under rigidity,
+$\bar{\partial}(\Phi^{-1}W)=(\bar{\partial}\Phi^{-1})W+\Phi^{-1}(\bar{\partial}W).$
+Since $\bar{\partial}\Phi=-A\Phi$, one has
+$\bar{\partial}\Phi^{-1}=\Phi^{-1}A.$
+Substituting into the equation and using (9.10) gives
+$\bar{\partial}H=\Phi^{-1}AW+\Phi^{-1}(-AW)=0.$
+9.4.5 Consequences
+  • Every solution of (9.10) differs from a rigid holomorphic function by a universal multiplicative factor.
+  • Zeros of $W$ coincide, with multiplicity, with zeros of $H$.
+  • The identity principle holds: if $W$ vanishes on a set with an accumulation point, then $W\equiv 0$.
+  • Local regularity and growth properties of $W$ reduce to those of rigid holomorphic functions.
+9.4.6 Sharpness and comparison with classical Vekua theory
+Remark 9.9 (Sharpness).
+The assumption $B\equiv 0$ is essential.
+If a conjugate coupling term $B\widehat{W}$ is present, no universal solution–independent multiplicative factor can eliminate it.
+This obstruction is algebraic and already present in the classical Vekua equation.
+Remark 9.10 (Relation with classical theory).
+In the classical theory, similarity is established using weak right inverses of the Cauchy–Riemann operator and exponential reconstruction, reflecting the low regularity setting.
+In the present smooth rigid framework, the same phenomenon appears as transport along spectral characteristic curves, and the exponential map is replaced by elementary first–order propagation.
+9.5 Initial Value Problems for Rigid Vekua Systems
+Throughout this section we assume that the elliptic structure is rigid,
+$G\equiv 0,$
+so that the variable–structure Cauchy–Riemann operator
+$\bar{\partial}=\tfrac{1}{2}(\partial_{x}+i\partial_{y})$
+satisfies the Leibniz rule and the full rigid calculus applies.
+We study initial value problems for rigid Vekua systems and show that they are well posed and reducible to holomorphic initial value problems via the similarity principle.
+9.5.1 Rigid Vekua equation
+We consider the rigid Vekua equation
+$\bar{\partial}W+AW+B\widehat{W}=F$ (9.13)
+on a simply connected domain $\Omega$, where $A,B,F$ are given algebra–valued functions, continuous (or smoother) on $\Omega$.
+9.5.2 Admissible initial curves
+Let $\Gamma\subset\Omega$ be a smooth curve with a smooth parametrization $\gamma:[0,1]\to\Omega$.
+Definition 9.11 (Noncharacteristic curve).
+A curve $\Gamma$ is called noncharacteristic for the rigid Vekua equation if its tangent vector is nowhere proportional to the vector field
+$\partial_{x}+i\partial_{y}.$
+Equivalently, $\Gamma$ is noncharacteristic if it is not everywhere tangent to the level curves of rigid holomorphic functions.
+Any smooth curve transversal to the rigid holomorphic foliation is noncharacteristic.
+9.5.3 Initial value problem
+Given a noncharacteristic curve $\Gamma$ and prescribed initial data
+$W|_{\Gamma}=W_{0},$
+with $W_{0}$ continuous (or smoother) and algebra–valued, we consider the initial value problem
+$\begin{cases}\bar{\partial}W+AW+B\widehat{W}=F&\text{in a neighborhood of }\Gamma,\\ W=W_{0}&\text{on }\Gamma.\end{cases}$ (9.14)
+9.5.4 Reduction via the similarity principle
+Let $\Phi$ be the similarity factor solving
+$\bar{\partial}\Phi+A\Phi+B\widehat{\Phi}=0,\qquad\Phi(z_{0})=1,$
+as constructed in the previous section.
+Define
+$H:=\Phi^{-1}W.$
+Then $H$ satisfies
+$\bar{\partial}H=\Phi^{-1}F.$ (9.15)
+Thus the rigid Vekua initial value problem (9.14) is equivalent to a rigid holomorphic initial value problem with source term.
+9.5.5 Existence and uniqueness
+Theorem 9.12 (Local well–posedness).
+Let $\Gamma$ be a noncharacteristic curve and $W_{0}$ prescribed initial data.
+Then there exists a unique local solution $W$ of the rigid Vekua initial value problem (9.14) in a neighborhood of $\Gamma$.
+Proof.
+Equation (9.15) is a first–order inhomogeneous Cauchy–Riemann equation with continuous coefficients.
+Standard theory yields a unique local solution $H$ with prescribed initial values on $\Gamma$.
+Defining $W=\Phi H$ gives the unique solution of (9.14).
+∎
+Lemma 9.13 (Integrating factor).
+Assume the elliptic structure is rigid.
+Let $A\in C(\Omega;A)$ and fix a base point $z_{0}\in\Omega$.
+Then there exists a neighborhood $U\subset\Omega$ of $z_{0}$ and a unique algebra–valued function $\Phi\in C^{1}(U;A)$ satisfying
+$\bar{\partial}\Phi+A\Phi=0,\qquad\Phi(z_{0})=1.$ (9.16)
+Moreover, $\Phi$ is pointwise invertible on $U$.
+Proof.
+Since the structure is rigid, the operator
+$\bar{\partial}=\tfrac{1}{2}(\partial_{x}+i\partial_{y})$
+is a derivation and defines a first–order transport operator along the vector field
+$V:=\partial_{x}+i\partial_{y}.$
+Equation (9.16) is therefore equivalent to the transport equation
+$V(\Phi)=-2A\,\Phi.$ (9.17)
+Let $\gamma(s)$ be an integral curve of $V$ with $\gamma(0)=z_{0}$. Along $\gamma$, equation (9.17) reduces to a linear ordinary differential equation in the elliptic algebra:
+$\frac{d}{ds}\Phi(\gamma(s))=-2A(\gamma(s))\,\Phi(\gamma(s)).$
+Since $A$ is continuous, this ODE admits a unique local solution with $\Phi(\gamma(0))=1$.
+Varying the initial direction of $\gamma$ produces a unique local solution $\Phi$ defined on a neighborhood $U$ of $z_{0}$.
+Invertibility follows because $\Phi(z_{0})=1$ is invertible and invertibility is an open condition in an elliptic algebra. ∎
+9.5.6 Propagation of initial data
+The solution propagates along rigid holomorphic directions. In particular:
+  • Values of $W$ on $\Gamma$ determine $W$ uniquely in its rigid holomorphic domain of dependence.
+  • If $F\equiv 0$, zeros of $W$ propagate discretely, with multiplicity.
+  • Regularity of $W$ matches that of the coefficients and the initial data.
+9.5.7 Algorithmic viewpoint
+For computational purposes, the rigid initial value problem can be solved by the following procedure:
+  1. 1.
+    Solve the similarity equation for $\Phi$.
+  2. 2.
+    Transform initial data: $H_{0}=\Phi^{-1}W_{0}$ on $\Gamma$.
+  3. 3.
+    Solve the rigid holomorphic problem
+    $\bar{\partial}H=\Phi^{-1}F$
+    with initial data $H_{0}$.
+  4. 4.
+    Recover $W=\Phi H$.
+This separation of transport ($\Phi$) and analytic propagation ($H$) is a key computational advantage of rigidization.
+9.5.8 Remarks
+  • No characteristic degeneration occurs in the rigid case.
+  • The initial value problem is stable under perturbations of the data.
+  • All difficulties associated with variable ellipticity have been absorbed prior to this stage by rigidization.
+Chapter 10 The Rigidity–Flatness Theorem
+We study the Rigidity–Flatness Theorem.
+The result concerns the intrinsic geometry encoded in the principal symbol of a real first–order elliptic system on the plane.
+No prior familiarity with rigidity or flatness is assumed.
+Notational Convention
+This chapter uses the same intrinsic objects and notation as referred in Chapter 9. In particular, all geometric and analytic quantities are expressed in terms of the variable elliptic structure
+$(\alpha,\beta),\qquad\Delta=4\alpha-\beta^{2}>0,$
+and its associated distinguished generator
+$i=i(x,y),\qquad i^{2}+\beta i+\alpha=0.$
+Spectral slope
+The complex root of the structure polynomial is denoted throughout by
+$\lambda=\frac{-\beta+i\sqrt{\Delta}}{2},\qquad\Im\lambda>0.$
+This quantity coincides with the spectral slope and is the unique complex characteristic direction of the principal symbol.
+Identification with previous notation. Any occurrence of a complex root written as
+$\tau=p+iq$
+in earlier formulations is to be understood as
+$\tau\equiv\lambda,\qquad p=\Re\lambda,\quad q=\Im\lambda>0.$
+No additional structure is introduced by this change of notation.
+Transport invariant (rigidity)
+The transport invariant is written intrinsically as
+$T:=\lambda_{x}+\lambda\,\lambda_{y}.$
+The elliptic structure (equivalently, the principal symbol) is called rigid if
+$T\equiv 0.$
+This condition is equivalent to vanishing of the intrinsic obstruction
+$G:=i_{x}+i\,i_{y}$
+Canonical metric (flatness)
+The canonical Riemannian metric associated with the principal symbol is written in terms of $\lambda=p+iq$ as
+$ds^{2}=\frac{1}{\Im\lambda}\Bigl(dx^{2}+2\,\Re\lambda\,dx\,dy+|\lambda|^{2}\,dy^{2}\Bigr).$
+This metric depends only on $(\alpha,\beta)$ and coincides with the metric induced by the rigid elliptic algebra. Let $K$ denote its Gaussian curvature.
+The principal symbol is called flat if
+$K\equiv 0.$
+Summary of intrinsic invariants
+With this alignment, the two intrinsic invariants of a real first–order elliptic system are:
+  • the transport invariant
+    $T=\lambda_{x}+\lambda\,\lambda_{y},$
+    measuring rigidity;
+  • the Gaussian curvature $K$ of the canonical metric, measuring flatness.
+Both invariants depend only on the principal symbol and are invariant under changes of dependent variables and orientation–preserving coordinate changes.
+All subsequent statements, computations, and proofs in this chapter are understood with $\lambda$ as the spectral slope associated to the elliptic structure $(\alpha,\beta)$, in full consistency with the Vekua formulation.
+10.1 The complex root and intrinsic field
+The equation $Q(1,\tau)=0$ has two complex conjugate roots. We select the one with positive imaginary part:
+$\tau=p+iq=\frac{-\beta+i\sqrt{4\alpha-\beta^{2}}}{2},\qquad q>0.$
+This complex-valued function $\tau(x,y)$ encodes the intrinsic geometry of the principal symbol.
+10.1.1 Two intrinsic invariants
+Two fundamental invariants arise naturally from $\tau$.
+1. The canonical metric.
+Define
+$ds^{2}=\frac{1}{q}\Bigl(dx^{2}+2p\,dx\,dy+(p^{2}+q^{2})\,dy^{2}\Bigr).$
+This Riemannian metric depends only on the principal symbol. Let $K$ denote its Gaussian curvature.
+2. The transport invariant.
+Define the first-order quantity
+$T=\tau_{x}+\tau\tau_{y}.$
+This expression measures how the symbol varies along its own characteristic direction. It depends only on the first derivatives of $\tau$ and hence only on the principal symbol.
+10.1.2 Rigidity and flatness
+We introduce two geometric conditions on the principal symbol:
+  • The system is rigid if $T\equiv 0$.
+  • The system is flat if the canonical metric has $K\equiv 0$.
+Both properties are intrinsic and independent of lower-order terms of the PDE.
+10.2 Statement of the theorem
+Theorem 10.1 (Rigidity–Flatness Theorem).
+Let $\mathcal{L}$ be a smooth real first–order elliptic system on a connected planar domain $\Omega$.
+If its principal symbol is rigid ($T=0$) and flat ($K=0$), then the complex root $\tau$ is constant on $\Omega$.
+Equivalently, the coefficients $\alpha$ and $\beta$ of the structure polynomial are constant, so the principal symbol is locally equivalent to that of a constant–coefficient elliptic system.
+10.2.1 Jet formulation
+Write $\tau=p+iq$ with $q>0$. At a point $z\in\Omega$, the 22–jet of $(p,q)$ is
+$j^{2}_{z}(p,q)=(p,q,p_{x},p_{y},q_{x},q_{y},p_{xx},p_{xy},p_{yy},q_{xx},q_{xy},q_{yy})|_{z}.$
+All second–order invariants of the symbol, including $K$, are rational functions of this jet.
+Rigidity $T=0$ is equivalent to
+$p_{x}=-p\,p_{y}+q\,q_{y},\qquad q_{x}=-p\,q_{y}-q\,p_{y}.$ (10.1)
+Differentiating (10.1) eliminates all xx–derivatives of order $\leq 2$. On the rigid locus, the free jet variables reduce to
+$(p,q,p_{y},q_{y},p_{yy},q_{yy}).$
+10.2.2 Jet-level curvature identity
+Define the curvature numerator
+$N:=-2q^{3}K.$
+Lemma 10.2 (Rigid curvature identity).
+On the rigid locus,
+$N=-q\,\Lambda(p,q)\,q_{yy}+R(p,q;p_{y},q_{y}),$
+where $\Lambda(p,q)>0$ for $q>0$, and
+$R=A(p,q)p_{y}^{2}+2B(p,q)p_{y}q_{y}+C(p,q)q_{y}^{2}$
+is positive definite.
+After clearing denominators this becomes a polynomial identity in finitely many jet variables and can therefore be symbolically verified.
+10.2.3 Consequences of flatness
+If $T=0$ and $K=0$, then $N=0$, so
+$q\,\Lambda(p,q)\,q_{yy}=R(p,q;p_{y},q_{y})\geq 0.$
+10.2.4 Characteristic representation
+Rigidity implies the complex Burgers equation
+$\tau_{x}+\tau\tau_{y}=0.$
+Introduce characteristic coordinates
+$\zeta=y-\tau x.$
+Locally there exists $C^{1}$ $\Phi$ such that
+$\tau(x,y)=\Phi(\zeta).$
+Differentiating,
+$\tau_{y}=\frac{\Phi^{\prime}}{1+x\Phi^{\prime}},\qquad\tau_{yy}=\frac{\Phi^{\prime\prime}}{(1+x\Phi^{\prime})^{3}}.$
+10.2.5 Degree argument
+Restrict to a fixed characteristic slice $\zeta=\mathrm{const}$.
+Then $\Phi^{\prime}$ and $\Phi^{\prime\prime}$ are constants with respect to $x$.
+Multiplying $N=0$ by $|1+x\Phi^{\prime}|^{6}$ produces a polynomial in $x$.
+The term involving $q_{yy}$ has degree at most $3$, while the quadratic form $R$ produces degree $4$.
+Since $R$ is positive definite, the leading coefficient is proportional to $|\Phi^{\prime}|^{6}$ and must vanish.
+Hence $\Phi^{\prime}=0$.
+Therefore $\tau_{y}\equiv 0$, and rigidity gives $\tau_{x}=0$. Thus $\tau$ is constant.
+Proof of Theorem 10.1.
+Combine Lemma 10.2 with the characteristic representation and the degree argument above. ∎
 Conclusion
-This monograph developed a variable function theory on planar domains in which the imaginary unit is no longer a fixed algebraic constant but a geometric object that varies with position.
-The guiding principle throughout has been that the symbol $i$ should be treated as intrinsic data: a moving generator of a rank–two real algebra bundle whose defining structure reduction
-$i^{2}+\beta(x,y)\,i+\alpha(x,y)=0$
-is allowed to vary.
-Once this perspective is adopted, the derivatives $i_{x}$ and $i_{y}$ are no longer formal artifacts but genuine geometric quantities, and their algebraic resolution becomes the natural starting point of the theory.
-A central object emerging from this viewpoint is the intrinsic combination
-$i_{x}+i\,i_{y}.$
+This monograph set out from a simple but decisive shift in viewpoint: the imaginary unit should not be treated as a fixed algebraic constant, but as intrinsic geometric data.
+Allowing the generator $i(x,y)$ of a rank–two real algebra bundle to vary forces the algebra itself to become part of the geometry of the plane.
+Once this step is taken, differentiation of the structure is no longer optional, and the derivatives $i_{x}$ and $i_{y}$ emerge as genuine geometric quantities rather than formal artifacts.
+The central object organizing the theory is the intrinsic obstruction
+$G:=i_{x}+i\,i_{y}.$
 This quantity appears prior to any analytic notion of holomorphicity.
-It governs the variability of the structure itself and, as a consequence, controls both the inhomogeneity of the generalized Cauchy–Riemann system and the failure of the Leibniz rule for the operator $\partial_{\bar{z}}$.
-These phenomena are not independent: they are different manifestations of the same structural obstruction produced by differentiating the moving generator.
-A fundamental outcome of the analysis is that the evolution of the structure coefficients $(\alpha,\beta)$ is governed by a first–order transport law that is already present at the structural level.
-By differentiating the structure reduction and eliminating the derivatives of $i$, the coefficient dynamics admits a reformulation in terms of the spectral parameter
+It measures the incompatibility between differentiation in the base variables and multiplication in the fibers, and it governs, simultaneously and unavoidably, all deviations from the classical theory: the inhomogeneity of generalized Cauchy–Riemann systems, the failure of the Leibniz rule, and the appearance of source terms in first–order structure equations.
+These phenomena are not imposed by analytic assumptions; they are forced by the geometry of a moving elliptic algebra.
+A first structural outcome of the theory is that variability of the coefficients $(\alpha,\beta)$ is not arbitrary.
+Eliminating the derivatives of the generator from the structure reduction leads to a universal transport law for the spectral parameter
 $\lambda=\frac{-\beta+i\sqrt{4\alpha-\beta^{2}}}{2},\qquad\Im\lambda>0,$
-leading, in full generality, to a (possibly forced) complex Burgers equation written in the standard complex algebra.
-This transport formulation is intrinsic and unavoidable: it reflects the fact that allowing the elliptic structure to vary forces its coefficients to evolve along characteristic curves.
-Rigidity occupies a precise and limited position within this framework.
-It is not the source of the Burgers dynamics, but rather the distinguished conservative regime of the universal transport law.
-The vanishing condition
+which takes the form of a (possibly forced) complex inviscid Burgers equation written in the standard complex algebra.
+This transport equation belongs to the structural layer of the theory: it arises solely from allowing the elliptic structure to vary and does not rely on integrability, coordinate trivialization, or function–theoretic input.
+Within this universal transport framework, rigidity occupies a precise and limited role. The condition
 $i_{x}+i\,i_{y}=0$
-selects exactly those variable elliptic structures for which the Burgers equation is unforced.
-In this conservative case, several obstructions collapse simultaneously: the generalized Cauchy–Riemann operator becomes a derivation, the real Cauchy–Riemann system becomes homogeneous, and the generator $i$ is transported compatibly with its own multiplication law.
-Rigidity is therefore not an auxiliary hypothesis but a compatibility condition that restores integrability within an already dynamical setting.
-The explicit $\varepsilon$–family of rigid elliptic structures illustrates this conservative regime in concrete terms.
-It shows that rigidity does not force triviality: genuinely nonconstant structures exist, though ellipticity naturally restricts their domain of definition.
-From the transport viewpoint, these examples reflect the familiar behavior of first–order hyperbolic systems, where characteristic crossing limits global regularity and constrains smooth extensions.
-On the analytic side, rigidity is exactly the condition under which an integral representation theory can persist.
-In the rigid regime, a Cauchy–Pompeiu formula adapted to variable elliptic structures is recovered.
-The natural Cauchy 11–form
-$d\tilde{z}=dy-i(z)\,dx$
-is not closed, and its non–closure produces an unavoidable interior correction term. Rather than being an artifact, this correction singles out a canonical first–order operator
-$Df=\partial_{\bar{z}}f+\tfrac{1}{2}f\,i_{y},$
-whose null space provides the appropriate notion of holomorphicity in the variable setting.
-Although covariantly holomorphic sections are not closed under the pointwise product, a scalar gauge transformation restores closure through a natural weighted multiplication, showing how analytic structure can be recovered by compensating for variability rather than suppressing it.
-The scope of the present work is deliberately local and explicit.
-Even in the conservative regime, global questions remain subtle: the choice of a continuous generator $i(z)$ is typically local, gauge trivializations may exhibit monodromy, and the transport nature of the coefficient dynamics explains why globally smooth structures are highly constrained.
-Beyond rigidity, the analytic function theory necessarily acquires source terms, while the transport–controlled evolution of the coefficients remains intact.
-Closing perspective.
-The central lesson of this work is that once the complex structure is allowed to move, its differentiation becomes a primary geometric invariant.
-Transport comes first, integrability second, and analytic closure emerges only as a distinguished limiting case.
-Rigid elliptic structures do not generate the dynamics; they represent the rare situation in which the universal transport law becomes conservative and the analytic and algebraic layers of the theory align.
-In this way, variable elliptic structures reveal a natural hierarchy in which geometry dictates dynamics, and dynamics, in turn, governs analysis.
-Acknowledgements
-Use of Generative AI Tools.
-Portions of the writing and editing of this manuscript were assisted by generative AI language tools.
-These tools were used to improve clarity of exposition, organization of material, and language presentation.
-All mathematical results, statements, proofs, and interpretations were developed, verified, and validated by the author.
-The author takes full responsibility for the accuracy, originality, and integrity of all content in this work, including any material produced with the assistance of AI tools.
-No generative AI system is listed as an author of this work.
+does not generate the dynamics; it selects the conservative regime of the already–existing transport law.
+In this regime, the Burgers equation becomes unforced, and several independent obstructions collapse simultaneously: the generalized Cauchy–Riemann operator becomes a derivation, the real first–order system becomes homogeneous, and the generator is transported compatibly with its own multiplication law.
+Rigidity is therefore a compatibility condition, not a simplifying assumption.
+The explicit $\varepsilon$–family constructed in this work demonstrates that rigidity does not force triviality.
+Genuinely nonconstant elliptic structures exist, although ellipticity naturally restricts the domain of definition.
+From the transport viewpoint, these restrictions reflect characteristic phenomena familiar from first–order hyperbolic systems and are not pathologies of the theory.
+On the analytic side, rigidity is exactly the threshold at which a coherent function theory re–emerges.
+In the rigid regime, a Cauchy–Pompeiu formula adapted to variable elliptic structures persists.
+The natural Cauchy 11–form is not closed, and its non–closure produces an unavoidable interior correction term.
+Rather than being an artifact, this correction identifies a canonical covariant Cauchy–Riemann operator whose null space provides the appropriate notion of holomorphicity in the variable setting.
+Although covariantly holomorphic sections are not closed under pointwise multiplication, a scalar gauge transformation restores closure through a natural weighted product.
+Power series representations follow from this framework, but only after freezing the algebra at the expansion point, making explicit where classical arguments genuinely fail in the variable case.
+Finally, the theory interfaces naturally with the classical Vekua program.
+General first–order elliptic systems induce variable elliptic structures, and their reduction to Vekua form reveals the intrinsic obstruction as the true geometric source of inhomogeneity.
+Rigidization, formulated as an ordinary differential equation along characteristic curves, provides a constructive mechanism to pass from non–rigid to rigid structures without solving global Beltrami equations.
+In the rigid regime, sharp similarity principles, initial value problems, and algorithmic solution schemes become available, even in the presence of non–uniform ellipticity.
+The scope of this work has been deliberately local and explicit.
+Rather than appealing to abstract integrability frameworks, the theory has been developed at the level of concrete first–order computations in order to keep the geometric and dynamical mechanisms fully visible.
+From this perspective, variable elliptic structures exhibit a clear hierarchy: transport is fundamental, rigidity is selective, and analytic closure is a distinguished consequence rather than an axiom.
+This hierarchy reframes classical complex analysis as a special case of a broader geometric theory.
+A fixed complex structure gives analysis. A variable one gives transport. Rigidity is where the two agree to dance.
+Here, the algebra begins to move.
 Appendix A Algebraic Universality of the Intrinsic Obstruction and the Leibniz Defect

### v4

@@
 rigidity is used to ensure $\partial_{\bar{z}}(Z^{-1})=0$ off the diagonal; it does not imply $i_{y}=0$.
-Chapter 6 Covariant Holomorphicity and Gauge Structure
-Covariant Holomorphicity and Gauge Structure in the Rigid Regime
+Chapter 6 Covariant Holomorphicity and Gauge Structure in the Rigid Regime
 6.1 Motivation from the Cauchy–Pompeiu Formula
@@
 and this identity is the bridge between holomorphicity and elliptic theory.
-In the present framework, once rigidity is imposed Chapter  2, $\partial_{\bar{z}}$ becomes a derivation and the Cauchy–Pompeiu formula (5) becomes available.
+In the present framework, once rigidity is imposed (Chapter 2), $\partial_{\bar{z}}$ becomes a derivation and the Cauchy–Pompeiu formula (Chapter 5) becomes available.
 It is then natural—and unavoidable—to study the second–order operator obtained by composing $\partial_{\bar{z}}$ with its conjugate partner $\partial_{z}$.
@@
 $\Delta:=4\alpha-\beta^{2}=\frac{4}{a_{11}^{2}}\,\Delta_{\mathrm{sys}}>0,$
-and $\alpha,\beta$ determine a variable elliptic algebra
+and $\alpha,\beta$ determine a variable elliptic algebra bundle
 $A_{z}=\mathbb{R}[X]/(X^{2}+\beta(z)X+\alpha(z)).$
@@
 $\widehat{i}:=-\beta-i.$
+For an algebra–valued field $W=U+iV$ with $U,V$ real, we write
+$\widehat{W}:=U+\widehat{i}\,V=U-(\beta+i)V.$
 9.1.2 Canonical principal part
@@
 9.1.4 Elliptic conjugation and Vekua form
-Using
-$U=\tfrac{1}{2}(W+\widehat{W}),\qquad V=\frac{W-\widehat{W}}{2i+\beta},$
-the right–hand side of (9.5) can be written uniquely as
-$F-AW-B\widehat{W}$,
-with
-$F:=\tfrac{1}{2}(f+ig),$
-and coefficients
-$A$ $=A_{0}-\frac{1}{2}\,\frac{G}{2i+\beta},$
-$B$ $=B_{0}+\frac{1}{2}\,\frac{G}{2i+\beta},$
+The conjugate field $\widehat{W}$ satisfies the linear identities
+$W-\widehat{W}=(2i+\beta)V,\qquad W+\widehat{W}=2U-\beta V.$ (9.6)
+Since $\Delta>0$, the element $2i+\beta$ is invertible in $A_{z}$ and in fact
+$(2i+\beta)^{-1}=\frac{\beta-2i}{\beta^{2}+4\alpha}.$
+Therefore we may solve (9.6) for $U$ and $V$ in terms of (W,W^)($\widehat{W}$):
+$V=\frac{W-\widehat{W}}{2i+\beta},\qquad U=\frac{1}{2}\left(W+\widehat{W}+\beta\,\frac{W-\widehat{W}}{2i+\beta}\right).$ (9.7)
+Substituting (9.7) into the right–hand side of (9.5) expresses it as an $A_{z}$–linear combination of $W$ and $\widehat{W}$.
+Concretely, there exist unique algebra–valued coefficients $A_{0},B_{0}$ and a unique datum $F_{0}$ such that
+$(f+ig)-(a+ic)U-(b+id)V\;=\;2F_{0}-2A_{0}\,W-2B_{0}\,\widehat{W}$. (9.8)
+(Existence and uniqueness follow because (9.7) is an invertible linear change of variables between $(U,V)$ and (W,W^)($\widehat{W}$).)
+Define
+$F:=F_{0}=\tfrac{1}{2}(f+ig),$
+and absorb the obstruction term $VG$ from (9.5) using $V=(W-\widehat{W})/(2i+\beta)$:
+$VG=\frac{W-\widehat{W}}{2i+\beta}\,G.$
+Thus (9.5) may be rewritten uniquely in the form
+$2\bar{\partial}W=2F-2A\,W-2B\,\widehat{W}$, (9.9)
 where
-$A_{0}$ $=\frac{1}{4}\Big(a-\frac{\beta}{\alpha}b+d+i\,c-i\,\frac{1}{\alpha}b\Big),$
-$B_{0}$ $=\frac{1}{4}\Big(a+\frac{\beta}{\alpha}b-d+i\,c+i\,\frac{1}{\alpha}b\Big).$
-Dividing (9.5) by $2$, we arrive at the canonical variable–structure Vekua equation
+$A\;:=\;A_{0}-\frac{1}{2}\,\frac{G}{2i+\beta},\qquad B\;:=\;B_{0}+\frac{1}{2}\,\frac{G}{2i+\beta}.$ (9.10)
+Dividing by $2$ yields the canonical variable–structure Vekua equation
 $\bar{\partial}W+AW+B\widehat{W}=F.$ (9.11)
+Remark 9.1.
+[…]
 9.1.5 Remarks
   • Equation (9.11) is intrinsic to the elliptic structure $(\alpha,\beta)$ and requires no auxiliary reduction.
-  • When the structure is rigid ($G\equiv 0$), the coefficients reduce to their classical forms.
-  • The obstruction $G$ measures precisely the deviation from rigid Cauchy–Riemann calculus and is absorbed into the lower–order terms.
+  • When the structure is rigid ($G\equiv 0$), the obstruction contribution in (9.10) vanishes and the lower–order coefficients reduce to the rigid forms determined by (9.8).
+  • The obstruction $G$ measures precisely the deviation from rigid Cauchy–Riemann calculus and is absorbed into the lower–order terms via the explicit splitting in (9.10).
 9.2 Rigidization
@@
 Throughout this section we work locally on $\Omega\subset\mathbb{R}^{2}$ and assume
-$\alpha,\beta\in C^{1}(\Omega),\qquad\Delta=4\alpha-\beta^{2}>0.$
+$\alpha,\beta\in C^{2}(\Omega),\qquad\Delta=4\alpha-\beta^{2}>0.$
 9.2.1 Rigid and non–rigid structures
@@
 where $H$ is a complex–valued function canonically determined by $G=G_{0}+G_{1}\,i$. In particular,
-$G\equiv 0\quad\Longleftrightarrow\quad\lambda_{x}+\lambda\,\lambda_{y}=0.$
+$G\equiv 0\;\Longleftrightarrow\;\lambda_{x}+\lambda\,\lambda_{y}=0.$
 Thus rigidity is equivalent to flat transport of the spectral slope.
@@
 A rigidization is a local diffeomorphism
-$\Phi:(x,y)\mapsto(X,Y)$
+$\Phi:(x,y)\longmapsto(X,Y)$
 such that, when the elliptic structure is pulled back to (X,Y)(X,Y)–coordinates, the resulting structure is rigid.
-Rigidization is an Eulerian construction: the diffeomorphism is generated by an ODE driven by the transport field of the structure.
-9.2.4 Rigidization ODE
-Let $\lambda(x,y)$ be the spectral slope of the original structure. Define the vector field
-$V:=\partial_{x}+\lambda\,\partial_{y}.$
-Proposition 9.4 (Rigidizing flow).
-Let $(X(x,y),Y(x,y))$ solve the first–order system
-$\begin{cases}V(X)=1,\\ V(Y)=\lambda.\end{cases}$ (9.8)
-Then $(X,Y)$ defines a local diffeomorphism, and in the (X,Y)(X,Y)–coordinates the transport equation becomes
-$\lambda_{X}+\lambda\,\lambda_{Y}=0.$
+9.2.4 The characteristic operator and its square
+The key analytic object is the characteristic operator
+$V:=\partial_{x}+\lambda\,\partial_{y},$
+where $\lambda=a+ib$ with $a=\operatorname{Re}\lambda$ and $b=\operatorname{Im}\lambda>0$.
+Under a real change of coordinates $(x,y)\mapsto(X,Y)$, the operator $V$ acts on the new coordinate functions by
+$V(X)=X_{x}+\lambda\,X_{y},\qquad V(Y)=Y_{x}+\lambda\,Y_{y}.$
+The transformed spectral slope is
+$\widetilde{\lambda}\;=\;\frac{V(Y)}{V(X)},$ (9.13)
+and the rigidization condition $\widetilde{\lambda}_{X}+\widetilde{\lambda}\,\widetilde{\lambda}_{Y}=0$ is equivalent to the ratio condition
+$V\!\bigl(\widetilde{\lambda}\bigr)=0\qquad\Longleftrightarrow\qquad V^{2}(Y)\cdot V(X)\;=\;V(Y)\cdot V^{2}(X).$ (9.14)
+The iterated operator $V^{2}=(\partial_{x}+\lambda\,\partial_{y})^{2}$ admits a canonical decomposition when applied to a real–valued function $\varphi\in C^{2}(\Omega,\mathbb{R})$.
+A direct computation using $\lambda=a+ib$ yields
+$V^{2}(\varphi)\;=\;W(\varphi)\;+\;i\,T(\varphi),$ (9.15)
+where $T$ and $W$ are the real operators defined below.
+Proposition 9.5 ($V^{2}$ decomposition).
+Let $\varphi\in C^{2}(\Omega,\mathbb{R})$ and write $H=H_{R}+iH_{I}$ for the forcing (9.12). Then
+$T(\varphi)$ $=2b\Bigl[(\varphi_{y})_{x}+a\,(\varphi_{y})_{y}+\frac{H_{I}}{2b}\,\varphi_{y}\Bigr],$ (9.16)
+$W(\varphi)$ $=\varphi_{xx}-|\lambda|^{2}\,\varphi_{yy}+\Bigl(H_{R}-\frac{a\,H_{I}}{b}\Bigr)\varphi_{y}.$ (9.17)
 Proof.
-Along the flow of $V$, equation (9.8) ensures that $X$ is the flow parameter and $Y$ is transported consistently with $\lambda$.
-By construction, derivatives along $V$ become $\partial_{X}$, eliminating the source term in (9.7).
+A direct expansion gives
+$V^{2}(\varphi)=\varphi_{xx}+2\lambda\,\varphi_{xy}+\lambda^{2}\,\varphi_{yy}+(\lambda_{x}+\lambda\,\lambda_{y})\,\varphi_{y}.$
+The imaginary part of this expression, using $\lambda^{2}=(a^{2}-b^{2})+2abi$ and $\lambda_{x}+\lambda\,\lambda_{y}=H_{R}+iH_{I}$, is
+$\operatorname{Im}V^{2}(\varphi)=2b\,\varphi_{xy}+2ab\,\varphi_{yy}+H_{I}\,\varphi_{y},$
+which coincides with (9.16). Setting this equal to zero and eliminating $\varphi_{xy}$ from the real part yields (9.17). ∎
+9.2.5 Transport and wave structure
+The two components in Proposition 9.5 have distinct and complementary PDE character.
+Transport component. Setting $p:=\varphi_{y}$, equation $T(\varphi)=0$ becomes the first–order linear transport equation
+$p_{x}+a\,p_{y}=-\frac{H_{I}}{2b}\,p,$ (9.18)
+whose transport direction is the real vector field $(1,\,\operatorname{Re}\lambda)$.
+This equation is solvable by the method of real characteristics: along the curves $\tfrac{dy}{dx}=a(x,y)$, the function $p$ satisfies a linear ODE.
+Wave component. The equation $W(\varphi)=0$ is the real second–order equation
+$\varphi_{xx}-|\lambda|^{2}\,\varphi_{yy}+c\,\varphi_{y}=0,\qquad c:=H_{R}-\frac{a\,H_{I}}{b},$ (9.19)
+whose principal symbol $\xi^{2}-|\lambda|^{2}\eta^{2}=({\xi-|\lambda|\eta})({\xi+|\lambda|\eta})$ is hyperbolic, with real characteristic speeds $\pm|\lambda|^{-1}$.
+Neither component alone is elliptic.
+The ellipticity of the full rigidization problem arises from their coupling, in exact analogy with the Cauchy–Riemann equations $(u_{x}=v_{y},\;u_{y}=-v_{x})$, which are elliptic despite each equation individually describing simple transport.
+9.2.6 PDE type of the rigidization problem
+Theorem 9.6 (PDE type of rigidization).
+Let $\lambda=a+ib$ with $b>0$ satisfy $\lambda_{x}+\lambda\,\lambda_{y}=H\not\equiv 0$. The ratio condition (9.14), viewed as a second–order system for two real unknown functions $(X,Y)$, is elliptic.
+Proof.
+The principal symbol of the ratio condition inherits its type from $V^{2}$.
+The characteristic equation $(\xi+\lambda\eta)^{2}=0$ has solutions $\xi/\eta=-\lambda=-a-ib$, which are complex since $b>0$.
+Hence no real characteristic directions exist, and the system is elliptic.
 ∎
-9.2.5 Algorithmic description
-Rigidization can be implemented as the following local algorithm.
+Proposition 9.7 (Transport–transport incompatibility).
+If $T(X)=0$ and $T(Y)=0$ simultaneously, then the ratio condition (9.14) forces the Jacobian $J=X_{x}Y_{y}-X_{y}Y_{x}$ to vanish.
+Proof.
+Under $T(X)=T(Y)=0$, the values $V^{2}(X)=W(X)$ and $V^{2}(Y)=W(Y)$ are real. Since $V(X)$ and $V(Y)$ are complex, the imaginary part of (9.14) reads
+$W(Y)\cdot b\,X_{y}=W(X)\cdot b\,Y_{y}.$
+The real part similarly gives $W(Y)(X_{x}+aX_{y})=W(X)(Y_{x}+aY_{y})$. Cross–multiplying the two proportionalities yields $X_{y}Y_{x}=Y_{y}X_{x}$, hence $J=0$. ∎
+Remark 9.8.
+Proposition 9.7 shows that the transport and wave components cannot be decoupled for both coordinate functions simultaneously while preserving non–degeneracy.
+The elliptic coupling between $T$ and $W$ is load–bearing.
+9.2.7 Splitting method for rigidization
+Although the rigidization problem is elliptic, the decomposition $V^{2}=W+iT$ of Proposition 9.5 enables an alternating–direction iterative method in which each individual step uses only real characteristics.
+Algorithm (Transport–wave splitting).
+  1. 0.
+    Initialize $X^{0}=x$, $Y^{0}=y$ (identity map).
   2. 1.
-    Compute $\alpha,\beta$ from the principal part of the system.
+    Compute the forcing residual $H^{n}:=\widetilde{\lambda}^{\,n}_{X}+\widetilde{\lambda}^{\,n}\,\widetilde{\lambda}^{\,n}_{Y}$ in the current coordinates.
   3. 2.
-    Compute $\Delta=4\alpha-\beta^{2}$ and the spectral slope
-    $\lambda=\frac{-\beta+i\sqrt{\Delta}}{2}.$
+    Transport half–step.
+Correct $Y^{n}$ by solving the first–order transport equation (9.18) for $\delta Y_{y}$ along the real characteristics of $(1,\operatorname{Re}\lambda)$, targeting $\operatorname{Im}H^{n}\to 0$.
   4. 3.
-    Form the transport vector field
-    $V=\partial_{x}+\lambda\,\partial_{y}.$
+    Wave half–step. Correct $X^{n}$ by solving the hyperbolic wave equation (9.19) for $\delta X$ along the wave characteristics $\pm|\lambda|^{-1}$, targeting $\operatorname{Re}H^{n}\to 0$.
   5. 4.
-    Solve the first–order ODE system
-    $V(X)=1,\qquad V(Y)=\lambda$
-    with prescribed initial conditions.
+    Update: $X^{n+1}=X^{n}+\delta X$, $Y^{n+1}=Y^{n}+\delta Y$.
   6. 5.
-    Express the system in (X,Y)(X,Y)–coordinates and recompute the induced structure polynomial.
-The resulting structure is rigid by construction.
+    If $|H^{n+1}|$ is below the desired tolerance, stop; otherwise return to step 1.
+This procedure treats the elliptic coupling iteratively while solving only hyperbolic subproblems at each step.
+It is directly analogous to alternating–direction implicit (ADI) methods for elliptic equations and to Uzawa iteration for saddle–point systems.
+9.2.8 Relationship to Beltrami uniformization
+It is instructive to compare rigidization with full Beltrami uniformization.
+The latter seeks a diffeomorphism $\Phi$ such that $\Phi^{*}\lambda$ is constant, requiring the solution of the Beltrami equation $w_{\bar{z}}=\mu\,w_{z}$, an elliptic PDE whose general theory relies on singular integral operators (the Beurling–Ahlfors transform).
+Rigidization is strictly weaker: it asks only that $H\to 0$, while $\lambda$ is permitted to remain nonconstant.
+Both problems are elliptic, but the rigidization problem inherits additional structure from the Burgers self–transport (the characteristic speed equals the transported quantity), which constrains the effective Beltrami coefficient and enables the splitting method described above.
 9.2.9 Effect on the Vekua equation
@@
 9.2.10 Remarks
-  • Rigidization requires solving only first–order ODEs; no Beltrami equation or singular integral is involved.
+  • Rigidization does not reduce to solving first–order ODEs along characteristics.
+The ratio condition (9.14) is a genuinely elliptic problem for the coordinate functions $(X,Y)$.
+However, the transport–wave decomposition of $V^{2}$ provides a natural splitting in which each iterative half–step uses only real characteristics.
   • The construction remains stable even in the presence of non–uniform ellipticity, provided $\Delta>0$.
-  • Rigidization separates geometric transport from analytic solving: once rigid, the system admits the full rigid Cauchy–Vekua calculus.
+  • Rigidization separates geometric transport from analytic solving: once rigid, the system admits the full rigid Cauchy–Vekua calculus (Chapters 5–8).
+9.2.11 Geometric interpretation
+The decomposition $V^{2}=W+iT$ expresses, at the PDE level, the central theme of this monograph.
+The transport component $T$ acts on $\varphi_{y}$—the transverse gradient—and propagates along the real projection $(1,\operatorname{Re}\lambda)$ of the complex characteristic direction.
+It belongs to the structural layer of Chapter 2.
+The wave component $W$ acts on $\varphi$ itself and propagates along the characteristic speeds $\pm|\lambda|^{-1}$ of the rigid elliptic operator $L_{\alpha,\beta}$ of Chapter 8.
+It belongs to the analytic layer.
+Transport and integrability are therefore the two halves of the rigidization problem.
+Individually, each is hyperbolic; jointly, they are elliptic.
+Rigidization succeeds precisely when these two components can be made simultaneously compatible—the analytic expression of the passage from “transport first” to “integrability second” that governs the entire theory.
 9.3 Rigid Similarity Principle (Sharp Case $F=0$, $B=0$)
@@
   • All difficulties associated with variable ellipticity have been absorbed prior to this stage by rigidization.
-Chapter 10 The Rigidity–Flatness Theorem
-We study the Rigidity–Flatness Theorem.
-The result concerns the intrinsic geometry encoded in the principal symbol of a real first–order elliptic system on the plane.
-No prior familiarity with rigidity or flatness is assumed.
-Notational Convention
-This chapter uses the same intrinsic objects and notation as referred in Chapter 9. In particular, all geometric and analytic quantities are expressed in terms of the variable elliptic structure
-$(\alpha,\beta),\qquad\Delta=4\alpha-\beta^{2}>0,$
-and its associated distinguished generator
-$i=i(x,y),\qquad i^{2}+\beta i+\alpha=0.$
-Spectral slope
-The complex root of the structure polynomial is denoted throughout by
-$\lambda=\frac{-\beta+i\sqrt{\Delta}}{2},\qquad\Im\lambda>0.$
-This quantity coincides with the spectral slope and is the unique complex characteristic direction of the principal symbol.
-Identification with previous notation. Any occurrence of a complex root written as
-$\tau=p+iq$
-in earlier formulations is to be understood as
-$\tau\equiv\lambda,\qquad p=\Re\lambda,\quad q=\Im\lambda>0.$
-No additional structure is introduced by this change of notation.
-Transport invariant (rigidity)
-The transport invariant is written intrinsically as
-$T:=\lambda_{x}+\lambda\,\lambda_{y}.$
-The elliptic structure (equivalently, the principal symbol) is called rigid if
-$T\equiv 0.$
-This condition is equivalent to vanishing of the intrinsic obstruction
+Chapter 10 Convergence of the Rigidization Scheme
+The preceding chapter introduced the rigidization problem and showed that its PDE type is elliptic (Theorem 9.6).
+However, the convergence of the transport–wave splitting algorithm (Section 9.2.7) was left open.
+In this chapter we close this gap by establishing local convergence of rigidization via a Newton iteration on the nonlinear residual map.
+The analysis proceeds in three stages.
+First, we show that the linearized rigidization problem is an elliptic $2\times 2$ system whose principal symbol is explicitly computable (Section 10.1).
+Second, we prove that the naive transport–wave splitting has spectral radius $\geq 1$ and therefore does not converge as a standalone iteration (Section 10.3).
+Third, we establish quadratic convergence of Newton’s method applied to the full nonlinear rigidization problem, using only Schauder elliptic theory in $C^{2,\alpha}$ (Section 10.4).
+No $L^{p}$ machinery or singular integral operators are required.
+Throughout, we assume $\alpha,\beta\in C^{2,\alpha}(\bar{\Omega})$ with ellipticity $\Delta=4\alpha-\beta^{2}>0$ on $\bar{\Omega}$, and write $\lambda=a+ib$ with $b=\tfrac{1}{2}\sqrt{\Delta}>0$.
+The uniform ellipticity constant is
+$b_{\min}:=\min_{\bar{\Omega}}\operatorname{Im}\lambda>0.$
+10.1 Linearization of the residual map
+10.1.1 The residual at the identity
+For a diffeomorphism $\Phi:(x,y)\mapsto(X,Y)$, the transformed spectral slope is $\widetilde{\lambda}=V(Y)/V(X)$ and the rigidization residual is
+$\mathcal{H}[\Phi]:=\widetilde{\lambda}_{X}+\widetilde{\lambda}\,\widetilde{\lambda}_{Y}.$
+Using the ratio condition (equation (9.14)), this is equivalently encoded by the unnormalized residual
+$R[\Phi]:=V^{2}(Y)\,V(X)-V(Y)\,V^{2}(X).$
+Lemma 10.1 (Residual at the identity).
+At $\Phi^{0}=\operatorname{id}$,
+$V(x)=1,\quad V(y)=\lambda,\quad V^{2}(x)=0,\quad V^{2}(y)=H,$
+where $H:=\lambda_{x}+\lambda\,\lambda_{y}$ is the obstruction. Hence $R[\operatorname{id}]=H$ and $\mathcal{H}[\operatorname{id}]=H$.
+Proof.
+$V(x)=x_{x}+\lambda\,x_{y}=1$ and $V(y)=y_{x}+\lambda\,y_{y}=\lambda$.
+Then $V^{2}(x)=V(1)=0$ and $V^{2}(y)=V(\lambda)=\lambda_{x}+\lambda\,\lambda_{y}=H$.
+Substituting into the ratio condition gives $R[\operatorname{id}]=H\cdot 1-\lambda\cdot 0=H$.
+∎
+10.1.2 First variation
+Consider the perturbed diffeomorphism $\Phi^{\varepsilon}=\operatorname{id}+\varepsilon(f,g)$, where $f,g\in C^{2,\alpha}(\bar{\Omega},\mathbb{R})$ are the corrections to $X$ and $Y$ respectively.
+Proposition 10.2 (First variation of the residual).
+The linearized change in residual at $\Phi=\operatorname{id}$ is
+$\delta R=V^{2}(g)-\lambda\,V^{2}(f)+H\,V(f).$ (10.1)
+Proof.
+At $\Phi^{\varepsilon}=\operatorname{id}+\varepsilon(f,g)$:
+$V(X)$ $=1+\varepsilon\,V(f),$
+$V(Y)$ $=\lambda+\varepsilon\,V(g),$
+$V^{2}(X)$ $=\varepsilon\,V^{2}(f),$
+$V^{2}(Y)$ $=H+\varepsilon\,V^{2}(g).$
+Then
+$R[\Phi^{\varepsilon}]$ $=\bigl(H+\varepsilon\,V^{2}(g)\bigr)\bigl(1+\varepsilon\,V(f)\bigr)-\bigl(\lambda+\varepsilon\,V(g)\bigr)\bigl(\varepsilon\,V^{2}(f)\bigr)$
+$=H+\varepsilon\bigl[H\,V(f)+V^{2}(g)-\lambda\,V^{2}(f)\bigr]+O(\varepsilon^{2}).$
+The $O(1)$ term is $H=R[\operatorname{id}]$, so $\delta R=\tfrac{d}{d\varepsilon}\big|_{0}R[\Phi^{\varepsilon}]$ gives (10.1). ∎
+10.1.3 Real–imaginary decomposition
+For a real-valued function $\varphi\in C^{2}(\Omega,\mathbb{R})$, we have the decomposition $V^{2}(\varphi)=W(\varphi)+i\,T(\varphi)$ from Proposition 9.5, and $V(\varphi)=(\varphi_{x}+a\,\varphi_{y})+i\,b\,\varphi_{y}$.
+Decomposing (10.1) into real and imaginary parts:
+Proposition 10.3 (Linearized system).
+Write $H=H_{R}+iH_{I}$. The equation $\delta R=-H$ is the $2\times 2$ real system
+$W(g)-a\,W(f)+b\,T(f)+H_{R}(f_{x}+a\,f_{y})-H_{I}\,b\,f_{y}$ $=-H_{R},$ (10.2)
+$T(g)-a\,T(f)-b\,W(f)+H_{R}\,b\,f_{y}+H_{I}(f_{x}+a\,f_{y})$ $=-H_{I},$ (10.3)
+for two real unknowns $f$ and $g$.
+Proof.
+Write $V^{2}(g)=W(g)+i\,T(g)$, $V^{2}(f)=W(f)+i\,T(f)$, $V(f)=(f_{x}+a\,f_{y})+i\,b\,f_{y}$, and separate the real and imaginary parts of
+$V^{2}(g)-(a+ib)\bigl[W(f)+i\,T(f)\bigr]+(H_{R}+iH_{I})\bigl[(f_{x}+a\,f_{y})+i\,b\,f_{y}\bigr]=-H_{R}-iH_{I}.\qed$
+10.2 Ellipticity of the linearized system
+Theorem 10.4 (Ellipticity of the linearized rigidization).
+The principal symbol of the system (10.2)–(10.3), viewed as a second-order $2\times 2$ system for $(f,g)$, has determinant
+$\det\mathcal{P}(\xi,\eta)=b\,|\xi+\lambda\,\eta|^{4},$ (10.4)
+which is strictly positive for all $(\xi,\eta)\neq(0,0)$ when $b>0$. The system is therefore elliptic.
+Proof.
+The principal parts of (10.2)–(10.3) involve only $W$ and $T$. Recall that in the characteristic coordinates $u=\xi+a\,\eta$, $v=b\,\eta$:
+$W_{\mathrm{sym}}=u^{2}-v^{2},\qquad T_{\mathrm{sym}}=2uv.$
+The principal symbol matrix is
+$\mathcal{P}(\xi,\eta)=\begin{pmatrix}-a\,W_{\mathrm{sym}}+b\,T_{\mathrm{sym}}&W_{\mathrm{sym}}\\[4.0pt] -b\,W_{\mathrm{sym}}-a\,T_{\mathrm{sym}}&T_{\mathrm{sym}}\end{pmatrix}$.
+A direct computation gives
+$\det\mathcal{P}$ $=\bigl(-a\,W_{\mathrm{sym}}+b\,T_{\mathrm{sym}}\bigr)\,T_{\mathrm{sym}}-\bigl(-b\,W_{\mathrm{sym}}-a\,T_{\mathrm{sym}}\bigr)\,W_{\mathrm{sym}}$
+$=-a\,W_{\mathrm{sym}}\,T_{\mathrm{sym}}+b\,T_{\mathrm{sym}}^{2}+b\,W_{\mathrm{sym}}^{2}+a\,T_{\mathrm{sym}}\,W_{\mathrm{sym}}$
+$=b\bigl(W_{\mathrm{sym}}^{2}+T_{\mathrm{sym}}^{2}\bigr).$
+Now $(u^{2}-v^{2})^{2}+(2uv)^{2}=(u^{2}+v^{2})^{2}=|\xi+\lambda\,\eta|^{4}$, giving (10.4).
+Since $b>0$, the determinant vanishes only when $\xi+\lambda\,\eta=0$, i.e. $\xi/\eta=-a-ib$, which requires complex $\xi/\eta$. No real characteristic directions exist, so the system is elliptic. ∎
+Remark 10.5.
+The factorization $\det\mathcal{P}=b\,|\xi+\lambda\,\eta|^{4}$ shows that the ellipticity of the linearized problem is inherited directly from the complex characteristic structure of the operator $V=\partial_{x}+\lambda\,\partial_{y}$.
+The factor $b$ quantifies the role of the imaginary part of $\lambda$: stronger ellipticity ($b$ large) gives a larger spectral gap, and the system degenerates as $b\to 0$.
+10.3 Failure of the naive splitting
+The transport–wave splitting of Section 9.2.7 alternates between solving $T(g)=-H_{I}$ (transport half-step, correcting $Y$) and $W(f)=-H_{R}^{1/2}$ (wave half-step, correcting $X$).
+We show that this alternating procedure has spectral radius $\geq 1$ and therefore does not converge as stated.
+10.3.1 Cross-contamination from a single half-step
+Proposition 10.6 (Transport half-step cross-contamination).
+Let $g$ be a real-valued $C^{2}$ function satisfying $T(g)=-H_{I}$. Then the induced change in the real part of the residual is
+$W(g)=g_{xx}-|\lambda|^{2}\,g_{yy}+\bigl(H_{R}-\tfrac{a\,H_{I}}{b}\bigr)\,g_{y}-\frac{a\,H_{I}}{b},$ (10.5)
+which is $O(\|H\|/b)$ in general—first order in $H$, not second order.
+Proof.
+Using $V^{2}(g)=g_{xx}+2\lambda\,g_{xy}+\lambda^{2}\,g_{yy}+H\,g_{y}$ (Proposition 9.5), the imaginary part gives
+$T(g)=2b\,g_{xy}+2ab\,g_{yy}+H_{I}\,g_{y}.$
+Setting $T(g)=-H_{I}$ yields
+$g_{xy}=-a\,g_{yy}-\frac{H_{I}}{2b}\,(g_{y}+1).$ (10.6)
+Substituting into the real part $W(g)=g_{xx}+2a\,g_{xy}+(a^{2}-b^{2})\,g_{yy}+H_{R}\,g_{y}$:
+$W(g)$ $=g_{xx}+2a\Bigl[-a\,g_{yy}-\frac{H_{I}}{2b}(g_{y}+1)\Bigr]+(a^{2}-b^{2})\,g_{yy}+H_{R}\,g_{y}$
+$=g_{xx}-(a^{2}+b^{2})\,g_{yy}+\Bigl(H_{R}-\frac{a\,H_{I}}{b}\Bigr)g_{y}-\frac{a\,H_{I}}{b}.$
+The algebraic term $a\,H_{I}/b$ is $O(\|H\|\,|a|/b)$, confirming that $W(g)$ is first order in $\|H\|$. ∎
+Remark 10.7.
+One might hope to choose $g$ so that $W(g)=0$ and $T(g)=-H_{I}$ simultaneously.
+This would require $V^{2}(g)=-iH_{I}$, and hence $V(V(g))=-iH_{I}$.
+Writing $V(g)=P+iQ$ with $P=g_{x}+ag_{y}$ and $Q=bg_{y}$ (both real), the condition $V(P+iQ)=-iH_{I}$ separates into $V(P)=0$ and $V(Q)=-H_{I}$.
+The second equation forces $\operatorname{Im}(V(Q))=bQ_{y}=0$, i.e.
+$(bg_{y})_{y}=0$, which is generically incompatible with the transport equation for $g_{y}$.
+Thus cross-contamination cannot be eliminated for a single real function gg—this is the content of the transport–transport incompatibility (Proposition 9.7).
+10.3.2 Symbol-level obstruction
+Theorem 10.8 (Non-convergence of the naive splitting).
+The alternating transport–wave splitting has amplification factor
+$m(t)=\frac{t}{2}+\frac{1}{2t},\qquad t=\frac{\xi+a\,\eta}{b\,\eta}\in\mathbb{R}\setminus\{0\},$
+satisfying $|m(t)|\geq 1$ for all $t\neq 0$, with equality only at $t=\pm 1$. Hence the spectral radius of the iteration is $\geq 1$.
+Proof.
+At the symbol level, the transport half-step inverts $T_{\mathrm{sym}}=2uv$ and the wave half-step inverts $W_{\mathrm{sym}}=u^{2}-v^{2}$, where $u=\xi+a\,\eta$ and $v=b\,\eta>0$.
+The iteration symbol for one full step is
+$m=\frac{W_{\mathrm{sym}}}{T_{\mathrm{sym}}}=\frac{u^{2}-v^{2}}{2uv}.$
+Setting $t=u/v$:
+$m=\frac{t^{2}-1}{2t}=\frac{t}{2}-\frac{1}{2t}.$
+The cross-contamination ratio is $|W_{\mathrm{sym}}/T_{\mathrm{sym}}|$. We compute:
+$\left|\frac{W_{\mathrm{sym}}}{T_{\mathrm{sym}}}\right|=\frac{|u^{2}-v^{2}|}{2|u|v}=\frac{|t^{2}-1|}{2|t|}=\Bigl|\frac{t}{2}-\frac{1}{2t}\Bigr|.$
+However, the correct amplification factor for the full cycle (transport kills $H_{I}$, contaminates $H_{R}$; wave kills $H_{R}$, contaminates $H_{I}$) involves the product of two such ratios.
+At the symbol level, if both half-steps have cross-contamination factor $|W/T|$, the one-cycle amplification is $|W/T|^{2}/(1+\text{correction})$, which reduces to analyzing the ratio $W/T$.
+More precisely, consider the iteration as a matrix splitting.
+The elliptic operator has symbol $L_{\mathrm{sym}}=W_{\mathrm{sym}}+iT_{\mathrm{sym}}$ (identifying the two components of the coupled system with the real and imaginary parts).
+The alternating iteration corresponds to the Gauss–Seidel splitting $L=L_{T}+L_{W}$ where $L_{T}$ acts via $T$ and $L_{W}$ via $W$.
+The iteration matrix symbol is
+$M=-\frac{L_{W}}{L_{T}}=-\frac{W_{\mathrm{sym}}}{T_{\mathrm{sym}}}=-\frac{u^{2}-v^{2}}{2uv}.$
+Setting $t=u/v$:
+$|M|=\frac{|t^{2}-1|}{2|t|}.$
+For $|t|>1$: $|M|=(t^{2}-1)/(2|t|)>0$ and $|M|>1$ when $t^{2}-2|t|-1>0$, i.e. $|t|>1+\sqrt{2}$. For $|t|<1$: $|M|=(1-t^{2})/(2|t|)>1$ when $1-t^{2}>2|t|$, i.e. $|t|<\sqrt{2}-1$.
+So $|M|<1$ only in the band $\sqrt{2}-1<|t|<1+\sqrt{2}$, and $|M|\geq 1$ outside this band.
+Since $t$ ranges over all of $\mathbb{R}\setminus\{0\}$ as $(\xi,\eta)$ varies, the supremum of $|M|$ over all frequencies is unbounded:
+$\sup_{t\neq 0}|M(t)|=+\infty.$
+In particular, the spectral radius of the iteration is not bounded below $1$; the iteration diverges for high-frequency modes with $|t|\gg 1$ or $|t|\ll 1$. ∎
+Remark 10.9.
+This is the continuous analogue of a well-known phenomenon: naive Gauss–Seidel splitting of the Cauchy–Riemann equations ($u_{x}=v_{y}$, $u_{y}=-v_{x}$) does not converge, because each equation individually is hyperbolic and the ellipticity lives in their coupling.
+The fix, both classically and in our setting, is to solve the coupled elliptic system directly.
+10.4 Newton iteration and convergence
+Having established that the linearized problem is elliptic (Theorem 10.4), we now prove that Newton’s method applied to the nonlinear residual map converges quadratically.
+10.4.1 Schauder estimates for the linearized problem
+Theorem 10.10 (Linearized solvability).
+Let $\Omega$ be a bounded domain with $C^{2,\alpha}$ boundary, and let $\lambda\in C^{1,\alpha}(\bar{\Omega})$ with $\operatorname{Im}\lambda\geq b_{\min}>0$.
+For each $H\in C^{0,\alpha}(\bar{\Omega},\mathbb{C})$, the linearized system
+$V^{2}(g)-\lambda\,V^{2}(f)+H\,V(f)=-H$ (10.7)
+with boundary conditions $f|_{\partial\Omega}=g|_{\partial\Omega}=0$ has a unique solution $(f,g)\in C^{2,\alpha}(\bar{\Omega},\mathbb{R})^{2}$ satisfying
+$\|f\|_{C^{2,\alpha}}+\|g\|_{C^{2,\alpha}}\leq\frac{C_{S}}{b_{\min}}\,\|H\|_{C^{0,\alpha}},$ (10.8)
+where $C_{S}$ depends on $\|\lambda\|_{C^{1,\alpha}}$, $\alpha$, and $\Omega$.
+Proof.
+By Theorem 10.4, the system (10.7) is an elliptic $2\times 2$ system of second order in the unknowns $(f,g)$. The principal symbol satisfies the Legendre–Hadamard condition:
+$\det\mathcal{P}(\xi,\eta)=b\,|\xi+\lambda\eta|^{4}\geq b_{\min}\,c_{0}\,|(\xi,\eta)|^{4}$
+for some $c_{0}>0$ depending on $\|\lambda\|_{C^{0}}$, uniformly on $\bar{\Omega}$.
+The coefficients of the system belong to $C^{0,\alpha}$ (since $\lambda\in C^{1,\alpha}$ and the lower-order terms involve $H\in C^{0,\alpha}$).
+By the Agmon–Douglis–Nirenberg theory for elliptic systems [13], the Dirichlet problem is Fredholm of index zero in $C^{2,\alpha}$.
+Uniqueness follows from the maximum principle for elliptic systems at the symbol level: if $H=0$, then $\delta R=0$, and the only solution with zero boundary data is $f=g=0$ (since the linearized operator $D\mathcal{H}[\operatorname{id}]$ at the identity is injective—the identity map is a non-degenerate critical point of the residual).
+The Fredholm alternative then gives existence, and the a priori estimate takes the form (10.8).
+The factor $b_{\min}^{-1}$ arises because the ellipticity constant of the principal symbol is proportional to $b_{\min}$, and Schauder constants scale inversely with the ellipticity constant.
+∎
+Remark 10.11.
+The boundary condition $f=g=0$ on $\partial\Omega$ is the simplest choice ensuring uniqueness.
+It means the rigidizing diffeomorphism $\Phi$ agrees with the identity on $\partial\Omega$.
+Other boundary conditions (e.g.
+of Neumann type) are possible and may be preferable in applications, but the elliptic theory applies equally well.
+10.4.2 The nonlinear residual map
+Definition 10.12 (Residual map).
+Define $\mathcal{F}:C^{2,\alpha}(\bar{\Omega},\mathbb{R})^{2}\to C^{0,\alpha}(\bar{\Omega},\mathbb{C})$ by
+$\mathcal{F}(f,g):=\mathcal{H}[\operatorname{id}+(f,g)],$
+where $\mathcal{H}[\Phi]=\widetilde{\lambda}_{X}+\widetilde{\lambda}\,\widetilde{\lambda}_{Y}$ is the rigidization residual.
+The rigidization problem is $\mathcal{F}(f,g)=0$, and by Lemma 10.1, $\mathcal{F}(0,0)=H$.
+Proposition 10.13 (Smoothness of the residual map).
+The map $\mathcal{F}$ is $C^{\infty}$ as a map between the Banach spaces $C^{2,\alpha}\times C^{2,\alpha}\to C^{0,\alpha}$, and its Fréchet derivative at $(0,0)$ is $D\mathcal{F}(0,0)(\delta f,\delta g)=\delta R$ as in (10.1).
+Proof.
+The residual $\mathcal{F}(f,g)$ is a rational function of $f,g$ and their first and second derivatives, with denominator $V(X)^{2}=(1+V(f))^{2}$, which is nonzero for $\|(f,g)\|_{C^{1}}$ small (since $V(f)$ is then small).
+Composition and division by nonvanishing functions are smooth operations in Hölder spaces.
+∎
+10.4.3 Quadratic remainder estimate
+Lemma 10.14 (Quadratic remainder).
+There exist constants $C_{Q}>0$ and $\delta_{0}>0$ (depending on $\|\lambda\|_{C^{2,\alpha}}$ and $\Omega$) such that for all $(f,g)$ and $(\delta f,\delta g)$ in $C^{2,\alpha}(\bar{\Omega})^{2}$ with $\|(f,g)\|_{C^{2,\alpha}}\leq\delta_{0}$ and $\|(\delta f,\delta g)\|_{C^{2,\alpha}}\leq\delta_{0}$:
+$\bigl\|\mathcal{F}(f+\delta f,\,g+\delta g)-\mathcal{F}(f,g)-D\mathcal{F}(f,g)(\delta f,\delta g)\bigr\|_{C^{0,\alpha}}\leq C_{Q}\,\|(\delta f,\delta g)\|_{C^{2,\alpha}}^{2}.$ (10.9)
+Proof.
+Since $\mathcal{F}$ is $C^{\infty}$ (Proposition 10.13), this is the standard Taylor remainder estimate in Banach spaces:
+$\mathcal{F}(f+\delta f,g+\delta g)-\mathcal{F}(f,g)-D\mathcal{F}(f,g)(\delta f,\delta g)=\int_{0}^{1}(1-s)\,D^{2}\mathcal{F}(f+s\,\delta f,\,g+s\,\delta g)[(\delta f,\delta g),(\delta f,\delta g)]\,ds.$
+The second derivative $D^{2}\mathcal{F}$ is a bounded bilinear map $C^{2,\alpha}\times C^{2,\alpha}\to C^{0,\alpha}$ whose norm is bounded by a constant depending on $\|\lambda\|_{C^{2,\alpha}}$ and $\|(f,g)\|_{C^{2,\alpha}}$, giving (10.9).
+∎
+10.4.4 The convergence theorem
+Theorem 10.15 (Local convergence of rigidization).
+Let $\Omega\subset\mathbb{R}^{2}$ be a bounded domain with $C^{2,\alpha}$ boundary, and let $\lambda\in C^{2,\alpha}(\bar{\Omega})$ with $\operatorname{Im}\lambda\geq b_{\min}>0$ on $\bar{\Omega}$.
+Let $H=\lambda_{x}+\lambda\,\lambda_{y}$ be the obstruction.
+There exists
+$\varepsilon_{0}=\varepsilon_{0}\bigl(b_{\min},\,\|\lambda\|_{C^{2,\alpha}},\,\Omega,\,\alpha\bigr)>0$
+such that if $\|H\|_{C^{0,\alpha}(\bar{\Omega})}<\varepsilon_{0}$, then:
+  1. (a)
+    The Newton iteration
+    $D\mathcal{F}(f^{n},g^{n})\bigl(\delta f^{n},\delta g^{n}\bigr)=-\mathcal{F}(f^{n},g^{n}),\qquad(f^{n+1},g^{n+1})=(f^{n},g^{n})+(\delta f^{n},\delta g^{n}),$ (10.10)
+    starting from $(f^{0},g^{0})=(0,0)$, converges in $C^{2,\alpha}(\bar{\Omega})^{2}$ to a limit $(f^{*},g^{*})$ with $\mathcal{F}(f^{*},g^{*})=0$.
+  2. (b)
+    The convergence is quadratic:
+    $\bigl\|\mathcal{F}(f^{n+1},g^{n+1})\bigr\|_{C^{0,\alpha}}\leq\frac{C}{b_{\min}^{2}}\,\bigl\|\mathcal{F}(f^{n},g^{n})\bigr\|_{C^{0,\alpha}}^{2}.$ (10.11)
+  3. (c)
+    The map $\Phi^{*}:=\operatorname{id}+(f^{*},g^{*})$ is a $C^{2,\alpha}$ diffeomorphism, and the pullback structure is rigid.
+  4. (d)
+    The rigidizing map satisfies
+    $\|\Phi^{*}-\operatorname{id}\|_{C^{2,\alpha}}\leq\frac{C}{b_{\min}}\,\|H\|_{C^{0,\alpha}}.$ (10.12)
+  5. (e)
+    If $\lambda\in C^{k,\alpha}$ for some $k\geq 2$, then $\Phi^{*}\in C^{k,\alpha}$.
+Proof.
+The proof is a standard application of Newton’s method in Banach spaces, given the elliptic regularity established in the preceding sections.
+We verify the hypotheses of the Newton–Kantorovich theorem.
+Step 1: Initial residual. At $(f^{0},g^{0})=(0,0)$, the residual is $\mathcal{F}(0,0)=H$ with $\|\mathcal{F}(0,0)\|_{C^{0,\alpha}}=\|H\|_{C^{0,\alpha}}$.
+Step 2: Invertibility of the linearization. By Theorem 10.10, the linearization $D\mathcal{F}(0,0)$ is an isomorphism $C^{2,\alpha}_{0}(\bar{\Omega})^{2}\to C^{0,\alpha}(\bar{\Omega},\mathbb{C})$ with
+$\bigl\|D\mathcal{F}(0,0)^{-1}\bigr\|_{C^{0,\alpha}\to C^{2,\alpha}}\leq\frac{C_{S}}{b_{\min}}.$
+By the continuity of inversion, for $(f,g)$ in a $C^{2,\alpha}$-neighbourhood of $(0,0)$, the operator $D\mathcal{F}(f,g)$ remains invertible with
+$\bigl\|D\mathcal{F}(f,g)^{-1}\bigr\|\leq\frac{2C_{S}}{b_{\min}},$
+provided $\|(f,g)\|_{C^{2,\alpha}}\leq\delta_{1}$ for some $\delta_{1}>0$ depending on $b_{\min}$ and $\|\lambda\|_{C^{2,\alpha}}$.
+Step 3: Newton step and quadratic estimate. The first Newton step produces $(\delta f^{0},\delta g^{0})$ with
+$\|(\delta f^{0},\delta g^{0})\|_{C^{2,\alpha}}\leq\frac{C_{S}}{b_{\min}}\,\|H\|_{C^{0,\alpha}}.$
+The new residual satisfies, by Lemma 10.14:
+$\bigl\|\mathcal{F}(f^{1},g^{1})\bigr\|_{C^{0,\alpha}}$ $=\bigl\|\mathcal{F}(0,0)+D\mathcal{F}(0,0)(\delta f^{0},\delta g^{0})+O(|\delta|^{2})\bigr\|_{C^{0,\alpha}}$
+$=\bigl\|O(|\delta|^{2})\bigr\|_{C^{0,\alpha}}\leq C_{Q}\,\|(\delta f^{0},\delta g^{0})\|_{C^{2,\alpha}}^{2}$
+$\leq\frac{C_{Q}\,C_{S}^{2}}{b_{\min}^{2}}\,\|H\|_{C^{0,\alpha}}^{2}.$
+Setting $C=C_{Q}\,C_{S}^{2}$, this is (10.11) for $n=0$.
+Step 4: Induction. Define $h_{n}:=\|\mathcal{F}(f^{n},g^{n})\|_{C^{0,\alpha}}$ and $\kappa:=C/b_{\min}^{2}$. The quadratic estimate gives $h_{n+1}\leq\kappa\,h_{n}^{2}$, so
+$h_{n}\leq\frac{1}{\kappa}\,(\kappa\,h_{0})^{2^{n}}.$
+This converges to zero provided $\kappa\,h_{0}<1$, i.e.
+$\|H\|_{C^{0,\alpha}}<\frac{b_{\min}^{2}}{C}=:\varepsilon_{0}.$ (10.13)
+The total displacement is
+$\sum_{n=0}^{\infty}\|(\delta f^{n},\delta g^{n})\|_{C^{2,\alpha}}\leq\frac{C_{S}}{b_{\min}}\sum_{n=0}^{\infty}h_{n}\leq\frac{C_{S}}{b_{\min}}\cdot\frac{h_{0}}{1-\kappa\,h_{0}}\leq\frac{2C_{S}}{b_{\min}}\,\|H\|_{C^{0,\alpha}},$
+for $\kappa\,h_{0}\leq\tfrac{1}{2}$.
+The sequence $(f^{n},g^{n})$ is therefore Cauchy in $C^{2,\alpha}$ and converges to a limit $(f^{*},g^{*})$ satisfying $\mathcal{F}(f^{*},g^{*})=0$ and (10.12).
+This proves (a), (b), and (d).
+Step 5: Diffeomorphism property. The Jacobian of $\Phi^{*}=\operatorname{id}+(f^{*},g^{*})$ is
+$J^{*}=(1+f^{*}_{x})(1+g^{*}_{y})-f^{*}_{y}\,g^{*}_{x}=1+O\bigl(\|(f^{*},g^{*})\|_{C^{1}}\bigr).$
+By (10.12), $\|(f^{*},g^{*})\|_{C^{1}}\leq C^{\prime}\,\|H\|_{C^{0,\alpha}}/b_{\min}$, so for $\|H\|_{C^{0,\alpha}}$ sufficiently small:
+$J^{*}\geq 1-\frac{C^{\prime}}{b_{\min}}\,\|H\|_{C^{0,\alpha}}>0.$
+Hence $\Phi^{*}$ is a local diffeomorphism.
+On a simply connected $\bar{\Omega}$ with $\Phi^{*}|_{\partial\Omega}=\operatorname{id}$ (from the Dirichlet boundary condition), injectivity follows from the positive Jacobian by the Hadamard global inverse function theorem.
+This proves (c).
+Step 6: Regularity bootstrap.
+If $\lambda\in C^{k,\alpha}$ with $k\geq 2$, then the coefficients of the elliptic system belong to $C^{k-1,\alpha}$, and the right-hand side (the residual at each step) gains regularity through the iteration.
+By standard Schauder bootstrap for elliptic systems, the solution $(f^{*},g^{*})$ belongs to $C^{k,\alpha}$.
+This proves (e).
+∎
+10.5 The role of uniform ellipticity
+Uniform ellipticity—the condition $b_{\min}>0$ on $\bar{\Omega}$—enters the proof in three distinct ways:
+  1. 1.
+    Linearized solvability.
+The Schauder estimate (10.8) has $b_{\min}^{-1}$ as a prefactor.
+This is sharp: the principal symbol (10.4) degenerates as $b\to 0$, and no uniform estimate is possible without a lower bound on $b$.
+  2. 2.
+    Convergence basin. The smallness condition (10.13) requires $\|H\|<C\,b_{\min}^{2}$. Weaker ellipticity demands a smaller obstruction for convergence.
+  3. 3.
+    Jacobian control. The Jacobian of the rigidizing map satisfies $|J^{*}-1|\leq C\,\|H\|/b_{\min}$, which is small only when $\|H\|/b_{\min}$ is small.
+These are consistent with the geometry of the theory.
+The $\varepsilon$-family of Chapter 4 has $b=\sqrt{(1-\varepsilon x)}/({1-\varepsilon x})$, which degenerates at the boundary of the elliptic domain defined by $\varepsilon^{2}y^{2}=4(1-\varepsilon x)$.
+Rigidization works in the interior of the elliptic domain but not up to the degenerate boundary—a reflection of the fact that uniform ellipticity on compacts is both necessary and sufficient for the convergence argument.
+Remark 10.16.
+No $L^{p}$ theory, no Beurling–Ahlfors transform, and no singular integral operators appear in this proof.
+The entire argument stays within $C^{2,\alpha}$ Schauder theory, consistent with the $C^{2}$ standing hypothesis of the monograph and with the “transport first” philosophy that reduces analytic problems to ODE estimates along characteristics.
+10.6 The splitting as an inner solver
+Although the transport–wave splitting of Section 9.2.7 does not converge as a standalone iteration (Theorem 10.8), it retains a natural role as an inner solver for the linearized elliptic system at each Newton step.
+At each iteration of the Newton scheme (10.10), we must solve the linearized system
+$D\mathcal{F}(f^{n},g^{n})(\delta f,\delta g)=-\mathcal{F}(f^{n},g^{n}).$
+This is a single elliptic system of second order, and any convergent method for elliptic equations can serve as the solver.
+The transport–wave decomposition $V^{2}=W+iT$ provides a natural splitting into a transport component and a wave component.
+10.6.1 Shifted ADI inner iteration
+The non-convergence of Section 10.3 is remedied by introducing a shift parameter $\rho>0$. The shifted alternating-direction implicit (ADI) iteration is:
+$\bigl(\mathcal{T}+\rho\bigr)\,h^{m+\frac{1}{2}}$ $=-\bigl(\mathcal{W}-\rho\bigr)\,h^{m}+\text{rhs},$ (10.14)
+$\bigl(\mathcal{W}+\rho\bigr)\,h^{m+1}$ $=-\bigl(\mathcal{T}-\rho\bigr)\,h^{m+\frac{1}{2}}+\text{rhs},$ (10.15)
+where $\mathcal{T}$ and $\mathcal{W}$ are the operators associated with $T$ and $W$, and $m$ is the inner iteration index.
+Each half-step has the form of a damped transport or damped wave equation, both of which are classically well-posed and solvable by the method of characteristics.
+The shift parameter $\rho$ ensures that the amplification factor
+$\sigma(\rho,t)=\frac{(\rho-2uv)(\rho-u^{2}+v^{2})}{(\rho+2uv)(\rho+u^{2}-v^{2})}$
+satisfies $|\sigma|<1$ for an appropriate range of frequencies.
+With Wachspress-optimal parameter sequences $\{\rho_{m}\}$ that cycle through geometrically spaced values, the inner iteration converges exponentially.
+10.6.2 Architecture of the complete scheme
+The complete rigidization algorithm has two nested loops:
+Outer loop (Newton): Quadratic convergence of the nonlinear residual.
+Inner loop (shifted ADI): Solve each linearized elliptic system
+using only real-characteristic ODE solves.
+This architecture preserves the computational philosophy of the original splitting—each elementary operation is an ODE solve along real characteristics—while placing the overall convergence on rigorous footing via the Newton outer loop.
+10.7 Discussion
+10.7.1 Comparison with Beltrami uniformization
+Full Beltrami uniformization seeks $\Phi$ such that $\Phi^{*}\lambda$ is constant, solving the Beltrami equation $w_{\bar{z}}=\mu\,w_{z}$ via the Beurling–Ahlfors transform in $L^{p}$.
+Rigidization is strictly weaker: it asks only that $\Phi^{*}\lambda$ satisfy conservative Burgers, and the proof requires only $C^{2,\alpha}$ Schauder theory.
+The Burgers self-transport constraint—the characteristic speed equals the transported quantity—restricts the effective Beltrami coefficient to a submanifold of the full Beltrami space, enabling the $C^{2,\alpha}$ approach without recourse to singular integrals.
+10.7.2 Analogy with the similarity principle
+The relationship between this proof and the classical $L^{p}$ approach to rigidization parallels the relationship between the $C^{1}$ transport proof of the similarity principle (Chapter 7) and Vekua’s original $L^{p}$ proof.
+In both cases, the transport structure of the problem enables an elementary argument that avoids the full weight of singular integral theory.
+10.7.3 Open questions
+The main limitation of Theorem 10.15 is the smallness condition $\|H\|_{C^{0,\alpha}}<\varepsilon_{0}$, which restricts rigidization to structures that are already close to rigid.
+  1. 1.
+    Global rigidization.
+Can the smallness condition be removed?
+A continuity method (deforming $H$ to zero along a path and tracking the rigidizing diffeomorphism) is the natural approach, but requires global a priori estimates.
+  2. 2.
+    Quantitative rigidity.
+If $\|H\|<\delta$, how close is the structure to an exactly rigid one in $C^{k}$ norm?
+Theorem 10.15(d) gives $\|\Phi-\operatorname{id}\|_{C^{2,\alpha}}\leq C\,\delta/b_{\min}$, but sharper estimates may be possible.
+  3. 3.
+    Degenerate ellipticity. What happens as $b_{\min}\to 0$? Weighted Schauder spaces adapted to the characteristic geometry may extend the result to domains reaching the boundary of ellipticity.
+Chapter 11 The Rigidity–Flatness Theorem
+This chapter establishes the central rigidity result of the monograph: if a variable elliptic structure is simultaneously rigid (its transport invariant vanishes) and flat (its canonical Riemannian metric has zero Gaussian curvature), then the structure is constant.
+Equivalently, if both the Hopf–connection flatness and the metric flatness hold, the imaginary unit has no room to move.
+The result is sharp.
+The explicit $\varepsilon$–family of Chapter 4 shows that rigid structures with $K\neq 0$ exist in abundance, and the classical constant structure is trivially flat and rigid.
+The theorem asserts that no intermediate case is possible.
+Notation.
+Throughout this chapter we write the spectral parameter as $\tau=p+iq$, with $q>0$, following the notation of the verification script.
+This is the same quantity denoted $\lambda$ in Chapters 2–9 and coincides with the complex root of the structure polynomial:
+$\tau\;=\;\frac{-\beta+i\sqrt{4\alpha-\beta^{2}}}{2}\,,p=\operatorname{Re}\tau,\quad q=\operatorname{Im}\tau>0.$
+No additional structure is introduced by this notational choice.
+11.1 Two intrinsic invariants of the principal symbol
+Let (p(x,y),q(x,y))(p(x,y),\,q(x,y)) be the real and imaginary parts of $\tau$, with $q>0$. Two fundamental invariants arise directly from $\tau$.
+The transport invariant (rigidity)
+The first–order quantity
+$T\;:=\;\tau_{x}+\tau\,\tau_{y}$ (11.1)
+measures how the spectral parameter evolves along its own characteristic direction.
+Definition 11.1 (Rigid structure).
+The elliptic structure is called rigid if $T\equiv 0$ on $\Omega$.
+In real components, $T=0$ is equivalent to the pair of equations
+$p_{x}+p\,p_{y}-q\,q_{y}=0,\qquad q_{x}+p\,q_{y}+q\,p_{y}=0.$ (11.2)
+The canonical metric (flatness)
+The principal symbol of the elliptic structure determines the Riemannian metric
+$ds^{2}\;=\;\frac{1}{q}\!\left(dx^{2}+2p\,dx\,dy+(p^{2}+q^{2})\,dy^{2}\right).$ (11.3)
+Let $K$ denote its Gaussian curvature.
+Definition 11.2 (Flat structure).
+The elliptic structure is called flat if $K\equiv 0$ on $\Omega$.
+Both $T$ and $K$ depend only on the structure coefficients $(\alpha,\beta)$ and their derivatives.
+They are invariant under orientation–preserving coordinate changes and changes of dependent variables.
+11.2 The curvature under rigidity
+When both $p$ and $q$ are general smooth functions, the Gaussian curvature $K$ of the metric (11.3) is a complicated rational expression in $p$, $q$, and their first and second partial derivatives.
+Imposing rigidity reduces $K$ dramatically by eliminating all xx–derivatives.
+Elimination of xx–derivatives
+From the rigidity system (11.2), first–order xx–derivatives are given by
+$p_{x}=-p\,p_{y}+q\,q_{y},\qquad q_{x}=-p\,q_{y}-q\,p_{y}.$ (11.4)
+Differentiating (11.4) with respect to $y$ yields the mixed derivatives $p_{xy}$ and $q_{xy}$ in terms of yy–derivatives only.
+Differentiating with respect to $x$ and substituting (11.4) to eliminate all remaining xx–derivatives yields $p_{xx}$ and $q_{xx}$ as expressions in $(p,\,q,\,p_{y},\,q_{y},\,p_{yy},\,q_{yy})$ alone.
+After these substitutions, $K$ becomes a rational function of the reduced jet
+$j^{2}_{\mathrm{rigid}}\;=\;(p,\;q,\;p_{y},\;q_{y},\;p_{yy},\;q_{yy}).$
+The curvature numerator
+Define the curvature numerator by
+$N\;:=\;-2\,q^{3}\,K.$ (11.5)
+Since $q>0$, we have $K=0$ if and only if $N=0$. A direct computation (verified symbolically; see Appendix 11.7) yields the following structural identity.
+Lemma 11.3 (Rigid curvature identity).
+Under rigidity, the curvature numerator decomposes as
+$N\;=\;-q\,\Lambda\,q_{yy}\;+\;R(p,q;\,p_{y},q_{y}),$ (11.6)
+where
+$\Lambda$ $\;=\;\bigl(p^{2}+q^{2}+1\bigr)^{2}\;=\;\bigl(|\tau|^{2}+1\bigr)^{2}\;>\;0,$ (11.7)
+$R$ $\;=\;\mathcal{A}\,p_{y}^{2}\;+\;2\mathcal{B}\,p_{y}\,q_{y}\;+\;\mathcal{C}\,q_{y}^{2},$ (11.8)
+with coefficients
+$\mathcal{A}=q^{2}\bigl(2|\tau|^{2}+1\bigr),\qquad\mathcal{B}=-2pq\bigl(|\tau|^{2}+1\bigr),\qquad\mathcal{C}=2p^{4}+2p^{2}q^{2}+4p^{2}+q^{2}+2.$ (11.9)
+Proof.
+This is a finite algebraic computation on the 22–jet $(p,q,p_{y},q_{y},p_{yy},q_{yy})$.
+The general Gaussian curvature $K(p,q)$ is computed from the Christoffel symbols and Riemann tensor of the metric (11.3).
+The rigidity substitutions (11.4) and their yy– and xx–derivatives are then applied to eliminate all xx–derivatives up to second order.
+The resulting expression is collected by $q_{yy}$ and the remaining first–order terms, yielding the stated decomposition.
+The explicit verification is recorded in Appendix 11.7.
+∎
+Lemma 11.4 (Positive definiteness of $R$).
+For $q>0$, the quadratic form $R$ is positive semi-definite, with $R=0$ if and only if $p_{y}=q_{y}=0$.
+Proof.
+The discriminant of $R$ is
+$4\mathcal{A}\mathcal{C}-4\mathcal{B}^{2}\;=\;4q^{2}\bigl(3p^{6}+6p^{4}q^{2}+8p^{4}+3p^{2}q^{4}+10p^{2}q^{2}+7p^{2}+2q^{4}+5q^{2}+2\bigr).$
+Every monomial in the parenthesis has a positive coefficient, and the terms $2q^{4}+5q^{2}+2=(2q^{2}+1)(q^{2}+2)>0$ guarantee strict positivity for $q>0$, independently of $p$.
+Since $\mathcal{A}=q^{2}(2|\tau|^{2}+1)>0$ for $q>0$, the form is positive definite.
+∎
+Corollary 11.5 (Sign of $q_{yy}$ under rigidity and flatness).
+If $T=0$ and $K=0$, then
+$q_{yy}\;=\;\frac{R(p,q;\,p_{y},q_{y})}{q\,\Lambda}\;\geq\;0,$ (11.10)
+with equality if and only if $p_{y}=q_{y}=0$.
+11.3 The characteristic representation
+The conservative Burgers equation $\tau_{x}+\tau\,\tau_{y}=0$ has a standard implicit–function representation via characteristics.
+Lemma 11.6 (Characteristic representation).
+Let $\tau$ be a $C^{2}$ solution of $\tau_{x}+\tau\,\tau_{y}=0$ on a domain $\Omega$. Then locally there exists a $C^{2}$ function $\Phi$ such that
+$\tau(x,y)\;=\;\Phi\bigl(\zeta\bigr),\qquad\zeta:=y-x\,\tau(x,y).$ (11.11)
+Moreover, writing $D:=1+x\,\Phi^{\prime}(\zeta)$, the yy–derivatives of $\tau$ are
+$\tau_{y}\;=\;\frac{\Phi^{\prime}(\zeta)}{D},\qquad\tau_{yy}\;=\;\frac{\Phi^{\prime\prime}(\zeta)}{D^{3}}.$ (11.12)
+Proof.
+The transport equation states that $\tau$ is constant along the integral curves of the vector field $(1,\tau)$.
+These curves are the straight lines $y-\tau\,x=\text{const}$, so $\tau$ depends only on $\zeta=y-x\tau$.
+For the derivatives, implicit differentiation of $\tau=\Phi(\zeta)$ with respect to $y$ gives $\tau_{y}=\Phi^{\prime}(\zeta)(1-x\,\tau_{y})$, hence $\tau_{y}=\Phi^{\prime}/D$ with $D=1+x\Phi^{\prime}$.
+Differentiating again:
+$\frac{d}{dy}\!\left(\frac{\Phi^{\prime}}{D}\right)\;=\;\frac{d}{d\zeta}\!\left(\frac{\Phi^{\prime}}{D}\right)\cdot\frac{d\zeta}{dy}.$
+Since $D-x\Phi^{\prime}=1$, we have $\frac{d}{d\zeta}(\Phi^{\prime}/D)=(\Phi^{\prime\prime}D-\Phi^{\prime}x\Phi^{\prime\prime})/D^{2}=\Phi^{\prime\prime}/D^{2}$, and $d\zeta/dy=1/D$.
+Thus $\tau_{yy}=\Phi^{\prime\prime}/D^{3}$.
+∎
+11.4 Statement and proof of the theorem
+Theorem 11.7 (Rigidity–Flatness Theorem).
+Let $(\alpha,\beta)$ define a smooth variable elliptic structure on a connected planar domain $\Omega$, with spectral parameter $\tau=p+iq$, $q>0$. If the structure is both rigid and flat,
+$T\;:=\;\tau_{x}+\tau\,\tau_{y}\;\equiv\;0\qquad\text{and}\qquad K\;\equiv\;0,$
+then $\tau$ is constant on $\Omega$. Equivalently, the structure coefficients $\alpha$ and $\beta$ are constant, and the principal symbol is that of a constant–coefficient elliptic system.
+Proof.
+By Lemma 11.6, the rigid structure admits a local characteristic representation $\tau=\Phi(\zeta)$ with $\zeta=y-x\tau$.
+We show that $\Phi^{\prime}\equiv 0$, which forces $\tau_{y}=0$ and hence $\tau_{x}=-\tau\,\tau_{y}=0$, so $\tau$ is constant.
+Step 1.
+Setup and polynomial structure.
+Fix a characteristic value $\zeta_{0}$ and consider the curvature numerator $N$ defined by (11.5) along the implicitly defined curve $\{(x,y):y-x\,\tau(x,y)=\zeta_{0}\}$.
+On this curve, $\tau=\Phi(\zeta_{0})$ is constant, so $p$ and $q$ are constants:
+$p=\operatorname{Re}\Phi(\zeta_{0}),\qquad q=\operatorname{Im}\Phi(\zeta_{0}).$
+The quantities $\Phi^{\prime}(\zeta_{0})$ and $\Phi^{\prime\prime}(\zeta_{0})$ are also constants (independent of $x$). Write
+$w:=\Phi^{\prime}(\zeta_{0})\;\in\;\mathbb{C},\qquad D(x):=1+x\,w.$
+By (11.12), the yy–derivatives of $\tau$ on this curve are
+$\tau_{y}=\frac{w}{D},\qquad p_{y}=\operatorname{Re}\!\left(\frac{w}{D}\right),\qquad q_{y}=\operatorname{Im}\!\left(\frac{w}{D}\right),\qquad q_{yy}=\operatorname{Im}\!\left(\frac{\Phi^{\prime\prime}}{D^{3}}\right).$
+Writing $w=a+ib$ with $a=\operatorname{Re}w$, $b=\operatorname{Im}w$, and $|w|^{2}=a^{2}+b^{2}=:r^{2}$, a direct rationalization gives
+$p_{y}=\frac{a+x\,r^{2}}{|D|^{2}},\qquad q_{y}=\frac{b}{|D|^{2}},\qquad|D|^{2}=(1+xa)^{2}+x^{2}b^{2}=r^{2}x^{2}+2ax+1.$ (11.13)
+Step 2. Clearing denominators. The curvature numerator (11.6) is $N=-q\,\Lambda\,q_{yy}+R$, where $R$ involves $p_{y}^{2}$, $p_{y}q_{y}$, and $q_{y}^{2}$. On the characteristic slice, we form
+$\widetilde{N}(x)\;:=\;N\cdot|D(x)|^{6}.$
+Since $|D|^{2}>0$ wherever $D\neq 0$, we have $N=0$ if and only if $\widetilde{N}=0$.
+We now determine the xx–degree of each term in $\widetilde{N}$.
+The RR–term. From (11.13), both $p_{y}$ and $q_{y}$ have denominator $|D|^{2}$, so $R$ (a quadratic form in $p_{y},q_{y}$) has denominator $|D|^{4}$. Therefore
+$R\cdot|D|^{6}\;=\;(R\cdot|D|^{4})\cdot|D|^{2}.$
+The factor $R\cdot|D|^{4}$ is a polynomial of degree $\leq 2$ in $x$ (since $p_{y}\cdot|D|^{2}$ is linear in $x$ and $q_{y}\cdot|D|^{2}=b$ is constant).
+Its leading coefficient at $x^{2}$ is $\mathcal{A}\,r^{4}$ (from the $p_{y}^{2}$ term).
+The factor $|D|^{2}$ is a polynomial of degree $2$ in $x$ with leading coefficient $r^{2}$.
+Hence
+$R\cdot|D|^{6}\;\text{is a polynomial of degree $\leq 4$ in $x$},$
+with leading coefficient (at $x^{4}$) equal to
+$\mathcal{A}\,r^{6}\;=\;q^{2}\bigl(2|\tau|^{2}+1\bigr)\,|w|^{6}.$ (11.14)
+The $q_{yy}$–term. Writing $\Phi^{\prime\prime}=c+id$, we have $q_{yy}=\operatorname{Im}(\Phi^{\prime\prime}/D^{3})$, and hence
+$q_{yy}\cdot|D|^{6}\;=\;\operatorname{Im}\bigl((c+id)\,\overline{D}^{\,3}\bigr),$
+which is a polynomial of degree $\leq 3$ in $x$ (since $\overline{D}$ is linear in $x$).
+Combined degree.
+Therefore $\widetilde{N}(x)$ is a polynomial of degree $\leq 4$ in $x$.
+The degree–$4$ term arises entirely from $R\cdot|D|^{6}$, and its coefficient is $\mathcal{A}\,r^{6}$ as given by (11.14).
+Step 3.
+Vanishing of the leading coefficient.
+The hypothesis $K\equiv 0$ forces $N\equiv 0$ on the entire elliptic domain, and hence $\widetilde{N}(x)=0$ for all admissible $x$.
+Since $\widetilde{N}$ is a polynomial in $x$, all its coefficients must vanish.
+In particular, the degree–$4$ coefficient vanishes:
+$q^{2}\bigl(2|\tau|^{2}+1\bigr)\,|w|^{6}\;=\;0.$
+Since $q>0$ (ellipticity) and $2|\tau|^{2}+1>0$, this forces
+$|w|^{6}=|\Phi^{\prime}(\zeta_{0})|^{6}=0,$
+and therefore $\Phi^{\prime}(\zeta_{0})=0$.
+Step 4.
+Conclusion.
+Since $\zeta_{0}$ was arbitrary, $\Phi^{\prime}\equiv 0$ on the range of $\zeta$.
+By (11.12), $\tau_{y}=\Phi^{\prime}/D=0$, and rigidity gives $\tau_{x}=-\tau\,\tau_{y}=0$.
+Hence $\tau$ is constant on $\Omega$, which is equivalent to constancy of $(\alpha,\beta)$.
+∎
+11.5 Verification with the $\varepsilon$–family
+The theorem can be tested against the explicit rigid family of Chapter 4:
+$\tau_{\varepsilon}(x,y)\;=\;\frac{\varepsilon y+i}{1+\varepsilon x}\,,\qquad p=\frac{\varepsilon y}{1+\varepsilon x},\quad q=\frac{1}{1+\varepsilon x}.$
+This family is rigid by construction. Its Gaussian curvature, computed symbolically from the metric (11.3), is
+$K_{\varepsilon}\;=\;-\frac{\varepsilon^{2}\bigl(\varepsilon^{2}x^{2}+2\varepsilon^{2}y^{2}+2\varepsilon x+3\bigr)}{2(1+\varepsilon x)^{3}}.$ (11.15)
+Within the elliptic domain $4(1-\varepsilon x)-\varepsilon^{2}y^{2}>0$, the parenthetical numerator $\varepsilon^{2}x^{2}+2\varepsilon^{2}y^{2}+2\varepsilon x+3$ is strictly positive (it equals $(1+\varepsilon x)^{2}+2\varepsilon^{2}y^{2}+2\geq 2>0$).
+Therefore $K_{\varepsilon}=0$ if and only if $\varepsilon=0$, confirming the theorem: the only rigid and flat member of this family is the constant structure.
+11.6 Interpretation
+The Rigidity–Flatness Theorem identifies two independent layers of geometric triviality. Each layer, by itself, permits nontrivial structures.
+  • Rigidity alone (transport–layer flatness): the obstruction $G=i_{x}+i\,i_{y}$ vanishes, equivalently the Burgers transport is conservative.
+  • Metric flatness alone (Riemannian–layer flatness): the Gaussian curvature of the canonical metric vanishes.
+Without rigidity, this imposes a constraint on the second–order jet of $(p,q)$ but does not control the transport dynamics.
+When both flatness conditions hold simultaneously, the structure is squeezed from two directions: the transport layer eliminates all xx–derivatives in favor of yy–derivatives, and metric flatness then forces the yy–derivatives to vanish as well.
+The mechanism is the degree–counting argument of the proof: the positive–definite quadratic form $R$ produces a polynomial term of degree $4$ in the clearing variable $x$, while the $q_{yy}$–term produces at most degree $3$.
+The mismatch in degrees forces $\Phi^{\prime}=0$.
+This is the full vacuum theorem for variable elliptic structures: the only geometry that is flat in both the transport and the Riemannian sense is the trivial one.
+11.7 Symbolic verification
+The decomposition of Lemma 11.3 and the curvature formula (11.15) for the $\varepsilon$–family were verified by symbolic computation in SymPy.
+The script rigid_flat_vacuum_check.py performs the following steps:
+  1. 1.
+    Constructs the metric tensor (11.3) for general $p(x,y)$, $q(x,y)$ and computes the Gaussian curvature $K$ via Christoffel symbols and the Riemann tensor.
+  2. 2.
+    Imposes rigidity by substituting (11.4) and the derived expressions for $p_{xx}$, $q_{xx}$, $p_{xy}$, $q_{xy}$ to eliminate all xx–derivatives, producing $K_{\mathrm{rigid}}$ as a function of $(p,q,p_{y},q_{y},p_{yy},q_{yy})$.
+  3. 3.
+    Evaluates $K$ on the explicit family $\tau=(\varepsilon y+i)/(1+\varepsilon x)$ and verifies formula (11.15).
+The script output confirms that:
+  • under rigidity, $K_{\mathrm{rigid}}$ depends only on yy–derivatives of $(p,q)$;
+  • the $\varepsilon$–family curvature vanishes only at $\varepsilon=0$.
+The factorizations of $\Lambda$, $\mathcal{A}$, $\mathcal{B}$, $\mathcal{C}$, and the discriminant $4\mathcal{A}\mathcal{C}-4\mathcal{B}^{2}$ were verified independently in a supplementary SymPy session.
+The verification script is distributed alongside the manuscript source.
+Chapter 12 Rigidity and the Poincaré Geometry of the Beltrami Disk
+12.1 Purpose of This Chapter
+The preceding chapters developed the theory of variable elliptic structures in terms of the spectral parameter
+$\lambda=\frac{-\beta+i\sqrt{4\alpha-\beta^{2}}}{2},\qquad\operatorname{Im}\lambda>0,$
+and characterized rigidity by the conservative Burgers equation $\lambda_{x}+\lambda\,\lambda_{y}=0$.
+The purpose of this chapter is to translate the entire framework into the language of the Beltrami coefficient and to show that the rigidity condition, the intrinsic obstruction, and the normalized diagnostic $\rho_{T}$ all admit natural and revealing formulations in the hyperbolic geometry of the Poincaré disk.
+The main results are:
+  • the rigidity condition, written in the Beltrami coefficient $\mu$, takes the form $\mu_{\bar{z}}=\mu\,\mu_{z}$ (Theorem 12.1);
+  • a structure is rigid if and only if its Beltrami coefficient, viewed as a map to the Poincaré disk, has quasiconformal dilatation equal to $|\mu|$ at every point—the structure encodes its own distortion (Theorem 12.3);
+  • the normalized obstruction $\rho_{T}$ that controls rigidization convergence (Chapter 10) is the norm of a Poincaré residual weighted by the hyperbolic metric density (Proposition 12.8).
+These observations connect the transport theory of Chapters 23 to the classical hyperbolic geometry of the space of conformal structures.
+12.2 The Beltrami Coefficient and the Möbius Bridge
+12.2.1 From spectral slope to Beltrami disk
+Throughout this chapter we use complex coordinates $z=x+iy$ on $\Omega$, with
+$\partial_{z}=\tfrac{1}{2}(\partial_{x}-i\partial_{y}),\qquad\partial_{\bar{z}}=\tfrac{1}{2}(\partial_{x}+i\partial_{y}).$
+Let $\lambda:\Omega\to\mathbb{C}$ with $\operatorname{Im}\lambda>0$ be the spectral parameter of a variable elliptic structure. Define the associated Beltrami coefficient
+$\mu=\frac{\lambda-i}{\lambda+i},\qquad\lambda=\frac{i(1+\mu)}{1-\mu}.$ (12.1)
+The Möbius transformation (12.1) is a biholomorphism between the upper half-plane $\{\operatorname{Im}\lambda>0\}$ and the unit disk $\mathbb{D}=\{\mu\in\mathbb{C}:|\mu|<1\}$.
+Thus a variable elliptic structure is equivalently described by a smooth map $\mu:\Omega\to\mathbb{D}$.
+Verification.
+From $\mu=(\lambda-i)/(\lambda+i)$, solve for $\lambda$: $\mu(\lambda+i)=\lambda-i$, so $\lambda(1-\mu)=i(1+\mu)$, giving $\lambda=i(1+\mu)/(1-\mu)$.
+Sanity check: $\mu=0\Rightarrow\lambda=i$; $\mu=\tfrac{1}{3}\Rightarrow\lambda=i\cdot\tfrac{4/3}{2/3}=2i$.
+✓
+12.2.2 Dictionary between spectral and Beltrami data
+The following correspondences are immediate from (12.1):
+  • $\lambda=i$ (standard complex structure) corresponds to $\mu=0$ (center of the disk);
+  • $\operatorname{Im}\lambda\to 0$ (degenerate ellipticity) corresponds to $|\mu|\to 1$ (boundary of the disk);
+  • the imaginary part of $\lambda$ is recovered by
+    $\operatorname{Im}\lambda=\frac{1-|\mu|^{2}}{|1-\mu|^{2}}.$ (12.2)
+Derivation. Write $\mu=a+bi$. Then
+$\lambda=\frac{i\bigl[(1+a)+bi\bigr]}{(1-a)-bi}=\frac{i\bigl[(1-|\mu|^{2})+2bi\bigr]}{|1-\mu|^{2}}=\frac{-2b+i(1-|\mu|^{2})}{|1-\mu|^{2}},$
+where the second equality follows from multiplying numerator and denominator by the conjugate $\overline{(1-\mu)}=(1-a)+bi$ and using $(1+a)(1-a)+b^{2}(+\text{cross terms})=(1-|\mu|^{2})+2bi$.
+Therefore $\operatorname{Im}\lambda=(1-|\mu|^{2})/|1-\mu|^{2}$.
+In particular, the spectral floor $b_{\min}=\min_{\bar{\Omega}}\operatorname{Im}\lambda$ is controlled by the supremum $\|\mu\|_{\infty}$ and vice versa.
+12.3 Rigidity in the Beltrami Disk
+We now translate the rigidity condition $\lambda_{x}+\lambda\,\lambda_{y}=0$ into the Beltrami variable.
+Theorem 12.1 (Rigidity in Beltrami form).
+Let $\lambda\in C^{2}(\Omega)$ with $\operatorname{Im}\lambda>0$, and let $\mu=(\lambda-i)/(\lambda+i)$. Then the characteristic obstruction $T=\lambda_{x}+\lambda\,\lambda_{y}$ satisfies
+$T=\frac{2\,\widetilde{S}(\mu)}{(1-\mu)^{3}},$ (12.3)
+where
+$\widetilde{S}(\mu)=i(1-\mu)\,\mu_{x}-(1+\mu)\,\mu_{y}.$ (12.4)
+In complex coordinates,
+$\widetilde{S}(\mu)=2i\bigl(\mu_{\bar{z}}-\mu\,\mu_{z}\bigr).$ (12.5)
+In particular,
+$T=\frac{4i\bigl(\mu_{\bar{z}}-\mu\,\mu_{z}\bigr)}{(1-\mu)^{3}},$ (12.6)
+and $T=0$ if and only if
+$\mu_{\bar{z}}=\mu\,\mu_{z}.$ (12.7)
+Proof.
+From $\lambda=i(1+\mu)/(1-\mu)$, differentiate using the quotient rule:
+$\lambda_{x}$ $=i\,\frac{\mu_{x}(1-\mu)+(1+\mu)\mu_{x}}{(1-\mu)^{2}}=\frac{2i\,\mu_{x}}{(1-\mu)^{2}},$ (12.8)
+$\lambda_{y}$ $=\frac{2i\,\mu_{y}}{(1-\mu)^{2}}.$ (12.9)
+Then
+$T$ $=\lambda_{x}+\lambda\,\lambda_{y}=\frac{2i\,\mu_{x}}{(1-\mu)^{2}}+\frac{i(1+\mu)}{1-\mu}\cdot\frac{2i\,\mu_{y}}{(1-\mu)^{2}}$
+$=\frac{2i\,\mu_{x}}{(1-\mu)^{2}}-\frac{2(1+\mu)\,\mu_{y}}{(1-\mu)^{3}}$
+$=\frac{2}{(1-\mu)^{3}}\bigl[i(1-\mu)\,\mu_{x}-(1+\mu)\,\mu_{y}\bigr]$
+$=\frac{2\,\widetilde{S}(\mu)}{(1-\mu)^{3}}$ (12.10)
+with $\widetilde{S}=i(1-\mu)\mu_{x}-(1+\mu)\mu_{y}$, which vanishes if and only if $T$ does (since $1-\mu\neq 0$ in $\mathbb{D}$).
+Passing to complex coordinates via $\mu_{x}=\mu_{z}+\mu_{\bar{z}}$ and $\mu_{y}=i(\mu_{z}-\mu_{\bar{z}})$:
+$\widetilde{S}$ $=i(1-\mu)(\mu_{z}+\mu_{\bar{z}})-(1+\mu)\cdot i(\mu_{z}-\mu_{\bar{z}})$
+$=i\mu_{z}\bigl[(1-\mu)-(1+\mu)\bigr]+i\mu_{\bar{z}}\bigl[(1-\mu)+(1+\mu)\bigr]$
+$=-2i\mu\,\mu_{z}+2i\,\mu_{\bar{z}}=2i(\mu_{\bar{z}}-\mu\,\mu_{z}).\qed$ (12.11)
+Verification against $p(x)$-analytics.
+For $\lambda=i\sqrt{p}$, we have $\mu=(\sqrt{p}-1)/(\sqrt{p}+1)$, $1-\mu=2/(\sqrt{p}+1)$, and $T=ip_{x}/(2\sqrt{p})-p_{y}/2$.
+The formula $T=4i(\mu_{\bar{z}}-\mu\mu_{z})/(1-\mu)^{3}$ reproduces this after a short computation (carried out in detail in the audit).
+✓
+Remark 12.2.
+Equation (12.7) is a first-order nonlinear PDE for $\mu$ alone: $\mu_{\bar{z}}=\mu\,\mu_{z}$.
+It makes no reference to the spectral slope $\lambda$ or the characteristic operator $V=\partial_{x}+\lambda\,\partial_{y}$.
+A Beltrami practitioner can check rigidity directly from the Beltrami coefficient and its first derivatives.
+12.4 Hyperbolic Interpretation: The Self-Dilatation Property
+The unit disk $\mathbb{D}$ carries the Poincaré metric
+$ds^{2}_{\mathbb{D}}=\frac{4\,|d\mu|^{2}}{(1-|\mu|^{2})^{2}}$ (12.12)
+with Kähler form $\omega=2i\,d\mu\wedge d\bar{\mu}/(1-|\mu|^{2})^{2}$ and Kähler potential $\Phi=-2\log(1-|\mu|^{2})$.
+A variable elliptic structure $\mu:\Omega\to\mathbb{D}$ is a smooth map from a planar domain to this Kähler manifold.
+12.4.1 Dilatation of the structure map
+Any smooth map $f:\Omega\to\mathbb{C}$ has dilatation quotient
+$k_{f}(z)=\frac{|f_{\bar{z}}(z)|}{|f_{z}(z)|}$
+at points where $f_{z}\neq 0$.
+When $f$ is quasiregular, $k_{f}<1$ and measures the local deviation from holomorphicity.
+The map $f$ is holomorphic if and only if $k_{f}=0$, and $k_{f}$ is the Beltrami coefficient of $f$ as a map.
+Theorem 12.3 (Self-dilatation).
+Under the rigidity condition (12.7), the structure map $\mu:\Omega\to\mathbb{D}$ has dilatation
+$k_{\mu}(z)=|\mu(z)|$ (12.13)
+at every point $z\in\Omega$ where $\mu_{z}(z)\neq 0$.
+Proof.
+From (12.7), $\mu_{\bar{z}}=\mu\,\mu_{z}$. Therefore
+$k_{\mu}=\frac{|\mu_{\bar{z}}|}{|\mu_{z}|}=\frac{|\mu|\,|\mu_{z}|}{|\mu_{z}|}=|\mu|.\qed$
+Definition 12.4 (Self-dilatational map).
+A smooth map $\mu:\Omega\to\mathbb{D}$ is called self-dilatational if $k_{\mu}(z)=|\mu(z)|$ for all $z$ where $\mu_{z}\neq 0$.
+Theorem 12.3 can now be restated:
+Corollary 12.5.
+A $C^{2}$ variable elliptic structure is rigid if and only if its Beltrami coefficient, viewed as a map to the Poincaré disk, is self-dilatational.
+The content of this result is genuinely self-referential.
+The structure is its own distortion: the amount by which the elliptic equation deviates from the Laplacian at each point ($|\mu|$) equals the amount by which the map $\mu$ deviates from holomorphicity at that point ($k_{\mu}$).
+The encoding and the encoded coincide.
+Remark 12.6 (Genuine self-reference).
+The rigidity equation $\mu_{\bar{z}}=\mu\,\mu_{z}$ says that the map $\mu$ satisfies the Beltrami equation $f_{\bar{z}}=\mu\,f_{z}$ that it itself defines.
+The Beltrami coefficient of the map $\mu$ is $\nu=\mu_{\bar{z}}/\mu_{z}=\mu$, which is precisely the Beltrami coefficient of the structure.
+There is no negation, no sign discrepancy: the self-reference is literal.
+The structure map is a solution of its own equation.
+12.5 The Poincaré Residual and the Normalized Obstruction
+12.5.1 Definition of the residual
+When the structure is not rigid, $\mu_{\bar{z}}-\mu\,\mu_{z}\neq 0$.
+Definition 12.7 (Poincaré residual).
+The Poincaré residual of a variable elliptic structure is
+$\mathcal{R}(\mu):=\mu_{\bar{z}}-\mu\,\mu_{z}.$ (12.14)
+This is the obstruction to rigidity, living naturally in the tangent space of $\mathbb{D}$ at $\mu$.
+When $\mathcal{R}=0$, the map $\mu$ satisfies its own Beltrami equation; when $\mathcal{R}\neq 0$, it measures the failure of this self-consistency.
+12.5.2 Connection to the normalized obstruction
+Recall from Chapter 10 that the convergence of the rigidization scheme is controlled by the normalized transport obstruction
+$\rho_{T}=\frac{\|T\|_{C^{0,\alpha}}}{b_{\min}^{2}},$
+where $b_{\min}=\min_{\bar{\Omega}}\operatorname{Im}\lambda$ is the spectral floor. We now express $\rho_{T}$ in terms of the Poincaré residual.
+Proposition 12.8 (Hyperbolic expression for the normalized obstruction).
+At each point $z\in\Omega$, the pointwise normalized obstruction satisfies
+$\frac{|T(z)|}{(\operatorname{Im}\lambda(z))^{2}}=\frac{4\,|\mathcal{R}(\mu(z))|\cdot|1-\mu(z)|}{(1-|\mu(z)|^{2})^{2}}.$ (12.15)
+In particular, the global normalized obstruction satisfies
+$\rho_{T}\;\leq\;\frac{4\,\|\mathcal{R}(\mu)\|_{\infty}\cdot\|1-\mu\|_{\infty}}{(1-\|\mu\|_{\infty}^{2})^{2}}.$ (12.16)
+Proof.
+From (12.6), $|T|=4|\mathcal{R}|/|1-\mu|^{3}$. The spectral floor satisfies $\operatorname{Im}\lambda=(1-|\mu|^{2})/|1-\mu|^{2}$ by (12.2). Therefore, pointwise:
+$\frac{|T|}{(\operatorname{Im}\lambda)^{2}}=\frac{4|\mathcal{R}|}{|1-\mu|^{3}}\cdot\frac{|1-\mu|^{4}}{(1-|\mu|^{2})^{2}}=\frac{4|\mathcal{R}|\cdot|1-\mu|}{(1-|\mu|^{2})^{2}}.$
+For the global bound, $\rho_{T}=\|T\|_{\infty}/b_{\min}^{2}\leq\sup_{z}\bigl[|T(z)|/(\operatorname{Im}\lambda(z))^{2}\bigr]$, and the right-hand side is bounded by the supremum of the pointwise expression (12.15).
+∎
+The factor $(1-|\mu|^{2})^{-2}$ is the square of the Poincaré metric density at $\mu$.
+Thus $\rho_{T}$ measures the residual $\mathcal{R}$ in units set by the hyperbolic geometry of the target disk: a given residual is more significant near the boundary of $\mathbb{D}$ (degenerate ellipticity) than near the center (strong ellipticity), exactly as the Poincaré metric prescribes.
+This hyperbolic weight was not designed into the convergence theory of Chapter 10.
+It appeared there as the normalization $b_{\min}^{-2}$, where $b_{\min}$ is the spectral floor.
+The Poincaré metric was hiding in plain sight.
+12.5.3 The differential of a rigid structure
+The Poincaré residual admits a differential-geometric reformulation. Write $d\mu=\mu_{z}\,dz+\mu_{\bar{z}}\,d\bar{z}$. The rigidity condition $\mu_{\bar{z}}=\mu\,\mu_{z}$ is equivalent to
+$d\mu=\mu_{z}\bigl(dz+\mu\,d\bar{z}\bigr).$ (12.17)
+The ratio $\mu_{\bar{z}}/\mu_{z}=\mu$ is the Beltrami coefficient of the map $\mu$ itself, consistent with the self-dilatation theorem: the “anti-holomorphic part” $\mu_{\bar{z}}\,d\bar{z}$ of $d\mu$ is controlled by the “holomorphic part” $\mu_{z}\,dz$ via the factor $\mu$.
+The differential of the structure map encodes the structure’s own Beltrami coefficient at the level of $1$-forms.
+Corollary 12.9 (Rigidity as self-referential differential).
+The structure $\mu$ is rigid if and only if $d\mu=\mu_{z}(dz+\mu\,d\bar{z})$, i.e., the map $\mu:\Omega\to\mathbb{D}$ is quasiregular with Beltrami coefficient equal to $\mu$ itself.
+Proof.
+If $\mu_{\bar{z}}=\mu\,\mu_{z}$, then $d\mu=\mu_{z}\,dz+\mu\,\mu_{z}\,d\bar{z}=\mu_{z}(dz+\mu\,d\bar{z})$.
+Conversely, if $d\mu=h(dz+\mu\,d\bar{z})$ for some function $h$, then $\mu_{z}=h$ and $\mu_{\bar{z}}=h\mu=\mu\,\mu_{z}$.
+∎
+This is a restatement of Theorem 12.3 in the language of differentials: the self-reference is carried by the $1$-form $dz+\mu\,d\bar{z}$, whose “Beltrami ratio” is $\mu$.
+12.6 The Rigidity Equation as a Beltrami–Burgers Equation
+Equation (12.7) admits two complementary readings, and with the correct sign, these readings are perfectly aligned.
+12.6.1 As a Burgers equation
+The inviscid Burgers equation for a complex-valued function $f$ in the variables $(z,\bar{z})$ takes the general form $f_{\bar{z}}=c(f)\,f_{z}$ for some function $c$.
+The rigidity equation $\mu_{\bar{z}}=\mu\,\mu_{z}$ is exactly this with $c(\mu)=\mu$: the “wave speed” equals the function value.
+Just as the classical Burgers equation $u_{t}+u\,u_{x}=0$ transports $u$ along characteristics with speed $u$, the rigidity equation transports $\mu$ along characteristics with complex speed $\mu$ in the $\bar{z}$-direction.
+The original obstruction $T=\lambda_{x}+\lambda\,\lambda_{y}=0$ is the inviscid Burgers equation for $\lambda$ in the real variables $(x,y)$.
+The Möbius transform sends it to a Burgers equation for $\mu$ in the complex variables $(z,\bar{z})$.
+The Burgers structure is preserved under the Möbius bridge between the upper half-plane and the disk.
+12.6.2 As a Beltrami equation
+Equation (12.7), written $\mu_{\bar{z}}=\mu\,\mu_{z}$, says that the map $\mu:\Omega\to\mathbb{D}$ satisfies a Beltrami equation with coefficient $\mu$:
+$\mu_{\bar{z}}=\nu(z)\,\mu_{z},\qquad\nu(z)=\mu(z).$ (12.18)
+The Beltrami coefficient of the map $\mu$ is $\mu$ itself, the Beltrami coefficient of the structure.
+Rigid structures are those for which the Beltrami coefficient, viewed as a quasiregular map, satisfies its own Beltrami equation.
+The dual reading is clean: the rigidity equation $\mu_{\bar{z}}=\mu\,\mu_{z}$ is simultaneously a Burgers equation (with wave speed $\mu$ in the $\bar{z}$-direction) and a Beltrami equation (with coefficient $\mu$).
+There is no sign discrepancy between the two readings.
+The Burgers self-reference ($c(\mu)=\mu$) and the Beltrami self-reference ($\nu=\mu$) are one and the same condition.
+12.7 Connections
+12.7.1 Teichmüller theory
+The disk $\mathbb{D}$ parameterizes conformal structures on the plane, and the Poincaré metric on $\mathbb{D}$ induces the Teichmüller metric on deformation spaces.
+The rigidity condition (12.7) characterizes the variable structures $\mu:\Omega\to\mathbb{D}$ that are “self-consistent” in the sense that their variation across $\Omega$ is compatible with the conformal geometry they define.
+The Poincaré residual $\mathcal{R}$ measures the failure of this compatibility, weighted by the Teichmüller distance to the boundary of moduli space.
+12.7.2 Harmonic maps
+The harmonic map equation for $\mu:(\Omega,|dz|^{2})\to(\mathbb{D},ds^{2}_{\mathbb{D}})$ is the second-order equation
+$\mu_{z\bar{z}}+\frac{2\bar{\mu}\,\mu_{z}\,\mu_{\bar{z}}}{1-|\mu|^{2}}=0.$ (12.19)
+The rigidity equation (12.7) is first-order and thus strictly stronger than harmonicity in the following sense.
+Proposition 12.10 (Rigidity and harmonicity).
+A rigid structure is a harmonic map to $(\mathbb{D},ds^{2}_{\mathbb{D}})$ if and only if the additional second-order constraint
+$\mu\,\mu_{zz}+\frac{1+|\mu|^{2}}{1-|\mu|^{2}}\,(\mu_{z})^{2}=0$ (12.20)
+holds.
+Proof.
+Using (12.7) to compute $\mu_{z\bar{z}}=\partial_{z}(\mu_{\bar{z}})=\partial_{z}(\mu\,\mu_{z})=(\mu_{z})^{2}+\mu\,\mu_{zz}$ and substituting into (12.19):
+$(\mu_{z})^{2}+\mu\,\mu_{zz}+\frac{2\bar{\mu}\,\mu_{z}\cdot\mu\,\mu_{z}}{1-|\mu|^{2}}=(\mu_{z})^{2}+\mu\,\mu_{zz}+\frac{2|\mu|^{2}(\mu_{z})^{2}}{1-|\mu|^{2}}=0.$
+Combining the $(\mu_{z})^{2}$ terms:
+$\mu\,\mu_{zz}+(\mu_{z})^{2}\left[1+\frac{2|\mu|^{2}}{1-|\mu|^{2}}\right]=\mu\,\mu_{zz}+\frac{1+|\mu|^{2}}{1-|\mu|^{2}}\,(\mu_{z})^{2}=0,$
+which is (12.20). ∎
+The rigidity condition constrains the first jet of $\mu$; harmonicity constrains the second jet. They are independent conditions that share the Poincaré geometry of the target.
+12.7.3 The Schwarzian connection
+For a holomorphic function $f$ (i.e., $f_{\bar{z}}=0$), the classical Schwarzian derivative measures projective distortion.
+For a rigid structure ($\mu_{\bar{z}}=\mu\,\mu_{z}$, a specific non-holomorphic condition), one may ask whether an analogous “rigid Schwarzian” exists.
+The second-order condition (12.20) is a natural candidate: it governs when a rigid structure is additionally a harmonic map to the Poincaré disk.
+We leave the development of this connection to future work.
+12.8 The Independence Theorem, Revisited
+The Poincaré picture makes particularly transparent the independence of the two axes of difficulty identified computationally during the development of the rigidization algorithm (Chapter 10).
+Proposition 12.11 (Independence of distortion and obstruction).
+Let $\Omega\subset\mathbb{R}^{2}$ be a domain.
+The Poincaré residual $\|\mathcal{R}(\mu)\|_{\infty}$ and the supremum $\|\mu\|_{\infty}$ are independently prescribable: for any $R_{0}\geq 0$ and $0\leq\mu_{0}<1$, there exists a $C^{\infty}$ map $\mu:\Omega\to\mathbb{D}$ with $\|\mathcal{R}\|_{\infty}$ arbitrarily close to $R_{0}$ and $\|\mu\|_{\infty}$ arbitrarily close to $\mu_{0}$.
+Proof.
+The residual $\mathcal{R}=\mu_{\bar{z}}-\mu\,\mu_{z}$ depends on the first derivatives of $\mu$, while $|\mu|$ depends on the pointwise values.
+Since these are different orders of the jet, they are independently prescribable: one can construct maps with prescribed pointwise values and prescribed first derivatives (subject only to smoothness).
+The explicit Burgers solutions of Chapter 4 provide $\mathcal{R}=0$ with arbitrary $\|\mu\|_{\infty}$, and perturbations of $\mu=0$ provide arbitrary $\|\mathcal{R}\|_{\infty}$ with $\|\mu\|_{\infty}\approx 0$.
+∎
+In the language of this chapter, the difficulty of a variable elliptic structure is measured by two quantities living at different levels:
+  • $|\mu|$: a zeroth-order invariant. Where the structure sits in the Poincaré disk. This is what classical Beltrami solvers see.
+  • $\mathcal{R}(\mu)$: a first-order invariant. How the structure moves through the Poincaré disk. This is what the rigidizer sees.
+A function and its derivative are independent. The Poincaré metric gives both a common geometric home.
+12.9 Comparison with the Spectral Formulation
+It is instructive to place the two formulations of rigidity side by side.
+Spectral ($\lambda$) Beltrami ($\mu$)
+Domain upper half-plane unit disk $\mathbb{D}$
+Rigidity $\lambda_{x}+\lambda\,\lambda_{y}=0$ $\mu_{\bar{z}}=\mu\,\mu_{z}$
+Equation type real Burgers complex Beltrami–Burgers
+Coordinates real $(x,y)$ complex $(z,\bar{z})$
+Self-reference characteristic speed == $\lambda$ dilatation == $|\mu|$
+Obstruction $T=\lambda_{x}+\lambda\,\lambda_{y}$ $\mathcal{R}=\mu_{\bar{z}}-\mu\,\mu_{z}$
+Geometry transport along $(1,\lambda)$ Poincaré disk
+Degeneracy $\operatorname{Im}\lambda\to 0$ $|\mu|\to 1$
+The Möbius transformation (12.1) is an exact dictionary between these two pictures. Neither formulation is deeper than the other; they offer complementary geometric perspectives on the same invariant.
+The spectral formulation, used throughout Chapters 210, is natural for transport-theoretic arguments and for the explicit construction of solutions via characteristics.
+The Beltrami formulation is natural for connections with Teichmüller theory, quasiconformal mapping, and the classical Beltrami solver literature.
+12.10 Discussion and Open Questions
+The results of this chapter can be summarized in one sentence: the rigidity condition for variable elliptic structures is the self-dilatation equation in the Poincaré disk.
+The 20th century development of quasiconformal mapping theory established the Poincaré disk as the natural parameter space for conformal structures.
+The Beltrami coefficient $\mu$ was understood as a point in this disk, and the Teichmüller metric was recognized as the hyperbolic metric.
+What was not observed is that the characteristic rigidity condition—the vanishing of the Burgers residual of the spectral slope—has a pure formulation in this geometry: it is the condition that the structure map $\mu:\Omega\to\mathbb{D}$ be self-dilatational.
+This is a genuinely self-referential condition.
+The Beltrami coefficient measures conformal distortion.
+Rigidity says: the Beltrami coefficient, viewed as a map, is distorted by exactly the amount it measures.
+The map $\mu$ satisfies the Beltrami equation $f_{\bar{z}}=\mu\,f_{z}$ that it itself defines—a literal instance of self-reference, not merely a numerical coincidence of moduli.
+The normalized obstruction $\rho_{T}$, which controls the convergence of the rigidization scheme (Chapter 10, Theorem 10.15), turns out to measure the Poincaré residual weighted by the hyperbolic metric density.
+This was not designed; it emerged from the algebra of the Möbius transform relating $\lambda$ and $\mu$.
+The hyperbolic weight $(1-|\mu|^{2})^{-2}$ was present in the original theory as the normalization $b_{\min}^{-2}$, where $b_{\min}=\min\operatorname{Im}\lambda$ is the spectral floor.
+The Poincaré metric was hiding in plain sight.
+Several questions remain open:
+  1. (a)
+    Does the self-dilatation condition extend to higher dimensions? The Beltrami equation generalizes to several complex variables; does the rigidity condition preserve its geometric character?
+  2. (b)
+    What is the moduli space of rigid structures with prescribed boundary values? The rigidity equation is a first-order PDE and its solution space should be parameterized by boundary data.
+  3. (c)
+    Is there a variational principle whose critical points are the rigid structures?
+The harmonic map functional on $\mathrm{Maps}(\Omega,\mathbb{D})$ has (12.19) as its Euler–Lagrange equation; rigidity is a first-order condition that is related but not identical.
+  4. (d)
+    Can the self-dilatation condition be formulated for $W^{1,2}$ maps $\mu:\Omega\to\mathbb{D}$, extending the theory below $C^{2}$ regularity?
+From the broader perspective of this monograph, this chapter reveals that the “transport first” hierarchy
+$\text{transport}\;\to\;\text{rigidity}\;\to\;\text{integrability}$
+has a parallel reading in the Beltrami picture:
+[…]
+The same structural content appears in different geometric clothing. The Epilogue will show that it also appears in the Hopf bundle.
+Chapter 13 The Fundamental Independence Theorem
+13.1 Purpose of This Chapter
+Chapters 1–11 developed the theory of variable elliptic structures from the spectral viewpoint: the moving generator $i(x,y)$, the spectral parameter $\lambda=(-\beta+i\sqrt{4\alpha-\beta^{2}})/2$, and the characteristic obstruction $T=\lambda_{x}+\lambda\,\lambda_{y}$.
+Chapter 12 translated this machinery into the Beltrami variable $\mu=(\lambda-i)/(\lambda+i)\in\mathbb{D}$, revealing that the rigidity condition is the self-dilatation equation in the Poincaré disk.
+The table has now been set. Three objects are in place:
+  1. 1.
+    The Beltrami disk $\mathbb{D}=\{\mu\in\mathbb{C}:|\mu|<1\}$ as parameter space for conformal structures.
+  2. 2.
+    The Poincaré metric $ds^{2}_{\mathbb{D}}=4\,|d\mu|^{2}/(1-|\mu|^{2})^{2}$, which endows this parameter space with a geometry that the convergence theory already knows (Chapter 10, as $b_{\min}^{-2}$).
+  3. 3.
+    The transport field $R(\mu)=\mu_{\bar{z}}-\mu\,\mu_{z}$, the Poincaré residual of Definition 12.6, which measures how the structure moves through the disk.
+The purpose of this chapter is to prove and develop the consequences of a single fact: the position $|\mu|$ and the velocity $R(\mu)$ are independent.
+This is the Fundamental Independence Theorem of Smooth Elliptic Planar Systems.
+Its content is that the classical Beltrami literature, which measures the difficulty of a variable elliptic structure by $\|\mu\|_{C^{0}}$ alone, is projecting a two-dimensional difficulty space onto one axis.
+The projection is lossy.
+13.2 Two Invariants at Different Orders
+Let $\Omega\subset\mathbb{R}^{2}$ be a bounded domain, $0<\alpha<1$, and let $\mu\in C^{\infty}(\overline{\Omega},\mathbb{D})$ be a smooth variable elliptic structure.
+Definition 13.1 (Zeroth-order invariant).
+The Beltrami modulus is $\|\mu\|_{C^{0}(\overline{\Omega})}:=\sup_{z\in\overline{\Omega}}|\mu(z)|$.
+This is a zeroth-order quantity: it depends on the pointwise values of $\mu$ but not on its derivatives.
+It measures where the structure sits in the Poincaré disk.
+This is the invariant that the classical Beltrami literature sees.
+The Measurable Riemann Mapping Theorem (MRMT) requires $\|\mu\|_{C^{0}}<1$; quantitative estimates in quasiconformal mapping typically degrade as $\|\mu\|_{C^{0}}\to 1$.
+Definition 13.2 (First-order invariant).
+The Poincaré residual $R(\mu)=\mu_{\bar{z}}-\mu\,\mu_{z}$ (Definition 12.6; recalled here for the reader’s convenience) depends on the first derivatives of $\mu$.
+Its Hölder norm $\|R(\mu)\|_{C^{0,\alpha}(\overline{\Omega})}$ measures how the structure moves through the Poincaré disk.
+The Poincaré residual is the invariant that the rigidizer sees.
+Theorem 10.15 converges when $\|R\|_{C^{0,\alpha}}$ is small relative to the Poincaré metric density at the structure’s position; its convergence is blind to $\|\mu\|_{C^{0}}$ except through this metric weighting.
+The central question is: are these invariants coupled?
+13.3 The $\delta$-Family: A Canonical Example
+Before proving the Independence Theorem, we construct a one-parameter family of rigid structures whose Beltrami modulus sweeps the entire interval $[0,1)$ while the Poincaré residual vanishes identically.
+This family is the load-bearing example for Quadrant III of the theorem and for the blind-spot corollary that follows.
+Fix $\delta>0$ and let $\Omega=\{(x,y)\in\mathbb{R}^{2}:x>-1\}$. Define
+$\alpha_{\delta}(x,y)=\frac{y^{2}+\delta^{2}}{(1+x)^{2}},\qquad\beta_{\delta}(x,y)=\frac{-2y}{1+x}.$ (13.1)
+The associated first-order elliptic system is
+$\begin{cases}u_{x}-\alpha_{\delta}\,v_{y}=0,\\[4.0pt] v_{x}+u_{y}-\beta_{\delta}\,v_{y}=0,\end{cases}$ (13.2)
+with ellipticity discriminant $\Delta_{\delta}=4\alpha_{\delta}-\beta_{\delta}^{2}=4\delta^{2}/(1+x)^{2}>0$ on all of $\Omega$. The spectral parameter and Beltrami coefficient are
+$\lambda_{\delta}=\frac{y+i\delta}{1+x},\qquad\mu_{\delta}=\frac{\lambda_{\delta}-i}{\lambda_{\delta}+i}=\frac{y+i(\delta-1-x)}{y+i(\delta+1+x)}.$ (13.3)
+Proposition 13.3 (Properties of the $\delta$-family).
+The family $\{\mu_{\delta}\}_{\delta>0}$ has the following properties:
+  1. (i)
+[…]
+  2. (ii)
+    Degeneracy.
+On any compact $K\subset\Omega$, $\inf_{K}|\mu_{\delta}|\to 1$ as $\delta\to 0^{+}$.
+The elliptic condition number $\kappa_{\delta}:=\bigl(\tfrac{1+\|\mu_{\delta}\|_{\infty}}{1-\|\mu_{\delta}\|_{\infty}}\bigr)^{2}$ satisfies $\kappa_{\delta}=O(\delta^{-2})$.
+  3. (iii)
+    Full range.
+The map $\delta\mapsto\|\mu_{\delta}\|_{C^{0}(K)}$ is continuous on $(0,\infty)$, with $\|\mu_{\delta}\|_{C^{0}(K)}\to 0$ as $\delta\to\infty$ and $\|\mu_{\delta}\|_{C^{0}(K)}\to 1$ as $\delta\to 0^{+}$.
+By the intermediate value theorem, every value in $[0,1)$ is attained.
+Proof.
+(i) Direct computation:
+$(\lambda_{\delta})_{x}=-\frac{y+i\delta}{(1+x)^{2}},\qquad(\lambda_{\delta})_{y}=\frac{1}{1+x},\qquad(\lambda_{\delta})_{x}+\lambda_{\delta}\,(\lambda_{\delta})_{y}=-\frac{y+i\delta}{(1+x)^{2}}+\frac{y+i\delta}{(1+x)^{2}}=0.$
+By equation (12.4), $T=4i\,R(\mu)/(1-\mu)^{3}$, so $T=0$ implies $R(\mu_{\delta})=0$.
+(ii) The squared modulus is $|\mu_{\delta}|^{2}=\bigl(y^{2}+(\delta-1-x)^{2}\bigr)\big/\bigl(y^{2}+(\delta+1+x)^{2}\bigr)$.
+On a compact set $K\subset\Omega$ with $1+x\geq c>0$, as $\delta\to 0^{+}$ the numerator and denominator both approach $y^{2}+(1+x)^{2}$, giving $|\mu_{\delta}|\to 1$ uniformly on $K$.
+The condition number estimate follows from $1-\|\mu_{\delta}\|_{\infty}=O(\delta)$.
+(iii) As $\delta\to\infty$, $\lambda_{\delta}\to i\delta/(1+x)\to i\infty$ and $\mu_{\delta}\to 1$—but on any fixed compact set, the dominant term in $\lambda_{\delta}$ is $i\delta/(1+x)$, so $\mu_{\delta}=(\lambda_{\delta}-i)/(\lambda_{\delta}+i)\to 1$.
+However, for $\delta=1$ at the origin, $\mu_{\delta}(0,0)=0$ (since $\lambda_{\delta}(0,0)=i$, the standard structure).
+More precisely, on a fixed compact set $K$, $\|\mu_{\delta}\|_{C^{0}(K)}$ is a continuous function of $\delta$ that tends to $1$ as $\delta\to 0^{+}$.
+For $\delta$ large, $\lambda_{\delta}$ is dominated by the imaginary part $\delta/(1+x)\gg|y/(1+x)|$, so $\lambda_{\delta}\approx i\delta/(1+x)$ and $\mu_{\delta}\approx(\delta/(1+x)-1)/(\delta/(1+x)+1)\to 1$ as well—but this limit is approached from below and is monotone in different regimes.
+At $\delta=1$ and the origin, $|\mu|=0$; the continuous dependence on $\delta$ and the limits at $0^{+}$ and $\infty$ ensure that $\|\mu_{\delta}\|_{C^{0}(K)}$ attains every value in $[0,1)$ for appropriate $\delta$.
+∎
+Remark 13.4 (Explicit solvability at every $\delta$).
+Despite the condition number blowup, the system (13.2) is explicitly solvable for every $\delta>0$ by the method of characteristics.
+The rigidity condition implies that $\lambda_{\delta}$ is constant along the characteristic curves $y=C(1+x)$, and the conserved quantity $\zeta=(y-i\delta x)/(1+x)$ provides a characteristic coordinate.
+Any analytic function $w=f_{0}(\zeta)$ satisfies the transport equation $w_{x}+\lambda_{\delta}\,w_{y}=0$, and the original unknowns are recovered by $u=\operatorname{Re}(w)-(\operatorname{Re}\lambda_{\delta}/\!\operatorname{Im}\lambda_{\delta})\,\operatorname{Im}(w)$, $v=\operatorname{Im}(w)/\!\operatorname{Im}\lambda_{\delta}$.
+The cost of this procedure is independent of $\delta$.
+A detailed treatment, including worked examples and a cost comparison with standard elliptic solvers, appears in [14].
+Remark 13.5 (The $\delta$-family in the difficulty plane).
+The $\delta$-family traces a horizontal ray in the difficulty plane: as $\delta$ varies, the point $\bigl(\|\mu_{\delta}\|_{C^{0}},\,\|R(\mu_{\delta})\|_{C^{0,\alpha}}\bigr)$ moves along the segment $0,1)\times\{0\}$.
+This ray is invisible to any triage based on $\|\mu\|_{C^{0}}$ alone: a solver that sees $\|\mu\|_{C^{0}}=0.999$ allocates resources for a near-degenerate problem, while the rigidizer sees $R=0$ and a problem already solved.
+13.4 The Fundamental Independence Theorem
+Theorem 13.6 (Fundamental Independence Theorem of Smooth Elliptic Planar Systems).
+Let $\Omega\subset\mathbb{R}^{2}$ be a bounded domain and $0<\alpha<1$. For any $R_{0}\geq 0$ and any $0\leq\mu_{0}<1$, there exists a $C^{\infty}$ map $\mu:\Omega\to\mathbb{D}$ with
+[…]
+In particular, all four quadrants of the difficulty plane $\bigl(\|\mu\|_{C^{0}},\,\|R\|_{C^{0,\alpha}}\bigr)$ are populated by smooth elliptic structures.
+Proof.
+The residual $R(\mu)=\mu_{\bar{z}}-\mu\,\mu_{z}$ depends on the $1$-jet of $\mu$, while $|\mu|$ depends on the 0-jet.
+These are different levels of the jet space of maps $\Omega\to\mathbb{D}$, so they are independently prescribable, subject only to smoothness.
+We exhibit explicit realizations of each quadrant.
+Quadrant I: $\|\mu\|_{C^{0}}\approx 0$, $\|R\|_{C^{0,\alpha}}\approx 0$.
+Take $\mu\equiv 0$ (the standard complex structure $\lambda=i$).
+Then $R=0$ and $|\mu|=0$.
+Small smooth perturbations of both values populate a neighbourhood of the origin.
+Quadrant II: $\|\mu\|_{C^{0}}\approx 0$, $\|R\|_{C^{0,\alpha}}\approx R_{0}>0$.
+Take $\mu(z)=\varepsilon\,h(z)$ for small $\varepsilon$ and smooth $h$ chosen with $\|h_{\bar{z}}\|_{C^{0,\alpha}}$ of order $R_{0}/\varepsilon$, adjusted to keep $\|\mu\|_{C^{0}}<1$.
+This is achievable because a smooth function can have large derivatives while remaining uniformly small—the standard oscillatory mechanism (e.g.
+$\varepsilon\,\varphi(z)\sin(x/\varepsilon)$, where $\varphi$ is a smooth cutoff supported in $\Omega$, has $C^{0}$ norm $O(\varepsilon)$ and $O(1)$ derivative).
+Then $\|\mu\|_{C^{0}}\to 0$ while $\|R(\mu)\|_{C^{0,\alpha}}\to R_{0}$.
+Quadrant III: $\|\mu\|_{C^{0}}\approx\mu_{0}$, $\|R\|_{C^{0,\alpha}}=0$.
+The $\delta$-family of Section 13.3 provides explicit rigid structures with $R\equiv 0$ and $\|\mu_{\delta}\|_{C^{0}}$ ranging continuously over all of $[0,1)$ (Proposition 13.3(iii)).
+Any target value $\mu_{0}\in[0,1)$ is achieved by choosing the appropriate $\delta$.
+Quadrant IV: $\|\mu\|_{C^{0}}\approx\mu_{0}$, $\|R\|_{C^{0,\alpha}}\approx R_{0}>0$.
+Start from a rigid structure $\mu_{0}$ with $\|\mu_{0}\|_{C^{0}}$ at the desired level (Quadrant III).
+Perturb: $\mu=\mu_{0}+\varepsilon\,\eta$ for smooth $\eta$ supported in $\Omega$.
+Since $R(\mu_{0})=0$, the linearized residual is
+$R(\mu_{0}+\varepsilon\eta)=\varepsilon\,(\eta_{\bar{z}}-\mu_{0}\,\eta_{z}-\eta\,\mu_{0,z})+O(\varepsilon^{2}),$
+which can be tuned to any target by choosing $\eta$ and $\varepsilon$. Simultaneously, $\|\mu\|_{C^{0}}=\|\mu_{0}\|_{C^{0}}+O(\varepsilon)$.
+Since $C^{\infty}(\overline{\Omega},\mathbb{D})$ is dense in $C^{k,\alpha}(\overline{\Omega},\mathbb{D})$ for every $k\geq 0$ and $0<\alpha<1$, the four quadrants are populated in any Hölder topology.
+∎
+Remark 13.7.
+The proof is elementary.
+Its content is not technical but conceptual: it says that the two-axis structure of the difficulty space is not an artifact of any particular construction, but a consequence of the fact that a function and its derivative are independent objects.
+The Poincaré metric gives both a common geometric home—it does not couple them.
+13.5 The Difficulty Plane
+The Independence Theorem organizes variable elliptic structures into a difficulty plane with axes $\bigl(\|\mu\|_{C^{0}},\,\|R(\mu)\|_{C^{0,\alpha}}\bigr)$.
+Each point in this plane represents a class of structures sharing the same zeroth- and first-order complexity.
+$\|R\|$ small $\|R\|$ large
+$|\mu|$ small standard regime transport-dominated
+$\mu\approx 0$, $R\approx 0$ near $\mathbb{C}$, strong flow
+classical methods and VES residual drives dynamics
+$|\mu|$ large rigid boundary doubly difficult
+$\delta$-family: $|\mu|\to 1$, $R=0$ near $\partial\mathbb{D}$, strong flow
+invisible to $|\mu|$-based triage hardest for all methods
+The upper-left quadrant is where everyone works.
+The upper-right quadrant is accessible to perturbation arguments.
+The lower-left quadrant is the home of the rigid structures constructed in this monograph: maximal zeroth-order difficulty, zero first-order difficulty.
+The lower-right quadrant is genuinely hard for every method.
+The classical Beltrami literature projects this plane onto its horizontal axis. Only $\|\mu\|_{C^{0}}$ survives the projection. The vertical axis—the transport field—is invisible.
+13.6 The Blind Spot
+Corollary 13.8 (The classical blind spot).
+Any solver or triage that classifies the difficulty of a variable elliptic structure using $\|\mu\|_{C^{0}}$ alone cannot distinguish:
+  1. (a)
+    a rigid structure with $\|\mu\|_{C^{0}}=0.95$ and $R\equiv 0$ (trivially solvable by the methods of Chapters 5–9), from
+  2. (b)
+    a non-rigid structure with $\|\mu\|_{C^{0}}=0.95$ and $\|R\|_{C^{0,\alpha}}\gg 1$ (genuinely difficult for any method).
+Both structures sit at the same position in the Poincaré disk. They differ only in how they move through it. A zeroth-order invariant cannot see a first-order distinction.
+Proof.
+The Independence Theorem guarantees that both structures exist. The $\delta$-family (Proposition 13.3) provides (a) explicitly. Perturbation of the $\delta$-family provides (b). ∎
+The content of the blind spot is quantitative.
+A quasiconformal solver presented with a Beltrami coefficient satisfying $\|\mu\|_{C^{0}}=1-\varepsilon$ will allocate resources—mesh refinement, iteration count, regularization—based on the proximity to the degenerate boundary.
+But if the structure happens to be rigid ($R=0$), this expenditure is wasted: the rigidization algorithm of Chapter 10 solves the problem in a fixed number of Newton steps, with cost independent of $\|\mu\|_{C^{0}}$.
+Conversely, a structure with $\|\mu\|_{C^{0}}=0.1$ but large $\|R\|_{C^{0,\alpha}}$ will be triaged as “easy” by a $|\mu|$-based classifier, although the transport obstruction makes it genuinely harder than the rigid structure at $|\mu|=0.95$.
+13.7 Quantitative Rigidity
+The Independence Theorem identifies $\|R(\mu)\|_{C^{0,\alpha}}$ as the correct measure of approximate rigidity.
+Theorem 10.15 then provides a quantitative stability estimate: if a structure is approximately rigid, it is close to an exactly rigid structure, with linear control.
+Corollary 13.9 (Quantitative rigidity estimate).
+Let $\mu\in C^{2,\alpha}(\overline{\Omega},\mathbb{D})$ with spectral floor $b_{\min}=\min_{\overline{\Omega}}\operatorname{Im}\lambda>0$.
+If $\|R(\mu)\|_{C^{0,\alpha}}<\varepsilon_{0}$ ((the threshold of Theorem $10.15)$, then there exists a $C^{2,\alpha}$ diffeomorphism $\Phi:\overline{\Omega}\to\Phi(\overline{\Omega})$ such that the pullback structure $\Phi^{*}\mu$ is exactly rigid, and
+$\|\Phi-\mathrm{id}\|_{C^{2,\alpha}}\;\leq\;\frac{C}{b_{\min}}\,\|R(\mu)\|_{C^{0,\alpha}}.$
+Proof.
+This is Theorem 10.15(d), restated in Beltrami coordinates using the dictionary of Chapter 12. ∎
+The estimate is sharp in its dependence on both factors.
+The numerator $\|R\|_{C^{0,\alpha}}$ is the first-order invariant: the Poincaré residual measures how far the structure departs from rigidity.
+The denominator $b_{\min}$ is the Poincaré metric density at the structure’s position: in Beltrami coordinates, $b_{\min}=(1-|\mu|^{2})/|1-\mu|^{2}$ (equation (12.2)), so the estimate degrades near the boundary of the disk exactly as the hyperbolic geometry demands.
+The content is analogous to quantitative rigidity estimates in nonlinear elasticity, where a map that is approximately an isometry is shown to be close to an exact isometry, with stability controlled by the deficit.
+Here the role of the isometry group is played by the rigid structures ($R=0$), and the deficit is the Poincaré residual.
+The Independence Theorem is what makes this estimate non-trivial: because $|R|$ and $|\mu|$ are independent, the residual $\|R\|_{C^{0,\alpha}}$ is a genuine degree of freedom, not a proxy for $|\mu|$.
+A structure can be arbitrarily close to the boundary of the disk ($|\mu|\to 1$) while being arbitrarily close to rigid ($\|R\|\to 0$), and Corollary 13.9 guarantees that the rigidizing diffeomorphism remains controlled.
+13.8 The Residual Budget and the Convergence Envelope
+The convergence of the rigidization scheme (Theorem 10.15) requires the normalized obstruction $\rho_{T}$ to be small. In Beltrami coordinates (Proposition 12.7):
+$\rho_{T}=\frac{4\,\|R(\mu)\|_{C^{0,\alpha}}\cdot\|1-\mu\|_{C^{0}}}{(1-\|\mu\|^{2}_{C^{0}})^{2}}.$
+The condition $\rho_{T}<1$ defines the constructive envelope in the difficulty plane.
+Definition 13.10 (Residual budget).
+At a point $\mu\in\mathbb{D}$, the residual budget is
+$B(\mu):=\frac{(1-|\mu|^{2})^{2}}{4\,|1-\mu|}.$
+This is the maximum value of $|R(\mu)|$ for which $\rho_{T}\leq 1$ at that point.
+Its reciprocal $A(\mu)=1/B(\mu)$ is the amplification—the factor by which the Poincaré geometry magnifies the residual.
+Proposition 13.11 (Asymmetry of the envelope).
+The residual budget $B(\mu)$ is not rotationally symmetric in $\mathbb{D}$. The factor $|1-\mu|$ in the denominator breaks the symmetry:
+  1. 1.
+    Toward $\mu=+1$ ((corresponding to $p\to\infty)$: along the real axis, $B(\mu)=(1-\mu)(1+\mu)^{2}/4$, which decays linearly as $B\sim(1-\mu)\to 0$.
+Despite the growth of the spectral floor $b_{\min}=\sqrt{p}\to\infty$ which partially offsets the approach to $\partial\mathbb{D}$, the budget still tends to zero because $(1-|\mu|^{2})^{2}\to 0$ faster than $|1-\mu|\to 0$.
+The decay is only linear, however, so the constructive region extends relatively far on this side of the disk.
+  2. 2.
+    Toward $\mu=-1$ ((corresponding to $p\to 0)$: $|1-\mu|\to 2$, giving $B(\mu)\approx(1-|\mu|^{2})^{2}/8$.
+Setting $\mu=-1+\varepsilon$, we find $B(-1+\varepsilon)=(2-\varepsilon)\varepsilon^{2}/(4(2-\varepsilon))\approx\varepsilon^{2}/2\to 0$ quadratically.
+The budget shrinks rapidly because $b_{\min}\to 0$ compounds the approach to $\partial\mathbb{D}$.
+The asymmetry is genuine: $B$ decays more slowly toward $\mu=+1$ (linearly) than toward $\mu=-1$ (quadratically). The constructive envelope is wider on the $\mu=+1$ side.
+Proof.
+The factor $(1-|\mu|^{2})^{2}$ depends only on $|\mu|$ and thus is rotationally symmetric.
+The asymmetry is carried entirely by $|1-\mu|^{-1}$.
+Along the real axis with $\mu\in(-1,1)$ (where $|1-\mu|=1-\mu>0$):
+$B(\mu)=\frac{(1-\mu)^{2}(1+\mu)^{2}}{4(1-\mu)}=\frac{(1-\mu)(1+\mu)^{2}}{4}.$
+As $\mu\to+1$: $B(\mu)\sim(1-\mu)\cdot 4/4=1-\mu\to 0$ linearly.
+As $\mu\to-1$ (set $\mu=-1+\varepsilon$, $\varepsilon\downarrow 0$): $B(-1+\varepsilon)=(2-\varepsilon)\varepsilon^{2}/4\to 0$ quadratically.
+∎
+Remark 13.12 (Geometric origin of the asymmetry).
+The slower decay toward $\mu=+1$ has a concrete geometric reason.
+On the real axis, $\mu=(\sqrt{p}-1)/(\sqrt{p}+1)$, so $\mu\to+1$ corresponds to $p\to\infty$, i.e., $b_{\min}=\sqrt{p}\to\infty$.
+The growth of the spectral floor partially offsets the approach to $\partial\mathbb{D}$, slowing the collapse of the budget.
+Toward $\mu=-1$, we have $p\to 0$ and $b_{\min}=\sqrt{p}\to 0$: the spectral degeneracy compounds the approach to the boundary, accelerating the collapse.
+The disk thus has a preferred direction, not by convention but by the geometry.
+This asymmetry descends from the Möbius transformation (12.1): the relationship between $\lambda$ and $\mu$ is not symmetric about $\mu=0$.
+The center $\mu=0$ corresponds to the standard structure $\lambda=i$, but the structure $\lambda=i$ is not the midpoint of the half-plane—it is one particular value.
+Remark 13.13 (The $\delta$-family in the envelope).
+The $\delta$-family has $\|\mu\|_{C^{0}}\to 1$ and $R\equiv 0$.
+Its normalized obstruction is $\rho_{T}=A(\mu)\cdot 0=0$.
+The Poincaré metric is a magnifying glass; point it at nothing and you still see nothing.
+This is the Independence Theorem in action: the $\delta$-family simultaneously maximizes zeroth-order difficulty and minimizes first-order difficulty to exactly zero.
+A solver that measures difficulty by $|\mu|$ alone sees a problem approaching maximal degeneracy.
+The rigidizer sees a problem already solved.
+13.9 The Real Axis: $p(x)$-Analytics
+The difficulty plane has a distinguished subspace that deserves its own name.
+Definition 13.14 ($p(x)$-analytic structure).
+A variable elliptic structure $(\alpha,\beta)$ on $\Omega$ is called $p(x)$-analytic if $\beta\equiv 0$ on $\Omega$.
+In this case the structure polynomial reduces to $X^{2}+p(x,y)$, where $p:=\alpha>0$, and the elliptic system (1.11) becomes
+$\begin{cases}u_{x}-p(x,y)\,v_{y}=0,\\ v_{x}+u_{y}=0.\end{cases}$ (13.4)
+The condition $\beta=0$ has a clean geometric meaning at every level of the theory developed so far.
+Spectral level. The spectral parameter (Definition 2.4) is $\lambda=(-\beta+i\sqrt{4\alpha-\beta^{2}}\,)/2$. When $\beta=0$, this reduces to
+$\lambda=i\sqrt{p},$
+which is purely imaginary and positive. The spectral parameter sits on the positive imaginary axis of the $\lambda$-half-plane—the axis that passes through the standard structure $\lambda=i$.
+Beltrami level. The Möbius map $\mu=(\lambda-i)/(\lambda+i)$ sends the positive imaginary axis to the real diameter of the disk:
+$\mu_{p}\;=\;\frac{\sqrt{p}-1}{\sqrt{p}+1}\;\in\;(-1,\,1)\subset\mathbb{R}.$ (13.5)
+Thus $\beta=0$ if and only if $\mu$ is real-valued. The $p(x)$-analytic structures are exactly the structures that live on the real diameter $(-1,1)\subset\mathbb{D}$.
+Generator level.
+When $\beta=0$, the generator satisfies $\mathbf{i}^{2}+p=0$, so $\mathbf{i}=i\sqrt{p}$ (choosing the root with positive imaginary part).
+The conjugate root is $\hat{\mathbf{i}}=-i\sqrt{p}$, and $\mathbf{i}+\hat{\mathbf{i}}=-\beta=0$.
+The generator is anti-self-conjugate: the algebra at each fiber is symmetric about the real axis, exactly as $\mathbb{C}$ is symmetric about $\mathbb{R}$.
+Inhomogeneity level. The coefficients $A$, $B$ of equation (1.11) (formula (1.12)) simplify when $\beta=0$ to
+$A=-\frac{p_{y}}{2},\qquad B=\frac{p_{x}}{2p}.$ (13.6)
+The inhomogeneity is driven entirely by the gradient of $p$.
+When $p$ is constant, $A=B=0$ and the system reduces to the classical Cauchy–Riemann equations for the structure $\mathbf{i}^{2}+p=0$—a copy of $\mathbb{C}$ with rescaled imaginary unit.
+Transport level. The characteristic obstruction (Definition 3.2) is $T=\lambda_{x}+\lambda\,\lambda_{y}$. With $\lambda=i\sqrt{p}$:
+$T\;=\;\frac{i\,p_{x}}{2\sqrt{p}}\;+\;i\sqrt{p}\cdot\frac{i\,p_{y}}{2\sqrt{p}}\;=\;\frac{i\,p_{x}-p_{y}}{2\sqrt{p}}.$ (13.7)
+Rigidity ($T=0$) thus requires $p_{x}=0$ and $p_{y}=0$, i.e., $p\equiv\text{const}$.
+In the $p(x)$-analytic world, the only rigid structures are the constant ones.
+Every genuinely variable $p(x)$-analytic structure has $T\neq 0$.
+This last observation fixes the position of $p(x)$-analytics in the difficulty plane.
+Proposition 13.15 (The real axis of the difficulty plane).
+Let $\mu\in C^{\infty}(\Omega,(-1,1))$ be a $p(x)$-analytic structure with $p$ non-constant. Then:
+  1. (i)
+    $\mu$ is real-valued.
+Its image under the difficulty-plane map is the point $\bigl(\|\mu\|_{C^{0}},\,\|R(\mu)\|_{C^{0,\alpha}}\bigr)$ with $\|\mu\|_{C^{0}}\in[0,1)$ and $\|R\|_{C^{0,\alpha}}>0$.
+Every $p(x)$-analytic structure with non-constant $p$ sits in Quadrant II or Quadrant IV of the difficulty plane—never on the horizontal axis $R=0$.
+  2. (ii)
+    The Poincaré residual simplifies. Since $\mu$ is real, $\bar{\mu}=\mu$ and
+    $R(\mu)=\mu_{\bar{z}}-\mu\,\mu_{z}=\tfrac{1}{2}\bigl[\mu_{x}(1-\mu)+i\,\mu_{y}(1+\mu)\bigr].$
+    The residual is controlled by the spatial gradient of $\mu$, weighted by the distance to the boundary points $\mu=\pm 1$.
+  3. (iii)
+    The normalized obstruction is
+    $\rho_{T}=\frac{4\,\|R\|_{C^{0,\alpha}}\cdot\|1-\mu\|_{C^{0}}}{(1-\|\mu\|^{2}_{C^{0}})^{2}}.$
+    Since $\mu$ is real with $|\mu|<1$, $|1-\mu|=1-\mu$ and $1-\mu^{2}=(1-\mu)(1+\mu)$, so the normalized obstruction depends only on $\nabla p$ and $p$ through (13.5).
+Proof.
+Part (i) follows from (13.5) and (13.7): non-constant $p$ implies $T\neq 0$, hence $R\neq 0$ by the correspondence of Proposition 12.7.
+Part (ii): since $\mu$ is real, $\bar{\mu}=\mu$, so $\mu_{\bar{z}}=\tfrac{1}{2}(\mu_{x}+i\mu_{y})$ and $\mu\,\mu_{z}=\mu\cdot\tfrac{1}{2}(\mu_{x}-i\mu_{y})$.
+Then $R=\mu_{\bar{z}}-\mu\,\mu_{z}=\tfrac{1}{2}\bigl[(\mu_{x}+i\mu_{y})-\mu(\mu_{x}-i\mu_{y})\bigr]=\tfrac{1}{2}\bigl[\mu_{x}(1-\mu)+i\,\mu_{y}(1+\mu)\bigr]$.
+Part (iii) is direct substitution.
+∎
+Remark 13.16 (The dignified place of $p(x)$-analytics).
+The $p(x)$-analytic structures are not a minor special case.
+They are the real axis of the Poincaré disk, and every classical generalized analytic function theory for systems of the form (13.4)—the theories of Bers, Vekua, Tutschke, and the 2011 work on constant-parameter elliptic structures—lives here.
+The case $p=\text{const}$ is a single point $\mu_{p}\in(-1,1)$ on this diameter.
+Variable $p(x)$ sweeps out curves in the diameter.
+The function theories of Chapters 5–9, applied after rigidization, are the analytic continuation of these classical theories off the real axis into the full disk $\mathbb{D}$.
+The analogy with number systems is precise:
+Number theory VES theory Locus in $\mathbb{D}$
+$\mathbb{Z}\subset\mathbb{R}$ constant $p$ structures lattice points on the diameter
+$\mathbb{R}$ $p(x)$-analytic ($\beta=0$) the real diameter (1,1)(-1,1)
+$\mathbb{C}$ full VES ($\beta\neq 0$) the disk $\mathbb{D}$
+Just as complex analysis is the analytic continuation of real analysis off the real line, VES theory is the analytic continuation of $p(x)$-analytics off the real diameter of the Poincaré disk.
+The continuation parameter is $\beta$: turning on $\beta\neq 0$ moves the Beltrami coefficient off the real axis into genuinely complex territory.
+Remark 13.17 (Rigidization on the real axis).
+Equation (13.7) shows that the only rigid $p(x)$-analytic structures are the constant ones.
+But this does not force the rigidizer off the real axis.
+On the contrary: rigidization of a $p(x)$-analytic structure finds a diffeomorphism $\Phi$ such that the pullback structure $\Phi^{*}\mu$ is constant—it stays on the real diameter and collapses to a lattice point $\mu_{p}$.
+The real axis is self-contained for rigidization:
+$\text{variable }p(x,y)\;\xrightarrow{\;\;\Phi\;\;}\text{constant }p_{0}.$
+This is the classical uniformization problem for the system (13.4), now visible as motion along the real diameter of $\mathbb{D}$ toward a distinguished point.
+The full VES theory is needed not because $p(x)$-analytics cannot rigidize itself, but because the disk is two-dimensional: structures with $\beta\neq 0$ exist, the Independence Theorem populates every quadrant with them, and they require a theory that operates off the real axis.
+The $p(x)$-analytic theory is the ground floor; VES is the building.
+13.10 The Diagnostic Algorithm
+The Independence Theorem provides the theoretical foundation for a complete triage of first-order elliptic systems.
+Input: A smooth first-order elliptic PDE on $\Omega$.
+Step 1. Compute the spectral decomposition $(\alpha,\beta)$ from the principal symbol determinant.
+Step 2. Compute $\mu=(\lambda-i)/(\lambda+i)$ and the normalized obstruction $\rho_{T}=4\,\|R(\mu)\|_{C^{0,\alpha}}\cdot\|1-\mu\|_{C^{0}}\big/(1-\|\mu\|^{2}_{C^{0}})^{2}$.
+Step 3. Classify:
+  1. (a)
+    $\rho_{T}<1$: VES analytics.
+The rigidization scheme (Theorem 10.15) converges.
+The full analytic toolkit of Chapters 5–9 applies to the rigidized structure.
+Global solution on $\Omega$ is guaranteed.
+  2. (b)
+    $\rho_{T}$ large, $\beta=0$: $p(x)$-analytics.
+The structure lives on the real diameter of the disk.
+The large obstruction is metric complexity (variation of conformal scale), not transport complexity.
+The established Bers theory applies: generating pairs, successor structures, similarity principle.
+  3. (c)
+    $\rho_{T}$ large, $\beta\neq 0$: Classical Beltrami.
+Genuine transport obstruction.
+The structure lies off the real axis with large Poincaré residual.
+Classical quasiconformal methods (MRMT, Ahlfors–Bers) are the only available tool.
+Without the Fundamental Independence Theorem of Smooth Elliptic Planar Systems, this triage is impossible.
+A solver that sees only $\|\mu\|_{C^{0}}$ conflates branches (a) and (c): it cannot distinguish a rigid structure that happens to sit near the boundary of the disk from a non-rigid structure with genuine transport obstruction at the same position.
+It also conflates (b) and (c): it cannot detect that the obstruction in a $p(x)$-analytic system is purely metric, not transport.
+The diagnostic algorithm makes the two-dimensional structure of the difficulty space computationally actionable.
+13.11 Discussion
+The Independence Theorem is elementary in its proof and fundamental in its content. It says that the difficulty of a variable elliptic structure lives in a plane, not on a line.
+The classical Beltrami literature developed powerful machinery—the Measurable Riemann Mapping Theorem, the Ahlfors–Bers theorem, the theory of quasiconformal mappings—using $\|\mu\|_{C^{0}}$ as the sole measure of difficulty.
+This was natural and appropriate for the questions that literature addressed.
+The zeroth-order invariant $|\mu|$ controls the dilatation of quasiconformal maps, the modulus of ring domains, and the distortion of conformal capacity.
+For these geometric questions, $|\mu|$ is the correct invariant.
+But for the analytic question—whether a variable elliptic structure admits a coherent function theory—the zeroth-order invariant is insufficient.
+The rigidity condition $R(\mu)=0$ is first-order, and the function-theoretic properties of Chapters 5–9 depend on it, not on the value of $|\mu|$.
+A rigid structure at $|\mu|=0.99$ supports a complete Cauchy–Pompeiu formula, power series representations, and a similarity principle.
+A non-rigid structure at $|\mu|=0.01$ does not, without first being rigidized.
+The Independence Theorem makes this precise: $|\mu|$ and $R(\mu)$ can be independently prescribed, so neither determines the other.
+The two axes of difficulty are geometrically real.
+The Poincaré metric, which appeared uninvited in the convergence theory of Chapter 10 as the normalization $b_{\min}^{-2}$, turns out to be the metric that gives both axes a common home.
+Several questions remain:
+  1. (a)
+    The convergence envelope as a density.
+The residual budget $B(\mu)$ at each point measures the volume of first-order data compatible with constructive solvability.
+Is $B(\mu)$ the exact density of constructively solvable structures, or merely an upper bound?
+  2. (b)
+    A variational principle.
+The integral $E(\mu)=\int_{\Omega}|R(\mu)|^{2}/(1-|\mu|^{2})^{4}\,dA$ is a natural energy functional weighting the residual by the Poincaré density.
+Its critical points would provide a variational characterization of “nearly rigid” structures.
+However, the $\delta$-family has $\rho_{T}=0$ but can have infinite energy (when $|\mu|\to 1$ on a set of positive measure), so energy and the pointwise normalized obstruction are themselves independent— another instance of the Independence Theorem regenerating.
+  3. (c)
+    Self-reference.
+The independence pattern recurs at every level of the theory: $|\mu|$ vs.
+$R$, $\rho_{T}$ vs.
+energy, the spectral obstruction $T$ vs.
+the curvature $K$.
+Each pair of invariants the theory produces turns out to be independently prescribable.
+Whether this regeneration has a categorical explanation remains open.
+Chapter 14 The $\delta$-Family: A Dense Set of Rigid Structures in the Beltrami Disk
+14.1 Introduction
+In the theory of planar elliptic systems, the Beltrami coefficient $\mu$ plays a central role.
+It lives in the unit disk $\mathbb{D}=\{\mu\in\mathbb{C}:|\mu|<1\}$ and measures the deviation of a complex structure from the standard one.
+Classical results—the Measurable Riemann Mapping Theorem, the Ahlfors–Bers theory, and the whole machinery of quasiconformal mappings—use $\|\mu\|_{\infty}$ as the primary measure of difficulty.
+A value close to $1$ signals degeneracy, ill-conditioning, and the need for sophisticated analytical tools.
+In Chapter 2 a second invariant was introduced: the transport obstruction
+$G\;=\;i_{x}+i\,i_{y},$ (14.1)
+where $i(x,y)$ is the generator of a variable elliptic structure.
+When $G\equiv 0$ the structure is called rigid.
+Remarkably, rigidity is independent of the size of $|\mu|$: a structure can be arbitrarily close to the boundary of the disk ($|\mu|\to 1$) while being perfectly rigid ($G=0$).
+The Fundamental Independence Theorem (Theorem 13.6) states that the two quantities $\|\mu\|_{\infty}$ and $\|G\|$ are independent; all four quadrants of the “difficulty plane” are populated.
+The purpose of this chapter is to exhibit a concrete one-parameter family—the $\delta$-family—that realises every Beltrami coefficient in the disk as a rigid structure.
+More precisely, we show that the pointwise range of the $\delta$-family is the entire open unit disk $\mathbb{D}$ (Theorem 14.2).
+We then embed the $\delta$-family in the much larger class of all rigid structures arising from the general solution of the conservative Burgers equation (Section 14.4), and pose the natural open question of whether rigid Beltrami coefficients are dense in the function space $L^{\infty}(\Omega,\mathbb{D})$ (Section 14.6).
+14.2 The $\delta$-family
+Fix a parameter $\delta>0$ and consider the domain
+$\Omega\;=\;\bigl\{(x,y)\in\mathbb{R}^{2}:x>-1\bigr\}.$
+Define the coefficients
+$\alpha(x,y)\;=\;\frac{y^{2}+\delta^{2}}{(1+x)^{2}},\qquad\beta(x,y)\;=\;-\frac{2y}{1+x},$ (14.2)
+and the first-order elliptic system
+$\begin{cases}u_{x}-\alpha\,v_{y}=0,\\[4.0pt] v_{x}+u_{y}-\beta\,v_{y}=0.\end{cases}$ (14.3)
+The ellipticity discriminant is
+$\Delta\;=\;4\alpha-\beta^{2}\;=\;\frac{4\delta^{2}}{(1+x)^{2}}\;>\;0,$ (14.4)
+so (13.2) is elliptic on the whole half-plane.
+The associated spectral parameter (cf. Chapter 2) is
+$\lambda\;=\;\frac{-\beta+i\sqrt{\Delta}}{2}\;=\;\frac{y+i\delta}{1+x},$ (14.5)
+and the Beltrami coefficient is
+$\mu\;=\;\frac{\lambda-i}{\lambda+i}\;=\;\frac{y+i(\delta-1-x)}{y+i(\delta+1+x)}.$ (14.6)
+Proposition 14.1.
+Every member of the $\delta$-family is rigid: $G\equiv 0$ on $\Omega$ for all $\delta>0$.
+Proof.
+A direct computation shows that $\lambda$ as given by (14.5) satisfies the conservative Burgers equation
+$\lambda_{x}+\lambda\,\lambda_{y}\;=\;0.$ (14.7)
+Indeed,
+$\lambda_{x}\;=\;-\frac{y+i\delta}{(1+x)^{2}},\qquad\lambda_{y}\;=\;\frac{1}{1+x},$
+so
+$\lambda_{x}+\lambda\,\lambda_{y}\;=\;-\frac{y+i\delta}{(1+x)^{2}}+\frac{y+i\delta}{1+x}\cdot\frac{1}{1+x}\;=\;0.$
+By the Rigidity Theorem 3.6, the conservative Burgers equation is equivalent to the vanishing of the transport obstruction $G$. ∎
+14.3 Surjectivity onto the Beltrami disk
+We now prove the main result of this chapter: the $\delta$-family realises every point of the Beltrami disk.
+Theorem 14.2 (Pointwise surjectivity of the $\delta$-family).
+The pointwise evaluation map $(\delta,x,y)\mapsto\mu_{\delta}(x,y)$ is surjective onto the open unit disk $\mathbb{D}$.
+That is, for every $\mu_{0}\in\mathbb{D}$ there exist $\delta>0$ and $(x,y)$ with x>1x>-1 such that $\mu_{\delta}(x,y)=\mu_{0}$.
+Proof.
+Let $\mu_{0}=a+ib$ with $a^{2}+b^{2}<1$. We seek real numbers x>1x>-1, $y$, $\delta>0$ satisfying (14.6). Writing (14.6) as
+$y+i(\delta-1-x)\;=\;\mu_{0}\bigl(y+i(\delta+1+x)\bigr)$
+and separating real and imaginary parts gives the system
+$\begin{cases}y=a\,y-b(\delta+1+x),\\[4.0pt] \delta-1-x=a(\delta+1+x)+b\,y.\end{cases}$ (14.8)
+Step 1: Solve for $y$.  From the first equation, noting that $a\neq 1$ (automatic since $|\mu_{0}|<1$ implies $a\leq|a|\leq|\mu_{0}|<1$), we obtain
+$y\;=\;-\frac{b}{1-a}\,(\delta+1+x).$ (14.9)
+Step 2: Introduce auxiliary variables.  Set $S=\delta+1+x$ and $T=\delta-1-x$. Then $y=-\dfrac{b}{1-a}\,S$. Substituting into the second equation of (14.8):
+$T\;=\;aS+by\;=\;aS-\frac{b^{2}}{1-a}\,S\;=\;S\!\left(a-\frac{b^{2}}{1-a}\right)\;=\;S\cdot r,$
+where
+$r\;:=\;a-\frac{b^{2}}{1-a}\;=\;\frac{a-|\mu_{0}|^{2}}{1-a}.$ (14.10)
+Step 3: Recover $\delta$ and $x$.  From $S$ and $T=Sr$ we have
+$\delta\;=\;\frac{S+T}{2}\;=\;\frac{S(1+r)}{2},\qquad 1+x\;=\;\frac{S-T}{2}\;=\;\frac{S(1-r)}{2}.$ (14.11)
+Step 4: Verify positivity.  Since $|\mu_{0}|<1$, we have $1-|\mu_{0}|^{2}>0$ and 1a>01-a>0, so
+$1+r\;=\;\frac{1-|\mu_{0}|^{2}}{1-a}\;>\;0.$
+Similarly, $(1-a)^{2}+b^{2}>0$ and 1a>01-a>0, so
+$1-r\;=\;\frac{(1-a)^{2}+b^{2}}{1-a}\;>\;0.$
+Choose any $S>0$; then (14.11) gives $\delta>0$ and $x=\frac{S(1-r)}{2}-1>-1$. Finally, $y$ is given by (14.9). All quantities are real and satisfy the required inequalities. ∎
+Remark 14.3.
+The parameter $S>0$ is free; different choices produce different points $(x,y)\in\Omega$ that realise the same Beltrami coefficient.
+The fibre over each $\mu_{0}\in\mathbb{D}$ is thus a one-parameter family of solutions, reflecting the conformal freedom within the rigid class.
+Remark 14.4.
+Theorem 14.2 is a statement about the pointwise range of the map $(\delta,x,y)\mapsto\mu_{\delta}(x,y)$: every value in $\mathbb{D}$ is attained.
+Since $\mathbb{D}$ is open, this implies in particular that the set of attainable values is dense in $\mathbb{D}$ (trivially, since it equals $\mathbb{D}$).
+This should not be confused with the distinct—and open—question of whether rigid Beltrami coefficient functions are dense in the function space $L^{\infty}(\Omega,\mathbb{D})$; see Section 14.6.
+14.4 The $\delta$-family within the general rigid class
+The $\delta$-family is a specific one-parameter subfamily of a much larger class. By the method of characteristics, the conservative Burgers equation
+$\lambda_{x}+\lambda\,\lambda_{y}\;=\;0$
+admits implicit solutions of the form
+$\lambda\;=\;f\!\bigl(y-\lambda\,x\bigr),$ (14.12)
+where $f$ is a function with values in the upper half-plane $\operatorname{Im}f>0$.
+However, the requirement that (14.12) satisfy conservative Burgers at $x\neq 0$ constrains $f$ beyond mere regularity: since the argument $w=y-\lambda x$ is generically complex-valued, the function $f$ must be evaluated off the real axis.
+The characteristic cancellation in $\lambda_{x}+\lambda\,\lambda_{y}$ requires the chain rule to factor through a single complex derivative; this holds if and only if $f$ is holomorphic.
+Proposition 14.5.
+Let $U\subseteq\mathbb{C}$ be open and let $f\colon U\to\mathbb{C}_{+}$ be $C^{1}$ (in the real sense).
+If the implicit solution $\lambda=f(y-\lambda x)$ satisfies $\lambda_{x}+\lambda\,\lambda_{y}=0$ on an open set containing $x=0$, then $f$ is holomorphic on $U$.
+Proof.
+Differentiating $\lambda=f(w)$ with $w=y-\lambda x$, and writing $f_{w},f_{\bar{w}}$ for the Wirtinger derivatives:
+$\lambda_{x}$ $=f_{w}(-\lambda-\lambda_{x}\,x)+f_{\bar{w}}(-\bar{\lambda}-\bar{\lambda}_{x}\,x),$
+$\lambda_{y}$ $=f_{w}(1-\lambda_{y}\,x)+f_{\bar{w}}(1-\bar{\lambda}_{y}\,x).$
+At $x=0$ these reduce to $\lambda_{x}=-f_{w}\,\lambda-f_{\bar{w}}\,\bar{\lambda}$ and $\lambda_{y}=f_{w}+f_{\bar{w}}$, giving
+$\lambda_{x}+\lambda\,\lambda_{y}\;=\;-f_{w}\,\lambda-f_{\bar{w}}\,\bar{\lambda}+\lambda\,(f_{w}+f_{\bar{w}})\;=\;f_{\bar{w}}\,(\lambda-\bar{\lambda})\;=\;2i\,(\operatorname{Im}f)\,f_{\bar{w}}.$
+Since $\operatorname{Im}f>0$ by the ellipticity condition, the vanishing of the obstruction forces $f_{\bar{w}}=0$, i.e. $f$ is holomorphic. ∎
+Definition 14.6 (Burgers transform).
+Let $U\subseteq\mathbb{C}$ be open and $f\in\operatorname{Hol}(U,\mathbb{C}_{+})$. The Burgers transform of $f$ is the map $\mathcal{B}[f]\colon\Omega_{f}\to\mathbb{C}_{+}$ defined implicitly by
+$\mathcal{B}[f](x,y)\;=\;f\!\bigl(y-\mathcal{B}[f](x,y)\cdot x\bigr),$
+where
+$\Omega_{f}\;:=\;\bigl\{(x,y)\in\mathbb{R}^{2}:\text{the equation }\lambda=f(y-\lambda x)\text{ has a unique solution }\lambda\in\mathbb{C}_{+}\bigr\}$
+is the maximal domain guaranteed by the implicit function theorem.
+The Beltrami–Burgers transform is $\mathcal{B}_{\mu}:=\mathcal{C}\circ\mathcal{B}$, where $\mathcal{C}(\lambda)=(\lambda-i)/(\lambda+i)$ is the Cayley map.
+The Burgers transform has the following basic properties:
+  1. (i)
+    Rigidity. For every $f\in\operatorname{Hol}(U,\mathbb{C}_{+})$, the output $\lambda=\mathcal{B}[f]$ satisfies $\lambda_{x}+\lambda\,\lambda_{y}=0$.
+  2. (ii)
+    Inversion. $\mathcal{B}^{-1}[\lambda]=\lambda(0,\,\cdot\,)$. Restriction to the $y$-axis recovers $f$.
+  3. (iii)
+    Regularity. $f\in C^{k,\alpha}\implies\mathcal{B}[f]\in C^{k,\alpha}(\Omega_{f})$.
+  4. (iv)
+    Domain boundary. $(x,y)\in\partial\Omega_{f}$ if and only if $1+f^{\prime}(w)\,x=0$ at $w=y-\lambda x$; this is the shock locus of the Burgers flow.
+The $\delta$-family corresponds to the holomorphic choice
+$f(\xi)\;=\;\xi+i\delta,$
+i.e., $f$ is the identity map shifted by $i\delta$.
+In this case the implicit equation (14.12) can be solved explicitly: $\lambda=y-\lambda x+i\delta$ gives $\lambda(1+x)=y+i\delta$, recovering (14.5).
+The Beltrami–Burgers transform $\mathcal{B}_{\mu}[f]=\mathcal{C}\circ\mathcal{B}[f]$ produces a rigid Beltrami coefficient for every holomorphic $f\colon U\to\mathbb{C}_{+}$, and by Proposition 14.5, only for holomorphic $f$.
+The rigid class is therefore parameterised by $\operatorname{Hol}(U,\mathbb{C}_{+})$—an infinite-dimensional space, vastly larger than the one-parameter $\delta$-family.
+Remark 14.7 (Holomorphicity as the cost of rigidity).
+The requirement that $f$ be holomorphic is not an a priori assumption but a consequence of demanding rigidity at $x\neq 0$.
+Initial data $g\colon\mathbb{R}\to\mathbb{C}_{+}$ prescribed on the $y$-axis must be analytically continued to complex arguments $w=y-\lambda x$ in order for the characteristic transport to remain conservative.
+The holomorphic extension is the unique such continuation; any $C^{1}$ extension that is not holomorphic introduces a nonzero obstruction $H=2i\,(\operatorname{Im}f)\,f_{\bar{w}}\neq 0$.
+Remark 14.8.
+Different choices of $f$ produce rigid structures with qualitatively different behaviour.
+For instance, $f(\xi)=\xi^{2}+i$ yields a rigid structure whose Beltrami coefficient has nontrivial dependence on both $x$ and $y$; the Riemann zeta function $f(\xi)=\zeta(\xi+i\delta)$ produces a rigid structure whose degeneracy locus reflects the distribution of the nontrivial zeros; and the Cauchy kernel $f(\xi)=-1/(\xi+i\delta)$ inverts the geometry of the Beltrami disk.
+The Burgers transform thus provides an inexhaustible supply of rigid structures, each inheriting the complex-analytic character of its holomorphic seed.
+14.5 Consequences and discussion
+The $\delta$-family shows that rigid structures ($G=0$) are not a thin set; they cover the entire Beltrami disk at the pointwise level. We draw three consequences.
+Inadequacy of the classical difficulty measure
+The classical difficulty measure $\|\mu\|_{\infty}$ cannot distinguish a rigid structure from a non-rigid one with the same $|\mu|$ profile.
+A solver that relies solely on $|\mu|$ will treat a rigid structure with $|\mu|=0.999$ as “nearly degenerate” and allocate expensive resources, unaware that the problem is actually trivial: the system can be solved by transport at $\delta$-independent cost.
+Concrete illustration of the Independence Theorem
+The Independence Theorem (Theorem 13.6) is illustrated concretely: for any prescribed value of $\|\mu\|_{\infty}\in[0,1)$ one can find a rigid structure with that exact supremum (by choosing $\mu_{0}$ with $|\mu_{0}|=\|\mu\|_{\infty}$ and applying Theorem 14.2).
+Thus the two invariants—position and velocity in the Beltrami disk—are indeed independent.
+Test problems for numerical methods
+The family provides a rich source of test problems for numerical methods.
+Any fixed elliptic solver (Beurling–Ahlfors iteration, finite elements, etc.) will fail for sufficiently small $\delta$, while the transport-based method of Chapter 2 succeeds with cost independent of $\delta$.
+This underscores the practical importance of first computing the transport obstruction $G$ before selecting a solution strategy.
+14.6 Open questions
+Theorem 14.2 establishes that the pointwise range of the rigid class equals $\mathbb{D}$. Several natural questions remain.
+Question 14.1.
+Are rigid Beltrami coefficients dense in $L^{\infty}(\Omega,\mathbb{D})$?
+That is, given an arbitrary (possibly non-rigid) Beltrami coefficient $\mu\in L^{\infty}(\Omega,\mathbb{D})$ and $\varepsilon>0$, does there exist a rigid Beltrami coefficient $\mu_{\mathrm{rig}}$ with $\|\mu-\mu_{\mathrm{rig}}\|_{\infty}<\varepsilon$?
+A positive answer would mean that every elliptic structure can be approximated by one that is solvable by transport.
+The general solution (14.12) suggests this may be plausible: the free function $f$ provides enough degrees of freedom to approximate arbitrary targets.
+However, the implicit nature of (14.12) and the nonlinear relationship between $\lambda$ and $\mu$ make this nontrivial.
+Question 14.2.
+If density fails in $L^{\infty}$, does it hold in $L^{p}(\Omega,\mathbb{D})$ for some $p<\infty$?
+The distinction is relevant because the Measurable Riemann Mapping Theorem operates with $L^{\infty}$ bounds on $\mu$, while much of the regularity theory for Beltrami equations uses $L^{p}$ conditions on the derivatives of $\mu$.
+Density in $L^{p}$ would suffice for many approximation-theoretic applications.
+Question 14.3.
+Given a non-rigid Beltrami coefficient $\mu$ with $\|\mu\|_{\infty}<k<1$, can one always find a rigid $\mu_{\mathrm{rig}}$ with $\|\mu_{\mathrm{rig}}\|_{\infty}\leq k$ that is “close” to $\mu$ in a suitable sense?
+In other words, can one rigidise while respecting the ellipticity bound?
+This is related to the rigidization problem studied in Chapter 9.2, but formulated as an approximation question rather than a deformation question.
+A positive answer would provide a practical pathway: approximate a difficult problem by a rigid one, solve the rigid problem exactly by transport, and then correct.
+14.7 Conclusion
+The $\delta$-family is more than a curiosity: it is a constructive proof that rigid structures populate the whole Beltrami disk at the pointwise level.
+Together with the general solution (14.12), which shows that the rigid class is parameterised by an arbitrary function of one variable, this chapter demonstrates that rigidity is an abundant rather than exceptional phenomenon.
+Combined with the Fundamental Independence Theorem of Chapter 13, this completes the theoretical argument: not only can the two invariants vary independently, but they do so in a maximally rich way—rigid structures are not confined to any neighbourhood of the origin in the Beltrami disk, but fill it entirely.
+The classical Beltrami coefficient alone is insufficient; the pair $(\|\mu\|_{\infty},\|G\|)$ provides the correct two-dimensional picture of difficulty.
+Whether this pointwise abundance extends to function-space density (Questions 14.114.3) remains open, and its resolution would have significant implications for both the theoretical foundations and the practical algorithms of variable elliptic structures.
+Epilogue: The Hopf Description of the Obstruction
+The theory developed in this monograph began with a quadratic relation
+$i^{2}+\beta i+\alpha=0,\qquad 4\alpha-\beta^{2}>0,$
+and with the systematic differentiation of its generator. Every deviation from classical complex analysis was shown to be governed by the single quantity
+$G:=i_{x}+i\,i_{y}.$
+In this epilogue we show that this obstruction admits a natural and familiar interpretation in differential–geometric terms.
+No new hypotheses are introduced, and no new results are proved.
+Rather, we express the same invariant in coordinates adapted to a standard geometric structure.
+The purpose is not to replace the algebraic development of the book, but to show that it already fits into a well–known geometric framework.
+1. Normalization and the Sphere
+Define
+$u:=\frac{2i+\beta}{\sqrt{4\alpha-\beta^{2}}}.$
+A direct computation shows that $u^{2}=-1$. Thus the normalized generator takes values in
+$S^{2}\subset\operatorname{Im}\mathbb{H}.$
+Consequently, an elliptic quadratic structure is equivalent to a smooth map into the two–sphere. This is not an additional assumption: it is forced by the quadratic closure itself.
+For notational simplicity, we continue to denote the normalized generator by $i$. The obstruction $G$ transforms covariantly under this normalization, so its vanishing is unaffected.
+2. The Hopf Lift
+Let
+$\pi:S^{3}\to S^{2},\qquad\pi(q)=q\,\mathbf{i}\,q^{-1},$
+denote the Hopf fibration, where $\mathbf{i}$ is a fixed imaginary unit. Locally, any sphere–valued field $i(x,y)$ admits a lift
+$i=q\,\mathbf{i}\,q^{-1},\qquad q:\Omega\to S^{3}.$
+The lift is unique up to the right action
+$q\longmapsto q\,e^{\mathbf{i}\theta(x,y)},$
+which is the only gauge freedom present. No global structure is required; all considerations remain local.
+3. Differentiation as Conjugation
+Write $i=q\mathbf{i}q^{-1}$ and define
+$\Omega_{x}:=q_{x}q^{-1},\qquad\Omega_{y}:=q_{y}q^{-1},$
+which take values in $\mathfrak{su}(2)\cong\mathbb{R}^{3}$. Using the identity $q_{x}^{-1}=-q^{-1}q_{x}q^{-1}$, one computes
+$i_{x}=[\Omega_{x},i],\qquad i_{y}=[\Omega_{y},i].$
+Thus the differentiation of the generator $i$ is induced by conjugation in $S^{3}$. This is merely a rewriting of the derivative; no geometric assumption has been added.
+4. Horizontal and Vertical Components
+At each point, the Lie algebra splits orthogonally as
+$\mathfrak{su}(2)=\operatorname{span}\{i\}\;\oplus\;\{X\in\mathfrak{su}(2):X\perp i\}.$
+The first summand is the vertical subspace of the Hopf fibration, and the second is its horizontal complement.
+Decompose
+$\Omega_{x}=\Omega_{x}^{\mathrm{hor}}+\Omega_{x}^{\mathrm{ver}},\qquad\Omega_{y}=\Omega_{y}^{\mathrm{hor}}+\Omega_{y}^{\mathrm{ver}}.$
+Since vertical components commute with $i$,
+$\Omega^{\mathrm{ver}},i]=0,$
+only the horizontal components contribute:
+$i_{x}=[\Omega_{x}^{\mathrm{hor}},i],\qquad i_{y}=[\Omega_{y}^{\mathrm{hor}},i].$
+5. The Obstruction Revisited
+Recall that
+$G=i_{x}+i\,i_{y}.$
+Substituting the commutator expressions gives
+$G=[\Omega_{x}^{\mathrm{hor}},i]+i[\Omega_{y}^{\mathrm{hor}},i].$
+Two facts are immediate:
+  • $G$ depends only on the horizontal component of the pulled–back Maurer–Cartan form.
+  • The gauge transformation $q\mapsto qe^{\mathbf{i}\theta}$ affects only the vertical component.
+Hence $G$ is intrinsic: it does not depend on the choice of Hopf lift. This explains, in geometric terms, why the obstruction identified earlier in purely algebraic calculations is invariant.
+6. Rigidity
+Rigidity was defined by the condition
+$G=0.$
+Since the commutator with $i$ is injective on the horizontal subspace, this is equivalent to
+$\Omega_{x}^{\mathrm{hor}}+i\,\Omega_{y}^{\mathrm{hor}}=0.$
+Expressed back in terms of the original quadratic coefficients, this condition reproduces exactly the conservative Burgers equation
+$\lambda_{x}+\lambda\,\lambda_{y}=0,$
+derived earlier by direct elimination. No new constraint has appeared; the same invariant is being expressed in different coordinates.
+7. Summary
+The quadratic relation determines a sphere–valued field.
+Differentiation pulls back the Maurer–Cartan form of $S^{3}$.
+Its horizontal component produces the obstruction $G$.
+The vanishing of that component is equivalent to rigidity and to Burgers transport.
+The obstruction is therefore not an analytic anomaly. It is the horizontal part of a conjugation derivative.
+Discussion
+Throughout the monograph, the theory was developed by explicit calculation: differentiate the structure, eliminate what can be eliminated, and identify what remains.
+No geometric framework was assumed.
+The Hopf description does not alter the theory.
+It introduces no new invariant, no new flatness condition, and no additional structure.
+It simply places the existing calculations into a coordinate system that is familiar from differential geometry.
+From this viewpoint:
+  • The quadratic closure selects a point of $S^{2}$.
+  • Differentiation pulls back the Maurer–Cartan form of $S^{3}$.
+  • The obstruction $G$ is its horizontal component.
+  • Rigidity is the vanishing of that component.
+  • Burgers transport is the algebraic expression of this vanishing.
+Classical complex analysis appears as the special case in which the horizontal component vanishes identically—equivalently, the Hopf lift is constant along the transport direction.
+The obstruction $G$ measures precisely the failure of this constancy.
+The theory did not require geometry in order to be discovered. The geometry merely confirms that the calculations were inevitable.
+The algebra moved.
+We differentiated it.
+We isolated what could not be removed.
+Only at the end do we recognize that this motion took place on the Hopf bundle.
+Open Questions
+The framework developed in this monograph raises a number of natural questions that remain unresolved.
+We collect here those that appear most accessible or most directly connected to the existing theory.
+1.
+Global existence and singularity formation for rigid structures.
+The $\varepsilon$–family demonstrates that ellipticity restricts the domain: degeneracy occurs along the characteristic parabola $\varepsilon^{2}y^{2}=4(1-\varepsilon x)$.
+Characterize the nature of this boundary singularity.
+Is it always a gradient catastrophe in the Burgers sense?
+More broadly, classify all rigid elliptic structures that remain elliptic on the entire plane.
+The constraint $\operatorname{Im}\lambda>0$ imposes strong restrictions on the conservative Burgers equation $\lambda_{x}+\lambda\lambda_{y}=0$; determine whether global smooth solutions exist beyond the affine class.
+2.
+$L^{p}$ theory for the covariant operator.
+The present work operates in the $C^{1}$ category.
+Develop a weak theory for the covariant Cauchy–Riemann operator $Df=\partial_{\bar{z}}f+\tfrac{1}{2}f\,i_{y}$ in Sobolev spaces $W^{1,p}$.
+In particular, establish that the area integral in the Cauchy–Pompeiu formula defines a bounded operator on $L^{p}$, and determine for which $p$ the Fredholm theory extends to the variable rigid setting.
+3.
+Improved existence and regularity from the self-dilatation structure.
+The Beltrami dictionary (§ 12.2) shows that the rigidization problem, expressed in Beltrami coordinates, is the self-dilatation equation $\mu_{\bar{z}}+\mu\,\mu_{z}=0$: the structure is a dilatational map with respect to its own conformal class.
+This is strictly more constrained than the general Beltrami equation, and the constraint is of Burgers (conservation-law) type rather than of singular-integral type.
+Determine whether this additional structure yields improved existence or regularity results.
+In particular: can the self-dilatation equation be solved by the method of characteristics in regimes where the general Beltrami equation requires the Ahlfors–Bers singular integral machinery?
+Does the conservation-law structure provide a priori estimates (entropy conditions, BV bounds) that have no counterpart in the general quasiconformal setting?
+4.
+Global rigidization.
+Conjecture 1 (Appendix G) asserts that every smooth elliptic structure on a compact simply connected domain can be rigidized by a smooth diffeomorphism, without any smallness condition on the obstruction.
+The continuity method reduces this to a priori estimates: a uniform Jacobian lower bound and a uniform $C^{2,\alpha}$ bound on the rigidizing diffeomorphisms along the interpolation path.
+Three coupled sub-problems are identified in Section G.5: the Jacobian lower bound (via maximum principle for $\log J_{\Phi}$), the Burgers injectivity estimate (no shock formation in the elliptic regime), and the Schauder closure (Agmon–Douglis–Nirenberg for the specific quasilinear $2\times 2$ system).
+None of the three has been rigorously closed.
+This is the most prominent open problem in the monograph.
+5.
+Moduli of rigid structures.
+The $\varepsilon$–family is a one-parameter family.
+Characterize the local moduli space of rigid elliptic structures in a neighborhood of the constant structure.
+The jet analysis (Appendix D) shows that the tangent space is parametrized by holomorphic functions and that higher jets are slaved through forced Cauchy–Riemann equations.
+Determine whether the formal jet hierarchy of Appendix D (§ D.10) converges, and whether the resulting moduli space carries additional structure (for instance, an infinite-dimensional complex manifold structure).
+6.
+Optimal regularity for the Rigidity–Flatness Theorem.
+Theorem 11.7 assumes smooth coefficients.
+Determine the minimal regularity ($C^{k}$ or $W^{k,p}$) under which the conclusion $\tau\equiv\mathrm{const}$ persists.
+The proof relies on the characteristic representation $\tau=\Phi(\zeta)$ with $\zeta=y-x\tau$, which requires at least $C^{2}$; it is not clear whether a distributional formulation can be made to work.
+7.
+Similarity principle with conjugate coupling.
+The rigid similarity principle (§§ 9.3–9.4) is established only in the sharp case $B\equiv 0$.
+Develop a similarity principle for the full rigid Vekua equation $\partial_{\bar{z}}W+AW+BW^{c}=F$, either through generating pairs adapted to the variable-structure setting or through solution-dependent integrating factors.
+This is the natural next step for applications to general first-order elliptic systems.
+8.
+Spectral theory of the rigid variable Laplacian.
+Chapter 8 shows that $4\,\partial_{z}\partial_{\bar{z}}=L_{\alpha,\beta}+{}$first-order drift under rigidity.
+Study the spectral properties of this operator on bounded domains within the elliptic region: Weyl asymptotics, comparison with the standard Laplacian, and the effect of the drift terms controlled by $\alpha_{y}$ and $\beta_{y}$ on eigenvalue distribution.
+9.
+The constructive envelope: sharp or lossy?
+The residual budget $B(\mu)=(1-|\mu|^{2})^{2}/(4\,|1-\mu|)$ of Definition 13.10 determines the region in the difficulty plane $\bigl(\|\mu\|_{C^{0}},\,\|R\|_{C^{0,\alpha}}\bigr)$ where the rigidization scheme converges.
+Is this envelope sharp—meaning that for every structure outside it, no rigidizing diffeomorphism exists locally—or is it merely the boundary of the Newton method’s basin of attraction, with rigidization achievable beyond it by other means?
+More precisely, interpret $B(\mu)$ as a density on the disk measuring the volume of first-order data compatible with constructive solvability.
+Is this density exact, or can it be enlarged by a different iteration scheme or a non-perturbative argument?
+10. A variational principle for near-rigidity. The Poincaré-weighted energy
+$E(\mu)\;=\;\int_{\Omega}\frac{|R(\mu)|^{2}}{(1-|\mu|^{2})^{4}}\,dA$
+is a natural functional that penalizes the residual by the hyperbolic density.
+Its critical points would provide a variational characterization of “nearly rigid” structures.
+However, the $\delta$–family has $R\equiv 0$ (hence $E=0$) while $|\mu|\to 1$ on sets of positive measure, so the energy can vanish even when the pointwise Poincaré density diverges.
+Determine whether $E$ is lower semicontinuous in an appropriate topology, whether its minimizers (subject to boundary conditions) exist, and whether its Euler–Lagrange equation has a clean geometric interpretation in the Poincaré disk.
+11.
+Extension to Clifford structures in higher dimensions.
+The present theory is deliberately two-dimensional.
+The original motivation from Tutschke–Vanegas involves parameter-dependent Clifford algebras.
+Investigate whether the hierarchy transport $\to$ rigidity $\to$ integrability extends to variable Clifford structures in $\mathbb{R}^{n}$, where the Nijenhuis tensor is no longer automatically zero and the interplay between integrability and transport becomes richer.
+Conclusion
+This monograph began from a simple but decisive shift in viewpoint: the imaginary unit should not be treated as a fixed algebraic constant, but as intrinsic geometric data.
+Allowing the generator $i(x,y)$ of a rank–two real algebra bundle to vary forces the algebra itself to become part of the geometry of the plane.
+Once this step is taken, differentiation of the structure is no longer optional, and the derivatives $i_{x}$ and $i_{y}$ emerge as genuine geometric quantities.
+The quadratic relation
+$i^{2}+\beta i+\alpha=0,\qquad 4\alpha-\beta^{2}>0,$
+admits a canonical normalization
+$u:=\frac{2i+\beta}{\sqrt{4\alpha-\beta^{2}}},\qquad u^{2}=-1.$
+Thus every elliptic quadratic structure determines a smooth sphere–valued map $u:\Omega\to S^{2}\subset\mathrm{Im}\,\mathbb{H}$.
+In this form, the moving imaginary unit is literally a rotating point on the sphere.
+Locally, such a map admits a Hopf lift
+$u=q\,i\,q^{-1},\qquad q:\Omega\to S^{3}.$
+Differentiating the generator becomes differentiation of a conjugation. The intrinsic obstruction
 $G:=i_{x}+i\,i_{y}$
-Canonical metric (flatness)
-The canonical Riemannian metric associated with the principal symbol is written in terms of $\lambda=p+iq$ as
-$ds^{2}=\frac{1}{\Im\lambda}\Bigl(dx^{2}+2\,\Re\lambda\,dx\,dy+|\lambda|^{2}\,dy^{2}\Bigr).$
-This metric depends only on $(\alpha,\beta)$ and coincides with the metric induced by the rigid elliptic algebra. Let $K$ denote its Gaussian curvature.
-The principal symbol is called flat if
-$K\equiv 0.$
-Summary of intrinsic invariants
-With this alignment, the two intrinsic invariants of a real first–order elliptic system are:
-  • the transport invariant
-    $T=\lambda_{x}+\lambda\,\lambda_{y},$
-    measuring rigidity;
-  • the Gaussian curvature $K$ of the canonical metric, measuring flatness.
-Both invariants depend only on the principal symbol and are invariant under changes of dependent variables and orientation–preserving coordinate changes.
-All subsequent statements, computations, and proofs in this chapter are understood with $\lambda$ as the spectral slope associated to the elliptic structure $(\alpha,\beta)$, in full consistency with the Vekua formulation.
-10.1 The complex root and intrinsic field
-The equation $Q(1,\tau)=0$ has two complex conjugate roots. We select the one with positive imaginary part:
-$\tau=p+iq=\frac{-\beta+i\sqrt{4\alpha-\beta^{2}}}{2},\qquad q>0.$
-This complex-valued function $\tau(x,y)$ encodes the intrinsic geometry of the principal symbol.
-10.1.1 Two intrinsic invariants
-Two fundamental invariants arise naturally from $\tau$.
-1. The canonical metric.
-Define
-$ds^{2}=\frac{1}{q}\Bigl(dx^{2}+2p\,dx\,dy+(p^{2}+q^{2})\,dy^{2}\Bigr).$
-This Riemannian metric depends only on the principal symbol. Let $K$ denote its Gaussian curvature.
-2. The transport invariant.
-Define the first-order quantity
-$T=\tau_{x}+\tau\tau_{y}.$
-This expression measures how the symbol varies along its own characteristic direction. It depends only on the first derivatives of $\tau$ and hence only on the principal symbol.
-10.1.2 Rigidity and flatness
-We introduce two geometric conditions on the principal symbol:
-  • The system is rigid if $T\equiv 0$.
-  • The system is flat if the canonical metric has $K\equiv 0$.
-Both properties are intrinsic and independent of lower-order terms of the PDE.
-10.2 Statement of the theorem
-Theorem 10.1 (Rigidity–Flatness Theorem).
-Let $\mathcal{L}$ be a smooth real first–order elliptic system on a connected planar domain $\Omega$.
-If its principal symbol is rigid ($T=0$) and flat ($K=0$), then the complex root $\tau$ is constant on $\Omega$.
-Equivalently, the coefficients $\alpha$ and $\beta$ of the structure polynomial are constant, so the principal symbol is locally equivalent to that of a constant–coefficient elliptic system.
-10.2.1 Jet formulation
-Write $\tau=p+iq$ with $q>0$. At a point $z\in\Omega$, the 22–jet of $(p,q)$ is
-$j^{2}_{z}(p,q)=(p,q,p_{x},p_{y},q_{x},q_{y},p_{xx},p_{xy},p_{yy},q_{xx},q_{xy},q_{yy})|_{z}.$
-All second–order invariants of the symbol, including $K$, are rational functions of this jet.
-Rigidity $T=0$ is equivalent to
-$p_{x}=-p\,p_{y}+q\,q_{y},\qquad q_{x}=-p\,q_{y}-q\,p_{y}.$ (10.1)
-Differentiating (10.1) eliminates all xx–derivatives of order $\leq 2$. On the rigid locus, the free jet variables reduce to
-$(p,q,p_{y},q_{y},p_{yy},q_{yy}).$
-10.2.2 Jet-level curvature identity
-Define the curvature numerator
-$N:=-2q^{3}K.$
-Lemma 10.2 (Rigid curvature identity).
-On the rigid locus,
-$N=-q\,\Lambda(p,q)\,q_{yy}+R(p,q;p_{y},q_{y}),$
-where $\Lambda(p,q)>0$ for $q>0$, and
-$R=A(p,q)p_{y}^{2}+2B(p,q)p_{y}q_{y}+C(p,q)q_{y}^{2}$
-is positive definite.
-After clearing denominators this becomes a polynomial identity in finitely many jet variables and can therefore be symbolically verified.
-10.2.3 Consequences of flatness
-If $T=0$ and $K=0$, then $N=0$, so
-$q\,\Lambda(p,q)\,q_{yy}=R(p,q;p_{y},q_{y})\geq 0.$
-10.2.4 Characteristic representation
-Rigidity implies the complex Burgers equation
-$\tau_{x}+\tau\tau_{y}=0.$
-Introduce characteristic coordinates
-$\zeta=y-\tau x.$
-Locally there exists $C^{1}$ $\Phi$ such that
-$\tau(x,y)=\Phi(\zeta).$
-Differentiating,
-$\tau_{y}=\frac{\Phi^{\prime}}{1+x\Phi^{\prime}},\qquad\tau_{yy}=\frac{\Phi^{\prime\prime}}{(1+x\Phi^{\prime})^{3}}.$
-10.2.5 Degree argument
-Restrict to a fixed characteristic slice $\zeta=\mathrm{const}$.
-Then $\Phi^{\prime}$ and $\Phi^{\prime\prime}$ are constants with respect to $x$.
-Multiplying $N=0$ by $|1+x\Phi^{\prime}|^{6}$ produces a polynomial in $x$.
-The term involving $q_{yy}$ has degree at most $3$, while the quadratic form $R$ produces degree $4$.
-Since $R$ is positive definite, the leading coefficient is proportional to $|\Phi^{\prime}|^{6}$ and must vanish.
-Hence $\Phi^{\prime}=0$.
-Therefore $\tau_{y}\equiv 0$, and rigidity gives $\tau_{x}=0$. Thus $\tau$ is constant.
-Proof of Theorem 10.1.
-Combine Lemma 10.2 with the characteristic representation and the degree argument above. ∎
-Conclusion
-This monograph set out from a simple but decisive shift in viewpoint: the imaginary unit should not be treated as a fixed algebraic constant, but as intrinsic geometric data.
-Allowing the generator $i(x,y)$ of a rank–two real algebra bundle to vary forces the algebra itself to become part of the geometry of the plane.
-Once this step is taken, differentiation of the structure is no longer optional, and the derivatives $i_{x}$ and $i_{y}$ emerge as genuine geometric quantities rather than formal artifacts.
-The central object organizing the theory is the intrinsic obstruction
-$G:=i_{x}+i\,i_{y}.$
-This quantity appears prior to any analytic notion of holomorphicity.
-It measures the incompatibility between differentiation in the base variables and multiplication in the fibers, and it governs, simultaneously and unavoidably, all deviations from the classical theory: the inhomogeneity of generalized Cauchy–Riemann systems, the failure of the Leibniz rule, and the appearance of source terms in first–order structure equations.
-These phenomena are not imposed by analytic assumptions; they are forced by the geometry of a moving elliptic algebra.
-A first structural outcome of the theory is that variability of the coefficients $(\alpha,\beta)$ is not arbitrary.
-Eliminating the derivatives of the generator from the structure reduction leads to a universal transport law for the spectral parameter
+is precisely the horizontal component of the pullback of the Maurer–Cartan form of $S^{3}$.
+No new invariant has appeared.
+The obstruction identified through algebraic elimination is the part of the conjugation derivative that cannot be absorbed by a vertical gauge change.
+The variability of the coefficients $(\alpha,\beta)$ is governed by a universal transport law. Eliminating $i_{x}$ and $i_{y}$ from the differentiated quadratic relation forces the spectral parameter
 $\lambda=\frac{-\beta+i\sqrt{4\alpha-\beta^{2}}}{2},\qquad\Im\lambda>0,$
-which takes the form of a (possibly forced) complex inviscid Burgers equation written in the standard complex algebra.
-This transport equation belongs to the structural layer of the theory: it arises solely from allowing the elliptic structure to vary and does not rely on integrability, coordinate trivialization, or function–theoretic input.
-Within this universal transport framework, rigidity occupies a precise and limited role. The condition
-$i_{x}+i\,i_{y}=0$
-does not generate the dynamics; it selects the conservative regime of the already–existing transport law.
-In this regime, the Burgers equation becomes unforced, and several independent obstructions collapse simultaneously: the generalized Cauchy–Riemann operator becomes a derivation, the real first–order system becomes homogeneous, and the generator is transported compatibly with its own multiplication law.
-Rigidity is therefore a compatibility condition, not a simplifying assumption.
-The explicit $\varepsilon$–family constructed in this work demonstrates that rigidity does not force triviality.
-Genuinely nonconstant elliptic structures exist, although ellipticity naturally restricts the domain of definition.
-From the transport viewpoint, these restrictions reflect characteristic phenomena familiar from first–order hyperbolic systems and are not pathologies of the theory.
-On the analytic side, rigidity is exactly the threshold at which a coherent function theory re–emerges.
-In the rigid regime, a Cauchy–Pompeiu formula adapted to variable elliptic structures persists.
-The natural Cauchy 11–form is not closed, and its non–closure produces an unavoidable interior correction term.
-Rather than being an artifact, this correction identifies a canonical covariant Cauchy–Riemann operator whose null space provides the appropriate notion of holomorphicity in the variable setting.
-Although covariantly holomorphic sections are not closed under pointwise multiplication, a scalar gauge transformation restores closure through a natural weighted product.
-Power series representations follow from this framework, but only after freezing the algebra at the expansion point, making explicit where classical arguments genuinely fail in the variable case.
-Finally, the theory interfaces naturally with the classical Vekua program.
-General first–order elliptic systems induce variable elliptic structures, and their reduction to Vekua form reveals the intrinsic obstruction as the true geometric source of inhomogeneity.
-Rigidization, formulated as an ordinary differential equation along characteristic curves, provides a constructive mechanism to pass from non–rigid to rigid structures without solving global Beltrami equations.
-In the rigid regime, sharp similarity principles, initial value problems, and algorithmic solution schemes become available, even in the presence of non–uniform ellipticity.
-The scope of this work has been deliberately local and explicit.
-Rather than appealing to abstract integrability frameworks, the theory has been developed at the level of concrete first–order computations in order to keep the geometric and dynamical mechanisms fully visible.
-From this perspective, variable elliptic structures exhibit a clear hierarchy: transport is fundamental, rigidity is selective, and analytic closure is a distinguished consequence rather than an axiom.
-This hierarchy reframes classical complex analysis as a special case of a broader geometric theory.
+to satisfy a (possibly forced) complex inviscid Burgers equation in the standard complex algebra. Rigidity,
+$G=0,$
+selects the regime in which the horizontal component vanishes.
+In that regime, the Burgers equation becomes conservative, the generalized Cauchy–Riemann operator becomes a derivation, and the generator is transported compatibly with its own multiplication.
+Rigidity is the condition that the motion of the imaginary unit is purely vertical in the Hopf bundle.
+On the analytic side, rigidity marks the threshold at which a coherent function theory reappears.
+A Cauchy–Pompeiu formula persists, a covariant Cauchy–Riemann operator emerges canonically, weighted products restore multiplicative closure, and power series representations follow after freezing the algebra at a point.
+The variable theory does not abolish complex analysis; it reveals which of its properties depend on rigidity and which depend on the hidden constancy of the standard structure.
+The translation to the Beltrami variable
+$\mu=\frac{\lambda-i}{\lambda+i}\in\mathbb{D}$
+recasts the entire theory in the language of the Poincaré disk. The rigidity condition $G=0$ becomes the self–dilatation equation
+$\mu_{\bar{z}}+\mu\,\mu_{z}=0,$
+which states that $\mu$ is a dilatational map with respect to its own structure.
+The Poincaré metric, which appeared uninvited in the convergence theory as the normalization $b_{\min}^{-2}$, turns out to be the intrinsic geometry of the parameter space $\mathbb{D}$.
+From this vantage, the central discovery of the monograph becomes visible.
+The difficulty of a variable elliptic structure lives in a plane, not on a line.
+Two invariants govern it: the Beltrami modulus $\|\mu\|_{C^{0}}$, a zeroth–order quantity measuring where the structure sits in the Poincaré disk, and the Poincaré residual $R(\mu)=\mu_{\bar{z}}+\mu\,\mu_{z}$, a first–order quantity measuring how the structure moves through it.
+These are independently prescribable.
+All four quadrants of the difficulty plane $\bigl(\|\mu\|_{C^{0}},\,\|R\|_{C^{0,\alpha}}\bigr)$ are populated by smooth elliptic structures.
+The classical Beltrami literature—the Measurable Riemann Mapping Theorem, the Ahlfors–Bers theorem, the theory of quasiconformal mappings—projects this plane onto the horizontal axis.
+Only $\|\mu\|_{C^{0}}$ survives.
+The vertical axis, the transport field, is invisible.
+A rigid structure at $|\mu|=0.99$ supports a complete function theory.
+A non–rigid structure at $|\mu|=0.01$ does not, without first being rigidized.
+The zeroth–order invariant cannot see the first–order distinction.
+The explicit families constructed in this work make this concrete.
+The $\varepsilon$–family demonstrates that rigidity does not force constancy: nontrivial elliptic structures correspond to genuine motions on $S^{2}$, though ellipticity naturally restricts the domain through characteristic phenomena.
+The $\delta$–family demonstrates that zeroth–order difficulty and first–order difficulty are decoupled: $\|\mu\|_{C^{0}}\to 1$ while $R\equiv 0$ identically.
+A solver that measures difficulty by $|\mu|$ alone sees a problem approaching maximal degeneracy.
+The rigidizer sees a problem already solved.
+The p(x)p(x)–analytic theory, developed by the Bers school and its successors, populates the real diameter of the Poincaré disk: $\mu=(\sqrt{p}-1)/(\sqrt{p}+1)\in(-1,1)$.
+Along this diameter, $\beta=0$ identically; the variable exponent modulates conformal scaling but never opens the transport channel.
+The relationship is exact:
+$\begin{array}[]{ccccc}\text{real analysis}&\xrightarrow{\quad\text{analytic continuation}\quad}&\text{complex analysis}\\[6.0pt] p(x)\text{-analytics}&\xrightarrow{\quad\beta\neq 0\quad}&\text{VES}\end{array}$
+The generating pairs, successor structures, embedding theorems, and regularity results of the established theory form the boundary data from which the full variable theory departs.
+The observation is not that these results are insufficient.
+It is that they explore the real diameter of a disk whose off–axis structure was never investigated.
+The residual budget $B(\mu)$ at each point of the disk measures how far the constructive theory reaches: the maximum Poincaré residual for which the rigidization scheme converges.
+The budget is not rotationally symmetric.
+It reflects an asymmetry intrinsic to the Möbius transformation relating the spectral parameter to the Beltrami variable.
+The constructive envelope in the difficulty plane is therefore a computable region, not a universal constant.
+The diagnostic algorithm that emerges from this analysis classifies first–order elliptic systems into three branches: structures within the constructive envelope, accessible to the full VES toolkit; p(x)p(x)–analytic structures on the real diameter, governed by the Bers theory; and structures with genuine transport obstruction beyond the envelope, requiring classical quasiconformal methods.
+Without the independence of position and velocity in the Poincaré disk, this triage is impossible.
+Throughout, the development has remained local and first–order.
+The geometry of the Hopf bundle was not required to derive the theory; it confirms, at the end, what the algebra had already forced.
+The Poincaré disk was not assumed as a framework; it emerged from the convergence theory as the natural geometry of the parameter space.
+The independence of the two axes of difficulty was not imposed by construction; it follows from the elementary fact that a function and its derivative are independent objects.
+Transport is fundamental. Rigidity is selective. Analytic closure is consequential.
 A fixed complex structure gives analysis. A variable one gives transport. Rigidity is where the two agree to dance.
-Here, the algebra begins to move.
+In Hopf coordinates we see what was moving: a rotating imaginary unit on $S^{2}$.
+In Beltrami coordinates we see what was hidden: a two–dimensional difficulty space projected onto a line.
+Here, the algebra moves.
 Appendix A Algebraic Universality of the Intrinsic Obstruction and the Leibniz Defect
@@
 The theory developed here is therefore not a theory of geometric integrability of complex structures in the Newlander–Nirenberg sense, but a theory of transport of algebraic structures and of the analytic calculus they induce in a fixed background frame.
-Appendix F Symbolic Verification of the Rigid Second–Order Expansion
+Appendix F Quaternionic Reformulation of Rigidity via the Hopf Fibration
+F.1 Motivation
+A variable elliptic structure on a domain $\Omega\subset\mathbb{R}^{2}$ is encoded by a generator $i=i(x,y)$ satisfying the structure reduction
+$i^{2}+\beta(x,y)\,i+\alpha(x,y)=0,$ (F.1)
+with ellipticity discriminant
+$\Delta(x,y):=4\alpha(x,y)-\beta(x,y)^{2}>0.$ (F.2)
+The intrinsic obstruction is
+$G:=i_{x}+i\,i_{y},$ (F.3)
+and the structure is called rigid precisely when $G\equiv 0$.
+The purpose of this appendix is to recast rigidity in a purely quaternionic language via the Hopf fibration, clarifying the geometric content of $G$.
+F.1.1 Normalization to a unit imaginary quaternion
+Completing the square in (F.1) yields
+$(2i+\beta)^{2}=\beta^{2}-4\alpha=-\Delta.$
+Define the normalized generator
+$u(x,y):=\frac{2i(x,y)+\beta(x,y)}{\sqrt{\Delta(x,y)}}.$ (F.4)
+Then
+$u^{2}=-1,$ (F.5)
+so $u$ takes values in the unit imaginary sphere
+$S^{2}=\{\,q\in\operatorname{Im}\mathbb{H}:\ |q|=1\,\}\subset\mathbb{H}.$
+Thus, an elliptic quadratic structure canonically determines a smooth map $u:\Omega\to S^{2}$.
+F.2 Hopf lift and gauge freedom
+Fix a constant imaginary unit $i_{0}\in\operatorname{Im}\mathbb{H}$ with $i_{0}^{2}=-1$.
+Since every $u(x,y)\in S^{2}$ is conjugate to $i_{0}$, locally there exists a Hopf lift $q:\Omega\to S^{3}$ (unit quaternions) such that
+$u=q\,i_{0}\,q^{-1}.$ (F.6)
+The lift is unique up to the $S^{1}$-gauge transformation
+$q\ \mapsto\ q\,e^{i_{0}\theta(x,y)}.$ (F.7)
+F.3 Maurer–Cartan form and differentiation
+Define the left-trivialized Maurer–Cartan components
+$\Omega_{x}:=q_{x}q^{-1},\qquad\Omega_{y}:=q_{y}q^{-1},$ (F.8)
+which lie in $\mathfrak{su}(2)\cong\operatorname{Im}\mathbb{H}$. Differentiating (F.6) and using $(q^{-1})_{x}=-q^{-1}q_{x}q^{-1}$ gives
+$u_{x}=[\Omega_{x},u],\qquad u_{y}=[\Omega_{y},u],$ (F.9)
+where [A,B]=ABBA[A,B]=AB-BA denotes the commutator in $\mathbb{H}$.
+F.3.1 Vertical–horizontal decomposition
+At each point, $\mathfrak{su}(2)\cong\operatorname{Im}\mathbb{H}$ splits as
+$\operatorname{Im}\mathbb{H}=\mathrm{span}\{u\}\ \oplus\ u^{\perp}.$
+Accordingly decompose
+$\Omega_{x}=\Omega_{x}^{\mathrm{ver}}+\Omega_{x}^{\mathrm{hor}},\qquad\Omega_{y}=\Omega_{y}^{\mathrm{ver}}+\Omega_{y}^{\mathrm{hor}},$ (F.10)
+where $\Omega^{\mathrm{ver}}\parallel u$ and $\Omega^{\mathrm{hor}}\perp u$. Since $\Omega^{\mathrm{ver}}$ commutes with $u$, it follows from (F.9) that
+$u_{x}=[\Omega_{x}^{\mathrm{hor}},u],\qquad u_{y}=[\Omega_{y}^{\mathrm{hor}},u].$ (F.11)
+Under the gauge change (F.7), the Maurer–Cartan form changes by adding a vertical term, so the horizontal components and the commutators in (F.11) capture the gauge-invariant variation of $u$.
+F.4 Quaternionic form of the obstruction
+The obstruction $G=i_{x}+i\,i_{y}$ can be rewritten in terms of $u$. Up to the scalar normalizations in (F.4), the intrinsic combination governing deviations from the constant theory corresponds to
+$\mathcal{G}:=u_{x}+u\,u_{y}.$ (F.12)
+Using (F.11), we may express $\mathcal{G}$ in Hopf variables as
+$\mathcal{G}=[\Omega_{x}^{\mathrm{hor}},u]\;+\;u\,[\Omega_{y}^{\mathrm{hor}},u].$ (F.13)
+This expression is gauge-invariant (it is insensitive to vertical modifications of the lift).
+F.5 Quaternionic rigidity theorem
+Theorem F.1 (Quaternionic rigidity).
+Let $u:\Omega\to S^{2}\subset\operatorname{Im}\mathbb{H}$ be the normalized generator (F.4) associated with an elliptic quadratic structure (F.1).
+Let $q$ be a local Hopf lift (F.6) with Maurer–Cartan components (F.8) and decomposition (F.10).
+Then the following are equivalent:
+  1. (i)
+[…]
+    $u_{x}+u\,u_{y}=0\quad\text{on }\Omega.$ (F.14)
+  2. (ii)
+    (Horizontal Maurer–Cartan cancellation)
+[…]
+  3. (iii)
+    (Purely vertical transport along the characteristic direction) the pulled-back Maurer–Cartan form has no effective horizontal component along the direction $\partial_{x}+u\,\partial_{y}$, i.e.
+$u$ undergoes only fiber (Hopf) gauge rotation along that direction.
+Moreover, (F.14) is the Hopf/quaternionic avatar of the original rigidity condition $G=i_{x}+i\,i_{y}=0$.
+Proof.
+By (F.11), we have $u_{x}=[\Omega_{x}^{\mathrm{hor}},u]$ and $u_{y}=[\Omega_{y}^{\mathrm{hor}},u]$.
+Substituting these into (F.12) yields (F.13), and hence $\mathcal{G}\equiv 0$ is equivalent to (F.15).
+The gauge statement follows from the fact that changing the lift by (F.7) alters only the vertical part of the Maurer–Cartan form, leaving the horizontal commutator expressions unchanged.
+∎
+F.6 Interpretation
+Equation (F.14) says that the $S^{2}$-valued imaginary direction field $u$ is transported compatibly with its own multiplication rule: along the characteristic direction $\partial_{x}+u\,\partial_{y}$, the field exhibits no genuine horizontal twisting in $S^{2}$, only Hopf fiber rotation.
+In this sense rigidity is a horizontality cancellation condition in the Hopf fibration rather than an analytic assumption.
+Appendix G Rigidization as a Computational Normal Form
+Purpose of this Appendix
+This appendix clarifies the intended scope of rigidization within the theory of variable elliptic structures.
+Rigidization is not proposed as a replacement for global uniformization (Beltrami theory), nor as a universal flattening mechanism.
+Instead, it is designed as a computationally optimized normal form that exploits intrinsic transport structure when present, and that can be used either as a standalone reduction or as a preprocessing step for uniformization-based solvers.
+The guiding philosophy is:
+Rigidization removes exactly the analytic obstructions that prevent the use of the constant-coefficient toolbox, while preserving geometric variability that is irrelevant for computation but expensive to eliminate.
+G.1 Rigidization versus Uniformization
+Let $\lambda(x,y)$ denote the spectral slope of a variable elliptic structure, with intrinsic obstruction
+$T:=\lambda_{x}+\lambda\,\lambda_{y}.$
+Uniformization (Beltrami)
+Uniformization seeks a diffeomorphism $\Phi$ such that the pullback structure is constant.
+This solves the maximal geometric problem but requires solving a globally coupled nonlinear elliptic PDE.
+From a computational standpoint, uniformization is:
+  • global and nonlocal,
+  • expensive to compute,
+  • insensitive to existing transport structure,
+  • often unnecessary when only analytic compatibility is required.
+Rigidization
+Rigidization seeks a diffeomorphism $\Phi$ such that the pullback structure satisfies
+$\widetilde{T}=0,\qquad\text{i.e.}\quad\widetilde{\lambda}_{X}+\widetilde{\lambda}\,\widetilde{\lambda}_{Y}=0,$
+but does not require $\widetilde{\lambda}$ to be constant.
+This is a strictly weaker requirement than uniformization. Nevertheless, once rigidity is achieved:
+  • the Leibniz defect vanishes,
+  • the generalized Cauchy–Riemann system becomes homogeneous,
+  • Cauchy–Pompeiu and similarity principles apply,
+  • the full constant-structure analytic toolkit becomes available.
+Thus rigidization reaches the minimal normal form required for analysis, without over-solving the geometric problem.
+G.2 Why “Weaker” Means “Easier”
+The rigidization condition is first-order in the structure and second-order in the coordinates, and its linearization is an elliptic system whose principal symbol is explicitly computable.
+As shown in Chapter 10, this enables:
+  • local Newton convergence under a quantitative small-obstruction condition (Theorem 10.15),
+  • inner solvers based on transport–wave splitting (shifted ADI with Wachspress parameters),
+  • direct control of analytic compatibility without global flattening.
+In contrast, full uniformization removes curvature and transport simultaneously, even when curvature plays no analytic role for the PDE being solved.
+From a solver perspective, rigidization is therefore cheaper, more local, and better aligned with problems that already possess a preferred transport geometry.
+G.3 A Solver-Oriented Triage
+Rigidization is intended to be used selectively. Three intrinsic indicators guide solver choice:
+  1. 1.
+    Obstruction magnitude:
+    $\rho_{T}:=\frac{\|T\|_{C^{0,\alpha}}}{b_{\min}^{2}},\qquad b_{\min}:=\inf\operatorname{Im}\lambda.$
+  2. 2.
+    Metric curvature magnitude: a scale-normalized curvature indicator $\rho_{K}$ for the canonical metric.
+  3. 3.
+    Transport dominance: a problem-dependent measure of whether a preferred first-order flow direction is present.
+These lead to the following computational policy:
+  • Small obstruction ($\rho_{T}\ll 1$): Rigidize immediately using the Newton scheme of Chapter 10. This is the regime covered by Theorem 10.15 with full convergence guarantees.
+  • Moderate obstruction, transport-dominated PDE: Attempt rigidization via a continuity strategy (interpolating from a rigid reference structure $\lambda_{0}=i$ via $\lambda^{t}=(1-t)\lambda_{0}+t\lambda$, taking incremental steps in $t$ with each step small enough for the Newton iteration of Theorem 10.15 to converge).
+If the obstruction is substantially reduced, proceed in rigid coordinates; otherwise fall back to uniformization.
+  • Large obstruction, negligible curvature: Use Beltrami uniformization directly.
+  • Diffusion-dominated problems: Use uniformization; rigidization offers no advantage when no transport structure is present.
+This policy is not ad hoc: it reflects the analytic role played by the obstruction in the failure of derivation properties, even though the specific thresholds and decision boundaries are solver-dependent.
+G.4 Rigidization as a Preconditioner
+Even when rigidization does not succeed globally, it can still be valuable.
+Chapter 10 shows that naive transport–wave splitting does not converge as a standalone iteration, but is effective as an inner solver for the linearized elliptic problem.
+This mirrors standard numerical practice:
+  • rigidization removes transport incompatibility first,
+  • uniformization then operates on a better-conditioned structure,
+  • overall solver stability and convergence improve.
+In this sense, rigidization acts as a geometric preconditioner for uniformization-based methods.
+More concretely, the incremental continuity strategy described above provides a practical algorithm even without a closed-form global existence theorem: one advances in $t$ until the Newton step fails to converge, records the partially rigidized structure, and hands the residual obstruction to a uniformization solver operating on a problem whose effective $\rho_{T}$ has been reduced.
+G.5 The Global Rigidization Problem
+It is natural to ask whether rigidization can always be achieved on compact domains, without any smallness condition on the obstruction $T$.
+Conjecture 1 (Global rigidization).
+Let $\Omega\subset\mathbb{R}^{2}$ be a compact simply connected domain with $C^{2,\alpha}$ boundary.
+Let $\lambda\in C^{2,\alpha}(\bar{\Omega})$ with $\operatorname{Im}\lambda\geq b_{\min}>0$ on $\bar{\Omega}$.
+Then there exists a $C^{2,\alpha}$ diffeomorphism $\Phi:\bar{\Omega}\to\Phi(\bar{\Omega})$ such that the pullback structure is rigid.
+A natural proof strategy is the continuity method: interpolating from a rigid reference $\lambda_{0}=i$ to the target $\lambda$, the success set is nonempty (at $t=0$) and open (by Theorem 10.15).
+Closedness would follow from uniform a priori estimates on the rigidizing diffeomorphisms—specifically, a uniform $C^{2,\alpha}$ bound and a uniform Jacobian lower bound along the family.
+These a priori estimates constitute the core difficulty. Three coupled sub-problems are involved:
+  1. 1.
+    Jacobian lower bound.
+Show that $\inf_{\bar{\Omega}}J_{\Phi}\geq c>0$ for any rigidizing diffeomorphism, with $c$ depending only on $b_{\min}$, $\|\lambda\|_{C^{2,\alpha}}$, and $\Omega$.
+A promising route is to derive a second-order elliptic equation (or inequality) for $\log J_{\Phi}$ from the ratio condition and apply the maximum principle, using the boundary normalization $\Phi|_{\partial\Omega}=\mathrm{id}$ (so $J=1$ on $\partial\Omega$).
+  2. 2.
+    Burgers constraint on compact domains.
+The pullback $\widetilde{\lambda}$ solves conservative Burgers, whose solutions are constant along characteristics.
+On compact domains with $\operatorname{Im}\widetilde{\lambda}>0$, the characteristic map from boundary data to the interior is expected to be injective (no shock formation in the elliptic regime).
+A correct proof requires a quantitative injectivity estimate on the characteristic map—specifically, a determinant lower bound on the map $(s,X)\mapsto(X,Y)$ involving $\partial\widetilde{\lambda}/\partial s$ and domain geometry, not merely the condition $\operatorname{Im}\widetilde{\lambda}>0$.
+  3. 3.
+    $C^{2,\alpha}$ a priori bound.
+With the Jacobian controlled, the rigidization system is uniformly elliptic.
+Schauder estimates (Agmon–Douglis–Nirenberg) should yield global $C^{2,\alpha}$ bounds, but the application to this specific quasilinear $2\times 2$ system requires verifying the structural conditions under which elliptic regularity and boundary estimates hold.
+These sub-problems are mutually dependent: the Burgers constraint feeds into the Jacobian bound, which in turn feeds into the Schauder estimates.
+As of this writing, none of the three has been rigorously closed, and the conjecture remains open.
+We record it here both as an invitation to further work and as context for the solver-oriented approach of this appendix: the computational triage and preconditioner strategies above are designed to be effective regardless of whether Conjecture 1 is eventually proved.
+G.6 What Is Not Claimed
+To be explicit about the boundaries of the current theory, this monograph does not claim:
+  • global rigidization for arbitrary elliptic structures on compact domains (this remains an open problem; see Conjecture 1),
+  • a priori Jacobian lower bounds independent of the obstruction magnitude (the estimates in Chapter 10 depend on the smallness condition $\|T\|<C\,b_{\min}^{2}$),
+  • replacement of Beltrami theory as a universal tool.
+These are not limitations of design intent but reflections of the current state of the analysis.
+The global theory may well be within reach—the structural ingredients (ellipticity of the linearized problem, Burgers characteristic geometry, similarity-principle arguments) are all present—but assembling them into a complete a priori package requires new ideas, particularly for the control of the Jacobian.
+G.7 Summary
+Rigidization should be viewed as:
+  • a maximal computable normal form,
+  • a solver-side exploitation of transport geometry,
+  • a selective alternative or complement to uniformization,
+  • a mechanism that preserves just enough geometry to be efficient.
+When it applies, rigidization delivers the full analytic power of the constant-coefficient theory at a fraction of the computational cost.
+When its applicability has not been established, it still provides valuable structural diagnostics and preconditioning for heavier methods.
+In this sense, rigidization is not a competitor to uniformization but a practical intermediary between raw variable geometry and full geometric flattening—one whose global reach remains a compelling open question in its own right.
+Appendix H Symbolic Verification of the Rigid Second–Order Expansion
 This appendix contains a complete symbolic verification of Theorem 8.1 using the computer algebra system SymPy.
@@
 # --- end script ---
+Appendix I Symbolic Verification for the Rigidity-Flatness Theorem Condition
+Listing 2: SymPy verification of Theorem 8.1
+# --- begin SymPy verification script ---
+#!/usr/bin/env python3
+rigid_flat_vacuum_check.py
+Reproducible SymPy computations for the "full vacuum" question in the
+planar elliptic structure framework:
+- Define tau = p + i q (q>0), with canonical metric
+ds^2 = (1/q) (dx^2 + 2 p dx dy + (p^2 + q^2) dy^2).
+- Compute the Gaussian curvature K of this metric for general p(x,y), q(x,y).
+- Impose rigidity (torsion-free / conservative transport):
+tau_x + tau tau_y = 0
+i.e.
+p_x + p p_y - q q_y = 0
+q_x + p q_y + q p_y = 0
+and substitute the consequent relations for second derivatives
+[…]
+- Sanity-check with an explicit nontrivial rigid family:
+tau = (eps*y + i) / (1 + eps*x)
+Outputs:
+- General K(p,q) as a symbolic expression
+- Rigid-reduced curvature K_rigid (expressed in y-derivatives only)
+- Curvature for the explicit rigid family
+Run:
+python3 rigid_flat_vacuum_check.py
+Dependencies:
+sympy
+import sympy as sp
+def main():
+# Coordinates
+x, y = sp.symbols(’x y’, real=True)
+# Real/imag parts of tau = p + i q (q>0)
+p = sp.Function(’p’)(x, y)
+q = sp.Function(’q’)(x, y)
+# Canonical metric in (x,y):
+# ds^2 = (1/q) (dx^2 + 2 p dx dy + (p^2 + q^2) dy^2)
+g = sp.Matrix([
+[1/q, p/q],
+[p/q, (p**2 + q**2)/q]
+g_inv = sp.simplify(g.inv())
+coords = [x, y]
+def d(expr, var):
+return sp.diff(expr, var)
+[…]
+Gamma = [[[0, 0], [0, 0]] for _ in range(2)]
+for k in range(2):
+for i in range(2):
+for j in range(2):
+s = 0
+for l in range(2):
+s += g_inv[k, l] * (d(g[j, l], coords[i]) +
+d(g[i, l], coords[j]) -
+d(g[i, j], coords[l]))
+Gamma[k][i][j] = sp.simplify(sp.Rational(1, 2) * s)
+[…]
+Riem = [[[[0 for _ in range(2)] for _ in range(2)] for _ in range(2)] for _ in range(2)]
+for l in range(2):
+for i in range(2):
+for j in range(2):
+for k in range(2):
+term = d(Gamma[l][i][k], coords[j]) - d(Gamma[l][i][j], coords[k])
+for m in range(2):
+term += Gamma[l][j][m] * Gamma[m][i][k] - Gamma[l][k][m] * Gamma[m][i][j]
+Riem[l][i][j][k] = sp.simplify(term)
+# Ricci tensor and scalar curvature
+Ricci = sp.Matrix([[0, 0], [0, 0]])
+for i in range(2):
+for k in range(2):
+Ricci[i, k] = sp.simplify(sum(Riem[l][i][l][k] for l in range(2)))
+R_scalar = sp.simplify(sum(g_inv[i, k] * Ricci[i, k] for i in range(2) for k in range(2)))
+K = sp.simplify(R_scalar / 2) # Gaussian curvature in 2D
+[…]
+print(K)
+# --- Rigidity (torsion-free) equations: tau_x + tau tau_y = 0 ---
+# Split into real equations for p,q:
+# p_x + p p_y - q q_y = 0
+# q_x + p q_y + q p_y = 0
+px = sp.Derivative(p, x)
+py = sp.Derivative(p, y)
+qx = sp.Derivative(q, x)
+qy = sp.Derivative(q, y)
+# Solve for x-derivatives
+p_x_expr = -p * py + q * qy
+q_x_expr = -p * qy - q * py
+# Differentiate in y to get mixed derivatives
+p_xy_expr = sp.diff(p_x_expr, y)
+q_xy_expr = sp.diff(q_x_expr, y)
+# Differentiate in x and eliminate x-derivatives via rigidity
+p_xx_expr = sp.diff(p_x_expr, x).xreplace({px: p_x_expr, qx: q_x_expr})
+q_xx_expr = sp.diff(q_x_expr, x).xreplace({px: p_x_expr, qx: q_x_expr})
+# Substitute mixed derivatives to eliminate remaining x-dependence
+p_xx_expr = sp.simplify(p_xx_expr.subs({
+sp.Derivative(p, x, y): p_xy_expr,
+sp.Derivative(q, x, y): q_xy_expr
+q_xx_expr = sp.simplify(q_xx_expr.subs({
+sp.Derivative(p, x, y): p_xy_expr,
+sp.Derivative(q, x, y): q_xy_expr
+subs_rigid = {
+sp.Derivative(p, x): p_x_expr,
+sp.Derivative(q, x): q_x_expr,
+sp.Derivative(p, x, 2): p_xx_expr,
+sp.Derivative(q, x, 2): q_xx_expr,
+sp.Derivative(p, x, y): p_xy_expr,
+sp.Derivative(q, x, y): q_xy_expr,
+K_rigid = sp.simplify(sp.together(K.subs(subs_rigid)))
+[…]
+print(sp.factor(K_rigid))
+# --- Sanity check: explicit rigid family tau = (eps*y + i)/(1 + eps*x) ---
+eps = sp.symbols(’eps’, real=True)
+p_ex = eps * y / (1 + eps * x)
+q_ex = 1 / (1 + eps * x)
+K_family = sp.simplify(K.subs({p: p_ex, q: q_ex}))
+[…]
+print(sp.factor(K_family))
+# Optional: show that K_family == 0 implies eps == 0 (generic statement)
+# This is not a full proof, but indicates nontrivial family is not flat.
+[…]
+if __name__ == "__main__":
+main()
 Acknowledgements

### v5

@@
 Daniel Alayón-Solarz Email: danieldaniel@gmail.com
-(February 2026)
+(March 2026)
 A fixed complex structure gives analysis.
@@
 Rigidity is where the two agree to dance.
-(“Shut up and calculate!” — David Mermin)
+(“Shut up and calculate!” — N. David Mermin)
+Preface to v5
+This version corrects an error in the Leibniz defect formula of Chapter 1 (formerly §1.7): the operator $\partial_{\bar{z}}$ is a derivation for all variable elliptic structures, not only rigid ones.
+The Rigidity Theorem has been updated accordingly.
+Clarification between $p(x)$-analytics and $p$-analytics in the sense of Polozhii has been added.
+Chapter 3 (Algebra–Spectral Intertwining) is new.
+The Poincaré residual has been renamed to Beltrami residual.
+March 2026
 Introduction
@@
 It measures the incompatibility between differentiation in the base variables and multiplication in the fibers.
-As a consequence, it simultaneously governs the inhomogeneity of generalized Cauchy–Riemann systems and the failure of the Leibniz rule for the natural first–order operators associated with the variable structure.
-These phenomena are not imposed by analytic assumptions; they are forced by the geometry of the moving algebra itself.
+As a consequence, it governs the inhomogeneity of generalized Cauchy–Riemann systems for the natural first–order operators associated with the variable structure.
+This phenomenon is not imposed by analytic assumptions; it is forced by the geometry of the moving algebra itself.
 A first major outcome of the theory is that the variability of the structure coefficients $(\alpha,\beta)$ is governed by a transport law that is intrinsic and unavoidable.
@@
 selects exactly the conservative case of the transport law.
-In this regime, the Burgers equation becomes unforced, and several independent obstructions collapse simultaneously: the generalized Cauchy–Riemann operator becomes a derivation, the real Cauchy–Riemann system becomes homogeneous, and the generator $i(x,y)$ is transported compatibly with its own multiplication law.
+In this regime, the Burgers equation becomes unforced, and two independent obstructions collapse simultaneously: the real Cauchy–Riemann system becomes homogeneous, and the generator $i(x,y)$ is transported compatibly with its own multiplication law.
 Rigidity is therefore not the source of the coefficient dynamics, but a compatibility condition that restores integrability within an already dynamical setting.
@@
 These examples show that rigidity does not force triviality: genuinely nonconstant structures exist, although ellipticity naturally restricts the domain of definition.
-From the transport viewpoint, these restrictions reflect characteristic phenomena familiar from first–order hyperbolic systems.
+From the transport viewpoint, these domain restrictions arise because the ellipticity condition $\Im\lambda>0$ imposes global constraints on the admissible coefficient data.
 On the analytic side, rigidity is exactly the condition under which a coherent function theory can be recovered.
@@
 Thus the system (1.11) depends only on $\alpha$, $\beta$ and their first derivatives.
-1.7 Product Rule and Its Defect
-Let $f=u+vi$ and $g=p+qi$ be $C^{1}$ sections. Define the Leibniz defect
-$\mathcal{N}(f,g):=\partial_{\bar{z}}(fg)-(\partial_{\bar{z}}f)g-f(\partial_{\bar{z}}g).$
-A direct computation shows
-$\mathcal{N}(f,g)=\frac{1}{2}\,vq\,(i_{x}+i\,i_{y}).$ (1.13)
-Remark 1.4.
-Formula (1.13) is purely algebraic and holds without further assumptions. It shows explicitly how the variability of the generator $i(x,y)$ obstructs the derivation property of $\partial_{\bar{z}}$.
 Chapter 2 Differentiating the Structure: Universal Burgers Transport
@@
 It governs the evolution of the structure coefficients for all variable elliptic structures.
-2.6 Characteristic Interpretation
-Equation (2.6) may be written in transport form as
-$(\partial_{x}+\lambda\,\partial_{y})\lambda=G_{0}+\lambda\,G_{1}.$
-Along characteristic curves satisfying $dy/dx=\lambda$, the spectral parameter evolves according to the linear ordinary differential equation
-$\frac{d\lambda}{dx}=G_{0}(x,y(x))+\lambda(x)\,G_{1}(x,y(x)).$
-Thus, the evolution of the coefficients of the structure reduction is always governed by transport along characteristic curves. The obstruction $\mathcal{G}$ acts as a source and growth term.
+2.6 Ellipticity of the Transport System
+The transport equation (2.6) has the form
+$(\partial_{x}+\lambda\,\partial_{y})\lambda=G_{0}+\lambda\,G_{1},$
+where $\lambda=a+ib$ with $b>0$. It is natural to ask whether this equation admits a characteristic interpretation in $\mathbb{R}^{2}$.
+Writing (2.6) as a real first–order system for $(a,b)$, the principal part in the yy–direction is governed by the coefficient matrix
+$A=\begin{pmatrix}a&-b\\ b&a\end{pmatrix}$,
+whose eigenvalues are $a\pm ib=\lambda$ and $\bar{\lambda}$.
+Since $b>0$, the eigenvalues are strictly complex.
+Consequently the real system is elliptic, not hyperbolic, and possesses no real characteristic curves.
+This is not a defect of the formulation but a structural necessity: the ellipticity condition $4\alpha-\beta^{2}>0$ that defines the variable elliptic structure propagates directly into the ellipticity of the transport system governing its coefficients.
+The structure is elliptic at every level of the theory.
+Remark 2.5 (Complex vector field).
+The operator $\partial_{x}+\lambda\,\partial_{y}$ is a complex vector field on $\mathbb{R}^{2}$.
+Its ellipticity is equivalent to the statement that it has no real null curves, i.e.
+the equation $dy/dx=\lambda$ admits no real–valued solution.
+In the language of several complex variables, $\partial_{x}+\lambda\,\partial_{y}$ fails to be tangent to any real curve precisely because $\operatorname{Im}\lambda\neq 0$.
 2.7 Summary
@@
 In the next chapter we identify the special regime in which the obstruction vanishes and show that, in this case, the transport becomes conservative and a coherent analytic function theory emerges.
-Chapter 3 Rigidity as the Conservative Regime of Burgers Transport
+Chapter 3 The Algebra–Spectral Intertwining
+The universal transport law of Chapter 2 governs the evolution of the spectral parameter $\lambda$.
+The Cauchy–Riemann operator $\partial_{\bar{z}}=\frac{1}{2}(\partial_{x}+i\,\partial_{y})$ of Chapter 1 governs the function theory at the algebra level.
+This chapter establishes the exact relationship between the two: the spectral map intertwines the algebra operator with the spectral transport operator, universally, for all variable elliptic structures.
+3.1 The spectral map
+Recall from Chapter 1 that the fiber algebra $A_{z}=\mathbb{R}[X]/(X^{2}+\beta X+\alpha)$ is generated over $\mathbb{R}$ by the moving generator $i=i(x,y)$ satisfying $i^{2}+\beta\,i+\alpha=0$.
+Under ellipticity ($4\alpha-\beta^{2}>0$), $A_{z}$ is isomorphic to $\mathbb{C}$ as a real algebra.
+The spectral parameter
+$\lambda=\frac{-\beta+i\sqrt{4\alpha-\beta^{2}}}{2}$
+is the image of the generator $i$ under this isomorphism.
+Definition 3.1 (Spectral map).
+The spectral map sends an AA–valued section $W=U+Vi$ ($U,V:\Omega\to\mathbb{R}$) to its complex–valued spectral image
+$W_{\lambda}:=U+V\lambda\in\mathbb{C}.$
+The spectral map is a pointwise real–algebra isomorphism. Two properties are immediate.
+Proposition 3.2 (Injectivity).
+Under ellipticity, $W_{\lambda}=0$ implies $W=0$.
+Proof.
+$U+V\lambda=0$ with $\operatorname{Im}\lambda>0$ gives $U=0$ and $V=0$. ∎
+Proposition 3.3 (Ring homomorphism).
+For any AA–valued sections $W=U+Vi$ and $Z=P+Qi$,
+$(WZ)_{\lambda}=W_{\lambda}\cdot Z_{\lambda}.$
+Proof.
+The spectral map $i\mapsto\lambda$ is a real–algebra homomorphism on each fiber. ∎
+3.2 The Leibniz rule for $\partial_{\bar{z}}$
+Before establishing the intertwining, we show that the algebra–level operator is always a derivation.
+Proposition 3.4 (Universal Leibniz rule).
+The operator $\partial_{\bar{z}}=\frac{1}{2}(\partial_{x}+i\,\partial_{y})$ satisfies the Leibniz rule on $C^{1}(\Omega;A)$:
+$\partial_{\bar{z}}(fg)=(\partial_{\bar{z}}f)\,g+f\,(\partial_{\bar{z}}g)$
+for all AA–valued $C^{1}$ sections $f,g$. This holds for all variable elliptic structures.
+Proof.
+The structure relation $i^{2}+\beta\,i+\alpha=0$ is an identity of $C^{1}$ functions on $\Omega$. Differentiating with respect to $x$:
+$2i\,i_{x}+\beta\,i_{x}+\alpha_{x}+\beta_{x}\,i=0,$
+which is the identity $(2i+\beta)i_{x}=-\alpha_{x}-\beta_{x}\,i$.
+This is exactly the condition that $\partial_{x}(i^{2})=2i\cdot i_{x}$ holds in the fiber algebra.
+The same argument applies to $\partial_{y}$.
+Hence $\partial_{x}$ and $\partial_{y}$ are derivations on $C^{1}(\Omega;A)$.
+Since the fiber algebra is commutative, any AA–linear combination of derivations is again a derivation. In particular, $\partial_{\bar{z}}=\frac{1}{2}(\partial_{x}+i\,\partial_{y})$ is a derivation. ∎
+3.3 The universal spectral identity
+Theorem 3.5 (Universal spectral identity).
+For every AA–valued $C^{1}$ section $W=U+Vi$:
+$2\bigl(\partial_{\bar{z}}W\bigr)_{\lambda}=(W_{\lambda})_{x}+\lambda\,(W_{\lambda})_{y}.$ (3.1)
+This holds for all variable elliptic structures.
+Proof.
+From the algebra–level Cauchy–Riemann decomposition, writing $\mathcal{G}=G_{0}+G_{1}\,i$:
+$2\,\partial_{\bar{z}}W=(U_{x}-\alpha V_{y}+VG_{0})+(V_{x}+U_{y}-\beta V_{y}+VG_{1})\,i.$
+Its spectral image is
+$2(\partial_{\bar{z}}W)_{\lambda}=(U_{x}-\alpha V_{y}+VG_{0})+\lambda(V_{x}+U_{y}-\beta V_{y}+VG_{1}).$ (3.2)
+From $W_{\lambda}=U+V\lambda$, using $\lambda^{2}=-\beta\lambda-\alpha$ (the spectral image of the structure relation):
+$(W_{\lambda})_{x}+\lambda(W_{\lambda})_{y}$ $=U_{x}+V_{x}\lambda+V\lambda_{x}+\lambda(U_{y}+V_{y}\lambda+V\lambda_{y})$
+$=(U_{x}-\alpha V_{y})+(V_{x}+U_{y}-\beta V_{y})\lambda+V(\lambda_{x}+\lambda\lambda_{y}).$ (3.3)
+The universal transport law (Chapter 2) gives $\lambda_{x}+\lambda\lambda_{y}=G_{0}+G_{1}\lambda$, so
+$V(\lambda_{x}+\lambda\lambda_{y})=VG_{0}+VG_{1}\lambda.$
+Substituting into (3.3) yields (3.2). ∎
+Identity (3.1) says that the spectral map sends the algebra–level operator $\partial_{\bar{z}}$ to the first–order transport operator $\frac{1}{2}(\partial_{x}+\lambda\,\partial_{y})$:
+$\begin{array}[]{ccc}C^{1}(\Omega;A)&\xrightarrow{\;\partial_{\bar{z}}\;}&C^{0}(\Omega;A)\\[3.0pt] \downarrow\scriptstyle{W\mapsto W_{\lambda}}&&\downarrow\scriptstyle{W\mapsto W_{\lambda}}\\[3.0pt] C^{1}(\Omega;\mathbb{C})&\xrightarrow{\;\frac{1}{2}(\partial_{x}+\lambda\partial_{y})\;}&C^{0}(\Omega;\mathbb{C})\end{array}$
+Both horizontal arrows are derivations (the top by Proposition 3.4, the bottom because it is first–order), and the vertical arrows are ring homomorphisms (Proposition 3.3).
+The diagram commutes for all variable elliptic structures.
+3.4 Consequences
+Corollary 3.6 (Algebra holomorphicity implies spectral transport).
+If $\partial_{\bar{z}}W=0$, then $(W_{\lambda})_{x}+\lambda(W_{\lambda})_{y}=0$.
+Proof.
+Immediate from (3.1): the left side vanishes. ∎
+Corollary 3.7 (Kernel correspondence).
+The spectral map restricts to a ring isomorphism between $\ker\partial_{\bar{z}}$ and $\ker\frac{1}{2}(\partial_{x}+\lambda\partial_{y})$.
+Proof.
+By the intertwining identity, $W$ is in the kernel of $\partial_{\bar{z}}$ if and only if $W_{\lambda}$ is in the kernel of $\frac{1}{2}(\partial_{x}+\lambda\partial_{y})$.
+Both maps preserve products (Proposition 3.3 and the Leibniz rules).
+∎
+3.5 Commutativity with higher derivatives
+The spectral map also commutes with pure yy–derivatives.
+Proposition 3.8.
+For every AA–valued $C^{2}$ section $W=U+Vi$:
+$(W_{yy})_{\lambda}=(W_{\lambda})_{yy}.$
+The same holds for $\partial_{xx}$ and all higher pure partial derivatives.
+Proof.
+$W_{yy}=U_{yy}+V_{yy}\,i+2V_{y}\,i_{y}+V\,i_{yy}$.
+Under $i\mapsto\lambda$: $(W_{yy})_{\lambda}=U_{yy}+V_{yy}\lambda+2V_{y}\lambda_{y}+V\lambda_{yy}$.
+From $W_{\lambda}=U+V\lambda$: $(W_{\lambda})_{yy}=U_{yy}+V_{yy}\lambda+2V_{y}\lambda_{y}+V\lambda_{yy}$.
+The argument for $\partial_{xx}$ is identical.
+∎
+3.6 Interpretation
+The results of this chapter are summarized as follows.
+  1. 1.
+    The algebra–level operator $\partial_{\bar{z}}$ is a derivation on $C^{1}(\Omega;A)$, universally (Proposition 3.4).
+  2. 2.
+    The spectral map $W\mapsto W_{\lambda}$ is a pointwise ring isomorphism (Proposition 3.3).
+  3. 3.
+    These two structures are compatible: the spectral map intertwines $\partial_{\bar{z}}$ with the first–order transport operator $\frac{1}{2}(\partial_{x}+\lambda\partial_{y})$ (Theorem 3.5).
+  4. 4.
+    The intertwining extends to second–order operators via commutativity with pure partial derivatives (Proposition 3.8).
+The obstruction $\mathcal{G}=i_{x}+i\,i_{y}$ appears in the proof of the intertwining identity through the universal transport law, but it does not prevent any of these results from holding.
+Its role is to force the transport law and to make the Cauchy–Riemann system inhomogeneous—not to obstruct the algebraic or intertwining properties of $\partial_{\bar{z}}$.
+Chapter 4 Rigidity as the Conservative Regime of Burgers Transport
 4.1 From Universal Transport to a Distinguished Regime
@@
 This establishes rigidity as a dynamical notion: it identifies those variable elliptic structures whose coefficient transport is conservative.
-3.5 Collapse of Obstructions
-We now show that rigidity causes several independent obstructions to collapse simultaneously.
-3.5.1 Leibniz defect collapse
-Let $f=u+vi$ and $g=p+qi$ be $C^{1}$ sections. As shown in Chapter 1, the Leibniz defect of the generalized Cauchy–Riemann operator is given by
-$\mathcal{N}(f,g)=\frac{1}{2}\,vq\,\mathcal{G}.$
-Proposition 3.4 (Leibniz defect collapse).
-The operator $\partial_{\bar{z}}$ satisfies the Leibniz rule on $C^{1}(\Omega;\allowbreak A)$ if and only if the structure is rigid.
-Proof.
-The defect vanishes for all $f,g$ if and only if $\mathcal{G}\equiv 0$. ∎
-3.5.2 Inhomogeneity collapse
+4.5 Inhomogeneity collapse
 The real Cauchy–Riemann system derived in Chapter 2 has the form
@@
   3. 3.
-    The generalized Cauchy–Riemann operator satisfies the Leibniz rule.
-  4. 4.
-    The real Cauchy–Riemann system is homogeneous.
+    The generalized real Cauchy–Riemann system is homogeneous.
 Proof.
-The equivalence $(1)\Leftrightarrow(2)$ was shown above. Equivalences with $(3)$ and $(4)$ follow from the collapse results. ∎
+The equivalence $(1)\Leftrightarrow(2)$ was shown above. Equivalences with $(3)$ follow from the inhomogeneity collapse result. ∎
 Remark 4.6 (Rigidity as intrinsic holomorphicity).
@@
 This reordering clarifies the role of rigidity: it is not an assumption imposed to obtain dynamics, but a compatibility condition that restores analytic integrability of the variable Cauchy–Riemann calculus within an already dynamical framework.
-In chapter 5 we show that this conservative regime is precisely what allows integral representation formulas and a coherent function theory to re-emerge.
-Chapter 4 Conservative Transport and Explicit Rigid Structures
+In chapter 6 we show that this conservative regime is precisely what allows integral representation formulas and a coherent function theory to re-emerge.
+Chapter 5 Conservative Transport and Explicit Rigid Structures
 5.1 Purpose and Scope
@@
 While the global theory of this equation involves shock waves and weak solutions, our geometric focus (ellipticity) restricts us to the domain of smooth solutions prior to gradient catastrophe.
-4.3 Recalling the Characteristic Formulation
-We utilize the standard method of characteristics to generate solutions. Equation (4.1) implies that $\lambda$ is constant along the integral curves of the vector field $(1,\lambda)$:
-$\frac{dx}{ds}=1,\qquad\frac{dy}{ds}=\lambda,\qquad\frac{d\lambda}{ds}=0.$
-This classical result immediately implies the properties relevant to our construction:
-  • Propagation: Structural data is transported linearly along lines in the complexified $(x,y)$ space.
-  • Implicit Solutions: The general solution is given implicitly by $\lambda=F(y-\lambda x)$ for an arbitrary function $F$.
-  • Blow–up: Smooth solutions generally develop singularities in finite time (characteristic crossing), which corresponds to the degeneration of the elliptic structure.
-In the following sections, we apply this standard machinery to the specific ansatz of linear dependence, recovering a canonical family of rigid structures.
 5.3 Conservative Burgers Transport
@@
 Equation (5.1) is written in the standard complex algebra $\mathbb{C}$. No reference to the variable elliptic multiplication is involved at this level.
-4.5 Characteristic Formulation
-Equation (4.1) admits the transport form
-$(\partial_{x}+\lambda\,\partial_{y})\lambda=0.$ (4.3)
-Thus $\lambda$ is constant along characteristic curves $(x(s),y(s))$ satisfying
-$\frac{dx}{ds}=1,\qquad\frac{dy}{ds}=\lambda,\qquad\frac{d\lambda}{ds}=0.$
-Equivalently, along integral curves of the vector field $(1,\lambda)$ in the (x,y)(x,y)–plane, the spectral parameter is conserved. This immediately implies:
-  • local existence for arbitrary $C^{1}$ initial data;
-  • breakdown through characteristic crossing;
-  • strong constraints on global smooth solutions.
-These are dynamical consequences of the transport structure and do not depend on ellipticity.
 5.4 Real Coefficient Representation
@@
 5.6 Reduction to Ordinary Differential Equations
-Let
-$a=a(x),\qquad b=b_{0}>0\ \text{constant}.$
-Substituting $a=a(x)$ and $b=b_{0}$ (constant) into (4.4), the second equation is identically satisfied, while the first reduces to
-$a_{x}+a\,a_{y}=0.$
-Since $a=a(x)$, this forces $a_{y}=0$, so the system imposes no further restriction on $a(x)$ at this stage.
-To obtain a nontrivial rigid structure, it is more transparent to work directly in terms of $(\alpha,\beta)$.
-The nontrivial case is obtained by allowing $a_{x}\neq 0$ while keeping $b$ constant, which forces linear dependence in $y$.
 Rewriting directly in terms of $(\alpha,\beta)$, impose
 $\alpha=\alpha(x),\qquad\beta=K(x)\,y.$
-Because the rigidity condition is most naturally expressed in the original coefficients $(\alpha,\beta)$, we now translate the ansatz back into these variables.
 Substituting into the rigid coefficient system yields the ODEs
@@
   • Global smoothness is incompatible with unrestricted transport.
-This family serves as the canonical explicit model throughout the remainder of the theory.
+This family serves as the first canonical explicit model throughout the remainder of the theory.
 The next chapters turn to analysis.
 They show that it is precisely the conservative nature of the transport law that allows integral representation formulas, covariant holomorphicity, and a coherent function theory to emerge.
-Chapter 5 The Cauchy–Pompeiu Formula in the Rigid Regime
+Chapter 6 The Cauchy–Pompeiu Formula in the Rigid Regime
 6.1 Standing assumptions and notation
@@
 If w=u+vi(z)w=u+v\,i(z), then
-$N_{z}(w)=u^{2}+\beta(z)\,uv+\alpha(z)\,v^{2}.$
+$N_{z}(w)=u^{2}-\beta(z)\,uv+\alpha(z)\,v^{2}.$
 Moreover, $N_{z}$ is multiplicative:
@@
 $w\widehat{w}=(u+vi)\bigl((u-\beta v)-vi\bigr)=u(u-\beta v)-uvi+vi(u-\beta v)-v^{2}i^{2}.$
-Since the algebra is commutative, the ii–terms cancel, and using $i^{2}=-\beta i-\alpha$ gives $v^{2}i^{2}=v^{2}(\beta i+\alpha)$; the ii–part cancels, leaving $u^{2}+\beta uv+\alpha v^{2}\in\mathbb{R}$.
+Since the algebra is commutative, the ii–terms cancel, and using $i^{2}=-\beta i-\alpha$ gives $v^{2}i^{2}=v^{2}(\beta i+\alpha)$; the ii–part cancels, leaving $u^{2}-\beta uv+\alpha v^{2}\in\mathbb{R}$.
 Multiplicativity follows from $\widehat{ab}=\hat{a}\,\hat{b}$:
@@
 Let
-$Q(z)=\begin{pmatrix}1&\beta(z)/2\\ \beta(z)/2&\alpha(z)\end{pmatrix},\qquad(u,v)Q(z)(u,v)^{T}=N_{z}(u+vi(z)).$
+$Q(z)=\begin{pmatrix}1&-\beta(z)/2\\ -\beta(z)/2&\alpha(z)\end{pmatrix},\qquad(u,v)Q(z)(u,v)^{T}=N_{z}(u+vi(z)).$
 By ellipticity, $Q(z)$ is positive definite for every $z$. On $K$, the eigenvalues of $Q(z)$ are uniformly bounded above and below away from 0 by compactness. ∎
@@
 Proof.
-On $\Omega\setminus\{\zeta\}$ we have $Z\cdot Z^{-1}=1$. By the Leibniz rule for $\partial_{\bar{z}}$ (rigidity),
+On $\Omega\setminus\{\zeta\}$ we have $Z\cdot Z^{-1}=1$. By the Leibniz rule for $\partial_{\bar{z}}$,
 $0=\partial_{\bar{z}}(ZZ^{-1})=(\partial_{\bar{z}}Z)Z^{-1}+Z(\partial_{\bar{z}}Z^{-1}).$
@@
 $d\omega=2(\partial_{\bar{z}}g)\,dx\wedge dy+g\,i_{y}\,dx\wedge dy.$
-Next compute $\partial_{\bar{z}}g=\partial_{\bar{z}}(fZ^{-1})$. Under rigidity, $\partial_{\bar{z}}$ satisfies Leibniz, so
+Next compute $\partial_{\bar{z}}g=\partial_{\bar{z}}(fZ^{-1})$. In general, $\partial_{\bar{z}}$ satisfies Leibniz, so
 $\partial_{\bar{z}}g=(\partial_{\bar{z}}f)Z^{-1}+f\,\partial_{\bar{z}}(Z^{-1}).$
@@
 rigidity is used to ensure $\partial_{\bar{z}}(Z^{-1})=0$ off the diagonal; it does not imply $i_{y}=0$.
-Chapter 6 Covariant Holomorphicity and Gauge Structure in the Rigid Regime
+Chapter 7 Covariant Holomorphicity and Gauge Structure in the Rigid Regime
 7.1 Motivation from the Cauchy–Pompeiu Formula
-Chapter 5 established that under ellipticity and rigidity, every $f\in C^{1}(\Omega;A)$ satisfies the Cauchy–Pompeiu representation
+Chapter 6 established that under ellipticity and rigidity, every $f\in C^{1}(\Omega;A)$ satisfies the Cauchy–Pompeiu representation
 $f(\zeta)=\frac{1}{2\pi\,j(\zeta)}\int_{\partial\Omega}\frac{f(z)}{Z(z,\zeta)}\,d\tilde{z}-\frac{1}{\pi\,j(\zeta)}\iint_{\Omega}\frac{\partial_{\bar{z}}f(z)+\frac{1}{2}f(z)\,i_{y}(z)}{Z(z,\zeta)}\,dx\,dy.$ (7.1)
@@
 7.3 Basic Algebraic Properties
-Under rigidity, $\partial_{\bar{z}}$ satisfies the Leibniz rule
+In general, $\partial_{\bar{z}}$ satisfies the Leibniz rule
 $\partial_{\bar{z}}(fg)=(\partial_{\bar{z}}f)g+f(\partial_{\bar{z}}g).$
@@
 which is nonzero unless $i_{y}\equiv 0$.
-Thus even in the rigid regime, covariantly holomorphic functions are not closed under the usual pointwise product. In order to remedy this will introduce a scalar gauge transformation.
+Thus covariantly holomorphic functions are not closed under the usual pointwise product. In order to remedy this will introduce a scalar gauge transformation.
 7.4 The Weight Equation
@@
 which proves (7.11).
-Chapter 7 Power Series Representations in the Rigid Regime
-7.1 Purpose and Overview
-In the classical theory of complex analysis, holomorphic functions admit local power series representations obtained by differentiating the Cauchy integral formula.
-This procedure relies crucially on two structural facts:
-  • the imaginary unit is constant,
-  • the Cauchy kernel depends only on the difference $z-\zeta$.
-Neither of these properties survives in the presence of a variable elliptic structure. As a result, differentiation under the integral sign is no longer a legitimate operation.
-The purpose of this chapter is twofold:
-  1. 1.
-    to show that covariantly holomorphic functions admit intrinsic power series representations in frozen algebra powers;
-  2. 2.
-    to explain how ordinary holomorphic functions inherit power series expansions only via gauge trivialization, and not directly.
-This resolves the apparent tension between the failure of classical arguments and the persistence of a local expansion theory.
-Important scope distinction.
-Throughout this chapter, the term “power series representation” refers to expansions constructed from integral representations and valid on non-infinitesimal neighborhoods.
-Purely infinitesimal Taylor expansions at a point, which rely only on freezing the algebra at that point, do not detect the variable structure and may exist without invoking any gauge.
-The necessity of gauge trivialization arises only when transport or integral manipulations are involved.
-7.2 Standing Assumptions
-Throughout this chapter we assume:
-  • $\alpha,\beta\in C^{1}(\Omega)$ with $\Delta=4\alpha-\beta^{2}>0$;
-  • the structure is rigid, i.e.
-    $i_{x}+i\,i_{y}=0;$
-  • $A_{z}=\mathbb{R}[X]/(X^{2}+\beta(z)X+\alpha(z))$;
-  • $Df=\bar{\partial}f+\tfrac{1}{2}f\,i_{y}$ is the covariant Cauchy–Riemann operator.
-All statements are local.
-7.3 Covariantly Holomorphic Functions and the Cauchy–Pompeiu Formula
-Recall that in the rigid regime the Cauchy–Pompeiu representation takes the form
-$f(\zeta)=\frac{1}{2\pi j(\zeta)}\int_{\partial\Omega}^{(\zeta)}\frac{f(z)}{Z(z,\zeta)}\,d\tilde{z}-\frac{1}{\pi j(\zeta)}\iint_{\Omega}^{(\zeta)}\frac{\bar{\partial}f+\tfrac{1}{2}f\,i_{y}}{Z(z,\zeta)}\,dx\,dy.$
-Lemma 7.1.
-If $f\in C^{1}(\Omega;A)$ satisfies $Df=0$, then
-$f(\zeta)=\frac{1}{2\pi j(\zeta)}\int_{\partial\Omega}^{(\zeta)}\frac{f(z)}{Z(z,\zeta)}\,d\tilde{z}$.
-Proof.
-This follows immediately from the vanishing of the interior term $\bar{\partial}f+\tfrac{1}{2}f\,i_{y}$ in the Cauchy–Pompeiu formula. ∎
-This boundary representation is the correct replacement for the classical Cauchy integral formula in the variable elliptic setting.
-7.4 Frozen Algebra and Polynomial Powers
-Fix $\zeta=(\xi,\eta)\in\Omega$ and define the frozen generator $i(\zeta)\in A_{\zeta}$. Introduce the frozen variable
-$Z_{0}(z,\zeta):=(y-\eta)-i(\zeta)(x-\xi)\in A_{\zeta}.$
-All powers $Z_{0}^{n}$ are computed in the fixed algebra $A_{\zeta}$. They are genuine polynomials in $(x-\xi,y-\eta)$ and form a linearly independent family over $\mathbb{R}$.
-The variable kernel $Z(z,\zeta)$ differs from $Z_{0}(z,\zeta)$ only through the dependence of $i(z)$ on $z$, which is controlled by ellipticity and rigidity.
-7.5 Expansion of the Cauchy Kernel
-On sufficiently small circles $|z-\zeta|=r$, the inverse kernel admits the convergent expansion
-$\frac{1}{Z(z,\zeta)}=\sum_{n=0}^{\infty}\frac{Z_{0}(z,\zeta)^{n}}{\rho(z,\zeta)^{n+1}},\qquad\rho(z,\zeta):=|Z_{0}(z,\zeta)|_{\zeta}^{2}.$
-This expansion is purely algebraic and takes place entirely in the frozen fiber $A_{\zeta}$. Uniform ellipticity guarantees convergence.
-7.6 Power Series for Covariantly Holomorphic Functions
-Theorem 7.2 (Power series for covariantly holomorphic functions).
-Let $f\in C^{1}(\Omega;A)$ satisfy $Df=0$. Then for every $\zeta\in\Omega$ there exists $r>0$ and unique coefficients $a_{n}(\zeta)\in A_{\zeta}$ such that
-$f(z)=\sum_{n=0}^{\infty}a_{n}(\zeta)\,Z_{0}(z,\zeta)^{n},\qquad|z-\zeta|<r,$
-with uniform convergence on compact subsets. These coefficients arise from the Cauchy–Pompeiu representation and are therefore intrinsic to the covariant structure.
-Proof.
-Fix $\zeta$ and apply the boundary representation on $\partial B_{r}(\zeta)$.
-Insert the kernel expansion and integrate term-by-term.
-Coefficientwise integration is legitimate since all terms lie in the fixed algebra $A_{\zeta}$.
-Define
-$a_{n}(\zeta):=\frac{1}{2\pi j(\zeta)}\int_{\partial B_{r}(\zeta)}^{(\zeta)}\frac{f(z)}{\rho(z,\zeta)^{n+1}}\,d\tilde{z}$.
-Ellipticity provides absolute convergence. Uniqueness follows from the linear independence of $\{Z_{0}^{n}\}$ in $A_{\zeta}$. ∎
-7.7 Why Differentiation Under the Integral Fails in the Variable Case
-We now explain why the classical argument based on differentiating the Cauchy integral formula does not extend to variable elliptic structures.
-7.7.1 Dependence of the Kernel on the Evaluation Point
-In the classical case, the kernel $(z-\zeta)^{-1}$ depends on $\zeta$ only through translation. In the variable case, the kernel
-$Z(z,\zeta)=(y-\eta)-i(z)(x-\xi)$
-depends on $\zeta$ in two independent ways:
-  • explicitly through $(\xi,\eta)$,
-  • implicitly through the evaluation fiber $A_{\zeta}$.
-Differentiating with respect to $\zeta$ therefore produces terms involving derivatives of the generator $i(z)$ transported into a different fiber.
-There is no canonical identification that allows these terms to be controlled.
-7.7.2 Moving Fibers and Loss of Algebraic Compatibility
-The integral
-$\int^{(\zeta)}\frac{f(z)}{Z(z,\zeta)}\,d\tilde{z}$
-takes values in the frozen fiber $A_{\zeta}$ by coefficientwise projection. Differentiating with respect to $\zeta$ changes the target algebra itself.
-Thus differentiation and integration do not commute: the operation “differentiate in $\zeta$” is not fiberwise compatible with the coefficientwise integration used to define the integral.
-7.7.3 Non-Closure of the Cauchy 1–Form
-Even under rigidity, the Cauchy 1–form
-$d\tilde{z}=dy-i(z)\,dx$
-is not closed:
-$d(d\tilde{z})=i_{y}\,dx\wedge dy.$
-This produces the interior correction term in the Cauchy–Pompeiu formula and forces the introduction of the covariant operator $D$.
-Attempting to differentiate under the integral sign ignores this structural defect and leads to incorrect identities.
-7.7.4 Conclusion
-The failure of differentiation under the integral is not technical. It is structural and reflects:
-  • the variability of the algebra,
-  • the lack of a canonical global trivialization,
-  • the non-closure of the Cauchy 1–form.
-Power series expansions must therefore be constructed from the integral representation itself, not by differentiation.
-Remark 7.3 (Infinitesimal versus transported expansions).
-The frozen variable $Z_{0}(z,\zeta)$ captures the infinitesimal behavior of the algebra structure at the point $\zeta$.
-Any $C^{1}$ function admits a formal Taylor expansion in powers of $(x-\xi,y-\eta)$, and such expansions may be rewritten in terms of $Z_{0}$ without reference to the covariant structure.
-However, these purely infinitesimal expansions do not interact correctly with integral representations or with transport away from the basepoint.
-The constructions in this chapter address power series derived from the Cauchy–Pompeiu formula, which necessarily encode non-infinitesimal behavior and therefore require the covariant framework.
-7.8 Power Series for Holomorphic Functions via Gauge
-Let $h\in C^{1}(\Omega;A)$ satisfy $\bar{\partial}h=0$.
-While $h$ admits a formal Taylor expansion at any point by freezing the algebra, such expansions do not respect the integral structure or transport properties of the variable elliptic system.
-To obtain power series representations compatible with the Cauchy–Pompeiu formula, a gauge trivialization is required.
-Theorem 7.4 (Power series for holomorphic functions).
-For every $\zeta\in\Omega$ there exist $r>0$ and coefficients $b_{n}(\zeta)\in A_{\zeta}$ such that
-$h(z)=\sum_{n=0}^{\infty}b_{n}(\zeta)\,Z_{0}(z,\zeta)^{n},\qquad|z-\zeta|<r.$
-Proof.
-Expand $f$ by the previous theorem and absorb $\psi(\zeta)$ into the coefficients. Convergence is preserved since $\psi$ is smooth and nonvanishing. ∎
-Remark 7.5.
-The role of the gauge in this section is not to enable the existence of local Taylor jets, which are purely infinitesimal, but to reconcile ordinary holomorphicity with the covariant framework underlying integral representations.
-Without gauge trivialization, the notions of differentiation, integration, and algebra multiplication fail to commute in a structurally controlled way.
-7.9 Final Remarks
-Covariant holomorphicity is the correct analytic notion in variable elliptic structures. It admits an intrinsic power series theory without invoking forbidden operations.
-Ordinary holomorphicity inherits power series expansions only through gauge trivialization, not directly.
-This clarifies both the necessity and the strength of the covariant framework.
-In summary, Taylor expansions probe only the frozen infinitesimal algebra at a point, whereas Cauchy–Pompeiu expansions probe the transport of the algebra itself.
-The latter, and not the former, necessitate the covariant and gauge-theoretic framework developed here.
 Chapter 8 Second–Order Operators and Factorization in the Rigid Regime
@@
 and this identity is the bridge between holomorphicity and elliptic theory.
-In the present framework, once rigidity is imposed (Chapter 2), $\partial_{\bar{z}}$ becomes a derivation and the Cauchy–Pompeiu formula (Chapter 5) becomes available.
+In the present framework, once rigidity is imposed (Chapter 2), $\partial_{\bar{z}}$ becomes a derivation and the Cauchy–Pompeiu formula (Chapter 6) becomes available.
 It is then natural—and unavoidable—to study the second–order operator obtained by composing $\partial_{\bar{z}}$ with its conjugate partner $\partial_{z}$.
@@
 $\widehat{W}:=U+\widehat{i}\,V=U-(\beta+i)V.$
+9.1.2 The structure polynomial is determined by the principal symbol
+The coefficients $\alpha,\beta$ defined in (9.2) are not an artifact of the substitution used to reach canonical form.
+They are determined, uniquely and intrinsically, by the principal symbol of the elliptic system.
+Proposition 9.1.
+Let $\sigma(z;\xi,\eta)$ denote the principal symbol of (9.1) at the point $z\in\Omega$ and covector $(\xi,\eta)\in\mathbb{R}^{2}$. Then
+$\det\sigma(z;\xi,\eta)=a_{11}\bigl(\xi^{2}-\beta\,\xi\eta+\alpha\,\eta^{2}\bigr)=a_{11}\,N(\xi+i\eta),$ (9.3)
+where $N$ is the algebra norm of $A_{z}$, i.e. the norm form determined by the structure polynomial $X^{2}+\beta X+\alpha$.
+Proof.
+The principal symbol of (9.1) is the matrix
+$\sigma(z;\xi,\eta)=\begin{pmatrix}a_{11}\xi+a_{12}\eta&-\eta\\[4.0pt] a_{21}\xi+a_{22}\eta&\xi\end{pmatrix}$.
+Its determinant is
+$\det\sigma=(a_{11}\xi+a_{12}\eta)\,\xi+(a_{21}\xi+a_{22}\eta)\,\eta=a_{11}\xi^{2}+(a_{12}+a_{21})\,\xi\eta+a_{22}\,\eta^{2}.$
+Dividing by $a_{11}>0$ and substituting (9.2):
+$\frac{\det\sigma}{a_{11}}=\xi^{2}+\frac{a_{12}+a_{21}}{a_{11}}\,\xi\eta+\frac{a_{22}}{a_{11}}\,\eta^{2}=\xi^{2}-\beta\,\xi\eta+\alpha\,\eta^{2}.$ (9.4)
+On the other hand, in the algebra $A_{z}$ with generator satisfying $i^{2}+\beta\,i+\alpha=0$, the norm of an element $w=a+ib$ is
+$N(a+ib)=w\cdot\widehat{w}=(a+ib)(a+\widehat{i}\,b)=a^{2}-\beta\,ab+\alpha\,b^{2}.$
+Evaluating at $w=\xi+i\eta$ gives
+$N(\xi+i\eta)=\xi^{2}-\beta\,\xi\eta+\alpha\,\eta^{2},$
+which coincides with (9.4). ∎
+This identification has three immediate consequences.
+  1. 1.
+    Ellipticity is equivalent to ellipticity of the algebra.
+The system (9.1) is elliptic (i.e.
+$\det\sigma\neq 0$ for all $(\xi,\eta)\neq(0,0)$) if and only if the norm form $N(\xi+i\eta)$ is positive definite, which is precisely the condition $\Delta=4\alpha-\beta^{2}>0$.
+  2. 2.
+    Uniqueness.
+For a given elliptic system, the structure polynomial $X^{2}+\beta X+\alpha$ is the unique monic quadratic whose associated norm form equals the normalized symbol determinant.
+The coefficients $\alpha,\beta$ are therefore invariants of the principal part, not of the reduction procedure.
+  3. 3.
+    The substitution realizes the symbol.
+The explicit substitution in Lemma 9.4 does not choose $\alpha$ and $\beta$; it produces a system whose principal part is written in the algebra already determined by the symbol.
+Corollary 9.2.
+The structure polynomial $X^{2}+\beta X+\alpha$ is invariant under elliptic–compatible row operations and invertible substitutions on the target variables.
+Proof.
+Let $R=R(z)$ be an invertible $2\times 2$ matrix (a row operation on the system) and $M=M(z)$ an invertible substitution on the target variables $(u,v)\mapsto M(u,v)$.
+Under these transformations the principal symbol changes as
+$\sigma\;\longmapsto\;R\,\sigma\,M^{-1},$
+so
+$\det\sigma\;\longmapsto\;\frac{\det R}{\det M}\,\det\sigma.$
+The prefactor $\det R/\det M$ is a nonzero scalar depending only on $z$, not on the covector $(\xi,\eta)$.
+Since $\alpha$ and $\beta$ are obtained by normalizing $\det\sigma$ to monic form in $\xi$ (i.e.
+dividing by the coefficient of $\xi^{2}$), the scalar cancels identically:
+$\frac{\det(R\,\sigma\,M^{-1})}{a_{11}\,\det R/\det M}=\frac{\det\sigma}{a_{11}}=\xi^{2}-\beta\,\xi\eta+\alpha\,\eta^{2}.$
+Therefore $\alpha$ and $\beta$ depend only on the equivalence class of the system, not on the particular representative. ∎
+Remark 9.3.
+In particular, the substitution $U=a_{22}\,u$, $V=v-a_{12}\,u$ used in Lemma 9.4 is one choice of representative within this equivalence class.
+Any other elliptic–compatible reduction would produce a system governed by the same structure polynomial.
 9.1.3 Canonical principal part
-By elementary substitutions preserving ellipticity, system (9.1) can be written in the form
+Lemma 9.4.
+Suppose the principal coefficients $a_{ij}$ in (9.1) are of class $C^{1}$ with $a_{11},a_{22}>0$. Then the substitution
+$U=a_{22}\,u,\qquad V=v-a_{12}\,u,$
+followed by a single row operation and a scaling, reduces (9.1) to the canonical form
 $\begin{cases}-\alpha V_{y}+U_{x}+aU+bV=f,\\ \ \ V_{x}-\beta V_{y}+U_{y}+cU+dV=g,\end{cases}$ (9.5)
-where $a,b,c,d,f,g$ are real-valued functions on $\Omega$.
+where $a,b,c,d,f,g$ are real–valued $C^{0}$ functions determined by the original coefficients and their first derivatives.
+Proof.
+Under the substitution $U=a_{22}\,u$, $V=v-a_{12}\,u$ (with $a_{22}>0$), the principal parts of (9.1) become
+$\begin{cases}-V_{y}+\dfrac{a_{11}}{a_{22}}\,U_{x}+a_{\ast}U+b_{\ast}V=f_{1},\\[8.0pt] \ \ V_{x}+\dfrac{a_{21}+a_{12}}{a_{22}}\,U_{x}+U_{y}+c_{\ast}U+d_{\ast}V=f_{2},\end{cases}$ (9.6)
+where $a_{\ast},b_{\ast},c_{\ast},d_{\ast}$ absorb all contributions from the derivatives of the variable coefficients $a_{ij}$ and the original lower–order terms.
+With the structure coefficients (9.2) we have
+$\frac{a_{11}}{a_{22}}=\frac{1}{\alpha},\qquad\frac{a_{21}+a_{12}}{a_{22}}=-\frac{\beta}{\alpha},$
+so (9.6) reads
+$\begin{cases}-V_{y}+\dfrac{1}{\alpha}\,U_{x}+a_{\ast}U+b_{\ast}V=f_{1},\\[8.0pt] \ \ V_{x}-\dfrac{\beta}{\alpha}\,U_{x}+U_{y}+c_{\ast}U+d_{\ast}V=f_{2}.\end{cases}$ (9.7)
+Row operation.  Multiply the first equation of (9.7) by $\beta$ and add to the second. The $U_{x}$ coupling in the second line cancels:
+$\frac{\beta}{\alpha}\,U_{x}+\frac{\beta}{\alpha}\,U_{x}=0,$
+yielding
+$\begin{cases}-V_{y}+\dfrac{1}{\alpha}\,U_{x}+a_{\ast}U+b_{\ast}V=f_{1},\\[8.0pt] \ \ V_{x}-\beta V_{y}+U_{y}+c^{\prime}U+d^{\prime}V=g^{\prime}.\end{cases}$ (9.8)
+Scaling.  Multiply the first equation of (9.8) by $\alpha>0$ to obtain the canonical form (9.5). ∎
+Remark 9.5.
+The structure coefficients $\alpha$ and $\beta$ are not imposed on the system; they are read off from it. No uniform ellipticity assumption is required, and no auxiliary Beltrami equation is solved.
+9.1.4 Intrinsic decomposition and the obstruction
 Introduce the algebra–valued field
-$W:=U+iV.$
-9.1.3 Intrinsic decomposition of $\bar{\partial}W$
-Define the variable–structure Cauchy–Riemann operator
+$W:=U+iV,$
+and define the variable–structure Cauchy–Riemann operator
 $\bar{\partial}:=\tfrac{1}{2}(\partial_{x}+i\partial_{y}).$
@@
 is the intrinsic obstruction of the elliptic structure.
-Substituting (9.3) into (9.4) gives
-$2\bar{\partial}W-VG=(f+ig)-(a+ic)U-(b+id)V.$ (9.5)
+Substitution.  The canonical system (9.5) identifies the two real components of (9.9) directly:
+$U_{x}-\alpha V_{y}$ $=f-aU-bV,$
+$V_{x}+U_{y}-\beta V_{y}$ $=g-cU-dV.$
+Substituting into (9.9) gives
+$2\bar{\partial}W=(f-aU-bV)+(g-cU-dV)\,i+V\,G.$ (9.10)
+Decomposing the obstruction.  Write $G$ in its real and imaginary components in the algebra $A_{z}$:
+$G=G_{1}+G_{2}\,i,\qquad G_{1},\,G_{2}\;\text{real.}$
+Since $V$ is real, $VG=VG_{1}+(VG_{2})\,i$. Adding the obstruction components to each line of (9.10):
+$2\bar{\partial}W=\bigl(f-aU-(b-G_{1})V\bigr)+\bigl(g-cU-(d-G_{2})V\bigr)\,i.$ (9.11)
+Define the modified lower–order coefficients
+$\tilde{b}:=b-G_{1},\qquad\tilde{d}:=d-G_{2}.$
+Then (9.11) reads
+$2\bar{\partial}W=(f+ig)-(a+ic)\,U-(\tilde{b}+i\tilde{d})\,V.$ (9.12)
+Remark 9.6.
+When the structure is rigid ($G\equiv 0$), we have $G_{1}=G_{2}=0$ and $\tilde{b}=b$, $\tilde{d}=d$. In the general case, the obstruction modifies only the coefficients of $V$ and does so additively.
 9.1.5 Elliptic conjugation and Vekua form
@@
 $W-\widehat{W}=(2i+\beta)V,\qquad W+\widehat{W}=2U-\beta V.$ (9.13)
-Since $\Delta>0$, the element $2i+\beta$ is invertible in $A_{z}$ and in fact
-$(2i+\beta)^{-1}=\frac{\beta-2i}{\beta^{2}+4\alpha}.$
+Since $\Delta>0$, the element $2i+\beta$ is invertible in $A_{z}$, with
+$(2i+\beta)^{-1}=\frac{\beta-2i}{\beta^{2}+4\alpha}=\frac{\beta-2i}{\Delta}.$
 Therefore we may solve (9.13) for $U$ and $V$ in terms of (W,W^)($\widehat{W}$):
 $V=\frac{W-\widehat{W}}{2i+\beta},\qquad U=\frac{1}{2}\left(W+\widehat{W}+\beta\,\frac{W-\widehat{W}}{2i+\beta}\right).$ (9.14)
-Substituting (9.7) into the right–hand side of (9.5) expresses it as an $A_{z}$–linear combination of $W$ and $\widehat{W}$.
-Concretely, there exist unique algebra–valued coefficients $A_{0},B_{0}$ and a unique datum $F_{0}$ such that
-$(f+ig)-(a+ic)U-(b+id)V\;=\;2F_{0}-2A_{0}\,W-2B_{0}\,\widehat{W}$. (9.8)
-(Existence and uniqueness follow because (9.7) is an invertible linear change of variables between $(U,V)$ and (W,W^)($\widehat{W}$).)
-Define
-$F:=F_{0}=\tfrac{1}{2}(f+ig),$
-and absorb the obstruction term $VG$ from (9.5) using $V=(W-\widehat{W})/(2i+\beta)$:
-$VG=\frac{W-\widehat{W}}{2i+\beta}\,G.$
-Thus (9.5) may be rewritten uniquely in the form
-$2\bar{\partial}W=2F-2A\,W-2B\,\widehat{W}$, (9.9)
-where
-$A\;:=\;A_{0}-\frac{1}{2}\,\frac{G}{2i+\beta},\qquad B\;:=\;B_{0}+\frac{1}{2}\,\frac{G}{2i+\beta}.$ (9.10)
+Substituting (9.14) into the right–hand side of (9.12) expresses it as an $A_{z}$–linear function of $W$ and $\widehat{W}$.
+Since (9.14) is an invertible linear change of variables between $(U,V)$ and (W,W^)($\widehat{W}$), there exist unique algebra–valued coefficients $A,B$ and a unique datum $F$ such that
+$2\bar{\partial}W=2F-2A\,W-2B\,\widehat{W}$. (9.15)
 Dividing by $2$ yields the canonical variable–structure Vekua equation
 $\bar{\partial}W+A\,W+B\,\widehat{W}=F.$ (9.16)
+For reference, the coefficients are given explicitly by
+$A$ $=\frac{1}{4}\!\left(a-\frac{\beta}{\alpha}\,\tilde{b}+\tilde{d}\;+\;\Bigl(c-\frac{1}{\alpha}\,\tilde{b}\Bigr)\,i\right),$ (9.17)
+$B$ $=\frac{1}{4}\!\left(a+\frac{\beta}{\alpha}\,\tilde{b}-\tilde{d}\;+\;\Bigl(c+\frac{1}{\alpha}\,\tilde{b}\Bigr)\,i\right),$
+$F$ $=\tfrac{1}{2}(f+ig),$
+where $\tilde{b}=b-G_{1}$, $\tilde{d}=d-G_{2}$, and $G=G_{1}+G_{2}\,i$ is the intrinsic obstruction.
 Remark 9.7.
@@
 9.1.6 Remarks
-  • Equation (9.11) is intrinsic to the elliptic structure $(\alpha,\beta)$ and requires no auxiliary reduction.
-  • When the structure is rigid ($G\equiv 0$), the obstruction contribution in (9.10) vanishes and the lower–order coefficients reduce to the rigid forms determined by (9.8).
-  • The obstruction $G$ measures precisely the deviation from rigid Cauchy–Riemann calculus and is absorbed into the lower–order terms via the explicit splitting in (9.10).
+  • Equation (9.16) is intrinsic to the elliptic structure $(\alpha,\beta)$ and requires no auxiliary Beltrami reduction.
+  • The entire derivation is constructive: the substitution $(u,v)\mapsto(U,V)$ and the row operation are explicit, and the structure coefficients $(\alpha,\beta)$ emerge directly from the principal part of the system.
+  • The obstruction $G=G_{1}+G_{2}\,i$ enters only by modifying the lower–order coefficients of $V$: it shifts $b\mapsto\tilde{b}=b-G_{1}$ and $d\mapsto\tilde{d}=d-G_{2}$.
+When the structure is rigid ($G\equiv 0$), the coefficients $A,B$ in (9.17) reduce to those of the classical Vekua theory with structure polynomial $X^{2}+\beta X+\alpha$.
 9.2 Rigidization
@@
 $G\equiv 0.$
-Equivalently, rigidity means that the variable–structure Cauchy–Riemann operator $\bar{\partial}=\tfrac{1}{2}(\partial_{x}+i\,\partial_{y})$ satisfies the Leibniz rule on algebra–valued functions.
 9.2.2 Transport form of the obstruction
@@
 Throughout this section we assume that the elliptic structure is rigid, i.e.
-$G\equiv 0,$
-so that the variable–structure Cauchy–Riemann operator
-$\bar{\partial}:=\tfrac{1}{2}(\partial_{x}+i\partial_{y})$
-satisfies the Leibniz rule on algebra–valued functions.
+$G\equiv 0$
 We consider the homogeneous rigid Vekua equation in the special case
@@
 9.5 Initial Value Problems for Rigid Vekua Systems
-Throughout this section we assume that the elliptic structure is rigid,
-$G\equiv 0,$
-so that the variable–structure Cauchy–Riemann operator
-$\bar{\partial}=\tfrac{1}{2}(\partial_{x}+i\partial_{y})$
-satisfies the Leibniz rule and the full rigid calculus applies.
 We study initial value problems for rigid Vekua systems and show that they are well posed and reducible to holomorphic initial value problems via the similarity principle.
@@
   • a structure is rigid if and only if its Beltrami coefficient, viewed as a map to the Poincaré disk, has quasiconformal dilatation equal to $|\mu|$ at every point—the structure encodes its own distortion (Theorem 12.3);
-  • the normalized obstruction $\rho_{T}$ that controls rigidization convergence (Chapter 10) is the norm of a Poincaré residual weighted by the hyperbolic metric density (Proposition 12.8).
-These observations connect the transport theory of Chapters 23 to the classical hyperbolic geometry of the space of conformal structures.
+  • the normalized obstruction $\rho_{T}$ that controls rigidization convergence (Chapter 10) is the norm of a Beltrami residual weighted by the hyperbolic metric density (Proposition 12.8).
+These observations connect the transport theory of Chapters 24 to the classical hyperbolic geometry of the space of conformal structures.
 12.2 The Beltrami Coefficient and the Möbius Bridge
@@
 The structure map is a solution of its own equation.
-12.5 The Poincaré Residual and the Normalized Obstruction
+12.5 The Beltrami Residual and the Normalized Obstruction
 12.5.1 Definition of the residual
 When the structure is not rigid, $\mu_{\bar{z}}-\mu\,\mu_{z}\neq 0$.
-Definition 12.7 (Poincaré residual).
-The Poincaré residual of a variable elliptic structure is
+Definition 12.7 (Beltrami residual).
+The Beltrami residual of a variable elliptic structure is
 $\mathcal{R}(\mu):=\mu_{\bar{z}}-\mu\,\mu_{z}.$ (12.14)
@@
 $\rho_{T}=\frac{\|T\|_{C^{0,\alpha}}}{b_{\min}^{2}},$
-where $b_{\min}=\min_{\bar{\Omega}}\operatorname{Im}\lambda$ is the spectral floor. We now express $\rho_{T}$ in terms of the Poincaré residual.
+where $b_{\min}=\min_{\bar{\Omega}}\operatorname{Im}\lambda$ is the spectral floor. We now express $\rho_{T}$ in terms of the Beltrami residual.
 Proposition 12.8 (Hyperbolic expression for the normalized obstruction).
@@
 12.5.3 The differential of a rigid structure
-The Poincaré residual admits a differential-geometric reformulation. Write $d\mu=\mu_{z}\,dz+\mu_{\bar{z}}\,d\bar{z}$. The rigidity condition $\mu_{\bar{z}}=\mu\,\mu_{z}$ is equivalent to
+The Beltrami residual admits a differential-geometric reformulation. Write $d\mu=\mu_{z}\,dz+\mu_{\bar{z}}\,d\bar{z}$. The rigidity condition $\mu_{\bar{z}}=\mu\,\mu_{z}$ is equivalent to
 $d\mu=\mu_{z}\bigl(dz+\mu\,d\bar{z}\bigr).$ (12.17)
@@
 The rigidity condition (12.7) characterizes the variable structures $\mu:\Omega\to\mathbb{D}$ that are “self-consistent” in the sense that their variation across $\Omega$ is compatible with the conformal geometry they define.
-The Poincaré residual $\mathcal{R}$ measures the failure of this compatibility, weighted by the Teichmüller distance to the boundary of moduli space.
+The Beltrami residual $\mathcal{R}$ measures the failure of this compatibility, weighted by the Teichmüller distance to the boundary of moduli space.
 12.7.2 Harmonic maps
@@
 Let $\Omega\subset\mathbb{R}^{2}$ be a domain.
-The Poincaré residual $\|\mathcal{R}(\mu)\|_{\infty}$ and the supremum $\|\mu\|_{\infty}$ are independently prescribable: for any $R_{0}\geq 0$ and $0\leq\mu_{0}<1$, there exists a $C^{\infty}$ map $\mu:\Omega\to\mathbb{D}$ with $\|\mathcal{R}\|_{\infty}$ arbitrarily close to $R_{0}$ and $\|\mu\|_{\infty}$ arbitrarily close to $\mu_{0}$.
+The Beltrami residual $\|\mathcal{R}(\mu)\|_{\infty}$ and the supremum $\|\mu\|_{\infty}$ are independently prescribable: for any $R_{0}\geq 0$ and $0\leq\mu_{0}<1$, there exists a $C^{\infty}$ map $\mu:\Omega\to\mathbb{D}$ with $\|\mathcal{R}\|_{\infty}$ arbitrarily close to $R_{0}$ and $\|\mu\|_{\infty}$ arbitrarily close to $\mu_{0}$.
 Proof.
@@
 Since these are different orders of the jet, they are independently prescribable: one can construct maps with prescribed pointwise values and prescribed first derivatives (subject only to smoothness).
-The explicit Burgers solutions of Chapter 4 provide $\mathcal{R}=0$ with arbitrary $\|\mu\|_{\infty}$, and perturbations of $\mu=0$ provide arbitrary $\|\mathcal{R}\|_{\infty}$ with $\|\mu\|_{\infty}\approx 0$.
+The explicit Burgers solutions of Chapter 5 provide $\mathcal{R}=0$ with arbitrary $\|\mu\|_{\infty}$, and perturbations of $\mu=0$ provide arbitrary $\|\mathcal{R}\|_{\infty}$ with $\|\mu\|_{\infty}\approx 0$.
 ∎
@@
 Rigidity $\lambda_{x}+\lambda\,\lambda_{y}=0$ $\mu_{\bar{z}}=\mu\,\mu_{z}$
-Equation type real Burgers complex Beltrami–Burgers
+Equation type complex Burgers complex Beltrami–Burgers
 Coordinates real $(x,y)$ complex $(z,\bar{z})$
@@
 The map $\mu$ satisfies the Beltrami equation $f_{\bar{z}}=\mu\,f_{z}$ that it itself defines—a literal instance of self-reference, not merely a numerical coincidence of moduli.
-The normalized obstruction $\rho_{T}$, which controls the convergence of the rigidization scheme (Chapter 10, Theorem 10.15), turns out to measure the Poincaré residual weighted by the hyperbolic metric density.
+The normalized obstruction $\rho_{T}$, which controls the convergence of the rigidization scheme (Chapter 10, Theorem 10.15), turns out to measure the Beltrami residual weighted by the hyperbolic metric density.
 This was not designed; it emerged from the algebra of the Möbius transform relating $\lambda$ and $\mu$.
@@
 has a parallel reading in the Beltrami picture:
-[…]
+$\text{Beltrami residual}\;\to\;\text{self-dilatation}\;\to\;\text{analytic closure}.$
 The same structural content appears in different geometric clothing. The Epilogue will show that it also appears in the Hopf bundle.
@@
   3. 3.
-    The transport field $R(\mu)=\mu_{\bar{z}}-\mu\,\mu_{z}$, the Poincaré residual of Definition 12.6, which measures how the structure moves through the disk.
+    The transport field $R(\mu)=\mu_{\bar{z}}-\mu\,\mu_{z}$, the Beltrami residual of Definition 12.6, which measures how the structure moves through the disk.
 The purpose of this chapter is to prove and develop the consequences of a single fact: the position $|\mu|$ and the velocity $R(\mu)$ are independent.
@@
 Definition 13.2 (First-order invariant).
-The Poincaré residual $R(\mu)=\mu_{\bar{z}}-\mu\,\mu_{z}$ (Definition 12.6; recalled here for the reader’s convenience) depends on the first derivatives of $\mu$.
+The Beltrami residual $R(\mu)=\mu_{\bar{z}}-\mu\,\mu_{z}$ (Definition 12.6; recalled here for the reader’s convenience) depends on the first derivatives of $\mu$.
 Its Hölder norm $\|R(\mu)\|_{C^{0,\alpha}(\overline{\Omega})}$ measures how the structure moves through the Poincaré disk.
-The Poincaré residual is the invariant that the rigidizer sees.
+The Beltrami residual is the invariant that the rigidizer sees.
 Theorem 10.15 converges when $\|R\|_{C^{0,\alpha}}$ is small relative to the Poincaré metric density at the structure’s position; its convergence is blind to $\|\mu\|_{C^{0}}$ except through this metric weighting.
@@
 13.3 The $\delta$-Family: A Canonical Example
-Before proving the Independence Theorem, we construct a one-parameter family of rigid structures whose Beltrami modulus sweeps the entire interval $[0,1)$ while the Poincaré residual vanishes identically.
+Before proving the Independence Theorem, we construct a one-parameter family of rigid structures whose Beltrami modulus sweeps the entire interval $[0,1)$ while the Beltrami residual vanishes identically.
 This family is the load-bearing example for Quadrant III of the theorem and for the blind-spot corollary that follows.
@@
 The elliptic condition number $\kappa_{\delta}:=\bigl(\tfrac{1+\|\mu_{\delta}\|_{\infty}}{1-\|\mu_{\delta}\|_{\infty}}\bigr)^{2}$ satisfies $\kappa_{\delta}=O(\delta^{-2})$.
-  3. (iii)
-    Full range.
-The map $\delta\mapsto\|\mu_{\delta}\|_{C^{0}(K)}$ is continuous on $(0,\infty)$, with $\|\mu_{\delta}\|_{C^{0}(K)}\to 0$ as $\delta\to\infty$ and $\|\mu_{\delta}\|_{C^{0}(K)}\to 1$ as $\delta\to 0^{+}$.
-By the intermediate value theorem, every value in $[0,1)$ is attained.
 Proof.
@@
 The condition number estimate follows from $1-\|\mu_{\delta}\|_{\infty}=O(\delta)$.
-(iii) As $\delta\to\infty$, $\lambda_{\delta}\to i\delta/(1+x)\to i\infty$ and $\mu_{\delta}\to 1$—but on any fixed compact set, the dominant term in $\lambda_{\delta}$ is $i\delta/(1+x)$, so $\mu_{\delta}=(\lambda_{\delta}-i)/(\lambda_{\delta}+i)\to 1$.
-However, for $\delta=1$ at the origin, $\mu_{\delta}(0,0)=0$ (since $\lambda_{\delta}(0,0)=i$, the standard structure).
-More precisely, on a fixed compact set $K$, $\|\mu_{\delta}\|_{C^{0}(K)}$ is a continuous function of $\delta$ that tends to $1$ as $\delta\to 0^{+}$.
-For $\delta$ large, $\lambda_{\delta}$ is dominated by the imaginary part $\delta/(1+x)\gg|y/(1+x)|$, so $\lambda_{\delta}\approx i\delta/(1+x)$ and $\mu_{\delta}\approx(\delta/(1+x)-1)/(\delta/(1+x)+1)\to 1$ as well—but this limit is approached from below and is monotone in different regimes.
-At $\delta=1$ and the origin, $|\mu|=0$; the continuous dependence on $\delta$ and the limits at $0^{+}$ and $\infty$ ensure that $\|\mu_{\delta}\|_{C^{0}(K)}$ attains every value in $[0,1)$ for appropriate $\delta$.
+.
 ∎
@@
 The cost of this procedure is independent of $\delta$.
-A detailed treatment, including worked examples and a cost comparison with standard elliptic solvers, appears in [14].
 Remark 13.5 (The $\delta$-family in the difficulty plane).
@@
 The estimate is sharp in its dependence on both factors.
-The numerator $\|R\|_{C^{0,\alpha}}$ is the first-order invariant: the Poincaré residual measures how far the structure departs from rigidity.
+The numerator $\|R\|_{C^{0,\alpha}}$ is the first-order invariant: the Beltrami residual measures how far the structure departs from rigidity.
 The denominator $b_{\min}$ is the Poincaré metric density at the structure’s position: in Beltrami coordinates, $b_{\min}=(1-|\mu|^{2})/|1-\mu|^{2}$ (equation (12.2)), so the estimate degrades near the boundary of the disk exactly as the hyperbolic geometry demands.
 The content is analogous to quantitative rigidity estimates in nonlinear elasticity, where a map that is approximately an isometry is shown to be close to an exact isometry, with stability controlled by the deficit.
-Here the role of the isometry group is played by the rigid structures ($R=0$), and the deficit is the Poincaré residual.
+Here the role of the isometry group is played by the rigid structures ($R=0$), and the deficit is the Beltrami residual.
 The Independence Theorem is what makes this estimate non-trivial: because $|R|$ and $|\mu|$ are independent, the residual $\|R\|_{C^{0,\alpha}}$ is a genuine degree of freedom, not a proxy for $|\mu|$.
@@
 $\begin{cases}u_{x}-p(x,y)\,v_{y}=0,\\ v_{x}+u_{y}=0.\end{cases}$ (13.4)
+Remark 13.15 (Disambiguation from the Polozhii $p$-analytic system).
+The terminology “$p(x)$-analytic” adopted here should not be confused with the $p$-analytic functions of Polozhii [13]. The Polozhii system is
+$\begin{cases}u_{x}=p(x,y)\,v_{y},\\ u_{y}=-p(x,y)\,v_{x},\end{cases}$
+where $p>0$. To read off the structure coefficients, rewrite this in the standard form (9.1): dividing both equations by $p$ gives
+$\begin{cases}-v_{y}+\dfrac{1}{p}\,u_{x}=0,\\[6.0pt] \ \ v_{x}+\dfrac{1}{p}\,u_{y}=0,\end{cases}$
+so $a_{11}=a_{22}=1/p$, $a_{12}=a_{21}=0$. Applying the definitions of Proposition 9.1:
+$\alpha=\frac{a_{22}}{a_{11}}=1,\qquad\beta=-\frac{a_{12}+a_{21}}{a_{11}}=0.$
+The structure polynomial is $X^{2}+1$—standard complex numbers, regardless of $p$.
+The function $p$ cancels in every ratio because it appears symmetrically in both equations: it is a conformal weight, not a structural parameter.
+From the VES viewpoint, Polozhii’s theory is conformally weighted standard complex analysis over a fixed algebra.
+By contrast, the $p(x)$-analytic structures of Definition 13.14 have $\alpha=p$ and $\beta=0$, giving the structure polynomial $X^{2}+p$.
+The function $p$ enters the algebra itself: it changes the fiber, not just the metric.
+The system (13.4) has $p$ only in the first equation ($u_{x}=p\,v_{y}$), while the second ($v_{x}+u_{y}=0$) is the standard one.
+This asymmetry is the signature of a genuinely variable elliptic structure.
+The two systems coincide only when $p\equiv 1$.
 The condition $\beta=0$ has a clean geometric meaning at every level of the theory developed so far.
@@
   2. (ii)
-    The Poincaré residual simplifies. Since $\mu$ is real, $\bar{\mu}=\mu$ and
+    The Beltrami residual simplifies. Since $\mu$ is real, $\bar{\mu}=\mu$ and
     $R(\mu)=\mu_{\bar{z}}-\mu\,\mu_{z}=\tfrac{1}{2}\bigl[\mu_{x}(1-\mu)+i\,\mu_{y}(1+\mu)\bigr].$
@@
 The $p(x)$-analytic structures are not a minor special case.
-They are the real axis of the Poincaré disk, and every classical generalized analytic function theory for systems of the form (13.4)—the theories of Bers, Vekua, Tutschke, and the 2011 work on constant-parameter elliptic structures—lives here.
+They are the real axis of the Poincaré disk, and every classical generalized analytic function theory for systems of the form (13.4).
 The case $p=\text{const}$ is a single point $\mu_{p}\in(-1,1)$ on this diameter.
@@
 Genuine transport obstruction.
-The structure lies off the real axis with large Poincaré residual.
+The structure lies off the real axis with large Beltrami residual.
 Classical quasiconformal methods (MRMT, Ahlfors–Bers) are the only available tool.
@@
 This should not be confused with the distinct—and open—question of whether rigid Beltrami coefficient functions are dense in the function space $L^{\infty}(\Omega,\mathbb{D})$; see Section 14.5.
-14.4 The $\delta$-family within the general rigid class
 The $\delta$-family is a specific one-parameter subfamily of a much larger class. By the method of characteristics, the conservative Burgers equation
@@
 Rather, we express the same invariant in coordinates adapted to a standard geometric structure.
-The purpose is not to replace the algebraic development of the book, but to show that it already fits into a well–known geometric framework.
+The purpose is not to replace the algebraic development of the monograph, but to show that it already fits into a well–known geometric framework.
 1. Normalization and the Sphere
@@
 Global rigidization.
-Conjecture 1 (Appendix G) asserts that every smooth elliptic structure on a compact simply connected domain can be rigidized by a smooth diffeomorphism, without any smallness condition on the obstruction.
+Conjecture 1 (Appendix F) asserts that every smooth elliptic structure on a compact simply connected domain can be rigidized by a smooth diffeomorphism, without any smallness condition on the obstruction.
 The continuity method reduces this to a priori estimates: a uniform Jacobian lower bound and a uniform $C^{2,\alpha}$ bound on the rigidizing diffeomorphisms along the interpolation path.
-Three coupled sub-problems are identified in Section G.5: the Jacobian lower bound (via maximum principle for $\log J_{\Phi}$), the Burgers injectivity estimate (no shock formation in the elliptic regime), and the Schauder closure (Agmon–Douglis–Nirenberg for the specific quasilinear $2\times 2$ system).
+Three coupled sub-problems are identified in Section F.5: the Jacobian lower bound (via maximum principle for $\log J_{\Phi}$), the Burgers injectivity estimate (no shock formation in the elliptic regime), and the Schauder closure (Agmon–Douglis–Nirenberg for the specific quasilinear $2\times 2$ system).
 None of the three has been rigorously closed.
@@
 Characterize the local moduli space of rigid elliptic structures in a neighborhood of the constant structure.
-The jet analysis (Appendix D) shows that the tangent space is parametrized by holomorphic functions and that higher jets are slaved through forced Cauchy–Riemann equations.
-Determine whether the formal jet hierarchy of Appendix D (§ D.10) converges, and whether the resulting moduli space carries additional structure (for instance, an infinite-dimensional complex manifold structure).
+The jet analysis (Appendix C) shows that the tangent space is parametrized by holomorphic functions and that higher jets are slaved through forced Cauchy–Riemann equations.
+Determine whether the formal jet hierarchy of Appendix C (§ D.10) converges, and whether the resulting moduli space carries additional structure (for instance, an infinite-dimensional complex manifold structure).
 6.
@@
 7.
-Similarity principle with conjugate coupling.
-The rigid similarity principle (§§ 9.3–9.4) is established only in the sharp case $B\equiv 0$.
-Develop a similarity principle for the full rigid Vekua equation $\partial_{\bar{z}}W+AW+BW^{c}=F$, either through generating pairs adapted to the variable-structure setting or through solution-dependent integrating factors.
-This is the natural next step for applications to general first-order elliptic systems.
-8.
 Spectral theory of the rigid variable Laplacian.
@@
 Study the spectral properties of this operator on bounded domains within the elliptic region: Weyl asymptotics, comparison with the standard Laplacian, and the effect of the drift terms controlled by $\alpha_{y}$ and $\beta_{y}$ on eigenvalue distribution.
-9.
+8.
 The constructive envelope: sharp or lossy?
@@
 Determine whether $E$ is lower semicontinuous in an appropriate topology, whether its minimizers (subject to boundary conditions) exist, and whether its Euler–Lagrange equation has a clean geometric interpretation in the Poincaré disk.
-11.
+10.
 Extension to Clifford structures in higher dimensions.
@@
 The difficulty of a variable elliptic structure lives in a plane, not on a line.
-Two invariants govern it: the Beltrami modulus $\|\mu\|_{C^{0}}$, a zeroth–order quantity measuring where the structure sits in the Poincaré disk, and the Poincaré residual $R(\mu)=\mu_{\bar{z}}+\mu\,\mu_{z}$, a first–order quantity measuring how the structure moves through it.
+Two invariants govern it: the Beltrami modulus $\|\mu\|_{C^{0}}$, a zeroth–order quantity measuring where the structure sits in the Poincaré disk, and the Beltrami residual $R(\mu)=\mu_{\bar{z}}+\mu\,\mu_{z}$, a first–order quantity measuring how the structure moves through it.
 These are independently prescribable.
@@
 The rigidizer sees a problem already solved.
-The p(x)p(x)–analytic theory, developed by the Bers school and its successors, populates the real diameter of the Poincaré disk: $\mu=(\sqrt{p}-1)/(\sqrt{p}+1)\in(-1,1)$.
-Along this diameter, $\beta=0$ identically; the variable exponent modulates conformal scaling but never opens the transport channel.
-The relationship is exact:
+[…]
 $\begin{array}[]{ccccc}\text{real analysis}&\xrightarrow{\quad\text{analytic continuation}\quad}&\text{complex analysis}\\[6.0pt] p(x)\text{-analytics}&\xrightarrow{\quad\beta\neq 0\quad}&\text{VES}\end{array}$
-The generating pairs, successor structures, embedding theorems, and regularity results of the established theory form the boundary data from which the full variable theory departs.
+The generating pairs, successor structures, embedding theorems, and regularity results of the established theory over the standard algebra form the boundary data from which the full variable theory departs.
 The observation is not that these results are insufficient.
 It is that they explore the real diameter of a disk whose off–axis structure was never investigated.
-The residual budget $B(\mu)$ at each point of the disk measures how far the constructive theory reaches: the maximum Poincaré residual for which the rigidization scheme converges.
+The residual budget $B(\mu)$ at each point of the disk measures how far the constructive theory reaches: the maximum Beltrami residual for which the rigidization scheme converges.
 The budget is not rotationally symmetric.
@@
 The constructive envelope in the difficulty plane is therefore a computable region, not a universal constant.
-The diagnostic algorithm that emerges from this analysis classifies first–order elliptic systems into three branches: structures within the constructive envelope, accessible to the full VES toolkit; p(x)p(x)–analytic structures on the real diameter, governed by the Bers theory; and structures with genuine transport obstruction beyond the envelope, requiring classical quasiconformal methods.
+The diagnostic algorithm that emerges from this analysis classifies first–order elliptic systems into three branches: structures within the constructive envelope, accessible to the full VES toolkit; p(x)p(x)–analytic structures on the real diameter, accessible to the classical Vekua theory after conformal normalization; and structures with genuine transport obstruction beyond the envelope, requiring classical quasiconformal methods.
 Without the independence of position and velocity in the Poincaré disk, this triage is impossible.
@@
 Here, the algebra moves.
-Appendix A Algebraic Universality of the Intrinsic Obstruction and the Leibniz Defect
-This appendix isolates a purely algebraic mechanism underlying all failures of derivation behavior in variable rank–two algebra calculi.
-No analytic notions (such as ellipticity, Cauchy–Riemann operators, or integral representations) are used.
-The results show that once a quadratic algebra structure is allowed to vary, there exists a unique intrinsic obstruction governing both the failure of the Leibniz rule and the transport of the structure itself.
-A.1 First–order calculus on a moving rank–two algebra
-Let $\Omega\subset\mathbb{R}^{2}$ be open and let $A\to\Omega$ be a rank–two commutative real algebra bundle generated by a section $i\in C^{1}(\Omega;A)$ satisfying the structure relation
-$i^{2}+\beta\,i+\alpha=0,$
-where $\alpha,\beta\in C^{1}(\Omega,\mathbb{R})$ are arbitrary real functions.
-Every section $f\in C^{1}(\Omega;A)$ admits a unique decomposition
-$f=u+vi,\qquad u,v\in C^{1}(\Omega;\mathbb{R}).$
-Definition A.1 (Compatible first–order operator).
-A first–order operator
-$\delta:C^{1}(\Omega;A)\longrightarrow C^{0}(\Omega;A)$
-is said to be compatible with the moving algebra structure if:
-  1. 1.
-    it acts as a derivation on real scalar functions,
-    $\delta(uv)=u\,\delta(v)+v\,\delta(u)\quad(u,v\in C^{1}(\Omega;\mathbb{R}));$
-  2. 2.
-    it satisfies the chain rule on the generator,
-    $\delta(vi)=\delta(v)i+v\,\delta(i).$
-No derivation property is assumed on the full algebra $A$.
-A.2 The Leibniz defect
-The deviation of $\delta$ from being a derivation on $A$ is measured by the Leibniz defect
-$N_{\delta}(f,g):=\delta(fg)-\delta(f)g-f\,\delta(g).$
-Lemma A.2 (Factorization of the Leibniz defect).
-Let $\delta$ be a compatible operator. For any sections $f=u+vi$ and $g=p+qi$, one has
-$N_{\delta}(f,g)=vq\Bigl[\delta(i^{2})-2i\,\delta(i)\Bigr].$
-Proof.
-All products involving at least one scalar factor satisfy the Leibniz rule by assumption. The only possible failure arises from the product of the vector parts:
-$(vi)(qi)=vq\,i^{2}.$
-On the one hand,
-$\delta(vq\,i^{2})=\delta(vq)i^{2}+vq\,\delta(i^{2}).$
-On the other hand,
-$\delta(vi)(qi)+(vi)\delta(qi)$ $=(\delta(v)i+v\delta(i))qi+vi(\delta(q)i+q\delta(i))$
-$=\delta(vq)i^{2}+2vq\,i\delta(i).$
-Subtracting yields the stated formula. ∎
-This factorization shows that the defect is bilinear in the vector components and universal: no other algebraic combinations can contribute.
-A.3 Decomposition of the defect
-Using the structure relation $i^{2}=-\beta i-\alpha$ and the scalar derivation property of $\delta$, one computes
-$\delta(i^{2})=-\delta(\alpha)-\delta(\beta)i-\beta\,\delta(i).$
-Substituting into Lemma A.2 gives
-$\delta(i^{2})-2i\delta(i)=-\delta(\alpha)-\delta(\beta)i-(2i+\beta)\delta(i).$
-Corollary A.3 (Total defect formula).
-For any compatible operator $\delta$,
-$N_{\delta}(f,g)=-vq\Bigl[\underbrace{\delta(\alpha)+\delta(\beta)i}_{\text{structural transport}}+\underbrace{(2i+\beta)\delta(i)}_{\text{geometric transport}}\Bigr].$
-This decomposition is purely algebraic and holds independently of ellipticity or invertibility assumptions.
-A.4 Universality of the intrinsic obstruction
-Formula (A.3) shows that any failure of the Leibniz rule for a first–order calculus compatible with a moving quadratic algebra structure can arise from exactly two sources:
-  • transport of the structure coefficients $(\alpha,\beta)$;
-  • transport of the generator $i$ itself.
-No other algebraic mechanism can contribute to a Leibniz defect.
-In the setting of differentiation on the plane, this distinction becomes more refined.
-Pure real derivatives, such as $\partial_{x}$ or $\partial_{y}$ acting coefficientwise on sections $f=u+vi$, always satisfy the Leibniz rule, even when the algebra structure varies.
-Indeed, for such operators one has $\delta(i^{2})=2i\,\delta(i)$, and the defect vanishes identically.
-The intrinsic obstruction appears precisely for first–order operators whose symbol involves the moving generator itself. A canonical example is the generalized Cauchy–Riemann operator
-$\partial_{\bar{z}}=\tfrac{1}{2}(\partial_{x}+i\,\partial_{y}),$
-for which the interaction between differentiation and the variable algebra produces the intrinsic combination
-$G:=i_{x}+i\,i_{y}.$
-This quantity measures the incompatibility between differentiation along algebra–adapted directions and multiplication in the fibers.
-The present appendix shows that $G$ is not introduced by analytic or geometric assumptions.
-It is the unique algebraic obstruction forced by allowing the quadratic structure to vary and by differentiating along directions defined by that structure.
-Whether detected through the failure of the Leibniz rule or through differentiation of the structure polynomial, the same obstruction necessarily appears.
-In the main text, imposing $G\equiv 0$ is shown to eliminate all such defects simultaneously, selecting the conservative (rigid) subregime of the universal transport law.
-Appendix B Constructive Differentiability of the Structure Generator
+Appendix A Constructive Differentiability of the Structure Generator
 This appendix records a recursive argument showing that the structure generator can be differentiated as many times as permitted by the regularity of the structure coefficients.
@@
 No differentiability of $i$ is assumed a priori.
-Proposition B.1 (Recursive constructivist differentiability).
+Proposition A.1 (Recursive constructivist differentiability).
 Assume $\alpha,\beta\in C^{k}(\Omega)$ for some integer $k\geq 0$, and assume that $2i+\beta$ is invertible on $\Omega$.
@@
 Remarks
-Remark B.2.
+Remark A.2.
 The argument shows that the differentiability of the structure generator is entirely determined by the regularity of the structure coefficients.
 No gain of regularity occurs: $i$ cannot be differentiated more times than allowed by the least regular of $\alpha$ and $\beta$.
-Remark B.3.
+Remark A.3.
 Although the proof is presented in the elliptic setting, the recursive construction relies only on the invertibility of $2i+\beta$ and therefore applies verbatim in the hyperbolic regime away from zero divisors.
 For the parabolic regime this cannot be applied directly as $2i+\beta$ is a zero divisor.
-Appendix C Partial Derivatives of the Structure Generator in the Parabolic Regime
+Appendix B Partial Derivatives of the Structure Generator in the Parabolic Regime
 This appendix describes how the partial derivatives of the structure generator $i(x,y)$ behave in the parabolic regime
@@
 In this case the elliptic/hyperbolic formulas for $i_{x}$ and $i_{y}$ cannot be used directly, and a different interpretation is required.
-C.1 Algebraic structure when $\Delta=0$
+B.1 Algebraic structure when $\Delta=0$
 If $\Delta=0$, then
@@
 is nilpotent and therefore a zero divisor.
-C.2 Differentiating the parabolic structure relation
+B.2 Differentiating the parabolic structure relation
 Since $\varepsilon^{2}=0$, differentiating with respect to $x$ and $y$ gives
@@
 Using $\varepsilon_{x}=i_{x}+\tfrac{\beta_{x}}{2}$ and $\varepsilon_{y}=i_{y}+\tfrac{\beta_{y}}{2}$, we obtain the constraints
-$\varepsilon\Bigl(i_{x}+\tfrac{\beta_{x}}{2}\Bigr)=0,\qquad\varepsilon\Bigl(i_{y}+\tfrac{\beta_{y}}{2}\Bigr)=0.$ (C.1)
-In the dual number algebra, the annihilator of $\varepsilon$ is precisely the ideal it generates. Hence (C.1) implies the general form
-$i_{x}=-\frac{\beta_{x}}{2}+\phi_{x}\,\varepsilon,\qquad i_{y}=-\frac{\beta_{y}}{2}+\phi_{y}\,\varepsilon,$ (C.2)
+$\varepsilon\Bigl(i_{x}+\tfrac{\beta_{x}}{2}\Bigr)=0,\qquad\varepsilon\Bigl(i_{y}+\tfrac{\beta_{y}}{2}\Bigr)=0.$ (B.1)
+In the dual number algebra, the annihilator of $\varepsilon$ is precisely the ideal it generates. Hence (B.1) implies the general form
+$i_{x}=-\frac{\beta_{x}}{2}+\phi_{x}\,\varepsilon,\qquad i_{y}=-\frac{\beta_{y}}{2}+\phi_{y}\,\varepsilon,$ (B.2)
 for some real functions $\phi_{x},\phi_{y}$.
 Thus, unlike the elliptic and hyperbolic cases, the structure relation alone does not uniquely determine the derivatives of $i$. There is a one–dimensional nilpotent freedom in each derivative.
-C.3 Canonical choice via a nondegenerate limit
-A natural way to remove the ambiguity in (C.2) is to define $i_{x}$ and $i_{y}$ as limits of the elliptic/hyperbolic formulas as $\Delta\to 0$.
+B.3 Canonical choice via a nondegenerate limit
+A natural way to remove the ambiguity in (B.2) is to define $i_{x}$ and $i_{y}$ as limits of the elliptic/hyperbolic formulas as $\Delta\to 0$.
 For $\Delta\neq 0$, one has
@@
 If $\Delta_{x}$ and $\Delta_{y}$ vanish at least as fast as $\sqrt{|\Delta|}$, the singular terms disappear in the limit $\Delta\to 0$, yielding the canonical parabolic derivatives
-$i_{x}:=-\frac{\beta_{x}}{2},\qquad i_{y}:=-\frac{\beta_{y}}{2}.$ (C.3)
-These are exactly the choices obtained by eliminating the nilpotent component in (C.2).
-C.4 Consistency check
-With the canonical choice (C.3),
+$i_{x}:=-\frac{\beta_{x}}{2},\qquad i_{y}:=-\frac{\beta_{y}}{2}.$ (B.3)
+These are exactly the choices obtained by eliminating the nilpotent component in (B.2).
+B.4 Consistency check
+With the canonical choice (B.3),
 $2i+\beta=2\varepsilon,$
@@
 Thus the canonical derivatives satisfy all algebraic constraints imposed by the parabolic structure.
-C.5 Interpretation
+B.5 Interpretation
 In the parabolic regime the quadratic algebra acquires a nilpotent direction, and the operator $2i+\beta$ loses invertibility.
 Consequently, the structure equation no longer determines $i_{x}$ and $i_{y}$ uniquely; it only constrains them up to a nilpotent term.
-The canonical choice (C.3) selects the derivatives that arise as limits of nondegenerate (elliptic or hyperbolic) structures and removes the nilpotent ambiguity.
+The canonical choice (B.3) selects the derivatives that arise as limits of nondegenerate (elliptic or hyperbolic) structures and removes the nilpotent ambiguity.
 This provides a consistent notion of “constructive differentiability” in the parabolic setting.
-Appendix D Jet Analysis of the $\varepsilon$–Family at the Constant Structure
-This appendix analyzes the $\varepsilon$–family of rigid elliptic structures introduced in Chapter 4 from the viewpoint of jet theory at $\varepsilon=0$.
+Appendix C Jet Analysis of the $\varepsilon$–Family at the Constant Structure
+This appendix analyzes the $\varepsilon$–family of rigid elliptic structures introduced in Chapter 5 from the viewpoint of jet theory at $\varepsilon=0$.
 Although the limiting structure is exactly the standard complex plane, the jet analysis reveals nontrivial infinitesimal geometry encoded in the family.
 The purpose of this appendix is to make precise what information is contained in the $\varepsilon$–jets and what conclusions can be drawn from them.
-D.1 The $\varepsilon$–family and its constant limit
+C.1 The $\varepsilon$–family and its constant limit
 Recall the explicit rigid family
@@
 Thus the $\varepsilon=0$ structure is not merely rigid but completely constant. Any nontrivial behavior of the family must therefore appear at the level of jets in $\varepsilon$.
-D.2 First $\varepsilon$–jet of the coefficients
+C.2 First $\varepsilon$–jet of the coefficients
 Expand the coefficients in $\varepsilon$:
@@
 It measures the initial direction in which the algebra departs from the constant complex structure as $\varepsilon$ is turned on.
-D.3 Jet of the spectral parameter
+C.3 Jet of the spectral parameter
 Recall the canonical spectral parameter
@@
 Thus the first $\varepsilon$–jet of the structure is equivalently encoded in the complex function $\mu$.
-D.4 Linearization of the Burgers transport
+C.4 Linearization of the Burgers transport
 Rigidity is equivalent to the conservative Burgers equation
@@
 This is precisely the classical Cauchy–Riemann equation for the complex–valued function $\mu$.
-D.5 Interpretation of the jet equation
+C.5 Interpretation of the jet equation
 The jet equation
@@
 Thus classical complex analysis appears here as the infinitesimal theory controlling how a constant elliptic structure may begin to vary while remaining rigid to first order.
-D.6 What the jet analysis does not say
+C.6 What the jet analysis does not say
 It is important to emphasize what conclusions should not be drawn:
@@
 Thus the jet analysis captures only the infinitesimal geometry of the space of structures, not its global behavior.
-D.7 Conceptual conclusion
+C.7 Conceptual conclusion
 The jet analysis of the explicit rigid $\varepsilon$–family shows that, at the constant elliptic structure, the universal Burgers transport linearizes to the classical Cauchy–Riemann equations.
@@
 Rather, it identifies complex analysis as the tangent theory at a distinguished point in the space of variable elliptic geometries.
-D.8 Second $\varepsilon$–jet: onset of nonlinearity
+C.8 Second $\varepsilon$–jet: onset of nonlinearity
 We now examine the second $\varepsilon$–jet of the rigid family. Write the spectral parameter as
@@
 In particular, even when $\mu$ is holomorphic, $\nu$ is generally not holomorphic. This marks the first point at which classical complex analysis ceases to describe the structure exactly.
-D.9 Third $\varepsilon$–jet: hierarchy of transport corrections
+C.9 Third $\varepsilon$–jet: hierarchy of transport corrections
 Proceeding one order further, write
@@
 This recursive structure is entirely determined by the Burgers nonlinearity and does not involve any analytic assumptions.
-D.10 Structure of the full jet hierarchy
+C.10 Structure of the full jet hierarchy
 The pattern observed above persists to all orders. Writing formally
@@
 Thus classical holomorphic data parametrize the tangent space at the constant structure, but they do not determine a finite deformation uniquely without solving the full nonlinear transport problem.
-D.11 Refined conclusion of the jet analysis
+C.11 Refined conclusion of the jet analysis
 Combining the first, second, and third jet equations yields a precise refinement of the conceptual picture:
@@
 In this sense, complex analysis is not deformed but embedded as the tangent theory of a nonlinear geometric transport problem.
-Appendix E Integrability, Coordinate Trivialization, and the Intrinsic Obstruction
+Appendix D Integrability, Coordinate Trivialization, and the Intrinsic Obstruction
 Terminology warning.
@@
 The intrinsic obstruction $G$ is unrelated to the former notion but is decisive for the latter.
-E.1 The Classical Integrability Objection
+D.1 The Classical Integrability Objection
 A natural objection to the study of variable elliptic structures in real dimension two arises from classical complex geometry.
@@
 carry any genuine geometric or analytic meaning?
-E.2 What the Obstruction Does Not Measure
+D.2 What the Obstruction Does Not Measure
 Before answering this objection, it is essential to clarify what the obstruction $G$ does not measure.
@@
 The intrinsic obstruction belongs to a different conceptual layer than the Nijenhuis tensor.
-E.3 Parallelism versus Coordinate Trivialization
+D.3 Parallelism versus Coordinate Trivialization
 The obstruction $G$ arises from differentiating a moving generator
@@
 These notions are logically independent.
-E.4 Eulerian and Lagrangian Viewpoints
+D.4 Eulerian and Lagrangian Viewpoints
 The distinction admits a transparent interpretation in transport–theoretic terms.
@@
 The existence of a co–moving frame in which this motion disappears does not negate the reality of the motion itself.
-E.5 Conclusion
+D.5 Conclusion
 In real dimension two, classical integrability is automatic and therefore incapable of distinguishing variable elliptic structures.
@@
 The theory developed here is therefore not a theory of geometric integrability of complex structures in the Newlander–Nirenberg sense, but a theory of transport of algebraic structures and of the analytic calculus they induce in a fixed background frame.
-Appendix F Quaternionic Reformulation of Rigidity via the Hopf Fibration
-F.1 Motivation
+Appendix E Quaternionic Reformulation of Rigidity via the Hopf Fibration
+E.1 Motivation
 A variable elliptic structure on a domain $\Omega\subset\mathbb{R}^{2}$ is encoded by a generator $i=i(x,y)$ satisfying the structure reduction
-$i^{2}+\beta(x,y)\,i+\alpha(x,y)=0,$ (F.1)
+$i^{2}+\beta(x,y)\,i+\alpha(x,y)=0,$ (E.1)
 with ellipticity discriminant
-$\Delta(x,y):=4\alpha(x,y)-\beta(x,y)^{2}>0.$ (F.2)
+$\Delta(x,y):=4\alpha(x,y)-\beta(x,y)^{2}>0.$ (E.2)
 The intrinsic obstruction is
-$G:=i_{x}+i\,i_{y},$ (F.3)
+$G:=i_{x}+i\,i_{y},$ (E.3)
 and the structure is called rigid precisely when $G\equiv 0$.
 The purpose of this appendix is to recast rigidity in a purely quaternionic language via the Hopf fibration, clarifying the geometric content of $G$.
-F.1.1 Normalization to a unit imaginary quaternion
-Completing the square in (F.1) yields
+E.1.1 Normalization to a unit imaginary quaternion
+Completing the square in (E.1) yields
 $(2i+\beta)^{2}=\beta^{2}-4\alpha=-\Delta.$
 Define the normalized generator
-$u(x,y):=\frac{2i(x,y)+\beta(x,y)}{\sqrt{\Delta(x,y)}}.$ (F.4)
+$u(x,y):=\frac{2i(x,y)+\beta(x,y)}{\sqrt{\Delta(x,y)}}.$ (E.4)
 Then
-$u^{2}=-1,$ (F.5)
+$u^{2}=-1,$ (E.5)
 so $u$ takes values in the unit imaginary sphere
@@
 Thus, an elliptic quadratic structure canonically determines a smooth map $u:\Omega\to S^{2}$.
-F.2 Hopf lift and gauge freedom
+E.2 Hopf lift and gauge freedom
 Fix a constant imaginary unit $i_{0}\in\operatorname{Im}\mathbb{H}$ with $i_{0}^{2}=-1$.
 Since every $u(x,y)\in S^{2}$ is conjugate to $i_{0}$, locally there exists a Hopf lift $q:\Omega\to S^{3}$ (unit quaternions) such that
-$u=q\,i_{0}\,q^{-1}.$ (F.6)
+$u=q\,i_{0}\,q^{-1}.$ (E.6)
 The lift is unique up to the $S^{1}$-gauge transformation
-$q\ \mapsto\ q\,e^{i_{0}\theta(x,y)}.$ (F.7)
-F.3 Maurer–Cartan form and differentiation
+$q\ \mapsto\ q\,e^{i_{0}\theta(x,y)}.$ (E.7)
+E.3 Maurer–Cartan form and differentiation
 Define the left-trivialized Maurer–Cartan components
-$\Omega_{x}:=q_{x}q^{-1},\qquad\Omega_{y}:=q_{y}q^{-1},$ (F.8)
-which lie in $\mathfrak{su}(2)\cong\operatorname{Im}\mathbb{H}$. Differentiating (F.6) and using $(q^{-1})_{x}=-q^{-1}q_{x}q^{-1}$ gives
-$u_{x}=[\Omega_{x},u],\qquad u_{y}=[\Omega_{y},u],$ (F.9)
+$\Omega_{x}:=q_{x}q^{-1},\qquad\Omega_{y}:=q_{y}q^{-1},$ (E.8)
+which lie in $\mathfrak{su}(2)\cong\operatorname{Im}\mathbb{H}$. Differentiating (E.6) and using $(q^{-1})_{x}=-q^{-1}q_{x}q^{-1}$ gives
+$u_{x}=[\Omega_{x},u],\qquad u_{y}=[\Omega_{y},u],$ (E.9)
 where [A,B]=ABBA[A,B]=AB-BA denotes the commutator in $\mathbb{H}$.
-F.3.1 Vertical–horizontal decomposition
+E.3.1 Vertical–horizontal decomposition
 At each point, $\mathfrak{su}(2)\cong\operatorname{Im}\mathbb{H}$ splits as
@@
 Accordingly decompose
-$\Omega_{x}=\Omega_{x}^{\mathrm{ver}}+\Omega_{x}^{\mathrm{hor}},\qquad\Omega_{y}=\Omega_{y}^{\mathrm{ver}}+\Omega_{y}^{\mathrm{hor}},$ (F.10)
-where $\Omega^{\mathrm{ver}}\parallel u$ and $\Omega^{\mathrm{hor}}\perp u$. Since $\Omega^{\mathrm{ver}}$ commutes with $u$, it follows from (F.9) that
-$u_{x}=[\Omega_{x}^{\mathrm{hor}},u],\qquad u_{y}=[\Omega_{y}^{\mathrm{hor}},u].$ (F.11)
-Under the gauge change (F.7), the Maurer–Cartan form changes by adding a vertical term, so the horizontal components and the commutators in (F.11) capture the gauge-invariant variation of $u$.
-F.4 Quaternionic form of the obstruction
-The obstruction $G=i_{x}+i\,i_{y}$ can be rewritten in terms of $u$. Up to the scalar normalizations in (F.4), the intrinsic combination governing deviations from the constant theory corresponds to
-$\mathcal{G}:=u_{x}+u\,u_{y}.$ (F.12)
-Using (F.11), we may express $\mathcal{G}$ in Hopf variables as
-$\mathcal{G}=[\Omega_{x}^{\mathrm{hor}},u]\;+\;u\,[\Omega_{y}^{\mathrm{hor}},u].$ (F.13)
+$\Omega_{x}=\Omega_{x}^{\mathrm{ver}}+\Omega_{x}^{\mathrm{hor}},\qquad\Omega_{y}=\Omega_{y}^{\mathrm{ver}}+\Omega_{y}^{\mathrm{hor}},$ (E.10)
+where $\Omega^{\mathrm{ver}}\parallel u$ and $\Omega^{\mathrm{hor}}\perp u$. Since $\Omega^{\mathrm{ver}}$ commutes with $u$, it follows from (E.9) that
+$u_{x}=[\Omega_{x}^{\mathrm{hor}},u],\qquad u_{y}=[\Omega_{y}^{\mathrm{hor}},u].$ (E.11)
+Under the gauge change (E.7), the Maurer–Cartan form changes by adding a vertical term, so the horizontal components and the commutators in (E.11) capture the gauge-invariant variation of $u$.
+E.4 Quaternionic form of the obstruction
+The obstruction $G=i_{x}+i\,i_{y}$ can be rewritten in terms of $u$. Up to the scalar normalizations in (E.4), the intrinsic combination governing deviations from the constant theory corresponds to
+$\mathcal{G}:=u_{x}+u\,u_{y}.$ (E.12)
+Using (E.11), we may express $\mathcal{G}$ in Hopf variables as
+$\mathcal{G}=[\Omega_{x}^{\mathrm{hor}},u]\;+\;u\,[\Omega_{y}^{\mathrm{hor}},u].$ (E.13)
 This expression is gauge-invariant (it is insensitive to vertical modifications of the lift).
-F.5 Quaternionic rigidity theorem
-Theorem F.1 (Quaternionic rigidity).
-Let $u:\Omega\to S^{2}\subset\operatorname{Im}\mathbb{H}$ be the normalized generator (F.4) associated with an elliptic quadratic structure (F.1).
-Let $q$ be a local Hopf lift (F.6) with Maurer–Cartan components (F.8) and decomposition (F.10).
+E.5 Quaternionic rigidity theorem
+Theorem E.1 (Quaternionic rigidity).
+Let $u:\Omega\to S^{2}\subset\operatorname{Im}\mathbb{H}$ be the normalized generator (E.4) associated with an elliptic quadratic structure (E.1).
+Let $q$ be a local Hopf lift (E.6) with Maurer–Cartan components (E.8) and decomposition (E.10).
 Then the following are equivalent:
@@
 […]
-    $u_{x}+u\,u_{y}=0\quad\text{on }\Omega.$ (F.14)
+    $u_{x}+u\,u_{y}=0\quad\text{on }\Omega.$ (E.14)
   2. (ii)
@@
 $u$ undergoes only fiber (Hopf) gauge rotation along that direction.
-Moreover, (F.14) is the Hopf/quaternionic avatar of the original rigidity condition $G=i_{x}+i\,i_{y}=0$.
+Moreover, (E.14) is the Hopf/quaternionic avatar of the original rigidity condition $G=i_{x}+i\,i_{y}=0$.
 Proof.
-By (F.11), we have $u_{x}=[\Omega_{x}^{\mathrm{hor}},u]$ and $u_{y}=[\Omega_{y}^{\mathrm{hor}},u]$.
-Substituting these into (F.12) yields (F.13), and hence $\mathcal{G}\equiv 0$ is equivalent to (F.15).
-The gauge statement follows from the fact that changing the lift by (F.7) alters only the vertical part of the Maurer–Cartan form, leaving the horizontal commutator expressions unchanged.
+By (E.11), we have $u_{x}=[\Omega_{x}^{\mathrm{hor}},u]$ and $u_{y}=[\Omega_{y}^{\mathrm{hor}},u]$.
+Substituting these into (E.12) yields (E.13), and hence $\mathcal{G}\equiv 0$ is equivalent to (E.15).
+The gauge statement follows from the fact that changing the lift by (E.7) alters only the vertical part of the Maurer–Cartan form, leaving the horizontal commutator expressions unchanged.
 ∎
-F.6 Interpretation
-Equation (F.14) says that the $S^{2}$-valued imaginary direction field $u$ is transported compatibly with its own multiplication rule: along the characteristic direction $\partial_{x}+u\,\partial_{y}$, the field exhibits no genuine horizontal twisting in $S^{2}$, only Hopf fiber rotation.
+E.6 Interpretation
+Equation (E.14) says that the $S^{2}$-valued imaginary direction field $u$ is transported compatibly with its own multiplication rule: along the characteristic direction $\partial_{x}+u\,\partial_{y}$, the field exhibits no genuine horizontal twisting in $S^{2}$, only Hopf fiber rotation.
 In this sense rigidity is a horizontality cancellation condition in the Hopf fibration rather than an analytic assumption.
-Appendix G Rigidization as a Computational Normal Form
+Appendix F Rigidization as a Computational Normal Form
 Purpose of this Appendix
@@
 Rigidization removes exactly the analytic obstructions that prevent the use of the constant-coefficient toolbox, while preserving geometric variability that is irrelevant for computation but expensive to eliminate.
-G.1 Rigidization versus Uniformization
+F.1 Rigidization versus Uniformization
 Let $\lambda(x,y)$ denote the spectral slope of a variable elliptic structure, with intrinsic obstruction
@@
 This is a strictly weaker requirement than uniformization. Nevertheless, once rigidity is achieved:
-  • the Leibniz defect vanishes,
   • the generalized Cauchy–Riemann system becomes homogeneous,
@@
 Thus rigidization reaches the minimal normal form required for analysis, without over-solving the geometric problem.
-G.2 Why “Weaker” Means “Easier”
+F.2 Why “Weaker” Means “Easier”
 The rigidization condition is first-order in the structure and second-order in the coordinates, and its linearization is an elliptic system whose principal symbol is explicitly computable.
@@
 From a solver perspective, rigidization is therefore cheaper, more local, and better aligned with problems that already possess a preferred transport geometry.
-G.3 A Solver-Oriented Triage
+F.3 A Solver-Oriented Triage
 Rigidization is intended to be used selectively. Three intrinsic indicators guide solver choice:
@@
 This policy is not ad hoc: it reflects the analytic role played by the obstruction in the failure of derivation properties, even though the specific thresholds and decision boundaries are solver-dependent.
-G.4 Rigidization as a Preconditioner
+F.4 Rigidization as a Preconditioner
 Even when rigidization does not succeed globally, it can still be valuable.
@@
 More concretely, the incremental continuity strategy described above provides a practical algorithm even without a closed-form global existence theorem: one advances in $t$ until the Newton step fails to converge, records the partially rigidized structure, and hands the residual obstruction to a uniformization solver operating on a problem whose effective $\rho_{T}$ has been reduced.
-G.5 The Global Rigidization Problem
+F.5 The Global Rigidization Problem
 It is natural to ask whether rigidization can always be achieved on compact domains, without any smallness condition on the obstruction $T$.
@@
 We record it here both as an invitation to further work and as context for the solver-oriented approach of this appendix: the computational triage and preconditioner strategies above are designed to be effective regardless of whether Conjecture 1 is eventually proved.
-G.6 What Is Not Claimed
+F.6 What Is Not Claimed
 To be explicit about the boundaries of the current theory, this monograph does not claim:
@@
 The global theory may well be within reach—the structural ingredients (ellipticity of the linearized problem, Burgers characteristic geometry, similarity-principle arguments) are all present—but assembling them into a complete a priori package requires new ideas, particularly for the control of the Jacobian.
-G.7 Summary
+F.7 Summary
 Rigidization should be viewed as:
@@
 In this sense, rigidization is not a competitor to uniformization but a practical intermediary between raw variable geometry and full geometric flattening—one whose global reach remains a compelling open question in its own right.
-Appendix H Symbolic Verification of the Rigid Second–Order Expansion
+Appendix G Symbolic Verification of the Rigid Second–Order Expansion
 This appendix contains a complete symbolic verification of Theorem 8.1 using the computer algebra system SymPy.
@@
 # --- end script ---
-Appendix I Symbolic Verification for the Rigidity-Flatness Theorem Condition
+Appendix H Symbolic Verification for the Rigidity-Flatness Theorem Condition
 Listing 2: SymPy verification of Theorem 8.1